### v1

# SLOPE and Designing Robust Studies for Generalization

Abstract
A popular task in generalization is to learn about a new, target population based on data from an existing, source population. This task relies on conditional exchangeability, which asserts that differences between the source and target populations are fully captured by observable characteristics of the two populations. Unfortunately, this assumption is often untenable in practice due to unobservable differences between the source and target populations. Worse, the assumption cannot be verified with data, warranting the need for robust data collection processes and study designs that are inherently less sensitive to violation of the assumption. In this paper, we propose SLOPE (Sensitivity of LOcal Perturbations from Exchangeability), a simple, intuitive, and novel measure that quantifies the sensitivity to local violation of conditional exchangeability. SLOPE combines ideas from sensitivity analysis in causal inference and derivative-based measure of robustness from Hampel, (1974). Among other properties, SLOPE can help investigators to choose (a) a robust source or target population or (b) a robust estimand. Also, we show an analytic relationship between SLOPE and influence functions (IFs), which investigators can use to derive SLOPE given an IF. We conclude with a re-analysis of a multi-national randomized experiment and illustrate the role of SLOPE in informing robust study designs for generalization.
Keywords: Generalizability; Conditional exchangeability; Sensitivity analysis; Causal inference; Influence Function; Exponential Tilting
1 Introduction
1.1 Motivation and Overview
There has been a growing interest in generalizing information from a source population to a new, target population under the assumption of conditional exchangeability, which posits that distributions in the two populations are identical after conditioning on a set of populations’ shared characteristics. Formally, suppose each datum is represented as random variables $(O,X)$ where $X$ denotes the shared characteristics between the two populations. The goal is to learn the target distribution $Q_{O,X}$ given (i) the “full” data $(O,X)$ from the source distribution $P_{O,X}$ and (ii) the “partial” data $X$ about the shared characteristics in the target distribution $Q_{X}$. Conditional exchangeability states
$Q_{O\mid X}(\cdot\mid X=x)=P_{O\mid X}(\cdot\mid X=x)\text{ almost everywhere in $Q_{X}$. }$ (1)
Under (1), we can learn the target distribution $Q_{O,X}$ via a heuristic yet simple equality
$Q_{O,X}=Q_{O\mid X}\times Q_{X}=P_{O\mid X}\times Q_{X}.$
The first equality is from the definition of conditional probability and the second equality is from (1). Also, when we are interested in a feature/functional of the target distribution $Q_{O,X}$, denoted as $\psi(Q_{O,X})$ and referred to as the target estimand (e.g., mean or median of $O$ in the target population), we can identify the estimand under (1) by $\psi(Q_{O,X})=\psi(P_{O\mid X}\times Q_{X})$. For concrete examples, see Section 2.1.
Unfortunately, recent works (Allcott,, 2015; Jin et al.,, 2024) have argued that conditional exchangeability is likely violated in practice, potentially threatening the validity of several empirical works on generalization and transfer learning. Worse, conditional exchangeability is inherently untestable because we never observe $O$ from the target distribution $Q$ (Dahabreh et al.,, 2023; Huang,, 2024); as mentioned above, we only know $P_{O,X}$ and $Q_{X}$. Taken together, before analyzing data for generalization, it’s essential to have robust data collection processes and study designs that are inherently less sensitive to violation of the assumption.
To this end, this paper proposes a simple, interpretable, and novel tool that helps gauge which study designs are robust to violations of conditional exchangeability and we present a high-level summary of the tool and its important caveats. Let $Q_{O\mid X}^{\gamma}$ be the distribution of $Q_{O\mid X}$ when conditional exchangeability is violated by a degree quantified by a sensitivity parameter $\gamma\in\mathbb{R}$ and $\gamma=0$ implies conditional exchangeability holds (i.e., $Q_{O\mid X}^{0}=P_{O\mid X}$). Then, we propose a metric called SLOPE, which stands for Sensitivity of LOcal Perturbations from Exchangeability:
${\rm SLOPE}(Q_{O,X}^{0},\psi)=\lim_{\gamma\to 0}\frac{\psi(Q_{O,X}^{\gamma})-\psi(Q_{O,X}^{0})}{\gamma},\quad{}\text{ where }Q_{O,X}^{\gamma}=Q_{O\mid X}^{\gamma}\times Q_{X}.$
Figure 1 provides a visual illustration of SLOPE. As the name hints, SLOPE is the slope of the tangent of the target estimand $\psi(Q_{O,X}^{\gamma})$ as a function of $\gamma$ at $\gamma=0$. In other words, SLOPE measures the change in the target estimand under $Q_{O,X}^{\gamma}$ (i.e., when conditional exchangeability doesn’t hold) and $Q_{O,X}^{0}$ (i.e., when conditional exchangeability holds) when there is a near-violation of conditional exchangeability. Generally speaking, a higher magnitude of SLOPE suggests that the estimand changes more dramatically to local violations of conditional exchangeability and a lower magnitude of SLOPE suggests otherwise. More broadly, SLOPE aligns with principles in robust statistics, where “robustness signifies insensitivity to small deviations from the assumptions” (Huber,, 1981, p. 1), where we added the emphasis on “small deviations”.
SLOPE depends on two quantities: (a) the target estimand of interest (i.e., $\psi$) and (b) the target distribution under conditional exchangeability (i.e., $Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$). Importantly, SLOPE does not depend on choices of any estimation procedure of $\psi$. To put it differently, SLOPE is an intrinsic, design-level characteristic about (a) the estimand/functional (i.e., $\psi$) and (b) the distributions that are involved in generalization (i.e., $P_{O\mid X}$ and $Q_{X}$) when there is a violation of conditional exchangeability.
The simplicity of SLOPE comes at some costs, which we briefly summarize here and discuss in detail in Section 3.1. First, as mentioned above, SLOPE is a local measure and for small deviations of $\gamma$ from 0, SLOPE is an accurate reflection about the change in $\psi$. For larger deviations (i.e., $\gamma$ is far from 0), SLOPE may still provide valuable intuition, subject to the usual limitations of linear approximations based of the tangent line. We also remark that some well-known robustness measures are local, including Hampel, (1974)’s influence function, which has taken a central role in robust statistics and estimation (e.g., for causal parameters). Second, the unit of SLOPE inherits the unit of the target estimand. While this fact is rather obvious from SLOPE’s definition since SLOPE measures the change in the target estimand, which generally has scientifically meaningful units, when interpreting the magnitude of SLOPE, we remind investigators to exercise the same level of caution as when they interpret the units of slopes in other areas of statistics (e.g., regression models). Third, SLOPE depends on the parametrization of $Q_{O\mid X}^{\gamma}$ or equivalently, the sensitivity model for conditional exchangeability. While our parametrization of $Q_{O\mid X}^{\gamma}$ has some technical and non-technical benefits, for instance (a) analytic derivation of SLOPE only requires basic tools from calculus, (b) the uniqueness of the tangent curve, especially compared to “bound-based” sensitivity models (see Remark 3), and (c) interpretability (see Section 3), it is not the only parametrization of $Q_{O\mid X}^{\gamma}$; see Rosenbaum, (1987), Tan, (2006), Ding and VanderWeele, (2016), Oster, (2019), and Cinelli and Hazlett, (2020) for other models of sensitivity analysis. Irrespective of the choice of the sensitivity model, we believe the high-level idea of SLOPE as a simple, derivative-based summary of a sensitivity analysis can provide new and important insights about designing robust studies for generalization.
Figure 1: Left: the data structure of our setup. Right: a visual illustration of SLOPE. The $y$-axis plots the target estimand $\psi(\cdot)$ and the x-axis plots $\gamma$ which represents the violation of conditional exchangeability; $\gamma=0$ indicates that conditional exchangeability holds.
1.2 Prior Works
As mentioned in our summary above, SLOPE and Hampel, (1974)’s influence function (IF) are related in that both SLOPE and IF measure local sensitivity (or robustness) and use derivatives to quantify sensitivity (robustness). But, there is an important difference between the two measures. Specifically, Hampel, (1974)’s IF measures a local change in an estimator due to contamination of a single data point whereas SLOPE measures a local change in an estimand due to violation of conditional exchangeability. Despite these differences, SLOPE and IF are analytically related. Section 4.1 formalizes this insightful relationship between the two local measures of robustness.
Our work fits into the broader literature on sensitivity analysis of conditional exchangeability for generalizability and transportability (Nguyen et al.,, 2017; Nie et al.,, 2021; Colnet et al.,, 2021; Dahabreh et al.,, 2022, 2023; Duong et al.,, 2023; Ek and Zachariah,, 2023; Huang,, 2024; Jin et al.,, 2024). Notably, our paper uses the exponentially tilting sensitivity model in Dahabreh et al., (2022, 2023), which was originally proposed by Robins et al., (2000) and used by others (e.g., Rotnitzky et al., (2001); Troxel et al., (2004); Birmingham et al., (2003); Franks et al., (2020); Nabi et al., (2024); Zhou and Yao, (2023)). What differentiates our paper from existing works on sensitivity analysis in this area is that existing works studied sensitivity of estimators of $\psi$. In contrast, our work studies sensitivity of study designs; as mentioned in our summary above, SLOPE is agnostic to how $\psi$ is estimated and is a population-level characteristic.
Outside of generalizability and transportability, there are several works on summary measures about the violation of key assumptions in their respective fields and we highlight some works. Gupta and Rothenhäusler, (2023) proposed the directional s-value which quantifies the minimum amount of covariate shift that alters the sign of the target parameter (usually the mean). Rosenbaum, (2004) proposed design sensitivity $\widetilde{\Gamma}\in(1,\infty)$, which is a scalar, odds-ratio based summary of a test statistic’s power under unmeasured confounding in an observational study. Andrews et al., (2017) considered local mis-specification of generalized method of moments (GMM) and their matrix-based measure, denoted as $\Lambda\in\mathbb{R}^{p\times p}$, quantified changes of the estimators for GMM parameters. In missing data, Troxel et al., (2004) proposed a measure, which quantifies the change in an estimator to a local violation of ignorable missingness. Troxel et al., (2004)’s measure is, loosely speaking, the “estimator” analog of SLOPE in missing data. Oster, (2019), Ding and VanderWeele, (2016), Cinelli and Hazlett, (2020), and Zhao, 2019b proposed scalar measures that summarize the impact of unmeasured confounding in observational studies. Specifically, each proposed metrics that quantify the “minimum unmeasured confounding bias” necessary to alter the original study conclusion under conditional exchangeability (e.g., the sensitivity bound does not contain the null effect). All these works, except for Gupta and Rothenhäusler, (2023) and Rosenbaum, (2004) to some extent, measure robustness of estimators or tests rather than that of estimands or study designs.
Finally, another related line of work assesses generalizability by comparing shared characteristics (i.e., $X$) of the two populations (Stuart et al.,, 2011; Tipton,, 2014; Chen et al., 2023b, ). These works are developed under the assumption of conditional exchangeability, whereas we assess robustness when the assumption is subject to violation. Also, while these works focus on generalizing the average treatment effect, our framework is applicable to a wide range of functionals/estimands that include the average treatment effect as a special case.
1.3 Organization and Notation
The rest of the manuscript is organized as follows. Section 2 introduces the setup and the sensitivity model. Section 3 formally introduces SLOPE, its properties, examples, and illustrations on using SLOPE for robust study designs. Section 4 discusses further insights and results, including the relationship between SLOPE and IF, estimation of SLOPE, and other sensitivity models. Section 5 demonstrates the application of SLOPE in robust designs through a re-analysis of a multi-site experiment by Banerjee et al., (2015). Section 6 concludes with a discussion on future directions and practical considerations. Technical proofs and auxiliary results are deferred to the Supplement.
We define the notations that we use throughout the paper. For a probability measure (i.e., a population) $P$ and random vectors $X_{1}$, $X_{2}$, we let $P_{X_{j}}$ be the marginal distribution of $X_{j}$ for $j=1,2$. Given $P_{X_{j}}$, we let $\textnormal{E}_{P_{X_{j}}}(\cdot)$ and $F_{P_{X_{j}}}$ be the expectation and cumulative distribution function (c.d.f.), respectively, under $P_{X_{j}}$. We let $P_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$ be the conditional distribution of $X_{1}$ given $X_{2}$. Similarly, we let $\textnormal{E}_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$, $F_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$, $f_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$ be the conditional distribution, conditional c.d.f., and conditional probability density function, respectively. We also let $\textnormal{E}_{X_{1}\mid X_{2}}(\cdot\mid x_{2})$ be the conditional expectation given a specific $X_{2}=x_{2}$. Throughout the paper, we assume sufficient regularity conditions for conditional distributions and conditional densities to exist; see Shao, (2008, Chapter 1) for the regularity conditions. Finally, for a probability distribution $Q$ defined on the same measurable space of $(X_{1},X_{2})$, we use $P_{X_{1}\mid X_{2}}\times Q_{X_{2}}$ to denote the joint distribution where $P_{X_{1}\mid X_{2}}\times Q_{X_{2}}(A\times B)=\int_{B}P_{X_{1}\mid X_{2}}(A\mid x_{2})dQ_{X_{2}}(x_{2})$ for $A\in\mathcal{B}$ and $B\in\mathcal{B}^{d}$ and $\mathcal{B}$ are Borel $\sigma$-fields with respect to the reals.
2 Setup
2.1 Goal and Key Assumptions in Generalization
Let $P_{O,X}$ and $Q_{O,X}$ be joint distributions of random variables $(O,X)$ from a source population $P$ and a target population $Q$, respectively, where $O$ is a scalar and $X$ is a vector. Suppose $\psi(\cdot)$ is a scalar-valued functional of distributions that extracts a quantity of scientific interest from the distribution. The goal is to learn the functional (i.e., $\psi$) of the target distribution $Q_{O,X}$, which we denote as $\psi(Q_{O,X})$ and refer to as the target estimand, from the (a) “full” data $(O,X)$ from the source distribution $P_{O,X}$ and (b) “partial” data $X$ from the target distribution $Q_{X}$. See Figure 1 for a visual illustration of the data setup.
We make two remarks on our setup. First, while our exposition considers a scalar functional $\psi$ (e.g., means, medians, average potential outcomes), most of our results extend to a vector valued $\psi$; see Section J.1 of the Supplement for a discussion. Moreover, with a user specified vector norm, the practical interpretation of SLOPE based on its magnitude will also extend; see Section 3.1. Second, almost all results below are agnostic to how the data were sampled within each population, for instance by simple random sampling, i.i.d. sampling, or even adaptive sampling. Specifically, our results remain at the population level until Section 4.2, where we propose an estimation procedure for SLOPE.
Under the setup, the two most popular assumptions to identify the target estimand $\psi(Q_{O,X})$ (e.g., Cole and Stuart, (2010); Tipton, (2014); Kern et al., (2016); Dahabreh et al., (2019); Huang et al., (2023); Zeng et al., (2023); Degtiar and Rose, (2023)) are as follows.
Assumption 1 (Overlap).
$Q_{X}$ is absolute continuous with respect to $P_{X}$.
Assumption 2 (Conditional Exchangeability).
$Q_{O\mid X}(\cdot\mid X)$ is absolute continuous with respect to $P_{O\mid X}(\cdot\mid X)$ and the Radon-Nikodym derivative satisfies $dQ_{O\mid X}(O,X)/dP_{O\mid X}(O,X)=1$ almost everywhere $P_{O\mid X}\times Q_{X}$.
Remark 1 (Alternative Formulation of Assumption 2).
We note that Assumption (2) is equivalent to equation (1) stated in Section 1. Further, when the density of $P_{O\mid X}$ exists and the density of $Q_{O\mid X}$ is assumed to exist, then Assumption 2 can be re-formulated as a condition on the density ratio, i.e., $f_{Q_{O\mid X}}(O,X)/f_{P_{O\mid X}}(O,X)=1$ almost everywhere $P_{O\mid X}\times Q_{X}$. For presentation purposes, we keep this density ratio representation in later sections, while keeping in mind that this can be formulated as a Radon-Nikodym derivative as in Assumption 2.
Assumption 1 states that the support of $Q_{X}$ is within the support of $P_{X}$. Assumption 2 enables the identification of $\psi(Q_{O,X})$ by replacing the unidentified $Q_{O\mid X}$ with the identified $P_{O\mid X}$. Under Assumptions 1 and 2, the target estimand can be identified as $\psi(Q_{X}\times P_{O\mid X})$. We provide some examples of target estimands and their identification below.
Example 1 (Mean).
Suppose we are interested in the mean of $O$ in the target distribution, denoted as $\psi^{{\textnormal{mean}}}(Q_{O})=\textnormal{E}_{Q_{O}}(O)$. Under Assumptions 1 and 2, the mean is identified via
$\psi^{{\textnormal{mean}}}(Q_{O})=\textnormal{E}_{Q_{O}}(O)=$ $\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{Q_{O\mid X}}(O)\right]=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O)\right].$
Example 2 (Median).
Suppose we are interested in the median of $O$ in the target distribution, denoted as $\psi^{{\textnormal{med}}}(Q_{O})=F_{Q_{O}}^{-1}(1/2)$, and $O$ is continuous. Under Assumptions 1 and 2, the median is identified as the solution to the following equation:
$\frac{1}{2}=\int\int_{-\infty}^{\psi^{{\textnormal{med}}}}dQ_{O\mid X}dQ_{X}=\int\int_{-\infty}^{\psi^{{\textnormal{med}}}}dP_{O\mid X}dQ_{X}.$
Example 3 (Z-Functionals).
Suppose the target estimand $\psi(Q_{O,X})$ is defined as the solution to the equation
$\textnormal{E}_{Q_{O,X}}\{s\left(O,X,\psi(Q_{O,X})\right)\}=0,$ (2)
where $s(O,X,\cdot)$ is a user-specified function, usually a score function, of the same dimension with $\psi$. Under Assumptions 1 and 2, $\psi$ is identified as the solution to the following equation:
$0=\textnormal{E}_{Q_{O,X}}\{s(O,X,\psi)\}=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{Q_{O\mid X}}\left\{s(O,X,\psi(Q_{O\mid X}\times Q_{X}))\right\}\right]=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\left\{s(O,X,\psi(Q_{O,X}^{0}))\right\}\right].$
Example 4 (Mean or Median of Potential Outcomes).
Suppose a randomized experiment has been conducted in the source population to measure the causal effect of treatment on an outcome. Let $Y(a)$ be the potential outcome if a study unit was assigned to treatment level $a\in\{0,1\}$ and $X$ be the pre-treatment covariates. The goal is to use this randomized experiment to learn about the causal effect in a new, target population based on the target population’s covariates $X$. Under Assumptions 1 and 2, the mean and the median of the potential outcome $Y(a)$ are identified via
$\psi^{{\textnormal{mean}}}(Q_{Y(a)})=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{Y(A)\mid X}}\{Y(A)\}\right],\text{ and }\quad{}\frac{1}{2}=\int\int_{-\infty}^{\psi^{{\textnormal{med}}}}dP_{Y(A)\mid X}dQ_{X}.$
We note that both equations involve potential outcomes (i.e., $P_{Y(a)\mid X}$), which is identified under a randomized experiment; see Section D in the Supplement for details.
2.2 Model for Sensitivity Analysis of Conditional Exchangeability
The identification of the target estimand depends on Assumptions 1 (i.e., overlap) and 2 (i.e., conditional exchangeability). While Assumption 1 can be empirically verified with observed data (i.e., $X$ from $P_{X}$ and $Q_{X}$), Assumption 2 cannot be verified since it concerns unobserved data, specifically $O$ in the target population $Q$. Faced with the inability to check the assumption, a sensitivity analysis seeks to study how conclusions from a study may change if Assumption 2 is violated by a certain amount. In this section, we describe a model for sensitivity analysis, which motivates how we define SLOPE.
Suppose conditional exchangeability (i.e., Assumption 2) no longer holds and instead, the distribution $Q_{O\mid X}$ is shifted away from $P_{O\mid X}$ by an exponential tilting term. Specifically, we consider Robins’ sensitivity model (Robins et al.,, 2000) where $Q_{O\mid X}^{\gamma}$ is the exponentially tilted distribution of $Q_{O\mid X}$. Suppose $Q_{O\mid X}^{\gamma}(\cdot\mid X)$ is absolute continuous with respect to $P_{O\mid X}(\cdot\mid X)$ almost everywhere $Q_{X}$ and their corresponding densities $f_{Q_{O\mid X}^{\gamma}}(O,X)$ and $f_{P_{O\mid X}}(O,X)$ satisfy
$\dfrac{f_{Q_{O\mid X}^{\gamma}}(O,X)}{f_{P_{O\mid X}}(O,X)}\propto\exp(\gamma\cdot O)\text{, almost everywhere }P_{O\mid X}\times Q_{X}.$ (3)
The notation $\propto$” means “proportional to” and $\int\exp(\gamma O)dP_{O\mid X}(O\mid X)<\infty$ almost everywhere $Q_{X}$. The term $\gamma$ is a scalar sensitivity parameter that quantifies the difference between $Q_{O\mid X}$ and $P_{O\mid X}$. If $\gamma=0$, the sensitivity model (3) reduces to Assumption 2 where the two distributions are identical. As $\gamma$ moves away from zero, Assumption 2 is violated by an amount quantified by $\gamma$.
Under Assumption 1 and the sensitivity model (3), given a value of $\gamma$, the target estimand can be identified as $\psi(Q_{O,X}^{\gamma})$ where $Q_{O,X}^{\gamma}=Q_{O\mid X}^{\gamma}\times Q_{X}$ represents the joint distribution of $Q_{O\mid X}^{\gamma}$ induced from (3) and $Q_{X}$. Since sensitivity parameters (e.g., $\gamma$) are inherently unidentifiable, in a sensitivity analysis, researchers specify a range of values for the sensitivity parameter $\gamma$ based on domain knowledge (Scharfstein et al.,, 1999; Rotnitzky et al.,, 2001; Nabi et al.,, 2024), benchmarking (Huang,, 2024) or calibration (Miao et al.,, 2024) and investigate how the target estimand changes between the case with $\gamma=0$ (i.e., $\psi(Q_{O,X}^{0})$ under conditional exchangeability) and cases where $\gamma\neq 0$ (i.e., $\psi(Q_{O,X}^{\gamma})$) with a specified range of $\gamma$ that violates conditional exchangeability.
We make some remarks about the sensitivity model (3). First, model (3) was proposed by Scharfstein et al., (1999) and Robins et al., (2000) as a non-parametric (just) identified model for describing selection bias in missing data. This sensitivity model was later referred to as an exponential tilting model (Rotnitzky et al.,, 2001; Birmingham et al.,, 2003) and an extrapolation-factorization model (Linero and Daniels,, 2018). The model was also used to conduct sensitivity analysis for unmeasured confounding in causal inference (Franks et al.,, 2020; Nabi et al.,, 2024) and sensitivity analysis for transportability (Dahabreh et al.,, 2022; Miao et al.,, 2024). Second, model (3) can be reformulated as a selection model or under some assumptions, the sensitivity parameter $\gamma$ can be reparameterized to pseudo-$R^{2}$ (Franks et al.,, 2020). Third, model (3) can be extended so that $\gamma$ depends on $X$ and $O$; see Remark 7 of the Supplement. Fourth, an important caveat of using model (3) is the non-collapsibility of the model with respect to $X$ as the model implies logistic selection model; see Section 7 of Scharfstein et al., (1999).
Lastly and importantly, we choose model (3) as the model (i) posits no testable implications on the data, (ii) makes our proposed measure SLOPE tractable in terms of having a unique tangent curve (see Section 4.3), and (iii) has an analytic connection to the influence function (see Section 4.1). In general, reflecting on the discussions by Robins, (2002) and Rosenbaum, (2002), we believe there is no “optimal” model for sensitivity analysis and depending on the scientific questions at hand, some models for sensitivity analysis are more useful than others. In our case, as we show below, model (3) is useful to define SLOPE and to inform robust study designs in generalizations. A further discussion on other sensitivity models is provided in Section 4.3.
3 SLOPE: Sensitivity to Local Perturbation from Exchangeability
3.1 Definition and Basic Properties
Under sensitivity model (3), $\psi(Q_{O,X}^{\gamma})$ can be viewed a function that maps the sensitivity parameter $\gamma$ to the target estimand. Therefore, the fluctuation of this function along $\gamma$ characterizes how much the target estimand changes when the conditional exchangeability is violated. This observation motivates us to introduce our metric SLOPE, which measures the change in the target estimand along $\gamma$ by taking a derivative with respect to $\gamma$.
Definition 1 (SLOPE).
The sensitivity to local perturbation from exchangeability (SLOPE) of a target functional $\psi(\cdot)$ with respect to the sensitivity model in (3) is defined as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\lim_{\gamma\to 0}\dfrac{\psi(Q_{O,X}^{\gamma})-\psi(Q_{O,X}^{0})}{\gamma},$ (4)
provided the limit exists.
[…]
When comparing the magnitudes of SLOPE, investigators should be aware that the unit of SLOPE inherits the unit of the target estimand/functional. For instance, in Examples 1 and 2 for the mean $\psi^{{\textnormal{mean}}}$ and the median $\psi^{{\textnormal{med}}}$, the SLOPE shares the unit of $O$. While this fact is rather obvious from SLOPE’s definition since it measures the change in the target estimand and the target estimand should carry scientifically meaningful units, we remind investigators to exercise the same level of caution when interpreting SLOPE’s magnitude akin as when investigators interpret the units of slopes in other areas of statistics (e.g., in regression models).
An important property about SLOPE is that it does not depend on any estimation procedure. Instead, SLOPE measures an intrinsic, robustness property about a study design when there is violation of conditional exchangeability. Specifically, SLOPE depends on two design level quantities: (a) the target functional/estimand $\psi(\cdot)$, and (b) the target distribution under conditional exchangeability ($Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$). For scientific researchers, SLOPE reflects their answers to two questions: (a) “what quantity would you like to learn?” and (b) “what type of dataset would you use to learn the target quantity?”. Changing either (a) or (b) can lead to a different value of SLOPE, reflecting a different level of robustness of the underlying study design. Consequently, SLOPE can be used to choose which estimand (e.g., the mean or the median), target distribution (e.g., $Q_{X}$), or the source distribution (e.g., $P_{O,X}$) is less sensitive for generalization, thereby informing choices of a robust study design. See Section 3.2 for an illustration and Section 5 for real data examples. Hereinafter, we may refer to the choices of target functionals (e.g., choice between $\psi^{{\textnormal{mean}}}$ and $\psi^{{\textnormal{med}}}$) as the choices of target estimands (e.g., choice between $\psi^{{\textnormal{mean}}}(Q_{O})$ and $\psi^{{\textnormal{med}}}(Q_{O})$) when the context is clear, since the term “estimand” is more familiar to practitioners.
When communicating the meaning of SLOPE, some audience may find it useful to interpret SLOPE as a measure of the “first-order change” of the estimand when conditional exchangeability is violated. Specifically, a first-order Taylor expansion of the target estimand yields
$\psi(Q_{O,X}^{\gamma})-\psi(Q_{O,X}^{0})\approx\gamma\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi).$ (5)
The Taylor expansion in (5) suggests that for a small $\gamma$ that is near zero, the SLOPE provides an accurate measure of the change in $\psi$ when conditional exchangeability is violated. When $\gamma$ is large in magnitude, SLOPE may still provide some intuition about the bias, with the usual limitations of first-order linear approximations. We remark, however, that SLOPE cannot identify the bias of an estimator of $\psi$ from violating conditional exchangeability since, as mentioned in Section 2.2, $\gamma$ and $Q_{O,X}^{\gamma}$ are not identified.
Finally, we remark that SLOPE does not always exist for every target estimand. For example, suppose $\psi$ is the sign of the mean of $O$ in the target population. Then, the limit in Definition 1 may not be well-defined when the sign changes near $\gamma=0$. One general condition for SLOPE to exist is to satisfy the conditions for the chain rule under Hadamard differentiability (Theorem 20.9 of Van der Vaart, (2000)); see Section A of the Supplement for the precise conditions. For Z-functionals defined through (2), a sufficient condition for the existence of SLOPE is guaranteed by the following smoothness and boundndess conditions that are typical to establish consistency of Z estimators.
Condition 1 (Existence of SLOPE for Z-Functionals).
(i) $\textnormal{E}_{Q_{O\mid X}^{\gamma}}[s(O,X,\psi(Q_{O,X}^{0})]$ is bounded by an integrable function almost surely $Q_{X}$ for $\gamma$ in a neighborhood of zero, and $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi(Q_{O,X}^{0})\left\{O-\mu(X)\right\}\right]$ exists; (ii) $s(O,X,\cdot)$ is differentiable almost everywhere and denote the derivative as $\dot{s}(O,X,\cdot)$; and (iii) $\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi(Q_{O,X}^{0}))\}$ exists and is nonsingular.
3.2 Example: SLOPE for the Mean and the Robustness of a Homogeneous Source Population
To gain intuition, we illustrate the SLOPE when the target estimand is the mean, i.e., $\psi^{{\textnormal{mean}}}$ in Example 1 with $\psi^{\textnormal{mean}}(Q_{O,X})=\textnormal{E}_{Q_{O}}(O)$. Given $\gamma$, sensitivity model (3) implies
$\psi^{{\textnormal{mean}}}(Q_{O,X}^{\gamma})=\textnormal{E}_{Q_{X}}\left[\dfrac{\textnormal{E}_{P_{O\mid X}}\left\{O\exp(\gamma O)\mid X\right\}}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\mid X\right\}}\right].$
Then SLOPE for the mean is the derivative of $\psi^{{\textnormal{mean}}}(Q_{O,X}^{\gamma})$ with respect to $\gamma$, evaluated at $\gamma=0$. Theorem 1 formally states the result under regularity conditions to make this derivative possible.
Theorem 1 (SLOPE for Mean).
Suppose Condition 1 holds with $s(O,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$. Then SLOPE for the mean is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\},$ (6)
where $\sigma^{2}(X)=\textnormal{Var}_{P_{O\mid X}}(O\mid X)$ is the conditional variance in the source population.
In words, the SLOPE of the mean is an average of $\sigma^{2}(X)$, the conditional variance of $O$ under $P_{O\mid X}$, over the target population (i.e., $Q_{X}$). When $\sigma^{2}(X)$ is constant across $X$ (i.e., homoskedastic), the mean’s SLOPE only depends on the marginal variance of $O$ in the source population. In other words, under homoskedasticity, SLOPE suggests that a source population with a small variance of $O$ is less sensitive to local violation of exchangeability.
This theoretical insight agrees with existing empirical recommendations for generalizing treatment effects, i.e., under the potential outcome framework in Example 4 with O=Y(1)Y(0)O=Y(1)-Y(0) representing an individual treatment effect. As noted from Tipton and Olsen, (2018), a more homogeneous individual treatment effect (say $O$ is constant in $P$) is generally more robust for generalization than a heterogeneous treatment effect, given all else is equal (i.e., sufficient overlap and covariate balance between the two populations). In Corollary 1, we present one simple mathematical formalization of the empirical insights from Tipton and Olsen, (2018) by using our proposed SLOPE.
Corollary 1 (SLOPE for Robust Source Distributions).
Fix the target distribution $Q_{X}$ and consider two source populations $P$ and $\widetilde{P}$ that satisfy $\sigma^{2}(X)\leq\widetilde{\sigma}^{2}(X)$ almost surely $Q_{X}$, where $\sigma^{2}(X)=\textnormal{Var}_{P_{O\mid X}}(O\mid X)$ and $\widetilde{\sigma}^{2}(X)=\textnormal{Var}_{\widetilde{P}_{O\mid X}}(O\mid X)$. Then
$\textnormal{SLOPE}(P_{O\mid X}\times Q_{X},\psi^{{\textnormal{mean}}})\leq\textnormal{SLOPE}(\widetilde{P}_{O\mid X}\times Q_{X},\psi^{{\textnormal{mean}}}),$
where the equality holds if and only if $\sigma^{2}(X)=\widetilde{\sigma}^{2}(X)$ almost surely $Q_{X}$.
To better illustrate Corollary 1, consider a toy example where $X$ is binary with values $x_{1}$ and $x_{2}$. Condition on $X$, suppose the source is normal with $P_{O\mid X=x_{1}}\sim N(\mu_{1},\sigma_{1}^{2})=N(0,0.5^{2})$ and $P_{O\mid X=x_{2}}\sim N(\mu_{2},\sigma_{2}^{2})$. Fixing $\mu_{2}=0.3$, top panels in Figure 2 show that as $\sigma_{2}^{2}$ increases (i.e., the conditional variance increases), the SLOPE for the mean increases (i.e., the sensitivity increases).
Figure 2: SLOPE of the mean in a toy example with two source distributions, $P_{O\mid X=x_{1}}\sim N(0,0.5^{2})$ and $P_{O\mid X=x_{2}}\sim N(\mu_{2},\sigma_{2}^{2})$. The two left panels fix $\mu_{2}=0.3$ and vary $\sigma_{2}$ and the right panels fix $\sigma_{2}=0.6$ and vary $\mu_{2}$, each with two target populations. Each panel plots the SLOPE in the y-axis against the data generating parameter ($\sigma_{2}$ or $\mu_{2}$) in the x-axis, for SLOPEs of the mean (circles on solid lines) and the median (crossmarks on dotted lines).
Similarly to Corollary 1, for a fixed source population, Corollary 2 states that one target population is more robust than another if the former has a lower measure on covariate values that correspond to a high conditional variance. More concretely, consider again the toy example with $\sigma_{2}=0.6>0.5=\sigma_{1}$. Then $x_{1}$ and $x_{2}$ can be viewed as two subgroups where units with $X=x_{2}$ are more heterogeneous. Consider two target populations $Q$ and $\widetilde{Q}$ where $Q_{X}(x_{2})=0.5$ and $\widetilde{Q}_{X}(x_{2})=0.8$. Since $_{X}(x_{2})<\widetilde{Q}_{X}(x_{2}), QQ$ has a lower proportion of units from the more heterogeneous subgroup than $\widetilde{Q}$, and thus the SLOPE under $Q$ is lower than the SLOPE under $\widetilde{Q}$. As illustrated in Figure 2, when $\sigma_{2}>\sigma_{1}$, the top left panel (i.e., with $Q$) has a smaller SLOPE than the top right panel (i.e., with $\widetilde{Q}$), and vice versa.
Corollary 2 (SLOPE for Robust Target Distributions).
Fix the source distribution $P_{O,X}$ and consider two target distributions $Q_{X}$ and $\widetilde{Q}_{X}$ over a common support $\mathcal{S}_{X}$ such that there exists a subset $\mathcal{S}_{X,1}\in\mathcal{S}_{X}$ that satisfies $Q_{X}(\mathcal{S}_{X,1})\leq\widetilde{Q}_{X}(\mathcal{S}_{X,1})$. If there exists a constant $c$ such that $\sigma^{2}(X)<c$ for $X\in\mathcal{S}_{X,1}$ and $\sigma^{2}(x)>c$ for $x\not\in\mathcal{S}_{X,1}$, then
$\textnormal{SLOPE}(P_{O\mid X}\times Q_{X},\psi^{{\textnormal{mean}}})\leq\textnormal{SLOPE}(P_{O\mid X}\times\widetilde{Q}_{X},\psi^{{\textnormal{mean}}}),$
where the equality holds if and only if $Q_{X}(\mathcal{S}_{X,1})=\widetilde{Q}_{X}(\mathcal{S}_{X,1})$.
3.3 Example: SLOPE for Median
Similar to the SLOPE of the mean, the SLOPE of the median also depends on the dispersion of the underlying distribution $Q_{O,X}^{0}$. We provide its general form in Theorem 2 along with two special cases.
Theorem 2 (SLOPE for Median).
Suppose Condition 1 holds with $s(O,X,\psi^{{\textnormal{med}}})=\mathds{1}(O\leq\psi^{{\textnormal{med}}})-\mathds{1}(O>\psi^{{\textnormal{med}}})$. Then SLOPE for the median is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=$ $\dfrac{\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{1/2}\mid X)\mu(X)\right]-\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{1/2})\right]}{f_{Q_{O}^{0}}(m_{1/2})}.$ (7)
(i) If $P_{O\mid X}$ is symmetric and $\mu(X)=m_{1/2}$ almost surely $Q_{X}$, then (7) simplifies to
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})=\dfrac{m_{1/2}-\textnormal{E}_{Q_{O}^{0}}\left[O\mid O\leq m_{1/2}\right]}{2f_{Q_{O}^{0}}(m_{1/2})}.$ (8)
(ii) If $P_{O\mid X}$ is Gaussian, i.e., $P_{O\mid X}\sim N\left(\mu(X),\sigma^{2}(X)\right)$, then (7) simplifies to
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=\textnormal{E}_{Q_{X}}\left[\sigma^{2}(X)\cdot\dfrac{f_{P_{O\mid X}}(m_{1/2})}{\textnormal{E}_{Q_{X}}\left\{f_{P_{O\mid X}}(m_{1/2})\right\}}\right].$
Compared with the mean, the SLOPE for the median crucially depends on the marginal median $m_{1/2}$ and the tail behavior. Therefore, it reflects the spread of $Q_{O,X}^{0}$ in ways that are more complicated than $\sigma^{2}(X)$. For example, in part (i) which induces symmetry, the SLOPE of the median in (8) is the product of two quantities: (a) the difference between the mean and the truncated mean that is lower than the median divided by two, $\left\{m_{1/2}-\textnormal{E}_{Q_{O}^{0}}\left[O\mid O\leq m_{1/2}\right]\right\}/2$, which measures the spread of $Q_{O}^{0}$ and (b) the inverse of the marginal density at the median, $1/f_{P_{O\mid X}}(m_{1/2})$, which is called Tukey’s sparsity (Tukey,, 1965) and measures the inverse of the concentration of $Q_{O}^{0}$. In part (ii) when $P_{O\mid X}$ is Gaussian, the SLOPE of the median becomes a weighted average of the conditional variance $\sigma^{2}(X)$ where the weight is determined by the conditional density $f_{O\mid X}$ evaluated at the median $m_{1/2}$. This form of the median’s SLOPE resembles the mean’s SLOPE (6), which is an (unweighted) average of $\sigma^{2}(X)$.
As the top row of Figure 2 shows, SLOPE for the median also increases when (a) the variance $\sigma_{2}^{2}$ increases, or (b) when the target population assigns a higher proportion for the subgroup with a higher variance. In general for the median, the choice of a robust distribution is more complex since $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})$ depends on $Q_{O,X}^{0}$ not only through the conditional variance $\sigma^{2}(X)$, but also the shape and tail behavior of the conditional density $f_{P_{O\mid X}}$.
4 Further Insights and Results
4.1 Relationship Between SLOPE and IF
Another measure of robustness that precedes and complements our work is Hampel, (1974)’s celebrated influence function (IF). In this section, we show how SLOPE is related to IF.
Briefly, the IF of $\psi(\cdot)$ under $Q_{O,X}^{0}$ is a derivative-based local measure of robustness that quantifies the effect of an infinitesimal contamination at the point $(o,x)$ on the estimand $\psi(Q_{O,X}^{0})$,
$\textnormal{IF}(o,x,\psi(Q_{O,X}^{0}))=$ $\lim_{t\downarrow 0}\dfrac{\psi\left((1-t)Q_{O,X}^{0}+t\delta_{o,x}\right)-\psi(Q_{O,X}^{0})}{t}.$ (9)
The term $\delta_{o,x}$ represents the dirac delta function at $(o,x)$. The IF was originally proposed by Hampel, (1974) as the “influence curve” and later called “influence function” in view of its generalization to higher dimensions (Hampel et al.,, 2011).
Both SLOPE and IFs are local measures of robustness. The main difference is that, IFs measure the local change of the target estimand due to contamination of a single data point whereas SLOPE measures the local change of the target estimand due to violation of conditional exchangeability from all data points. Nevertheless, since both can be viewed as evaluations of directional derivatives, SLOPE is equal to a “tilting” of the IF onto a particular direction as the following Theorem shows.
Theorem 3 (Connection between SLOPE and IF).
Let $\mu(X)=\textnormal{E}_{P_{O\mid X}}(O\mid X)$ be the conditional mean of $O$ given $X$ on the source population. Then SLOPE can be written as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left\{O-\mu(X)\right\}\mid X\right]\right),$ (10)
whenever either (i) $\psi$ is a Z-functional defined in (2) and Condition 1 hold, or (ii) Conditions 2-3 in the Supplement hold.
Similar to Condition 1, condition (ii) in Theorem 3 ensures the chain rule under Hadamard differentiability holds. In words, equation (10) states that SLOPE is the conditional covariance of the IF and $O-\mu(X)$. As shown in Remark 7 of the Supplement, the term $O-\mu(X)$ is a result of the sensitivity model (3) and can be interpreted as the “direction” of violating conditional exchangeability. From this perspective, SLOPE is the projection of the IF onto the direction in which conditional exchangeability is violated. If the IF is nearly orthogonal to the subspace based on violating conditional exchangeability (or more specifically, $O-\mu(X)$), SLOPE will be close to zero.
In addition to the geometric interpretation of SLOPE, Theorem 3 provides a general formula to derive the SLOPE given an IF. Sections 3.2 and 3.3 derive SLOPEs for the mean and the median using their respective IFs. Also, Section C of the Supplement derives SLOPEs for scale parameters, ordinary least squares (OLS) coefficients, Pearson correlation, L-functionals and some other Z-functionals.
4.2 Estimation
We briefly discuss two estimators of SLOPE, the weighting estimator and the regression estimator. In short, estimation of SLOPE follows from the theory of M-estimation (e.g., Chapter 5 of Van der Vaart, (2000)) and to avoid overloaded notations, we focus on estimating the SLOPE of Z-functionals defined in equation (2), and defer the implementation details, theoretical results on the asymptotic properties, and simulations to Sections E and G of the Supplement.
Suppose we have $n_{q}$ independent and identically distributed (i.i.d.) samples $X_{i}\sim Q_{X}$ with $i=1,\cdots n_{q}$ and another independent, $n_{p}$ i.i.d. samples $(X_{i},O_{i})\sim P_{O,X}$ with $i=n_{q}+1,\cdots,n_{p}+n_{q}=n$. Under part (i) of Condition 1, the SLOPE of a Z-functional is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=$ $-\left\{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\dot{s}(O,X,\psi)\right]\right)\right\}^{-1}\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}s\left(O,X,\psi\right)\right]\right).$ (11)
Let $\omega(X)=f_{Q_{X}}(X)/f_{P_{X}}(X)$ be the density ratio of the shared characteristic $X$ between the two populations. The weighting estimator of (11) estimates the two parts of (11) by re-weighing the source population to match the target distribution $Q_{X}$:
[…]
Here, $\widehat{Q}_{O,X}^{0}$ represents samples from the source and target distributions, and $\widehat{\omega}$, $\widehat{\mu}$, and $\widehat{\psi}$ are estimates of $\omega$, $\mu$, and $\psi(Q_{O,X}^{0})$, respectively, say by parametric or semi-parametric methods. We note that the estimation of the density ratio $\omega$ is often based on balancing approaches or selection models; see Section E.7 of the Supplement for examples.
The regression estimator of (11) works by estimating the corresponding conditional expectations $\textnormal{E}_{P_{O\mid X}}$, denoted as $\widehat{\textnormal{E}}_{P_{O\mid X}}$, using parametric or semi-parametric methods:
[…]
For example, for the target mean $\psi^{{\textnormal{mean}}}$ with $s(O,X,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$, the weighting estimator and the regression estimator for its SLOPE are
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})$ $=\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{O_{i}-\widehat{\mu}(X_{i})\}(O_{i}-\widehat{\psi}^{{\textnormal{mean}}}),$
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})$ $=\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i}).$
Here, $\hat{\psi}^{{\textnormal{mean}}}$ is the estimate of the target mean under conditional exchangeability and $\widehat{\sigma}^{2}(x)$ is an estimate of $\sigma^{2}(x)$. A classic way is to estimate $\widehat{\mu}(X)$ with least squares and $\widehat{\sigma}^{2}(X)$ with a log-variance linear model from weighted least squares (Harvey,, 1976; Carroll and Ruppert,, 1982; Davidian and Carroll,, 1987).
4.3 “Sensitivity” of Our Results to Sensitivity Model (3)
Our development of SLOPE has been based on the sensitivity model (3). While our reasons of choosing this model have been discussed in Section 2.2, we further discuss potential extensions of SLOPE to other sensitivity models.
To begin with, we remark on extending the exponential tilting shift, i.e., the right hand side of equation (3), to more general forms. As shown in Remark 7, this extension is straightforward; SLOPE can be still well defined and the connection between SLOPE and IF still holds in a more general form.
Remark 2 (The Direction of Sensitivity).
We consider a broader class of sensitivity models that tilts the density ratio by a given non-negative function $\rho(O,X,\gamma)$ and incorporates the exponential tilting model (3) as a special case with $\rho(O,X,\gamma)=\exp(\gamma\cdot O)$:
$\dfrac{f_{Q_{O\mid X}^{\gamma}}(O,X)}{f_{P_{O\mid X}}(O,X)}$ $\propto\rho(O,X,\gamma),$
where $\rho(O,X,0)=1$ and $\int\rho(O,X,\gamma)dP_{O\mid X}\in(0,\infty)$. Then the result in Theorem 3 becomes:
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{O,X}^{0}}\left(\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\}\right]\right),$
where $\dot{\rho}$ is the derivative of $\rho$ with respect to $\gamma$. Here, $\dot{\rho}$ characterizes the belief on how violations to conditional exchangeability may change the conditional distribution $Q_{O\mid X}^{\gamma}$. For example, when $\rho(O,X,\gamma)=\exp(\gamma\cdot O)$ as in (3), the derivative is $\dot{\rho}(O,X,0)=O$. Therefore, the term $\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\}\right]$ (i.e., $O-\mu(X)$ in (10)) can be viewed as the direction of sensitivity that’s unexplained by the observed variable $X$.
Next, we note that SLOPE under model (3) is based on the shift in a conditional density ratio. One important future direction is to generalize the concept of SLOPE, i.e., a derivative-based robustness measure that quantifies design-level robustness, to other types of sensitivity models. For example, in contrast to specifying $Q_{O\mid X}/P_{O\mid X}$ (semi-)parametrically, one may bound the difference between $Q_{O\mid X}$ and $P_{O\mid X}$ by some specified degree, which elicits a set identification instead of a point identification. Such extension is non-trivial due to the difficulty in (a) generalizing the notion of derivative from a point to a set, and (b) using the derivative as a metric to guide applications. We briefly illustrate the challenges in Remark 3 and defer details to Section J.5 of the Supplement.
Remark 3 (Challenges in Defining SLOPE for Bound-Based Models).
To fix ideas, consider Zeng et al., (2023)’s sensitivity model for $\psi^{\textnormal{mean}}$, which assumes the target conditional distribution, denoted as $Q_{O\mid X}^{\textnormal{bias}}$, deviate from $P_{O\mid X}$ by at most $\gamma$ in conditional means:
$-\gamma+\textnormal{E}_{P_{O\mid X}}(O\mid X)\leq\textnormal{E}_{Q_{O\mid X}^{\textnormal{bias}}}(O\mid X)\leq\gamma+\textnormal{E}_{P_{O\mid X}}(O\mid X).$ (14)
Under (68), Zeng et al., (2023) showed that the target estimand is sharply bounded below and above by $\gamma+\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})$ and $\gamma+\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})$, respectively. Then, an analogous definition of SLOPE where we take the derivative of the upper and lower bounds with respect to $\gamma$ yields 1-1 and $1$, respectively. Since these two numbers disagree, the two-sided limit in Definition 1 no longer exists and the corresponding derivative is not well-defined. Moreover, even if we take the maximum magnitude of derivatives for the target estimand in (68), the resulting value (i.e., $1$) cannot be naturally interpreted by intrinsic properties of the design.
5 Application
5.1 Data Background
We illustrate the use of SLOPE through a re-analysis of Banerjee et al., (2015)’s multi-national experiment that assesses the Graduation program in six countries (Ethiopia, Ghana, Honduras, India, Pakistan, and Peru). The Graduation program provides a holistic set of services, including asset transfers, consumption support and other career and health services, to poor households, with the purpose of improving their livelihoods. Between 2007 and 2014, in each village, eligible households were randomized to the intervention or control group by local non-profit organizations. The experiment lasted for 24 months, and data were collected before and after the program.
We adopt the potential outcome notation stated in Example 4 where $Y(a)$ denotes the potential outcome under treatment $a$, with $a=1$ denoting participation in the program and $a=0$ denoting otherwise. In subsections below, we use SLOPE to study the violation of conditional exchangeability in transporting the average potential outcome under treatment, i.e., $O=Y(1)$, from one country (i.e., the source population $P$) to another country (i.e., the target population $Q$). Since $O=Y(1)$ is a potential outcome in contrast to an observed variable, identification of the SLOPE requires additional causal assumptions, in specific stable unit treatment variable assumption (SUTVA) and strong ignorability, on the source population. See Section D in the Supplement for these assumptions and Section F.1 for their validity in this dataset.
We focus on two types of outcome variables, the per capita consumption and the physical health index, respectively, in Sections 5.2 and 5.3. For each outcome, the baseline covariate corresponds to the same variable measured prior to intervention. Our analyses are based on the complete data with overlapped baseline measurement (i.e., Assumption 1 holds). To harmonize the baseline covariate across countries and for simplicity, we discretize the measurement into coarse categorical variables as $X$ with consistent definitions. Note that due to the randomization of $A$, transformations of $X$ will not affect the plausibility of strong ignorability of $A$. For estimation, we have assumed that the conditional expectation of the outcome is linear in the baseline $X$ and the village where the household locates. Moreover, for SLOPE for the median in Section 5.2, the residual of the linear model is assumed to be normal. For details on data-preprocessing, estimation, and auxiliary results for Sections 5.2 and 5.3, see Sections F.3 and F.4, respectively, in the Supplement.
5.2 Which Source Country Is More Robust?
In this section, we study the SLOPE of transporting the per capita consumption across countries. The outcome variable is the per capita consumption averaged over measurements taken at two time points: immediately after the program and one year later. We adopted the logarithm transformation and after transformation, the outcome ranges from 1.4 to 6.4. We analyze complete data (83% of the sample) with overlapped baseline measurement (99% of the complete data) in five out of six countries. One country (Ethiopia) was excluded from this analysis because their consumption support, as part of the services during intervention, is substantially different from other countries. Sample sizes range from 771 to 2359 households across countries, with an average treatment rate of $35\%$. For each pair of countries, we treat one as the target and estimate the SLOPE of transporting the mean and median (i.e., $\psi^{\rm{\textnormal{mean}}}$ and $\psi^{{\textnormal{med}}}$) from the other (i.e., source).
Results are shown in Table 1. Given a target country and an estimand, the SLOPE is primarily determined by the data distribution in the source country. In most cases, using India or Peru as the source yields a lower SLOPE (i.e., a lower sensitivity) compared with using other source countries while using Pakistan yields a higher SLOPE (i.e., a higher sensitivity). This pattern of SLOPE remains consistent across estimands, i.e., source countries with a lower SLOPE of the mean usually also has a lower SLOPE of the median (e.g., India and Peru) and vice versa (e.g., Ghana). Overall, SLOPE suggests that among the six countries, India and Peru are more robust and Pakistan is least robust to generalize from. To understand the ranking in robustness across source countries, we recall from Sections 3.2 and 3.3 that the SLOPE for mean and median depends on the heterogeneity/dispersion of the source distribution. As shown by the box plot of $Y(1)$ in Figure 3, source countries with a higher SLOPE often have a uniformly higher spread of $Y(1)$ conditional on each category of $X$. Conversely, source countries with a lower SLOPE, for example India and Peru, have more concentrated values of $Y(1)$. A deeper contextual explanation for the robustness of India and Peru is beyond the scope of this work; we present one hypothesis in Section F.3 of the Supplement based on our understanding of the design of Banerjee et al., (2015).
Table 1: The estimated SLOPEs for transporting the counterfactual log-transformed per capita consumption under treatment from a source country (by rows) to a target country (by columns). Bootstrap standard errors are in the parentheses. For each target country and each estimand, the lowest SLOPE values among source countries are bold faced.
Estimand ($\psi$) Source ($P_{O\mid X}$) Target ($Q_{X}$)
Ghana Honduras India Pakistan Peru
Mean Ghana 0.24 (0.01) 0.24 (0.01) 0.24 (0.02) 0.24 (0.02)
Honduras 0.24 (0.01) 0.24 (0.01) 0.25 (0.02) 0.25 (0.02)
India 0.14 (0.01) 0.13 (0.01) 0.20 (0.04) 0.20 (0.04)
Pakistan 0.21 (0.03) 0.21 (0.03) 0.21 (0.04) 0.20 (0.01)
Peru 0.15 (0.01) 0.15 (0.02) 0.15 (0.02) 0.15 (0.01)
Median Ghana 0.24 (0.01) 0.24 (0.01) 0.24 (0.02) 0.24 (0.02)
Honduras 0.24 (0.01) 0.24 (0.01) 0.25 (0.02) 0.25 (0.02)
India 0.13 (0.01) 0.12 (0.01) 0.19 (0.03) 0.19 (0.03)
Pakistan 0.20 (0.03) 0.21 (0.03) 0.21 (0.04) 0.20 (0.01)
Peru 0.15 (0.02) 0.15 (0.02) 0.14 (0.02) 0.15 (0.01)
In addition to comparing robustness across source countries, SLOPEs in Table 1 also present the comparison in robustness between the mean and the median. SLOPEs for mean and median are close and no one is dominating the other. In most cases the median has a smaller SLOPE, with exceptions where the its SLOPE is higher than the mean’s by no greater than 0.01. In addition, Section F.3.2 in the Supplement shows that the median and mean themselves (not their SLOPEs) have comparable estimated variances. Therefore, generally, we recommend using the median in transporting the average of per capita consumption.
Figure 3: Boxplots of $Y(1)$ across countries (in y-axis) and categories of $X$ (in panels).
Finally, since the outcome data are measured in the target country, we can assess the bias introduced by conditional exchangeability (i.e., “oracle bias”) empirically. This bias can also be approximated with SLOPE and an estimate of $\gamma$, in light of the first-order approximation in (5). Although bias approximation is not the designed/recommended usage of SLOPE, for completeness, we check the bias approximation with SLOPE by comparing it with the oracle bias. As shown by Figure 7 of the Supplement, the approximation works reasonably well. Additional details are in Section F.3.3 of the Supplement.
5.3 Which Health Index Is More Robust for Generalization?
Banerjee et al., (2015) constructed a physical health index to capture the variation in overall physical health of individuals in a household. Specifically, the index is an (equally weighted) average of three standardized variables (z-scores): did not miss work due to illness ($Y_{\text{notMiss}}(1)$), activities of daily living score ($Y_{\text{act}}(1)$), and perception of health status ($Y_{\text{perc}}(1)$). In this section, we ask whether there is another way to define a physical health index using SLOPE so that it’s less sensitive for generalization.
Formally, suppose we rewrite the physical health index as a weighted average of three z-scores:
$\alpha_{\text{notMiss}}Y_{\text{notMiss}}(1)+\alpha_{\text{act}}Y_{\text{act}}(1)+\alpha_{\text{perc}}Y_{\text{perc}}(1).$
In the original analysis, the weights $\alpha_{\text{notMiss}}$, $\alpha_{\text{act}}$, and $\alpha_{\text{perc}}$ were set to $1/3$ (i.e., equally weighted). Our goal is to find a new vector of weights $\alpha=[\alpha_{\text{notMiss}},\alpha_{\text{act}},\alpha_{\text{perc}}]^{\intercal}$ in a simplex that minimizes the SLOPE of the mean, i.e., $\psi^{{\textnormal{mean}}}=\textnormal{E}_{Q_{O}}(O)$.
We focus on households in three countries (India, Ethiopia, Peru) where all three variables were measured prior to and after the treatment assignment. Our analysis is based on complete data (84% of all samples) with overlapped baseline measurement (80% of the complete data), which results in 730 households in India, 785 in Ethiopia, and 1307 in Peru with an average treatment rate of 43%.
Figure 4 presents the SLOPE across different weight combinations where the target country is Ethiopia and the source country is (left) India or (right) Peru. When the source country is India, the SLOPE is minimized at $\alpha_{\text{notMiss}}=0.10$, $\alpha_{\text{act}}=0.55$, $\alpha_{\text{perc}}=0.35$. Upon closer examination, the variable notMiss has the highest variance compared to the other two variables and thus, putting a low weight on it minimizes the overall variance of $O$. Also, when the source country is Peru, the SLOPE is minimized with the weight representing the left bottom vertex where $\alpha_{\text{notMiss}}=1$. Again, upon closer examination, this was because the variance of $Z_{\text{notMiss}}$ in Peru is substantially lower than the other two variables. More broadly, similar to the theoretical discussion from Section 3.2 and Corollary 1, SLOPE prefers distributions of the physical index that are more homogeneous and less variable in order to be less sensitive to generalization. While this phenomenon is intuitive, some readers may be concerned that generalizing from a homogeneous distribution may not be “useful”. We will address such concerns and further clarify the role of SLOPE in Section 6.
Figure 4: SLOPE for the mean of the health summary across weighting schemes. In each panel, edges of the triangle represent the weight in percentage scale of three z-scores. The contours represent the SLOPE where a lighter color means a lower SLOPE, which is more favorable. The star represents the case where the SLOPE is lowest in the simplex.
6 Discussion
This work proposes a measure, Sensitivity to LOcal Perturbation from Exchangeability (SLOPE), to quantify the change in the target estimand with respect to local violations of the conditional exchangeability assumption in transportability and generalizability. It reflects several design-level quantities of a generalization problem and can thus guide the design for generalization, including choices of source/target distributions and the target estimand. In particular, when interpreted with care, notably with respect to units (see Section 3), SLOPE can provide a simple and intuitive metric for choosing robust study designs for generalization under violations of conditional exchangeability. With SLOPE, under the validity of Assumption 1, some observations are useful in choosing robust study designs in practice:
  1. 1.
    With a fixed target population, a source population that’s more homogeneous in $P_{O\mid X}$ is more robust to violation of the conditional exchangeability. For instance, the mean’s SLOPE becomes smaller when $\sigma^{2}(X)$ reduces (Corollary 1), which is in agreement to Tipton and Olsen, (2018). As another example, in Section 5.2, countries with a smaller SLOPE (e.g., India and Peru as shown in Table 1) also have a less variant distribution of $O$ given $X$ (Figure 3).
  2. 2.
    With a fixed source population, a target population that contains more “homogeneous units” with respect to $P_{O\mid X}$ is more robust to generalize to. For example, in the left panel of Figure 2, SLOPE increases as the target population incorporates a higher proportion of units from a subgroup (defined with $X$) that corresponds to a more diverse $O$.
  3. 3.
    When holding populations fixed, a target functional is more robust when its influence function projects more (in proportion) onto the space of shared variable $X$. In a trivial example, as implied by Theorem 3, a target functional that concerns $X$ only, i.e., $\psi(Q_{O,X})=\psi(Q_{X})$, has zero SLOPE. Additionally, as shown in Section 5.3, for a weighted average of several physical health variables, increasing weights to variables that are better explained by $X$ will also reduce the magnitude of SLOPE, thereby improving robustness.
We re-emphasize that SLOPE is developed under the validity of the overlap assumption (i.e., Assumption 1), which means that the source population is already sufficiently large compared with the target population, in terms of $P_{X}$ and $Q_{X}$. In addition, while SLOPE is proposed under situations when conditional exchangeability is unlikely to hold, we echo Tipton and Olsen, (2018) and Degtiar and Rose, (2023) on the importance of guaranteeing conditional exchangeability through domain expertise and careful data collection (e.g., by collecting a rich $X$). However, when it is unrealistic to meet the conditional exchangeability with the observed set of $X$, SLOPE is a useful tool to assess sensitivity/robustness of the underlying design, and to guide designs in future generalization problems.
Finally, we highlight some extensions and future directions. First, while this paper focuses on the conditional exchangeability assumption in generalizations, by defining $P$ and $Q$ differently, SLOPE naturally extends to measuring the sensitivity for other types of conditional exchangeability assumptions in causal inference and missing data; see Section J.4 of the Supplement. Second, an interesting example in Section J.6 of the Supplement connects SLOPE mathematically to the marginal interventional effect proposed by Zhou and Opacic, (2022) with the incremental propensity score intervention (Kennedy,, 2019), highlighting the potential analogies of SLOPE in incremental treatment effects. Lastly, while our result has been based on sensitivity model (3), the high level idea of SLOPE as a derivative-based robustness measure for study designs is “generalizable”. As stated in Section 4.3, an important direction is to extend SLOPE from the specific model (3) to other sensitivity models that posit less parametric assumptions.
Acknowledgment
The authors gratefully acknowledge Guanhua Chen, Melody Huang, Edward Kennedy, Jae-Kwang Kim, Qingyuan Zhao, and participants of the 2024 International Conference on Statistics and Data Science, 2024 Joint Statistical Meetings, 2025 American Causal Inference Conference, and the statistics seminar at Seoul National University for their valuable feedback.
Appendix
The Appendix is organized as follows. Section A discusses the existence of SLOPE, including the definition of SLOPE through Hadamard differentiability and some regularity conditions deferred from the main text. Section B exemplifies the use of SLOPE in choosing a robust location estimand. Section C discusses SLOPE for other estimands, including risks, quantiles, trimmed mean, OLS coefficients, some scale parameters, Pearson correlation coefficient, and general formulas of SLOPEs for L-functionals and Z-functionals. Section D discusses transporting functionals of potential outcomes. Next, Section E and Section F supplement Section 4.2 (estimation) and Section 5 (data application) in the main text, respectively, by providing details, auxiliary results, and some deferred discussions. Section G includes a simulation study that verifies the asymptotic properties of proposed estimators. Afterwards, Sections H and I provide proof for the derivation of the SLOPE (at population level) and the estimation of the SLOPE, respectively. Finally, Section J details some extended remarks, including SLOPE with a vector valued $\psi(\cdot)$, SLOPE for other types of conditional exchangeability assumptions, the challenge of extending SLOPE to bound-based sensitivity models, and the mathematical connection between SLOPE and the marginal interventional effect.
Appendix A Existence of SLOPE
A.1 Notation
We introduce some notation for normed spaces and functional analysis (Van Der Vaart and Wellner,, 1996). For an arbitrary set $T$, let $l^{\infty}(T)$ be the set of all uniformly bounded, real functions on $T$: the set with elements satisfying $z:T\to\mathbb{R}$ such that $\lVert z\rVert_{T}:=\sup_{t\in T}|z(t)|<\infty$. For two normed spaces $\mathbb{D}$ and $\mathbb{E}$, a map $\phi:\mathbb{D}_{\phi}\subset\mathbb{D}\to\mathbb{E}$ is Hadamard differentiable at $\theta\in\mathbb{D}_{\phi}$ tagentially to $\mathbb{D}_{0}$ if there exists a continuous linear map $\phi_{\theta}^{\prime}:\mathbb{D}\to\mathbb{E}$ such that $\{\phi(\theta+t_{n}h_{n})-\phi(\theta)\}/t_{n}\to\phi_{\theta}^{\prime}(h)$ as $n\to\infty$, for all converging sequences $t_{n}\to 0$ and $h_{n}\to h$ such that $\theta+t_{n}h_{n}\in\mathbb{D}_{\phi}$ and $h\in\mathbb{D}_{0}$, where $\mathbb{D}_{\phi}$ and $\mathbb{D}_{0}$ are two subsets of $\mathbb{D}$.
We introduce notations for the supports under $Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$. Let $\mathcal{S}_{O,X}$ be the support under $Q_{O,X}^{0}$ and $\mathcal{S}_{X}$ and $\mathcal{S}_{O}$ be the supports for the marginals $Q_{X}$ and $Q_{O}^{0}$, respectively.
A.2 SLOPE Through Hadamard Differentiability.
We define SLOPE through the derivative of a composite of two functionals. First, define the map from $\gamma$ to $Q_{O,X}^{\gamma}$ as $\phi:\left(\mathbb{R},l^{\infty}(\mathcal{S}_{O,X})\right)\to l^{\infty}(\mathcal{S}_{O,X})$ such that $(\gamma,Q_{O,X}^{0})\mapsto Q_{O,X}^{\gamma}$, with domain $\mathbb{D}_{\phi}=\left([-\varepsilon,\varepsilon],Q_{O,X}^{0}\right)\subset\left(\mathbb{R},l^{\infty}(\mathcal{S}_{O,X})\right)$ for some $\varepsilon>0$. Next, consider the functional $\psi$ as $l^{\infty}(\mathcal{S}_{O,X})\to\mathbb{R}$ with domain $\mathbb{D}_{\psi}\subset l^{\infty}(\mathcal{S}_{O,X})$ which contains probability distributions on $\mathcal{S}_{O,X}$. Then holding $Q_{O,X}^{0}$ fixed, SLOPE as defined in (4) is the Hadamard derivative of the composite function $\psi\circ\phi$ with respect to $\gamma$ at zero. It exists under Condition 2, the standard condition that enables the chain rule of Hadamard differentiability. For Z-functionals, SLOPE exists under Condition 1 of the main text. Part (i) ensures exchanging integration and differentiation and the existence of SLOPE as an integral; parts (ii) and (iii) are analogous to regularity conditions that ensures the existence of the IF of $\psi$ (see Section 4.1).
Condition 2 (Existence of SLOPE).
Suppose $\phi$ as a function of $\gamma$ is Hadamard differentiable at 0 and $\psi$ is Hadamard differentiable at $\phi(0,Q_{O,X}^{0})=Q_{O,X}^{0}$ tangentially to $\phi^{\prime}_{0}(\{0\})$.
Condition 3 is a regularity condition that ensures IF is an evaluation of the Hadamard derivative and therefore its connection to SLOPE follows from the linearity of Hadamard differentiability. We note that Hadamard differentiability (i.e., Condition 3) is stronger than the directional differentiability (i.e., (9) in the main text) since the former requires the limit to exist for every sequence of directions that converges to $\delta_{o,x}-Q_{O,X}^{0}$ whereas the latter only requires the limit to exist in this single direction.
Condition 3 (IF as an Evaluation of a Hadamard Derivative).
Suppose $\psi$ is Hadamard differentiable at $Q_{O,X}^{0}$ tangentially to $\delta_{o,x}-Q_{O,X}^{0}$.
A.3 Regularity Conditions
Condition 4 (Existence of SLOPE for the Mean).
Suppose $\textnormal{E}_{P_{O\mid X}}\{O\exp(\gamma O)\mid X\}/\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}$ is uniformly bounded by an integrable function under $Q_{X}$ for $\gamma$ in a neighborhood of zero, and $\sigma^{2}(X)<\infty$ almost surely on $Q_{X}$.
Condition 5 (Existence of SLOPE for the Median).
(i) Suppose $F_{Q_{O}^{0}}$ is differentiable at $m_{1/2}$ with a positive derivative.
(ii) Suppose $\textnormal{E}_{P_{O\mid X}}\left[O\mathds{1}(O\leq m_{1/2})\exp(\gamma O)\mid X\right]/\textnormal{E}_{P_{O\mid X}}(O\mid X)$ is uniformly bounded by an integrable function under $Q_{X}$ for $\gamma$ in a neighborhood of zero, and $\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{1/2}\mid X)\mu(X)\right]$ and $\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{1/2})\right]$ exist.
Condition 6 (Existence of SLOPE for the $q$-th Quantile).
(i) Suppose $F_{Q_{O}^{0}}$ is differentiable at $m_{q}$ with a positive derivative.
(ii) Suppose $\textnormal{E}_{P_{O\mid X}}\left[\mathds{1}(O\leq m_{q})O\exp(\gamma O)\mid X\right]/\textnormal{E}_{P_{O\mid X}}(O\mid X)$ is uniformly bounded by an integrable function under $Q_{X}$ for $\gamma$ in a neighborhood of zero, and $\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{q}\mid X)\mu(X)\right]$ and $\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{q})\right]$ exist.
Appendix B Using SLOPE to Choose A Robust Location Parameter
In this section, we illustrate how SLOPE can guide the choice of a robust location parameter. We focus on comparing two estimands, the mean with functional $\psi^{\textnormal{mean}}$ and the median with functional $\psi^{\textnormal{med}}$. Their SLOPEs have been stated in Theorems 1 and 2 in the main text. When $P_{O\mid X}$ is normal, we recall that in this case the both SLOPEs are weighted average of the conditional variance $\sigma^{2}(X)$. In this case, Theorem 4 lists some sufficient conditions where the SLOPE for the mean is larger than (or equal to) the SLOPE for the median.
Theorem 4 (Comparison in SLOPEs of Mean and Median).
Suppose $P_{O\mid X}\sim N\left(\mu(X),\sigma^{2}(X)\right)$.
(a) If $\sigma^{2}(X)=\sigma^{2}$ almost surely $Q_{X}$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=\sigma^{2}$.
(b) If $\mu(X)=\mu$ almost surely $Q_{X}$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})\geq\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$.
(c) More generally, $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{mean}})\geq\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})$ if and only if
$\textnormal{Corr}_{Q_{X}}[\sigma^{2}(X),f_{P_{O\mid X}}(m_{1/2})]\leq 0,\text{ where }\textnormal{Corr}\text{ represents correlation}.$
In part (a) of Theorem 4 where the conditional variance is constant, the SLOPEs for the median and the mean are identical. From part (b) of Theorem 4, roughly speaking, when the conditional means of $P_{O\mid X}$ are sufficiently uniform, the median is more robust than or equally robust with the mean in terms of violation of conditional exchangeability. Intuitively, this is because the conditional density at the median, i.e., $f_{P_{O\mid X=x}}(m_{1/2})$, should be higher for $x$’s with a lower conditional variance ($\sigma^{2}(x)$). Then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$, as a weighted average of $\sigma^{2}(x)$, will assign a higher weight to $x$’s with a lower $\sigma^{2}(x)$. This makes the SLOPE for the median no larger than the SLOPE for the mean, which is an (unweighted) average of $\sigma^{2}(x)$.
While Theorem 4 lists some sufficient conditions for the median to be more robust than the mean, in general, the relationship can be reversed. For example, suppose the target population contains two subgroups indicated by $X\in\{x_{1},x_{2}\}$. In each subgroup, the source distribution is normal with $P_{O\mid X=x_{1}}\sim N(\mu_{1},\sigma_{1}^{2})$ and $P_{O\mid X=x_{2}}\sim N(\mu_{2},\sigma_{2}^{2})$, where we set $\mu_{1}=0$, $\sigma_{1}=0.5$, and $\sigma_{2}=0.6$. Then SLOPE of the median becomes
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})=(1-w_{2})\sigma_{1}^{2}+w_{2}\sigma_{2}^{2},$
where $w_{2}$ is the weight for the more heterogeneous subgroup ($X=x_{2}$) with the higher variance $\sigma_{2}^{2}$. As shown in Lemma 2, once $\mu_{1}$ and $\sigma_{2}>\sigma_{1}$ are fixed, $w_{2}$ (as a function of $\mu_{2}-\mu_{1}$) is monotonically increasing with $\mu_{2}-\mu_{1}$, the mean difference between the two subgroups. Therefore, as $w_{2}$ increases, the SLOPE for the median assigns an increasing weight for the heterogeneous subgroup, which will eventually exceed the SLOPE for the mean. Intuitively, as $\mu_{2}$ increases, the marginal target distribution under exchangeability $Q_{O}^{0}\sim 0.2N(0,0.5^{2})+0.8N(\mu_{2},0.6^{2})$ becomes more asymmetric and less uni-modal, and thus the usual understanding on median’s robustness no longer holds. Figure 5 provides a visual illustration of different underlying marginal distributions $Q_{O}^{0}$ and how they correspond to different SLOPEs for the median and the mean.
Figure 5: Some toy examples where the SLOPE of the median is more, equal, or less than the SLOPE of the mean.
Appendix C SLOPE for Other Estimands
In this section we provide examples of SLOPE with some other target functionals $\psi(\cdot)$ that were not included in the main text.
C.1 SLOPE for Expectations
Suppose the target estimand is the expectation of a known function $\xi(O,X)$,
$\psi^{\xi}(Q_{O,X})$ $=\textnormal{E}_{Q_{O,X}}\{\xi(O,X)\}.$ (15)
Lemma 1 shows that SLOPE is the expectation (under $Q_{X}$) of the conditional covariance (under $P_{O\mid X})$ of $O$ and $\xi(O,X)$.
Lemma 1.
Suppose Condition 1 holds with $s(O,\psi^{\xi})=\psi^{\xi}-\xi(O,X)$. Then SLOPE for the functional $\psi^{\xi}$ defined in (15) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\xi})$ $=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}\left[O,\xi(O,X)\right]\right\}.$
For example, if $\xi(O,X)=O$, then $\psi^{\xi}=\psi^{{\textnormal{mean}}}$ and the SLOPE is the expectation of the conditional variance of $O$. Additional examples include the centered moment in Example 5, risks and excess risks in Section C.2.
Example 5 (Centered Moments).
Consider $\psi(Q_{O,X})=\textnormal{E}_{Q_{X}}[\{O-\mu(X)\}^{k}]$ for a positive integer $k$. Then the SLOPE is $\textnormal{E}_{Q_{O,X}^{0}}[\{O-\mu(X)\}^{k+1}]$ provided that the expectation exists.
C.2 SLOPE for Risks
Suppose $\delta(X)$ is a (fixed) decision rule that maps from $X$ to the support of $O$. For a loss function $L(O,\cdot)$, define the risk of $\delta(X)$ on the target distribution as $R(\delta)=\textnormal{E}_{Q_{O,X}}[L(O,\delta(X))]$. In this section we consider SLOPEs for risks and excess risks. Note that they can be viewed as special cases of the SLOPE for expectations in Lemma 1 with different choices of $\xi$.
C.2.1 SLOPE for Mean Squared Error
Consider the squared loss with $L(O,\delta(X))=\{O-\delta(X)\}^{2}$. The target functional is the mean squared error of the decision $\delta(X)$ on the target population, i.e., $\psi=\textnormal{E}_{Q_{O,X}}[\{O-\delta(X)\}^{2}]$. Then the SLOPE is $\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=\textnormal{E}_{Q_{O,X}^{0}}\left[\{\delta(X)-O\}^{2}\{O-\mu(X)\}\right]$.
C.2.2 SLOPE for Excess Risk under Squared Loss
Consider again the squared loss $L(O,\delta(X))=\{O-\delta(X)\}^{2}$. Then $\mu(X)$ minimizes the corresponding risk. For any (fixed) $\delta(X)$, let the target functional be the excess risk, i.e., $\psi(Q_{O,X})=R(\delta)-R(\mu)=\textnormal{E}_{Q_{O,X}}\left[\{\delta(X)-O\}^{2}-\{\mu(X)-O\}^{2}\right]$. SLOPE for this excess risk under squared loss is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=2\textnormal{E}_{Q_{X}}\left[\{\mu(X)-\delta(X)\}\sigma^{2}(X)\right].$
C.2.3 SLOPE for Excess Risk under 0-1 Loss
Suppose $O\in\{1,-1\}$ and let $\eta(X)=P(O=1\mid X)$. Consider the 0-1 loss, $L(O,\delta(X))=\mathds{1}\{\delta(X)\neq O\}$, and the corresponding risk
$R(\delta)=\textnormal{E}_{Q_{O,X}}\{L(O,\delta(X)\}=\textnormal{E}_{Q_{O,X}}\left[\mathds{1}\{\delta(X)\neq 1\}\{2\eta(X)-1\}+1-\eta(X)\right].$
Then one of the Bayes classifiers is $\delta^{*}(X)=\mathds{1}\{\eta(X)\geq 1/2\}-\mathds{1}\{\eta(X)<1/2\}$. For any fixed $\delta(X)$, let the target functional be the excess risk, $\psi(Q_{O,X})=R(\delta)-R(\delta^{*})$. Then the SLOPE for the excess risk under 0-1 loss is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=\textnormal{E}_{Q_{X}}\left[\sigma^{2}(X)\mathds{1}\{\delta(X)\neq\delta^{*}(X)\}\textnormal{sign}\{\eta(X)-1/2\}\right],$
where $\textnormal{sign}(\cdot)$ is the sign function such that $\textnormal{sign}(t)=1$ if $t>0$, $\textnormal{sign}(t)=0$ if $t=0$, and $\textnormal{sign}(t)=-1$ if $t<0$.
C.3 SLOPE for Quantiles
For $q\in(0,1)$, let $\psi(Q_{O})=F_{Q_{O}}^{-1}(q)$ be the $q$ quantile of the marginal distribution $Q_{O}$; for example when $q=1/2$, $\psi=\psi^{{\textnormal{med}}}$. Then SLOPE for the quantile, as a general case of the SLOPE for the median (Theorem 2), is presented in Theorem 5.
Theorem 5 (SLOPE for Quantiles).
Suppose Condition 6 holds. Then SLOPE for the $q$-th quantile is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=$ $\dfrac{\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{q}\mid X)\mu(X)\right]-\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{q})\right]}{f_{Q_{O}^{0}}(m_{q})},$
where $m_{q}$ satisfies $\int_{-\infty}^{m_{q}}dQ_{O}^{0}=q$.
C.4 SLOPE for $\alpha$-Trimmed Mean
Let the target functional be the $\alpha$-trimmed mean, $\psi^{{\alpha\textnormal{-trim}}}$, which is the mean of $Q_{O}$ after trimming off the lower and upper $\alpha$ quantiles. Specifically, the $\psi^{{\alpha\textnormal{-trim}}}$ is defined as
$\psi^{{\alpha\textnormal{-trim}}}(Q_{O})=\dfrac{1}{1-2\alpha}\int_{\alpha}^{1-\alpha}F_{Q_{O}}^{-1}(p)dp,$ (16)
for $\alpha\in(0,1/4)$.
Theorem 6 (SLOPE for $\alpha$-Trimmed Mean).
SLOPE for the $\alpha$-trimmed mean $\psi^{{\alpha\textnormal{-trim}}}$ defined in (16) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\alpha\textnormal{-trim}}})=$ $\frac{F^{-1}_{Q_{O}^{0}}(1-\alpha)}{1-2\alpha}\textnormal{E}_{Q_{O,X}^{0}}\left\{F_{P_{O\mid X}}(F_{Q_{O}^{0}}^{-1}(1-\alpha))\mu(X)-O\mathds{1}_{(-\infty,F_{Q_{O}^{0}}^{-1}(1-\alpha)]}(O)\right\}$
$-\frac{F^{-1}_{Q_{O}^{0}}(\alpha)}{1-2\alpha}\textnormal{E}_{Q_{O,X}^{0}}\left\{F_{P_{O\mid X}}(F_{Q_{O}^{0}}^{-1}(\alpha))\mu(X)-O\mathds{1}_{(-\infty,F_{Q_{O}^{0}}^{-1}(\alpha)]}(O)\right\}$
$+\dfrac{1}{1-2\alpha}\textnormal{E}_{Q_{O,X}^{0}}\left[O\{O-\mu(X)\}\mathds{1}_{\left(F_{Q_{O}^{0}}^{-1}(\alpha),F_{Q_{O}^{0}}^{-1}(1-\alpha)\right)}(O)\right].$
Moreover, if $P_{O\mid X}\sim N\left(\mu(X),\sigma^{2}(X)\right)$, then the SLOPE is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\alpha\textnormal{-trim}}})=\dfrac{1}{1-2\alpha}\textnormal{E}_{Q_{X}}\left[\sigma^{2}(X)\left\{\Phi\left(\frac{F_{Q_{O}^{0}}^{-1}(1-\alpha)-\mu(X)}{\sigma(X)}\right)-\Phi\left(\frac{F_{Q_{O}^{0}}^{-1}(\alpha)-\mu(X)}{\sigma(X)}\right)\right\}\right],$
where $\Phi(\cdot)$ is the cumulative distribution function of standard normal.
C.5 SLOPE for OLS Coefficients
Suppose $O=Y$ is an outcome variable and $X$ is a vector of covariates that includes one. We are interested in the OLS coefficient of regressing $Y$ on $X$ in the target distribution, i.e., $\psi^{{\textnormal{OLS}}}(Q_{Y,X})$ such that
$\textnormal{E}_{Q_{Y,X}}\left[XX^{\intercal}\psi^{{\textnormal{OLS}}}-XY\right]=0.$ (17)
C.5.1 SLOPE for OLS Coefficients with Omitted Variables
We consider transferring OLS coefficients $\psi^{{\textnormal{OLS}}}$ in (17) where $X$ is a vector of covariates and contains one as the intercept, $O=Y$ is a scalar outcome variable. The target estimand $\psi^{{\textnormal{OLS}}}$ can be equivalently defined as the solution to the following least squares problem,
$\psi^{{\textnormal{OLS}}}=\arg\min_{\psi^{{\textnormal{OLS}}}}\textnormal{E}_{Q_{O,X}}\left[\left(Y-X^{\intercal}\psi^{{\textnormal{OLS}}}\right)^{2}\right]$ (18)
For generality, we also consider the OLS coefficient where only a subset $X_{\textnormal{Sub}}\subset X$ that contains the intercept is being modeled:
$\psi^{\textnormal{OLS,Sub}}=\arg\min_{\psi^{\textnormal{OLS,Sub}}}\textnormal{E}_{Q_{O,X}}\left[\left(Y-X_{\textnormal{Sub}}^{\intercal}\psi^{\textnormal{OLS,Sub}}\right)^{2}\right].$ (19)
Note that (19) contains (18) as a special case by setting $X_{\textnormal{Sub}}=X$.
Theorem 7 (SLOPE for OLS Coefficient).
Suppose Condition 1 holds with $s(Y,\psi^{\textnormal{OLS,Sub}})=X_{\textnormal{Sub}}X_{\textnormal{Sub}}^{\intercal}\psi^{\textnormal{OLS,Sub}}-X_{\textnormal{Sub}}Y$. SLOPE for the OLS coefficient with a subset of covariates (19) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{OLS,Sub}})=\left\{\textnormal{E}_{Q_{X}}\left(X_{\textnormal{Sub}}X_{\textnormal{Sub}}^{\intercal}\right)\right\}^{-1}\textnormal{E}_{Q_{X}}\left\{X_{\textnormal{Sub}}\sigma^{2}(X)\right\}.$
In the special case where $X_{\textnormal{Sub}}=1$, coefficient $\psi^{\textnormal{OLS,Sub}}$ becomes the outcome mean (Theorem 1). As expected, SLOPE in Theorem 7 becomes $\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}$, which is identical to the SLOPE for the mean. On the contrary, in another special case where $X_{\textnormal{Sub}}=X$, the SLOPE becomes $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})=\left\{\textnormal{E}_{Q_{X}}\left(XX^{\intercal}\right)\right\}^{-1}\textnormal{E}_{Q_{X}}\left\{X\sigma^{2}(X)\right\}$.
C.5.2 SLOPE in ANCOVA
Suppose $X=[1,A,L^{\intercal}]^{\intercal}$ where $A$ is a binary treatment and $L$ contains pre-treatment covariates. We are interested in the effect of $A$ on $Y$ in two models of analysis of covariance (ANCOVA) in which (20) adjusts for covariates $L$ and (21) does not:
$(\alpha_{\textnormal{a}},\tau_{\textnormal{a}},\beta_{\textnormal{a}})$ $=\arg\min_{(\textnormal{W}\alpha_{\textnormal{a}},\textnormal{W}\tau_{\textnormal{a}},\textnormal{W}\beta_{\textnormal{a}})}\textnormal{E}_{Q_{O,X}^{0}}\left[\left(Y-\textnormal{W}\alpha_{\textnormal{a}}-\textnormal{W}\tau_{\textnormal{a}}A-\textnormal{W}\beta_{\textnormal{a}}^{\intercal}L\right)^{2}\right],$ (20)
$(\alpha_{\textnormal{u}},\tau_{\textnormal{u}})$ $=\arg\min_{(\textnormal{W}\alpha_{\textnormal{u}},\textnormal{W}\tau_{\textnormal{u}})}\textnormal{E}_{Q_{O,X}^{0}}\left[\left(Y-\textnormal{W}\alpha_{\textnormal{u}}-\textnormal{W}\tau_{\textnormal{u}}A\right)^{2}\right].$ (21)
Specifically, consider SLOPE for the regression slopes $\tau_{\textnormal{a}}$ and $\tau_{\textnormal{u}}$. We will see that the conditional varaince $\sigma^{2}(X)$ still plays an important role. To reflect its dependency on the treatment $A$ and covariates $L$, we slightly abuse the notation and let $\sigma^{2}(L,A)=\textnormal{Var}_{P_{Y\mid X}}(Y\mid A,L)=\textnormal{Var}_{P_{Y\mid X}}(Y\mid X)$.
Remark 4 (Difference Between Transporting $Y\mid A$ and Transporting $Y(A)$).
Coefficients $\tau_{\textnormal{a}}$ and $\tau_{\textnormal{u}}$ can be interpreted as the (causal) treatment effect of $A$ on $Y$ under modeling assumptions. Nevertheless, they are different from the causal effects in Section D which are based on transporting potential outcomes, even under the same modeling assumptions.
In Section D, the transportation is for the potential outcome $O=Y(A)$, based on the commonly observed pre-treatment covariates in the two populations (i.e., $X$). The SLOPE depends on target population only through the distribution of pre-treatment covariates; notably, the treatment $A$ itself need not be well defined on the target population. The SLOPE depends on the source population only through the distribution of the potential outcome $Y(A)$ given pre-treatment covariates. It does not directly depend on how $A$ has been randomized on the source population, as longs as the identification is guaranteed. This setting is widely adopted in generalizing a causal effect from one population where an experimental has taken place to another population where only baseline covariates are collected.
In this Section, the transportation is for the observed outcome $O=Y$, based on the commonly observed intervention and pre-treatment covariates (i.e., $X$ contains both $A$ and pre-treatment covariates). The intervention $A$ is observed on the target population and the SLOPE depends on the target population not only through covariates, but also through $A$. This setting is less common than the previous one. One example of this setting is in surveys (or in general, missing data) where $P$ contains units whose outcome variables are observed and $Q$ contains units whose outcome variables are not observed.
The following Theorem 8 gives the SLOPE for regression slopes $\tau_{\textnormal{a}}$ and $\tau_{\textnormal{u}}$.
Theorem 8 (SLOPE in ANCOVA).
The SLOPEs for $\tau_{\textnormal{u}}$ and $\tau_{\textnormal{a}}$ are
$\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{u}})$ $=\dfrac{\textnormal{Cov}_{Q_{L,A}}[A,\sigma^{2}(L,A)]}{\textnormal{Var}_{Q_{A}}(A)},$ (22)
$\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})$ $=\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{u}})+\delta^{\intercal}V^{-1}\delta\textnormal{Cov}_{Q_{L,A}}[A,\sigma^{2}(L,A)]-\delta^{\intercal}V^{-1}\textnormal{Cov}_{Q_{L,A}}[Z,\sigma^{2}(L,A)],$ (23)
respectively, where $\delta=\textnormal{Cov}_{Q_{X}}[L,A]/\textnormal{Var}_{Q_{A}}[A]$ and $V=\textnormal{Cov}_{Q_{Z}}(L)-\delta\delta^{\intercal}\textnormal{Var}_{Q_{A}}(A)$.
From Theorem 8, if $\textnormal{Cov}_{Q_{L,A}}(Z,A)=0$ almost surely, e.g., in cases when the intervention $A$ has been randomized, then by definition $\delta=0$ and
$\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})=\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{u}})=\textnormal{Cov}_{Q_{L,A}}[A,\sigma^{2}(L,A)]/\textnormal{Var}_{Q_{A}}(A).$
C.6 SLOPE for Scale Parameters
C.6.1 SLOPE for Variance
Let the target functional be the variance of $O$, i.e., $\psi^{\text{var}}=\textnormal{Var}_{Q_{O}}(O)$. Then the SLOPE is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\text{var}})=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}[O^{2},O\mid X]\right\}-2\textnormal{E}_{Q_{O}^{0}}(O)\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}.$
C.6.2 SLOPE for MAD
Let the target functional be the median absolute deviation from the median (MAD), $\psi^{{\textnormal{MAD}}}(\cdot)$ such that
$Q_{O}\left(\left|O-\psi^{{\textnormal{med}}}(Q_{O})\right|\leq\psi^{{\textnormal{MAD}}}\right)=1/2,$ (24)
where $\psi^{{\textnormal{med}}}$ is the functional that maps to the marginal median which has been defined in Example 2 of the main text. Recall that $m_{1/2}=\psi^{{\textnormal{med}}}(Q_{O}^{0})$ and to ease notation, let ${\textnormal{MAD}}=\psi^{{\textnormal{MAD}}}(Q_{O}^{0})$ be the MAD of $Q_{O}^{0}$.
Theorem 9 (SLOPE for MAD).
Suppose Condition 2 holds for $\psi^{{\textnormal{MAD}}}$. Then SLOPE for the MAD defined through $\psi^{{\textnormal{MAD}}}$ in (24) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{MAD}}})=$ $\dfrac{f_{Q_{O}^{0}}\left(m_{1/2}-{\textnormal{MAD}}\right)-f_{Q_{O}^{0}}\left(m_{1/2}+{\textnormal{MAD}}\right)}{f_{Q_{O}^{0}}\left(m_{1/2}-{\textnormal{MAD}}\right)+f_{Q_{O}^{0}}\left(m_{1/2}+{\textnormal{MAD}}\right)}\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$
$-\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\mathds{1}_{[m_{1/2}-{\textnormal{MAD}},m_{1/2}+{\textnormal{MAD}}]}(O)\{O-\mu(X)\}\right]}{f_{Q_{O}^{0}}\left(m_{1/2}-{\textnormal{MAD}}\right)+f_{Q_{O}^{0}}\left(m_{1/2}+{\textnormal{MAD}}\right)},$
where $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$ is the SLOPE for median (Theorem 2).
C.6.3 SLOPE for $\alpha$-Quantile Range
The $\alpha$-quantile range of $Q_{O}$ is $\psi^{\alpha\text{-range}}(Q_{O})=F_{Q_{O}}^{-1}(1-\alpha)-F_{Q_{O}}^{-1}(\alpha)$ for a given $\alpha\in(0,1/2)$. A common choice is to let $\alpha=1/4$ and the $\alpha$-quantile range is the interquartile range. By the SLOPE for quantiles (Theorem 5), the SLOPE for the $\alpha$-quantile range is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\alpha\text{-range}})=$ $\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[(1-\alpha)O-\mathds{1}\left(O\leq F_{Q_{O}^{0}}^{-1}(1-\alpha)\right)O\right]}{f_{Q_{O}^{0}}\left(F_{Q_{O}^{0}}^{-1}(1-\alpha)\right)}$
$-\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\alpha O-\mathds{1}\left(O\leq F_{Q_{O}^{0}}^{-1}(\alpha)\right)O\right]}{f_{Q_{O}^{0}}\left(F_{Q_{O}^{0}}^{-1}(\alpha)\right)}.$
C.7 SLOPE for Pearson Correlation Coefficient
Suppose $O\in\mathbb{R}$ and let the target functional be the Pearson correlation between $O$ and $X$:
$\psi^{\textnormal{Corr}}(Q_{O,X})=\dfrac{\textnormal{E}_{Q_{O,X}}(XO)-\textnormal{E}_{Q_{X}}(X)\textnormal{E}_{Q_{O}}(O)}{\sqrt{\textnormal{Var}_{Q_{X}}(X)\textnormal{Var}_{Q_{O}}(O)}}.$ (25)
The SLOPE for the Pearson correlation coefficient is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{Corr}})$ $=\dfrac{\textnormal{Cov}_{Q_{X}}[X,\sigma^{2}(X)]}{\sqrt{\textnormal{Var}_{Q_{X}}(X)\textnormal{Var}_{Q_{O}^{0}}(O)}}-\dfrac{\textnormal{Corr}_{Q_{O,X}^{0}}[X,O]}{2\textnormal{Var}_{Q_{O}^{0}}(O)}\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{Var}}),$
where $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{Var}})$ is the SLOPE for the variance of $O$ (Section C.6.1). We also note that the relationship between SLOPE and IF (Theorem 3) can be easily verified, where the IF is
$\textnormal{IF}\left(O,X,\psi^{\textnormal{Corr}}(Q_{O,X}^{0})\right)=$ $\dfrac{X-\textnormal{E}_{Q_{X}}(X)}{\sqrt{\textnormal{Var}_{Q_{X}}(X)}}\cdot\dfrac{O-\textnormal{E}_{Q_{O}^{0}}(O)}{\sqrt{\textnormal{Var}_{Q_{O}^{0}}(O)}}$
$-\dfrac{\textnormal{Corr}_{Q_{O,X}^{0}}(X,O)}{2}\left\{\dfrac{\{X-\textnormal{E}_{Q_{X}}(X)\}^{2}}{\textnormal{Var}_{Q_{X}}(X)}+\dfrac{\{O-\textnormal{E}_{Q_{O}^{0}}(O)\}^{2}}{\textnormal{Var}_{Q_{O}^{0}}(O)}\right\}$
by Devlin et al., (1975).
C.8 SLOPE for L-Functionals
C.8.1 L-Functionals and Their SLOPEs
We discuss one important class of estimands whose sample counterparts correspond to L-estimates (Huber,, 1981, Section 3.3). Specifically, consider a one-dimensional functional $\psi(Q_{O,X})=\psi(Q_{O})$ defined on the marginal distribution of $O$:
$\psi(Q_{O})$ $=\int_{0}^{1}h(F_{Q_{O}}^{-1}(p))l(p)dp.$ (26)
Choosing a particular function $h(\cdot)$ and a density function $l$, both defined over the support $(0,1)$, determines a location parameter $\psi(Q_{O})$. For example, when $l(p)=1$, equation (26) reduces to the mean of $O$ in the target population, i.e., $\psi^{{\textnormal{mean}}}(Q_{O})$. When $l(p)=\mathds{1}_{\frac{1}{2}}(p)$ where $\mathds{1}_{C}(p)$ for a set $C$ is the indicator function such that $\mathds{1}_{C}(p)=1$ if $s\in C$ and $\mathds{1}_{C}(p)=0$ otherwise, equation (26) becomes the marginal median, $\psi^{{\textnormal{med}}}(Q_{O})$. When $l(p)=\mathds{1}_{[\alpha,1-\alpha]}(p)/(1-2\alpha)$, then the target functional becomes the marginal trimmed mean, $\psi^{{\alpha\textnormal{-trim}}}$.
Theorem 10 derives the SLOPE for estimands defined in (26).
Theorem 10 (SLOPE for L-Functionals).
Suppose Condition 2 holds with $\psi$ in (26). Then SLOPE of $\psi$ defined in (26) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$
$=$ $\int_{0}^{1}\dfrac{-\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(p))\right]+\textnormal{E}_{Q_{O,X}^{0}}\left[\mu(X)\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(p))\right]}{f_{Q_{O}^{0}}\left(F_{Q_{O}^{0}}^{-1}(p)\right)}h^{\prime}\left(F_{Q_{O}^{0}}^{-1}(p)\right)l(p)dp,$
where $h^{\prime}(\cdot)$ is the derivative of $h(\cdot)$.
C.8.2 Lemma 2
Lemma 2 (SLOPE of Median in Two-Component Gaussian Mixtures).
Suppose $X\in\{x_{1},x_{2}\}$ and $q_{1}=Q_{X}(x_{1})$ and $q_{2}=Q_{X}(x_{2})$, $P_{O\mid X=x_{1}}\sim N(\mu_{1},\sigma_{1}^{2})$ and $P_{O\mid X=x_{2}}\sim N(\mu_{2},\sigma_{2}^{2})$ with $\sigma_{2}>\sigma_{1}$ fixed. We denote the SLOPE for the median as a function of $\Delta=\mu_{2}-\mu_{1}$: $\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]$. Then the followings hold.
(i) If $0<q_{1}<q_{2}$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]$ is increasing with $\Delta$.
(ii) If $q_{1}>q_{2}>0$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]$ is decreasing with $\Delta$.
(iii) If $q_{1}=q_{2}=1/2$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]=\sigma_{1}\sigma_{2}$ does not depend on $\Delta$.
Proof of Lemma 2. We start with parts (i) and (ii). Without loss of generality we suppose $\mu_{1}=0$ and $\sigma_{1}=1$ and therefore $\mu_{2}=\Delta$ and $\sigma_{2}>1$.
The SLOPE of the median can be expressed as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]=w_{1}(\Delta)(\sigma_{1}^{2}-\sigma_{2}^{2})+\sigma_{2}^{2},$
The weight is
$w_{1}(\Delta)$ $=\dfrac{q_{1}f_{P_{Y\mid X=x_{1}}}(m_{1/2})}{q_{1}f_{Y\mid X=x_{1}}(m_{1/2})+q_{2}f_{P_{Y\mid X=x_{2}}}(m_{1/2})}$
$=\dfrac{1}{1+\dfrac{q_{2}}{q_{1}\sigma_{2}}\dfrac{\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)}{\varphi(m_{1/2}(\Delta))}}$
$=\dfrac{1}{1+\dfrac{q_{2}}{q_{1}\sigma_{2}}\exp\left\{\frac{1}{2}\left[\{m_{1/2}(\Delta)\}^{2}-\{m_{1/2}(\Delta)-\Delta\}^{2}/\sigma_{2}^{2}\right]\right\}}$
$=\dfrac{1}{1+\dfrac{q_{2}}{q_{1}\sigma_{2}}\exp\{h(\Delta)\}}$
where $m_{1/2}=m_{1/2}(\Delta)$ is the marginal median that satisfies
$\int_{-\infty}^{m_{1/2}(\Delta)}\left[q_{1}\varphi(y)+q_{2}\varphi\left(\frac{y-\Delta}{\sigma_{2}}\right)/\sigma_{2}\right]dy=1/2.$ (27)
and $h(\Delta)=\dfrac{1}{2}\left[\{m_{1/2}(\Delta)\}^{2}-\{m_{1/2}(\Delta)-\Delta\}^{2}/\sigma_{2}^{2}\right]$. Since $\dfrac{q_{2}}{q_{1}\sigma_{2}}>0$ and the function $\exp(\cdot)$ is increasing, to prove $w_{1}(\Delta)$ is monotonically decreasing (resp. increasing) with $\Delta$, it’s sufficient to prove $h(\Delta)$ is monotonically increasing (resp. decreasing) with $\Delta$.
We start with case (i) when $0<q_{1}<q_{2}$. From (27) we know that the derivative of $m(\Delta)$ is
$m_{1/2}^{\prime}(\Delta)=\dfrac{\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)}{q_{1}\varphi(m_{1/2}(\Delta))+\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)}$
and it satisfies $m_{1/2}^{\prime}(\Delta)\in(0,1)$. Then the derivative of $h(\Delta)$ is
$h^{\prime}(\Delta)=$ $m_{1/2}(\Delta)\cdot m_{1/2}^{\prime}(\Delta)-\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}^{2}}\left[m_{1/2}^{\prime}(\Delta)-1\right]$
$=$ $\left[q_{1}\varphi(m_{1/2}(\Delta))+\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)\right]\cdot$
$\left[m_{1/2}(\Delta)\dfrac{q_{2}}{\sigma_{2}}\varphi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)+\dfrac{1}{\sigma_{2}^{2}}\left\{m_{1/2}(\Delta)-\Delta\right\}q_{1}\varphi\left(m_{1/2}(\Delta)\right)\right]\cdot$
$=$ $\underbrace{\dfrac{1}{\sigma_{2}}\left[q_{1}\varphi(m_{1/2}(\Delta))+\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)\right]}_{>0}\cdot$
$\left[q_{2}m_{1/2}(\Delta)\varphi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)-\dfrac{q_{1}}{\sigma_{2}}\left\{\Delta-m_{1/2}(\Delta)\right\}\varphi\left(m_{1/2}(\Delta)\right)\right].$
Since the first term $\dfrac{1}{\sigma_{2}}\left[q_{1}\varphi(m_{1/2}(\Delta))+\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)\right]>0$, the sign of $h^{\prime}(\Delta)$ is determined by the sign of the second term. Hence, we have that
$h^{\prime}(\Delta)>0$ $\Longleftrightarrow q_{2}m_{1/2}(\Delta)\varphi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)-\dfrac{q_{1}}{\sigma_{2}}\left\{\Delta-m_{1/2}(\Delta)\right\}\varphi\left(m_{1/2}(\Delta)\right)>0$
$\Longleftrightarrow\dfrac{q_{2}m_{1/2}(\Delta)\varphi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)}{\dfrac{q_{1}}{\sigma_{2}}\left\{\Delta-m_{1/2}(\Delta)\right\}\varphi\left(m_{1/2}(\Delta)\right)}>1$
$\Longleftrightarrow\dfrac{q_{2}}{q_{1}}\cdot\dfrac{\varphi\left(\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}\right)/\left\{\Delta-m_{1/2}(\Delta)\right\}}{\varphi\left(m_{1/2}(\Delta)\right)/m_{1/2}(\Delta)}>1,$
where the second line follows from the fact that both the numerator and the denominator are positive, and the third line follows by re-organizing terms. In order to show $h^{\prime}(\Delta)>0$, it’s sufficient to show $m_{1/2}(\Delta)>\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}$, because
$h^{\prime}(\Delta)>0$ $\Longleftarrow\dfrac{\varphi\left(\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}\right)/\left\{\Delta-m_{1/2}(\Delta)\right\}}{\varphi\left(m_{1/2}(\Delta)\right)/m_{1/2}(\Delta)}>1$
$\Longleftarrow m_{1/2}(\Delta)>\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}},$
where the first line follows from the fact that $q_{2}/q_{1}>1$ and the second line follows since the function $\varphi(x)/x=\exp(-x^{2}/2)/x$ is decreasing when $x>0$.
Therefore, we are left to show $m_{1/2}(\Delta)>\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}$. Recall the definition of the median in (2), we have
$q_{1}\Phi\left(m_{1/2}(\Delta)\right)+q_{2}\Phi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)=1/2$
$\Longrightarrow$ $q_{1}\left[\Phi\left(m_{1/2}(\Delta)\right)-1/2\right]+q_{2}\left[\Phi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)-1/2\right]=0$
$\Longrightarrow$ $q_{1}\left[\Phi\left(m_{1/2}(\Delta)\right)-1/2\right]=-q_{2}\left[\Phi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)-1/2\right]=q_{2}\left[\Phi\left(\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}\right)-1/2\right]$
$\Longrightarrow$ $\dfrac{\Phi\left(m_{1/2}(\Delta)\right)-1/2}{\Phi\left(\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}\right)-1/2}=\dfrac{q_{2}}{q_{1}}>1$
$\Longrightarrow$ $m_{1/2}(\Delta)>\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}.$
From the above, we have $h^{\prime}(\Delta)>0$ and in hence $w_{1}^{\prime}(\Delta)<0$. The SLOPE $w_{1}(\Delta)(\sigma_{1}^{2}-\sigma_{2}^{2})+\sigma_{2}^{2}$ is therefore increasing with $\Delta$. We have proven part (i).
The proof of part (ii) follows with a similar argument. When $q_{2}<q_{1}$, we have $m_{1/2}(\Delta)<\left[\delta-m_{1/2}(\Delta)\right]/\sigma_{2}$, and therefore $h^{\prime}(\Delta)<0$. This in turn gives $w_{1}^{\prime}>0$ and the SLOPE is decreasing with $\Delta$.
Part (iii) follows by noticing the median is $m_{1/2}=(\mu_{1}\sigma_{2}+\mu_{2}\sigma_{1})/(\sigma_{1}+\sigma_{2})$.
$\square$
C.9 SLOPE for Z-Functionals
For a Z-functional defined through (2) in the main text, the corresponding SLOPE is presented in the Corollary 3 below, which is an immediate result of Theorem 3.
Corollary 3 (SLOPE for Z-Functionals).
Under Condition 3, the SLOPE for a Z-functional is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=-\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi(Q_{O,X}^{0}))\}^{-1}\textnormal{E}_{Q_{O,X}^{0}}\left[s\left(O,X,\psi(Q_{O,X}^{0})\right)\{O-\mu(X)\}\right].$ (28)
Appendix D SLOPE for Functionals of Potential Outcomes
In this section, we discuss how to use SLOPE to quantify the sensitivity of the conditional exchangeability assumption in generalizing parameters in causal inference. We provide a brief review on notation in causal inference in Section D.1 and discuss the SLOPE in Section D.2.
D.1 Notation for Causal Inference
Let $Y$ be the observed outcome variable, $A\in\{0,1\}$ denote the observed treatment, and $X$ denote pre-treatment covariates. Let $Y(a)$ denote the potential outcome if the study unit, contrary to fact, was assigned treatment value $a\in\{0,1\}$. Throughout the paper, we assume that we observe only one of the potential outcomes based on the observed treatment assignment, i.e., $Y=Y(A)$ almost surely (Rubin,, 1980). A central goal is in causal inference to identify some functional of $Y(a)$, for example, the mean of $Y(a)$, for $a\in\{0,1\}$.
D.2 Transporting Functionals of a Potential Outcome
Suppose $P$ is the source population where researchers have completed a randomized experiment to study the effect of treatment $A\in\{0,1\}$ on an outcome $Y$. We adopt notation in Section D.1 and let $Y(a)$ represent the potential outcome under treatment $a$. The goal is to estimate the “average” of the potential outcome $Y(a)$ on the target country, where only the marginal $Q_{X}$ can be identified. In this problem, $O=Y(a)$ is the potential outcome under the treated. To identify its average on the target distribution, in the literature it’s common to impose the overlap condition between $Q_{X}$ and $P_{X}$ and the conditional exchangeability condition concerns the counterfactual outcome between the source and target populations, i.e., Assumptions 1 and 2 with $O=Y(a)$. Specifically, Assumption 2 becomes: $Q_{Y(a),X}(\cdot\mid X)$ is absolute continuous with respect to $P_{Y(a)\mid X}(\cdot\mid X)$ and the Radon-Nikodym derivative satisfies
$dQ_{Y(a),X}(O,X)/dP_{Y(a)\mid X}(O,X)=1\text{ almost everywhere }P_{Y(a)\mid X}\times Q_{X}$ (29)
Accordingly, the sensitivity model (3) becomes
$\dfrac{f_{Q_{Y(a)\mid X}^{\gamma}}(O,X)}{f_{P_{Y(a)\mid X}}(O,X)}\propto\exp(\gamma\cdot O)\text{, almost surely }P_{Y(a)\mid X}\times Q_{X}.$ (30)
Following Theorems 1 and 2, the SLOPEs for the mean potential outcome and median potential outcome are
$\textnormal{SLOPE}(Q_{Y(a),X}^{0},\psi^{{\textnormal{mean}}})=$ $\textnormal{E}_{Q_{X}}\left\{\textnormal{E}_{P_{Y(a)\mid X}}\left[\left\{Y(1)-\textnormal{E}_{P_{Y(a)\mid X}}[Y(a)\mid X]\right\}^{2}\Bigg|X\right]\right\},\text{ and}$ (31)
$\textnormal{SLOPE}(Q_{Y(a),X}^{0},\psi^{{\textnormal{med}}})=$ $\dfrac{\textnormal{E}_{Q_{X}}\{F_{P_{Y(a)\mid X}}(m_{1/2})\mu(X)\}-\textnormal{E}_{Q_{Y(a)}^{0}}\{Y(a)\mathds{1}(Y(a)\leq m_{1/2})\}}{f_{Q_{Y(a)}^{0}}(m_{1/2})},$ (32)
respectively, where we recall that $m_{1/2}=F_{Q_{Y(a)}^{0}}(1/2)$ is the marginal median under conditional exchangeability and $\mu(X)=\textnormal{E}_{P_{Y(a)\mid X}}\{Y(a)\mid X\}$ is the conditional mean function.
We make some remarks on the SLOPEs in (31) and (32) compared with their counterparts for an observed outcome in Theorems 1 and 2. First, the SLOPE for functionals of a potential outcome is no different from in form from the SLOPE for functionals of an observed outcome, except for a change in notation. This is because the identification strategy in transporting/generalizing a causal quantity is based on the couterfactual outcome $Y(a)$ (i.e., conditional exchangeability in (29), and so as its violation (i.e., the sensitivity model (30)). In this perspective, the SLOPE for a potential outcome should exactly mimic the SLOPE for an observed outcome, as we have seen above.
Second, although the forms look identical, we remind the readers that working with potential outcomes requires more discretion because their SLOPEs (e.g., (31) and (32)) usually involve the counterfactual distribution $P_{Y(a)\mid X}$ which is un-identified without additional causal assumptions. Fortunately, if the randomization procedure in the source population is known, then the identification is straightforward with additional justifiable assumptions. We discuss these assumption in the followings.
Suppose the randomization is known to depend on $X$ and another variable $V$ which are both measured in the source population, then by design, the strong ignorability assumption (Assumption 3) holds on the source population, which in turn provides equivalence between $P_{Y(a)\mid X,V}$ and $P_{Y(a)\mid X,V,A=1}$. Further, under SUTVA (Assumption 4), $P_{Y(a)\mid X,V,A=1}$ equals to $P_{Y\mid X,V,A=1}$. Consequently, one can replace $P_{Y(1)\mid X}$ in the SLOPE with $\int P_{Y\mid X,V,A=1}dP_{V\mid X,A=1}$, which no longer involves counterfactual quantities and thus can be identified. Under these identification assumptions, the SLOPEs can be expressed in observed quantities; see (52) and (53) in Section F.1. In the special case when the randomization solely depends on $X$ (i.e., all confounders can been observed in both the source and the target), we have $P_{Y(1)\mid X}=P_{Y(1)\mid X,A=1}$.
Assumption 3 (Strong Ignorability on $P$; Rosenbaum and Rubin, (1983)).
$P_{X,V}(\cdot)$ is absolute continuous with respect to $P_{X,V\mid A=a}(\cdot)$ and $P_{Y(a)\mid X,V}(\cdot\mid x,v)=P_{Y(a)\mid X,V,A=a}(\cdot\mid x,v)$ almost everywhere $P_{X,V}$.
Assumption 4 (Stable Unit Treatment Variable Assumption (SUTVA)).
$Y=Y(a)$ if $A=a$ on the source population $P$.
Appendix E Supplementary Materials for Estimation
we provide supplementary results and some deferred discussions for the estimation of SLOPE presented in Section 4.2 of the main text. We start with defining some notations in Section E.1. Then we discuss regularity conditions and asymptotic properties of the weighting estimator and the regression estimator in Sections E.2 and E.3, respectively. Next, we detail estimators of SLOPEs for the mean, OLS coefficients, and the median in Sections E.4 to E.6, including the weighting and regression estimators presented in the main text and estimators based on the efficient influence function of the SLOPE (if exists). Then we discuss ways of estimating the nuisance function $\omega(X)$ in Section E.7. Finally, we present a general statement of the efficient influence function for the SLOPE of scalar valued Z-functionals in Section E.8.
E.1 Notation and Setup
As stated in the main, suppose we have i.i.d. samples from the target population with size $n_{q}$ and i.i.d. samples from the source population with size $n_{p}$. We pool the samples together and denote by $T_{i}=1$ if the sample comes from the target distribution $Q$ and $T_{i}=0$ otherwise. Therefore, we have a random sample $\{(X_{i},\textnormal{NA},T_{i}=1),i=1,\cdots,n_{q}\}\cup\{(X_{i},O_{i},T_{i}=0),i=n_{q}+1,\cdots,n\}$ from the pseudo population that combines $Q$ and $P$, where NA means unobserved. We will establish the asymptotic properties for the weighting and regression estimators by re-expressing these estimators as solutions to estimating equations, and then apply standard M-estimation theory (Van der Vaart,, 2000).
After introducing the pseudo population that combines $P$ and $Q$, in this section, we use $\textnormal{pr}(\cdot)$ and $\textnormal{E}(\cdot)$, $f(\cdot)$ without subscripts to denote probabilities, expectations, and densities on this pseudo population, and similarly for conditional quantities. For quantities on a single population (e.g., expectations on $P$ or $Q$ only), there are two equivalent notations. One is to use subscripts, e.g., $\textnormal{E}_{P_{O\mid X}}(O\mid X)$ and $f_{Q_{X}}(X)$, as adopted in the main text. The other is to conditioning on $T$, e.g., $\textnormal{E}(O\mid X,T=0)$ and $f(X\mid T=1)$, We prefer the first approach in order to be consistent with our convention of notation in the main text, while in this section we use the second approach if parts of Section E.7 when the context is clearer.
E.2 Asymptotic Properties for the Weighting Estimator
To start with, we express the SLOPE in equation (11) as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=-\eta_{2}^{-1}\eta_{1},\text{ where}$
$\eta_{2}=\left\{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\dot{s}(O,X,\psi)\right]\right)\right\},\text{ and }\eta_{1}=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}s\left(O,X,\psi\right)\right]\right).$
Similarly, we re-express the weighting estimator by the estimates of $\eta_{1}$ and $\eta_{2}$ as follows,
[…]
$\widehat{\eta}_{1}^{\textnormal{W}}$ $=\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{O_{i}-\widehat{\mu}(X_{i})\}s(O_{i},X_{i},\widehat{\psi}),\text{ and}$
[…]
In order to establish the consistency and asymptotic normality of the weighting estimator, we re-express $\widehat{\eta}_{1}^{\textnormal{W}}$ and $\widehat{\eta}_{2}^{\textnormal{W}}$ as solutions to some estimating equations and then apply standard M-estimation theory. Specifically, consider a vector of parameters $\eta^{\textnormal{W}}=[\eta_{1},\eta_{2},\eta_{3}^{\intercal},\eta_{4}^{\intercal},\eta_{5}^{\intercal}]^{\intercal}$, and suppose these parameters are estimated by solving estimating equations $G^{\textnormal{W}}=\left[\left(g_{1}^{\textnormal{W}}\right)^{\intercal},\left(g_{2}^{\textnormal{W}}\right)^{\intercal},\left(g_{3}^{\textnormal{W}}\right)^{\intercal},\left(g_{4}^{\textnormal{W}}\right)^{\intercal},\left(g_{5}^{\textnormal{W}}\right)^{\intercal}\right]^{\intercal}$, i.e., by
$\sum_{i=1}^{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=0$ (33)
Specifically, elements in $\eta^{\textnormal{W}}$ and estimating equations $G^{\textnormal{W}}$ are defined as follows. The first two parameters, $\widehat{\eta}_{1}^{\textnormal{W}}$ and $\widehat{\eta}_{2}^{\textnormal{W}}$, have been defined above, and their corresponding estimating equations are
$g_{1}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})$ $=\dfrac{1-T_{i}}{\textnormal{pr}(T_{i}=0)}\omega(X_{i},\eta_{5})s(O_{i},X_{i},\eta_{3})\{O_{i}-\mu(X_{i},\eta_{4})\}-\eta_{1}=0,\text{and}$
$g_{2}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})$ $=\dfrac{1-T_{i}}{\textnormal{pr}(T_{i}=0)}\omega(X_{i},\eta_{5})\dot{s}(O_{i},X_{i},\eta_{3})-\eta_{2}=0.$
Next, $\eta_{3}=\psi(Q_{O,X}^{0})$ is defined as the target estimand under conditional exchangeability, and is estimated through
$g_{3}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=\dfrac{1-T_{i}}{\textnormal{pr}(T_{i}=0)}\omega(X_{i},\eta_{5})s(O_{i},X_{i},\eta_{3}).$
Finally, suppose parametric models of $\mu(x)$ and $\omega(x)$ are posited by the researcher with parameters $\eta_{4}$ and $\eta_{5}$, respectively, and let $g_{4}^{\textnormal{W}}$ and $g_{5}^{\textnormal{W}}$ be the corresponding estimating equations (e.g., score functions). Denote the two nuisance functions as $\mu(x,\eta_{5})$ and $\omega(x,\eta_{4})$ and hence their estimates are $\widehat{\mu}(x)=\mu(x,\widehat{\eta}_{5})$ and $\widehat{\omega}(x)=\omega(x,\widehat{\eta}_{4})$. Suppose Let $g_{4}^{\textnormal{W}}$ and $g_{5}^{\textnormal{W}}$ are estimating equations (e.g., score functions) that correspond to $\mu(x)$ and $\omega(x)$, respectively, via parametric models posited by the researcher. To indicate the dependencies on nuisance parameters, we denote these two functions as $\mu(x,\eta_{5})$ and $\omega(x,\eta_{4})$, respectively, and hence their estimates are $\widehat{\mu}(x)=\mu(x,\widehat{\eta}_{5})$ and $\widehat{\omega}(x)=\omega(x,\widehat{\eta}_{4})$.
With $G^{W}$ defined as above, the asymptotic properties of the weighting estimator for SLOPE in (12) can be established under standard regularity conditions in M-estimation theory (Newey and McFadden,, 1994; Van der Vaart,, 2000). Below, Condition 7 parallels the condition in Theorem 2.6 of Newey and McFadden, (1994) for consistency, and Condition 8 parallels assumptions in Theorem 3.4 of Newey and McFadden, (1994) and Theorem 5.31 of Van der Vaart, (2000) for asymptotic normality.
Condition 7 (Regularity Conditions for Consistency for Weighting Estimator).
(i) $\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\}=0$ implies $\eta=\eta^{\textnormal{W}}$. (ii) $\eta^{\textnormal{W}}\in\Theta$ where $\Theta$ is compact. (iii) $G^{\textnormal{W}}$ is continuous at each $\eta\in\Theta$ with probability one and $\textnormal{E}\left\{\sup_{\eta\in\Theta}\lVert G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\rVert\right\}<\infty$. (iv) $\eta_{2}=\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O_{i},X_{i},\psi(Q_{O,X}^{0})\right]\neq 0$ with probability one.
Condition 8 (Regularity Conditions for Asymptotic Normality for Weighting Estimator).
(i) $\eta^{\textnormal{W}}$ lies in the interior of $\Theta$. (ii) $\textnormal{E}\lVert G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})\rVert^{2}<\infty$. (iii) The function class $\left\{G^{\textnormal{W}}:\lVert\eta-\eta^{\textnormal{W}}\rVert<\delta\right\}$ is Donsker for some $\delta>0$ and $\textnormal{E}\lVert G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)-G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})\rVert^{2}\to 0$ as $\eta\to\eta^{\textnormal{W}}$. (iv) The map $\eta\mapsto\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\}$ is differentiable at $\eta^{\textnormal{W}}$ with a non-singular derivative matrix $\Omega^{\textnormal{W}}$ with inverse matrix $V^{\textnormal{W}}$.
With Conditions 7 and 8, we establish the asymptotic properties of the weighting estimator.
Theorem 11 (Weighting Estimator).
Let $\widehat{\textnormal{SLOPE}}^{\textnormal{W}}$ be the weighting estimator in (12) where $\widehat{\eta}_{j}$’s are estimated with (33) and we drop the notation in parentheses of the SLOPE for ease of communication. Suppose Condition 7 holds, then $\widehat{\textnormal{SLOPE}}^{\textnormal{W}}$ converges to SLOPE in probability. Additionally suppose Condition 8 holds, then $\sqrt{n}(\widehat{\textnormal{SLOPE}}^{\textnormal{W}}-\textnormal{SLOPE})$ converges in distribution to a normal distribution with mean zero and variance $(\eta_{2})^{2}V_{11}^{\textnormal{W}}/(\eta_{1})^{4}+V_{22}^{\textnormal{W}}/(\eta_{1})^{2}-2\eta_{2}V^{W}_{12}/(\eta_{1})^{3}$, where $V^{\textnormal{W}}$ is inverse of the derivative matrix of $\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})\}$ with respect to $\eta^{\textnormal{W}}$, with $V^{\textnormal{W}}_{ij}$ denoting its entry at the $i$-th row and $j$-th column.
We note that depending on the target estimand, some nuisance parameters listed above may be trivial. For example, for the SLOPE for the mean, $\eta_{2}=-1$ need not be estimated since $\dot{s}=-1$ is constant. In addition, we provide example estimators for other estimands later in this section and discuss ways of estimating $\omega(x)$ in Section E.
Finally, we remark on the weighting estimator for SLOPE when the target functional $\psi(\cdot)$ is vector valued.
Remark 5 (Weighting Estimator with a Vector Valued $\psi$).
With a vector valued target functional $\psi(\cdot)$, the target estimand and the SLOPE become vectors of the same dimension, say $p$. The SLOPE formula derived from the IF (i.e., (11)) still holds and the weighting estimator (12) is still applicable. The difference is that $\eta_{1}$ becomes a vector of length $p$ and $\eta_{2}$ becomes a $p$ by $p$ matrix. To establish the asymptotic properties, we need to adjust the vector of nuisance parameter, $\eta^{\textnormal{W}}$, mainly to include a vectorization of $\eta_{2}$ instead of $\eta_{2}$ itself. More concretely, let $\eta^{\textnormal{W}}=[\eta_{1}^{\intercal},\{\vec{(}\eta_{2})\}^{\intercal},\eta_{3}^{\intercal},\eta_{4}^{\intercal},\eta_{5}^{\intercal}]^{\intercal}$, where $\vec{(}\cdot)$ is vectorization of a matrix, and let $G^{\textnormal{W}}$ be modified such that $g_{2}^{\textnormal{W}}$ corresponds to $\vec{(}\eta_{2})$. Then under the same regularity conditions on the updated $G^{\textnormal{W}}$ and $\eta^{\textnormal{W}}$, the weighting estimator is still consistent and asymptotically normal. The asymptotic variance is $A\Sigma^{\textnormal{W}}A^{\intercal}$, where $\Sigma^{\textnormal{W}}$ is the first $(p^{2}+p)$ diagonal matrix of $V^{\textnormal{W}}$, and $A=\left[-(\textnormal{SLOPE})^{\intercal}\otimes(\eta_{2})^{-1},\eta_{2}^{-1}\right]$.
E.3 Asymptotic Properties for the Regression Estimator
We re-express the regression estimator (13) with $\widehat{\eta}_{1}^{\textnormal{R}}$ and $\widehat{\eta}^{\textnormal{R}}_{2}$, the regression typed estimators for $\eta_{1}$ and $\eta_{2}$ respectively, as follows:
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}(\widehat{Q}_{O,X}^{0},\widehat{\psi})=-\left(\eta_{2}^{\textnormal{R}}\right)^{-1}\eta_{1}^{\textnormal{R}},$ where
$\eta_{1}=\sum_{i=1}^{n_{q}}\widehat{\textnormal{E}}_{P_{O\mid X}}\left[\{O_{i}-\widehat{\mu}(X_{i})\}s(O_{i},X_{i},\psi)\mid X_{i}\right],$ $\text{ and }\eta_{2}=\left[\sum_{i=1}^{n_{q}}\widehat{\textnormal{E}}_{P_{O\mid X}}\{\dot{s}(O_{i},X_{i},\psi)\mid X_{i}\}\right].$
Next, we establish the asymptotic properties of the regression estimator in a similar way as done for the weighting estimator. Specifically, we consider nuisance parameters $\eta^{\textnormal{R}}=[\eta_{1},\eta_{2},\eta_{3},\eta_{4}^{\intercal},\eta_{6}^{\intercal},\eta_{7}^{\intercal},\eta_{8}^{\intercal}]^{\intercal}$ estimated through
$\sum_{i=1}^{n}G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})=0$ (34)
with estimating equations $G^{\textnormal{R}}=\left[g_{1}^{\textnormal{R}},g_{2}^{\textnormal{R}},g_{3}^{\textnormal{R}},\left(g_{4}^{\textnormal{R}}\right)^{\intercal},\left(g_{6}^{\textnormal{R}}\right)^{\intercal},\left(g_{7}^{\textnormal{R}}\right)^{\intercal},\left(g_{8}^{\textnormal{R}}\right)^{\intercal}\right]^{\intercal}$. Specifically, $\eta_{1}$ to $\eta_{4}$ have been defined previously with these new estimating equations based on regressions on the target samples instead of weighting on the source samples,
$g_{1}^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})$ $=\dfrac{T_{i}}{\textnormal{pr}(T_{i}=1)}\textnormal{E}_{P_{O\mid X}}\left[s(O_{i},X_{i},\eta_{3})\{O_{i}-\mu(X_{i})\}\mid X_{i},\eta_{6}\right]-\eta_{1},$
$g_{2}^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})$ $=\dfrac{T_{i}}{\textnormal{pr}(T_{i}=1)}\textnormal{E}_{P_{O\mid X}}\left\{\dot{s}(O_{i},X_{i},\eta_{3})\mid X_{i},\eta_{7}\right\}-\eta_{2},$
$g_{3}^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})$ $=\dfrac{T_{i}}{\textnormal{pr}(T_{i}=1)}\textnormal{E}_{P_{O\mid X}}\{s(O_{i},X_{i},\eta_{3})\mid X_{i},\eta_{8}\},$
and $g_{4}^{\textnormal{R}}=g_{4}^{\textnormal{W}}$ is the estimating equation for the regression function $\mu(x)$ which is now denoted as $\mu(x,\eta_{4})$ as in Section E.2. As denoted in the proceeding formulas, the additional nuisance parameters, $\eta_{6}$, $\eta_{7}$, and $\eta_{8}$ represent the nuisance parameters in parametric models for $\textnormal{E}_{P_{O\mid X}}\{s(O_{i},X_{i},\eta_{3})\{O_{i}-\mu(X_{i})\mid X\}$, $\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O_{i},X_{i},\eta_{3})\mid X\}$ and $\textnormal{E}_{P_{O\mid X}}\{s(O_{i},X_{i},\eta_{3})\mid X\}$, respectively, with estimating equations $g_{6}^{\textnormal{R}}$, $g_{7}^{\textnormal{R}}$ and $g_{8}^{\textnormal{R}}$. Depending on the specific estimand, some nuisance parameters may be trivial and the estimation is simpler than what has been shown. That being said, using regression based estimator typically involves more nuisance functions/parameters than the weighting estimator.
The regularity conditions for the regression estimator are standard and are similar to the conditions for the weighting estimator. Condition 9 parallels the condition in Theorem 2.6 of Newey and McFadden, (1994) for consistency and Condition 10 parallels assumptions in Theorem 3.4 of Newey and McFadden, (1994) and Theorem 5.31 of Van der Vaart, (2000) for asymptotic normality.
Condition 9 (Regularity Conditions for Consistency for Regression Estimator).
(i) $\textnormal{E}\{G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta)\}=0$ implies $\eta=\eta^{\textnormal{R}}$. (ii) $\eta^{\textnormal{R}}\in\Theta$ where $\Theta$ is compact. (iii) $G^{\textnormal{R}}$ is continuous at each $\eta\in\Theta$ with probability one and $\textnormal{E}\left\{\sup_{\eta\in\Theta}\lVert G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta)\rVert\right\}<\infty$. (iv) $\eta_{2}=\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O_{i},X_{i},\psi(Q_{O,X}^{0})\right]\neq 0$ with probability one.
Condition 10 (Regularity Conditions for Asymptotic Normality for Regression Estimator).
(i) $\eta^{\textnormal{R}}$ lies in the interior of $\Theta$. (ii) $\textnormal{E}\lVert G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})\rVert^{2}<\infty$. (iii) The function class $\left\{G^{\textnormal{R}}:\lVert\eta-\eta^{\textnormal{R}}\rVert<\delta\right\}$ is Donsker for some $\delta>0$ and $\textnormal{E}\lVert G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta)-G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})\rVert^{2}\to 0$ as $\eta\to\eta^{\textnormal{R}}$. (iv) The map $\eta\mapsto\textnormal{E}\{G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta)\}$ is differentiable at $\eta^{\textnormal{R}}$ with a non-singular derivative matrix $\Omega^{\textnormal{R}}$ with inverse matrix $V^{\textnormal{R}}$.
Under regularity conditions listed above, the consistency and asymptotic normality of the regression estimator is presented as follows.
Theorem 12 (Regression Estimator).
Let $\widehat{\textnormal{SLOPE}}^{\textnormal{R}}$ be the regression estimator in (13) where $\eta_{j}$’s are estimated with (34) and we drop the notation in parentheses of the SLOPE for ease of communication. Suppose Condition 9 in the Supplement holds, then $\widehat{\textnormal{SLOPE}}^{\textnormal{R}}$ converges to SLOPE in probability. Additionally suppose Condition 10 in the Supplement holds, then $\sqrt{n}(\widehat{\textnormal{SLOPE}}^{\textnormal{R}}-\textnormal{SLOPE})$ converges in distribution to a Gaussian distribution with mean zero and variance $(\eta_{2})^{2}V_{11}^{\textnormal{R}}/(\eta_{1})^{4}+V_{22}^{\textnormal{R}}/(\eta_{1})^{2}-2\eta_{2}V^{\textnormal{R}}_{12}/(\eta_{1})^{3}$, where $V^{\textnormal{R}}$ is the derivative matrix of $\textnormal{E}\{G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})\}$ with respect to $\eta^{\textnormal{R}}$, with $V^{\textnormal{R}}_{ij}$ denoting its entry at the $i$-th row and $j$-th column.
Remark 6 (Regression Estimator with a Vector Valued $\psi$).
When the target functional is vector valued, the regression estimator is still applicable with a similar argument as in Remark 5 for the weighting estimator.
E.4 Estimating SLOPE for The Mean
We elaborate on the estimators of the SLOPE for the mean, $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})$. In Section 4.2, we have presented two estimators, a weighting estimator and a regression estimator,
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})=$ $\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{O_{i}-\widehat{\mu}(X_{i})\}(O_{i}-\widehat{\psi}^{{\textnormal{mean}}}),$ (35)
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})=$ $\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i}).$ (36)
To implement the two estimators, one can resort to Section E.7 for estimating $\omega(x)$ and any regression method for estimating $\mu(x)$. In addition, $\widehat{\psi}^{{\textnormal{mean}}}$ can be obtained by a weighted average of outcomes over source samples, i.e., $\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})O_{i}/{n_{p}}$, and $\widehat{\sigma}^{2}(x)$ can be obtained by regressing the squared residual, $\{O_{i}-\widehat{\mu}(X_{i})\}^{2}$ over $X_{i}$.
Next we motivate the estimator based on the efficient influence function. By expressing the SLOPE explicitly,
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}^{2}\mid X\right]\right),$
an alternative weighting estimator is naturally motivated:
$\widehat{\textnormal{SLOPE}}^{\text{W,Alt}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})=$ $\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{O_{i}-\widehat{\mu}(X_{i})\}^{2}.$ (37)
Moreover, a fourth estimator is motivated by combining properties of the alternative weighting estimator (37) with the regression estimator (36), i.e.,
$\widehat{\textnormal{SLOPE}}^{\text{EIF}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})=$ $\widehat{\textnormal{SLOPE}}^{\text{W,Alt}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})-\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{\sigma}^{2}(X_{i})+\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i})$
$=$ $\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left[\{O_{i}-\widehat{\mu}(X_{i})\}^{2}-\widehat{\sigma}^{2}(X_{i})\right]+\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i}).$ (38)
Since this estimator (38) can be naturally motivated from the efficient influence function (EIF) of the SLOPE (see Section E.8 below), we refer to it as the EIF-based estimator and denote it using the superscript “EIF”. Under standard regularity conditions, the EIF-based estimator is consistent and asymptotically normal,
$\sqrt{n}\left\{\widehat{\textnormal{SLOPE}}^{\text{EIF}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})-{\textnormal{SLOPE}}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})\right\}\to_{d}N\left(0,\textnormal{E}\{\textnormal{EIF}(T,O,X,\textnormal{SLOPE})\}\right),$
where $\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$ is the efficient influence function of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})$ and it takes the form of
$\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$ $=\dfrac{1-T}{\textnormal{pr}(T=0)}\omega(X)\{O-\mu(X)\}^{2}+\dfrac{T}{\textnormal{pr}(T=1)},$
where $\textnormal{pr}(T=1)$ is the probability limit of $n_{q}/n$ and $\textnormal{pr}(T=0)=1-\textnormal{pr}(T=1)$; see Section E.8 (in particular, Proposition 1) for a general formula of the EIF for the SLOPE. Therefore, the asymptotic variance can be estimated with
$\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\left\{\widehat{\sigma}^{2}(X_{i})-\widehat{\textnormal{SLOPE}}^{\text{EIF}}\right\}+\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left[\{O_{i}-\widehat{\mu}(X_{i})\}^{2}-\widehat{\sigma}^{2}(X_{i})\right],$ (39)
which is consistent under regularity conditions. Consequently, the standard error (SE) of the EIF-based estimator can be estimated with
$\dfrac{1}{\sqrt{n}}\sqrt{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\left\{\widehat{\sigma}^{2}(X_{i})-\widehat{\textnormal{SLOPE}}^{\text{EIF}}\right\}+\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left[\{O_{i}-\widehat{\mu}(X_{i})\}^{2}-\widehat{\sigma}^{2}(X_{i})\right]}.$
E.5 Estimating SLOPE for OLS Coefficients
Suppose $O$ is an outcome variable and $X$ is a vector of covariates that includes $1$ as the first component. We are interested in the OLS coefficient of regressing $O$ on $X$ in the target distribution, i.e., $\psi^{{\textnormal{OLS}}}(Q_{O,X})$ such that $\textnormal{E}_{Q_{O,X}}\left[XX^{\intercal}\psi^{{\textnormal{OLS}}}-XO\right]=0$. Then according to Theorem 7, the SLOPE for the OLS coefficient $\psi^{{\textnormal{OLS}}}$ is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$ $=\left\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\right\}^{-1}\textnormal{E}_{Q_{X}}\{X\sigma^{2}(X)\}.$
To estimate the SLOPE, we consider three estimators based on weighting, regression, and the efficient influence function:
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{OLS}}}\right)=\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{r}_{i}^{2}X_{i},$ (40)
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{OLS}}}\right)=\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i})X_{i}$ (41)
$\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{OLS}}}\right)=\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\left[\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{\widehat{r}_{i}^{2}-\widehat{\sigma}^{2}(X_{i})\}X_{i}+\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}\widehat{\sigma}^{2}(X_{i})\right],$ (42)
where $\widehat{r}_{i}=O_{i}-\widehat{\mu}(X_{i})$ is the regression residuals on the source data.
In addition, for the EIF-based estimator, the variance can be consistently estimated by
$\dfrac{1}{n}\sum_{i=1}^{n}\widehat{\textnormal{EIF}}\left(T_{i},O_{i},X_{i},\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}}\right)\left\{\widehat{\textnormal{EIF}}(T_{i},O_{i},X_{i},\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}})\right\}^{\intercal},$ (43)
where
$\widehat{\textnormal{EIF}}\left(T_{i},O_{i},X_{i},\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}}\right)=$ $\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\left[\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})X_{i}\{\widehat{r}_{i}^{2}-\widehat{\sigma}^{2}(X_{i})\}\right]$
$+\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\left[\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\left\{-X_{i}^{\intercal}\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}}+\widehat{\sigma}^{2}(X_{i})\right\}X_{i}\right].$
E.6 Estimating SLOPE for The Median
In this section, we consider estimating the SLOPE for the median. We start with a simpler case when $P_{O\mid X}$ is Gaussian, i.e., $P_{O\mid X}\sim N\left(\mu(X),\sigma^{2}(X)\right)$. Then by part (ii) of Theorem 2, the SLOPE for the median is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=\dfrac{\textnormal{E}_{Q_{X}}\left[\sigma(X)\varphi(\{m_{1/2}-\mu(X)\}/\sigma(X))\right]}{\textnormal{E}_{Q_{X}}\left[\varphi(\{m_{1/2}-\mu(X)\}/\sigma(X))/\sigma(X)\right]}.$
This motivates the following weighting and regression estimators:
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{\sigma}(X_{i})\varphi\left(\{\widehat{m}_{1/2}-\widehat{\mu}(X_{i})\}/\widehat{\sigma}(X_{i})\right)}{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\varphi\left(\{\widehat{m}_{1/2}-\widehat{\mu}(X_{i})\}/\widehat{\sigma}(X_{i})\right)/\widehat{\sigma}(X_{i})},$ (44)
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}(X_{i})\varphi\left(\{\widehat{m}_{1/2}-\widehat{\mu}(X_{i})\}/\widehat{\sigma}(X_{i})\right)}{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\varphi\left(\{\widehat{m}_{1/2}-\widehat{\mu}(X_{i})\}/\widehat{\sigma}(X_{i})\right)/\widehat{\sigma}(X_{i})},$ (45)
where $\varphi(\cdot)$ is the density of the standard normal distribution.
Next, we consider the general case with SLOPE given by (7). Since it involves conditional densities, the efficient influence function does not exist in general. We will consider the weighting estimator and the regression estimator only. Let $\widehat{F}_{P_{O\mid X}}$, $\widehat{f}_{P_{O\mid X}}$ and $\widehat{\textnormal{E}}_{P_{O\mid X}}\left\{O\mathds{1}(O\leq\widehat{m}_{1/2})\right\}$ be estimates of the c.d.f. $F_{P_{O\mid X}}$, the p.d.f. $f_{P_{O\mid X}}$, and the truncated mean $\textnormal{E}_{P_{O\mid X}}\left\{O\mathds{1}(O\leq\widehat{m}_{1/2})\right\}$, respectively, and $\widehat{m}_{1/2}$ be an estimate of $m_{1/2}$ where all estimates are based on parametric models. Then the weighting and regression estimators are
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left\{\widehat{F}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})\widehat{\mu}(X_{i})-O_{i}\mathds{1}(O_{i}\leq\widehat{m}_{1/2})\right\}}{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{f}_{P_{O\mid X}}(\widehat{m}_{1/2})},\text{ and}$
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{F}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})\widehat{\mu}(X_{i})-\widehat{\textnormal{E}}_{P_{O\mid X}}\left\{O_{i}\mathds{1}(O_{i}\leq\widehat{m}_{1/2})\mid X_{i}\right\}}{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{f}_{P_{O\mid X}}(\widehat{m}_{1/2})},$
respectively. In practice, one may impose parametric assumptions on $P_{O\mid X}$ and estimate $F_{P_{O\mid X}}$, $f_{P_{O\mid X}}$ and $\textnormal{E}_{P_{O\mid X}}\left\{O\mathds{1}(O\leq\widehat{m}_{1/2})\right\}$ accordingly. Then the marginal $m_{1/2}$ can be estimated by numerically solving either of the following equations using bisection,
$\sum_{i=1}^{n_{q}}\widehat{F}_{P_{O\mid X}}(m_{1/2}\mid X_{i})$ $=1/2,\text{ or}$
$\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{F}_{P_{O\mid X}}(m_{1/2}\mid X_{i})$ $=1/2.$
E.7 Example Estimating Equations for $\omega(x)$
In this section, we discuss some common approaches to estimating the density ratio $\omega(x)$, including the logistic regression, entropy balancing, and a method based on empirical distributions when the support of $Q_{X}$ (i.e., $\mathcal{S}_{X}$) is finite and discrete. Specifically, following the notation in Section E.2, we build estimating equations $g_{5}^{\textnormal{W}}$ for the nuisance parameter $\eta_{5}$ and after obtaining the estimate $\widehat{\eta}_{5}$, we present $\widehat{\omega}(x)$ in terms of $\widehat{\eta}_{5}$. For the sake of exposition, we take the convention that $X$ does not include one (i.e., a constant that can serve as an intercept in regression models).
E.7.1 Logistic Regression
Let $\textnormal{pr}(T=1\mid x)=\textnormal{pr}(T=1\mid X=x)$ be the probability of being included in the target sample. Since the density ratio $\omega(x)$ can be re-formulated in terms of $\textnormal{pr}(T=1\mid x)$ via Bayes rule,
$\omega(x)$ $=\dfrac{f(x\mid T=1)}{f(x\mid T=0)}$ (By definition)
$=\dfrac{\textnormal{pr}(T=1\mid x)}{\textnormal{pr}(T=0\mid x)}\dfrac{\textnormal{pr}(T=0)}{\textnormal{pr}(T=1)}$ $\text{ (By Bayes rule)},$
one common strategy to estimate $\omega(x)$ is to first estimate the conditional probability function $\textnormal{pr}(T=1\mid x)$ using some binary classification/regression model, and then plug in the estimates $\widehat{\textnormal{pr}}(T=1\mid x)$ to obtain $\widehat{\omega}(x)$,
$\widehat{\omega}(x)=\dfrac{\widehat{\textnormal{pr}}(T=1\mid x)}{\widehat{\textnormal{pr}}(T=0\mid x)}\dfrac{\widehat{\textnormal{pr}}(T=0)}{\widehat{\textnormal{pr}}(T=1)}=\dfrac{\widehat{\textnormal{pr}}(T=1\mid x)}{\widehat{\textnormal{pr}}(T=0\mid x)}\dfrac{n_{p}}{n_{q}},$ (46)
where $\widehat{\textnormal{pr}}(T=0)=n_{p}/n$ and $\widehat{\textnormal{pr}}(T=1)=n_{q}/n$.
Next, we list the estimation equations based on maximum likelihood estimation when $\textnormal{pr}(T=1\mid x)$ is modeled by logistic regression, one of the most popular binary regression models. This can be implemented in R using the built-in function glm. To demonstrate the estimating equation for the nuisance parameter $\eta_{5}$, we let $\eta_{5}=[\alpha_{\textnormal{LR}},\beta_{\textnormal{LR}}^{\intercal}]^{\intercal}$ where $\alpha_{\textnormal{LR}}$ and $\beta_{\textnormal{LR}}$ are the intercept and slope coefficients in the logistic regression model. Then the likelihood can be expressed as
$l_{\textnormal{LR}}(\alpha_{\textnormal{LR}},\beta_{\textnormal{LR}})$ $=\sum_{T_{i}=0}(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X_{i})-\sum_{i=1}^{n}\log\left(1+\exp\{\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X_{i}\}\right),$
By setting $\partial l_{\textnormal{LR}}/\partial\alpha_{\textnormal{LR}}=0$ and $\partial l_{\textnormal{LR}}/\partial\beta_{\textnormal{LR}}=0$ and checking the second-order conditions, parameters $\alpha_{\textnormal{LR}}$ and $\beta_{\textnormal{LR}}$ can be estimated by $\widehat{\alpha}_{\textnormal{LR}}$ and $\widehat{\beta}_{\textnormal{LR}}$ which are solutions to the following equations,
$n_{p}-\sum_{i=1}^{n}\dfrac{\exp(\widehat{\alpha}_{\textnormal{LR}}+\widehat{\beta}_{\textnormal{LR}}^{\intercal}X_{i})}{1+\exp(\widehat{\alpha}_{\textnormal{LR}}+\widehat{\beta}_{\textnormal{LR}}^{\intercal}X_{i})}=0,\text{ and}\sum_{T_{i}=0}X_{i}-\sum_{i=1}^{n}\dfrac{X_{i}\exp(\widehat{\alpha}_{\textnormal{LR}}+\widehat{\beta}_{\textnormal{LR}}^{\intercal}X_{i})}{1+\exp(\widehat{\alpha}_{\textnormal{LR}}+\widehat{\beta}_{\textnormal{LR}}^{\intercal}X_{i})}.$
Therefore, the estimating equation $g_{5}^{\textnormal{W}}$ is
$g_{5}^{\textnormal{W}}(T,O,X,\eta^{\textnormal{W}})$ $=\begin{bmatrix}\mathds{1}(T=0)-\dfrac{\exp(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X)}{1+\exp(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X)}\\ {}\\ \mathds{1}(T=0)X-\dfrac{X\exp(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X)}{1+\exp(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X)}\end{bmatrix}$,
and the estimate $\widehat{\eta}_{5}=[\widehat{\alpha}_{\textnormal{LR}},\widehat{\beta}_{\textnormal{LR}}^{\intercal}]^{\intercal}$ is obtained by setting $\sum_{i=1}^{n}g_{5}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=0$. After estimating these parameters, by (46), the resulting estimate for the density ratio is
$\widehat{\omega}(x)$ $=\exp\left\{-\left[\widehat{\alpha}_{\textnormal{LR}}-\log(n_{p}/n_{q})\right]-\widehat{\beta}_{\textnormal{LR}}^{\intercal}x\right\}.$
E.7.2 Entropy Balancing
Note that for any measurable function $h(x)$, the density ratio $\omega(x)$ satisfies that
$\textnormal{E}\{\omega(X)h(X)\mid T=0\}=\textnormal{E}\{h(X)\mid T=1\}.$
In other words, $\omega(x)$ reweighs the source population in order to match the target population. This property motivates estimating $\omega(x)$ by balancing functions (usually moments) of the source samples so that they match the target samples.
In this section, we introduce the method of entropy balancing and aim at balancing the first moments of $X$. Specifically, for source samples with $i=1+n_{q},\cdots,n$, suppose $\omega_{i}=\omega(X_{i})$ are the weights. Entropy balancing seeks for $\omega_{i}$’s that maximizes their entropy as well as balances the first moment of $X$:
$\begin{split}&\underset{w_{i}}{\rm argmin}\sum_{T_{i}=0}w_{i}\log(w_{i}),\quad{}\textnormal{s.t. }\dfrac{1}{n_{p}}\sum_{T_{i}=0}w_{i}X_{i}=\dfrac{1}{n_{q}}\sum_{T_{i}=1}X_{i}.\end{split}$ (47)
According to Lee et al., (2023); Chen et al., 2023a , solutions to (47), denoted as $\widehat{\omega}_{i}$, can be expressed by
$\widehat{\omega}_{i}=\widehat{\omega}(X_{i})=\exp\left(-\widehat{\alpha}_{\textnormal{EB}}-\widehat{\beta}_{\textnormal{EB}}^{\intercal}X_{i}\right),$ (48)
where $\widehat{\alpha}_{\textnormal{EB}}$ and $\widehat{\beta}_{\textnormal{EB}}$ satisfy
$\sum_{T_{i}=0}\exp(\widehat{\alpha}_{\textnormal{EB}}-\widehat{\beta}_{\textnormal{EB}}^{\intercal}X_{i})=n_{p},\quad\dfrac{1}{n_{p}}\sum_{T_{i}=0}\exp\left(\widehat{\alpha}_{\textnormal{EB}}-\widehat{\beta}_{\textnormal{EB}}^{\intercal}X_{i}\right)=\dfrac{1}{n_{q}}\sum_{T_{i}=1}X_{i}.$
Following the notation in the main text, let the nuisance parameter be $\eta_{5}=[\alpha_{\textnormal{EB}},\beta_{\textnormal{EB}}^{\intercal}]^{\intercal}$, then the estimating equation $g_{5}^{\textnormal{W}}$ is
$g_{5}^{\textnormal{W}}(T,O,X,\eta^{\textnormal{W}})=\begin{bmatrix}\mathds{1}(T=0)\exp(-\alpha_{\textnormal{EB}}-\beta_{\textnormal{EB}}^{\intercal}X)-\mathds{1}(T=0)\\ \mathds{1}(T=0)X_{i}\exp(-\alpha_{\textnormal{EB}}-\beta_{\textnormal{EB}}^{\intercal}X)-\dfrac{n_{p}}{n_{q}}\mathds{1}(T=1)X\end{bmatrix}$,
and the estimate $\widehat{\eta}_{5}=[\widehat{\alpha}_{\textnormal{EB}},\widehat{\beta}_{\textnormal{EB}}^{\intercal}]^{\intercal}$ is obtained by setting $\sum_{i=1}^{n}g_{5}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=0$. After estimating these parameters, by (48), the resulting estimate for the density ratio is
$\widehat{\omega}(x)$ $=\exp\left(\widehat{\alpha}_{\textnormal{EB}}-\widehat{\beta}_{\textnormal{EB}}^{\intercal}x\right).$
We note that the two methods, entropy balancing and logistic regression, are related in that when models are correctly specified, $\beta_{\textnormal{EB}}=\beta_{\textnormal{LR}}$ and $\alpha_{\textnormal{EB}}=\alpha_{\textnormal{LR}}-\log(\textnormal{pr}(T=0)/\textnormal{pr}(T=1))$, while their estimates are numerically different due to different first-order conditions (Zhao and Percival,, 2017). In practice, we recommend using entropy balancing over logistic regression since by enabling covariate balancing, entropy balancing is more robust when models are slightly mis-specified (Imai and Ratkovic,, 2014; Zhao, 2019a, ).
E.7.3 Estimation for Discrete Covariates
Suppose $X$ is discrete with a finite support, $\mathcal{S}_{X}=\{x_{1},x_{2},\cdots,x_{K}\}$ where $K$ is fixed. Then $\omega(x)$ can be estimated by the ratio of the empirical distributions of $f_{Q_{X}}(x)$ and $f_{P_{X}}(x)$ for $x\in\mathcal{S}_{X}$, i.e.,
$\widehat{\omega}(x)=\dfrac{\sum_{T_{i}=1}\mathds{1}(X_{i}=x)/n_{q}}{\sum_{T_{i}=0}\mathds{1}(X_{i}=x)/n_{p}}.$ (49)
Following the notation in Section E.2, let $\eta_{5}=[w_{1},w_{2},\cdots,w_{K}]^{\intercal}$ where $w_{k}=f_{Q_{X}}(x_{k})/f_{P_{X}}(x_{k})$ for $k=1,2,\cdots,K$. Then by defining the estimating equations as
$g_{5}^{\textnormal{W}}(T,O,X,\eta^{\textnormal{W}})$ $=\begin{bmatrix}\dfrac{T\mathds{1}(X=x_{1})}{n_{q}}-w_{1}\cdot\dfrac{(1-T)\mathds{1}(X=x_{1})}{n_{p}}\\ \ldots\\ \dfrac{T\mathds{1}(X=x_{K})}{n_{q}}-w_{K}\cdot\dfrac{(1-T)\mathds{1}(X=x_{K})}{n_{p}}\end{bmatrix}$,
the nuisance parameter can be estimated by $\widehat{\eta}_{5}=[\widehat{w}_{1},\widehat{w}_{2},\cdots,\widehat{w}_{5}]$ where $\widehat{w}_{k}$’s are estimated by setting $\sum_{i=1}^{n}g_{5}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=0$. Then $\widehat{\omega}(x_{k})$ can be estimated with $\widehat{w}_{k}$ for $x_{k}\in\mathcal{S}_{X}$.
E.8 General Statements for The Efficient Influence Function
In this section, we derive the efficient influence function (EIF) of SLOPE when the target functional (and therefore the SLOPE) is scalar valued. Let $\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$ be EIF for $\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$. It is provided by Proposition 1. Additionally, the next Proposition 2 gives the EIF of the target functional, denoted as $\textnormal{EIF}(T,O,X,\psi)$.
Proposition 1 (EIF of SLOPE).
Suppose $\psi$ and $s$ are one-dimensional. The efficient influence function of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ is
$\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\cdot\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$
$=$ $\dfrac{(1-T)\omega(X)}{\textnormal{pr}(T=0)}\left[-s(O,X,\psi)\{O-\mu(X)\}+\textnormal{Cov}_{P_{O\mid X}}[s(O,X,\psi),O\mid X]+\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\{O-\mu(X)\}\right]$
$-\dfrac{T}{\textnormal{pr}(T=1)}\textnormal{E}_{P_{O\mid X}}[s(O,X,\psi)\{O-\mu(X)\}\mid X]$
$-\textnormal{SLOPE}(Q_{O,X}^{0},\psi)\left(\dfrac{(1-T)\omega(X)}{\textnormal{pr}(T=0)}\left[\dot{s}(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}\right]-\dfrac{T}{\textnormal{pr}(T=1)}\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}\right)$
$-\textnormal{EIF}(T,O,X,\psi)\left(\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]+\textnormal{SLOPE}(Q_{O,X}^{0},\psi)\cdot\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}\right),$
where $\textnormal{EIF}(T,O,X,\psi)$ be the EIF of the target functional $\psi(Q_{O,X}^{0})$ (see Proposition 2).
Proposition 2 (EIF of The Target Functional).
The efficient influence function of $\psi(Q_{O,X}^{0})$ is
$\textnormal{EIF}(T,O,X,\psi)=$ $-\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\dfrac{(1-T)\omega(X)}{\textnormal{pr}(T=0)}\left[\{s(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\right]$
$-\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\dfrac{T}{\textnormal{pr}(T=1)}\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}.$
We elaborate these two propositions on a few examples. First, suppose the target functional is $\psi^{{\textnormal{mean}}}$ with $s(O,X,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$. Then $\dot{s}(O,X,\psi)=-1$ and $\ddot{s}(O,X,\psi)=0$. With Proposition 2, the EIF for the target functional is
$\textnormal{EIF}(T,O,X,\psi^{{\textnormal{mean}}})$ $=\dfrac{1-T}{\textnormal{pr}(T=0)}\omega(X)\{O-\mu(X)\}+\dfrac{T}{\textnormal{pr}(T=1)}\left\{\mu(X)-\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})\right\}.$ (50)
This is identical to the EIF derived by Zeng et al., (2023) in their special case when the source and target samples share the same set of covariates. Estimators for the mean based on this EIF were proposed by Dahabreh et al., (2020) and then by Zeng et al., (2023) in a more general setting.
Further, by Proposition 1, the EIF of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})$ is
$\begin{split}\textnormal{EIF}(T,O,X,\textnormal{SLOPE})=&\dfrac{1-T}{\textnormal{pr}(T=0)}\omega(X)\left[\{O-\mu(X)\}^{2}-\sigma^{2}(X)\right]\\ &+\dfrac{T}{\textnormal{pr}(T=1)}\left\{\sigma^{2}(X)-\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})\right\}.\end{split}$ (51)
(51) consists of two parts. The first part is indexed by 1T1-T and it is a weighted, mean-zero term involving the source data. The weight is the density ratio $\omega(X)$ that reweighs the source covariates to match the target population, and the mean zero part can be viewed as the residual of estimating $\sigma^{2}(X)=\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}^{2}\mid X\right]$. The second part is indexed by $T$ and it can be viewed as the conditional variance $\sigma^{2}(X)$ re-centered over the target population. By estimating nuisance functions $\mu(X)$, $\omega(X)$, and $\sigma^{2}(X)$ and setting the summation of the empirical EIFs to zero, we obtain an EIF-based estimator that was presented in (38).
For the second example, suppose the target functional is the OLS coefficient $\psi^{{\textnormal{OLS}}}$ where $s(O,X)=\textnormal{E}_{Q_{O,X}}\left[XX^{\intercal}\psi^{{\textnormal{OLS}}}-XO\right]=0$ with $X$ includes an intercept term as its first component. We notice that although Proposition 1 has been stated for one-dimensional parameters, it is also valid for the $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$ because the corresponding $\ddot{s}(O,X,\psi^{{\textnormal{OLS}}})=0$. Consequently, the EIF for $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$ is
$\textnormal{EIF}(T,O,X,\textnormal{SLOPE})=$ $\left\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\right\}^{-1}\left\{\dfrac{1-T}{\textnormal{pr}(T=0)}\omega(X)\left[\{O-\mu(X)\}^{2}-\sigma^{2}(X)\right]+\dfrac{T}{\textnormal{pr}(T=1)}\sigma^{2}(X)\right\}.$
After presenting these examples of the EIF, finally, we comment on some limitations for estimators based on the EIF. First, the EIF for the SLOPE does not always exist. For example, when the target estimand is the median, $\psi^{{\textnormal{med}}}$, the SLOPE involves the conditional density $f_{P_{O\mid X}}$, and hence the EIF does not exist. Second, unlike many well-known examples in the literature, the SLOPE estimators based on the EIF often do not enjoy the double robustness property. In the simplest case when the target functional is the mean, $\psi^{{\textnormal{mean}}}$, the EIF-based estimator (38) involves three nuisance functions, $\omega(X)$, $\mu(X)$, and $\sigma^{2}(X)$. Roughly speaking, this estimator is conditionally doubly robust in that as long as $\widehat{\mu}(x)$ is consistent, the estimator will be consistent when either $\widehat{\omega}(x)$ or $\widehat{\mu}(x)$ is consistent. A similar result for the conditional rate double robustness holds under standard regularity conditions or cross fitting procedures. Due to these limitations and the complexity of the EIF, in practice we recommend using weighting or regression estimators presented in Section 4.2 of the main text. Meanwhile, we implement all three estimators in the numeric simulations (Section G) and we recognize the development of robust estimators of the SLOPE as an import future direction.
Appendix F Supplementary Materials for Data Application
In this section, we provide supplementary results for the data application in Section 5 of the main text.
F.1 Causal Identification
On the source population, we assume SUTVA holds (Assumption 4). Let $X$ be the baseline measurement and $V$ be the village that a household belongs to. We assume the strong ignorability assumption holds with $X$ and $V$, i.e., Assumption 3. We note that this holds by the design of Banerjee et al., (2015). Under these identification assumptions, the SLOPE for the mean becomes
$\textnormal{SLOPE}(Q_{Y(a),X}^{0},\psi^{{\textnormal{mean}}})$
$=$ $\textnormal{E}_{Q_{X}}\left\{\textnormal{Var}_{P_{Y(a)\mid X}}[Y(a)\mid X]\right\}$
$=$ $\textnormal{E}_{Q_{X}}\textnormal{E}_{P_{V\mid X}}\left\{\textnormal{Var}_{P_{Y\mid X,V,A=a}}[Y\mid X,V,A=a]\mid X\right\}+\textnormal{Var}_{P_{V\mid X}}\left[\mu_{a}(X,V)\mid X\right],$ (52)
where we let $\mu_{a}(X,V)=\textnormal{E}_{P_{Y\mid X,V,A=a}}(Y\mid X,V,A=a)$.
The SLOPE for the median becomes
$\textnormal{SLOPE}(Q_{Y(a),X}^{0},\psi^{{\textnormal{med}}})$
$=$ $\dfrac{\textnormal{E}_{Q_{X}}\left[F_{P_{Y(a)\mid X}}(m_{1/2}\mid X)\mu(X)\right]-\textnormal{E}_{Q_{Y(a),X}^{0}}\left[Y(a)\mathds{1}\{Y(a)\leq m_{1/2}\}\right]}{f_{Q_{Y(a)}^{0}}(m_{1/2})}.$
$=$ $\dfrac{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{V\mid X}}\left[F_{P_{Y\mid X,V,A=a}}(m_{1/2}\mid X,V)\mu_{a}(X,V)-\textnormal{E}_{Y\mid X,V,A=a}\{Y\mathds{1}(Y\leq m_{1/2}\}\Bigg|X\right]\right)}{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{V\mid X}}\left[f_{Y\mid X,V,A=a}(m_{1/2}\mid X,V)\Bigg|X\right]\right)},$ (53)
where $m_{1/2}$ is the marginal median such that
$\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{V\mid X}}\left\{F_{P_{Y\mid X,V,A=a}}(m_{1/2}\mid X,V)\mid X\right\}\right]=1/2.$ (54)
F.2 Estimation of The SLOPE
With a slight abuse of notation, we let $\mu_{a}(X,V)=\textnormal{E}_{P_{Y\mid X,V,A=a}}(Y\mid X,V,A=a)$ and $\mu_{a}(X)=\textnormal{E}_{P_{V\mid X}}\{\mu_{a}(X,V)\mid X\}$. In notation we keep $a\in\{0,1\}$ for generality while in the main text we have focused on $a=1$. For estimation, we assume a linear model whereas for individual $i$ in treatment group $a$, i.e. $A_{i}=a$, we have
$\mu_{a}(x,v)=\alpha_{v}+\beta_{x}+\delta_{a}+(\beta\delta)_{xa},$ (55)
where $v$ ranges in all villages in the source country and $x\in\mathcal{S}_{X}$ ranges in the categories of the baseline variable, and regression coefficients are constrained to guarantee identification: $\beta_{1}=0$, $\delta_{0}=0$, $(\beta\delta)_{x0}=0$ for $x\in\mathcal{S}_{X}$ and $(\beta\delta)_{0a}=0$ for $a\in\{0,1\}$. In addition, we assume that the outcome variance depends on the baseline measurement $x$ as below.
$\sigma^{2}_{a}(x):=\textnormal{Var}[Y\mid X=x,V=v,A=a].$
In Sections 5.2 and 5.3, the SLOPE for the mean is estimated by a simple plug-in estimator of (52):
$\sum_{x\in\mathcal{S}_{X}}\left[\widehat{Q}_{X}(x)\left\{\widehat{\sigma}^{2}_{a}(x)+\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X=x}(v\mid x)\widehat{\mu}_{a}(x,v)\right\}\right],$
where $\widehat{\mu}_{a}(x,v)$ is estimated by the least squares estimator that regresses $Y$ on $X$ and $V$ as in (55), $\widehat{\sigma}^{2}_{a}(x)$ is estimated by the sample variance of regression residuals, $\widehat{Q}_{X}(x)$ and $\widehat{P}_{V\mid X=x}(v\mid x)$ are estimated by the empirical distribution of the corresponding distributions. Since $X$ is discrete, this plug-in estimator can alternatively be viewed as a weighting estimator with weights estimated by empirical distributions of $X$ in the source and target populations.
To estimate the SLOPE for the median in Section 5.2, we assume $Y_{i}-\mu_{a}(X_{i},V_{i})$ follows a normal distribution. Under this conditional normality assumption, the SLOPE in (53) can be estimated by a plug-in (or equivalently, weighting) estimator as follows:
$\sum_{x\in\mathcal{S}_{X}}\widehat{Q}_{X}(x)\left[\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X}(v\mid x)\Phi\left(\dfrac{\widehat{m}_{1/2}-\widehat{\mu}_{a}(x,v)}{\widehat{\sigma}_{a}^{2}(x)}\right)\right]\left[\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X}(v\mid x)\widehat{\mu}_{a}(x,v)\right]$
$-\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X}(v\mid x)\left[\widehat{\mu}_{a}(x,v)\Phi\left(\dfrac{\widehat{m}_{1/2}-\widehat{\mu}_{a}(x,v)}{\widehat{\sigma}_{a}^{2}(x)}\right)-\widehat{\sigma}_{a}(x)\varphi\left(\dfrac{\widehat{m}_{1/2}-\widehat{\mu}_{a}(x,v)}{\widehat{\sigma}_{a}^{2}(x)}\right)\right],$
where $\varphi(\cdot)$ and $\Phi(\cdot)$ are the probability density functional and cumulative distribution function of the standard normal distribution, and $\widehat{m}_{1/2}$ is the solution of the following equation using bisection search,
$\sum_{x\in\mathcal{S}_{X}}\widehat{Q}_{X}(x)\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X}(v\mid x)\Phi\left(\dfrac{\widehat{m}_{1/2}-\widehat{\mu}_{a}(x,v)}{\widehat{\sigma}_{a}^{2}(x)}\right)=1/2.$
F.3 Auxiliary Data Information and Results for Section 5.2
In this section, we provide additional data information and analysis results for Section 5.2 where the outcome variable is the log-transformed per capita consumption.
F.3.1 Auxiliary Data Information
Table 2 gives the distribution of the categorized baseline measurement of the log-transformed per-capita consumption across countries on the overlapped sample.
Table 2: Baseline Measurement $X$ of the log-transformed per-capita consumption.
Peru Pakistan India Honduras Ghana
Sample size 1768 829 771 2152 2379
Category of $X$
(-0.41,3.65] 97 ( 5.5%) 50 ( 6.0%) 450 (58.4%) 1026 (47.7%) 1001 (42.1%)
(3.65,4.24] 449 (25.4%) 170 (20.5%) 260 (33.7%) 817 (38.0%) 832 (35.0%)
(4.24,8] 1222 (69.1%) 609 (73.5%) 61 (7.9%) 309 (14.4%) 546 (23.0%)
Figure 6 provides the normal diagnostics for the conditional normal assumption imposed in Section 5.2 when estimating the SLOPE for the median. From these QQ-plots, the normality assumption is reasonable well for most countries, but less than perfect for India because of the tail behavior.
Figure 6: $Q$-plots for the residuals of the linear model (55) across source countries. Each plot is generated by qqplot in R where the straight line generated by qqline passes through the first and third quartiles.
F.3.2 Additional Results
First, we present one hypothesis on why SLOPE for the average per capita consumption is lower in India and Peru. For India, the experiment was located at West Bengal, an area abutting Bangladesh and shares a language and a culture. Therefore, the unique cultural and locational features may have caused the uniformity of the underlying population. For Peru, according to Banerjee et al., (2015), there has already been a consumption support program implemented on part of the households. This may has led to a higher homogeneity among the treated households.
Second, as mentioned in the main text, Table 3 presents the mean and median of the transported per-capita consumption.
Table 3: The estimated mean and median (not their SLOPEs) for transporting the counterfactual log-transformed per capita consumption under intervention from a source country (i.e., the rows of the table) to a target country (i.e., the columns of the table). Bootstrap standard errors are in the parentheses.
Estimand ($\psi$) Source ($P_{O\mid X}$) Target ($Q_{X}$)
Ghana Honduras India Pakistan Peru
Mean Ghana 3.47 (0.02) 3.44 (0.02) 3.67 (0.03) 3.67 (0.03)
Honduras 4.23 (0.02) 4.15 (0.02) 4.42 (0.04) 4.42 (0.04)
India 4.07 (0.03) 4.04 (0.02) 4.24 (0.07) 4.24 (0.07)
Pakistan 4.24 (0.05) 4.21 (0.05) 4.17 (0.06) 4.42 (0.02)
Peru 4.77 (0.03) 4.74 (0.03) 4.71 (0.04) 4.93 (0.02)
Median Ghana 3.48 (0.02) 3.44 (0.02) 3.68 (0.03) 3.67 (0.03)
Honduras 4.23 (0.02) 4.16 (0.02) 4.42 (0.04) 4.41 (0.04)
India 4.05 (0.02) 4.03 (0.02) 4.23 (0.07) 4.22 (0.06)
Pakistan 4.24 (0.04) 4.21 (0.05) 4.18 (0.06) 4.42 (0.02)
Peru 4.76 (0.03) 4.73 (0.03) 4.70 (0.04) 4.94 (0.02)
F.3.3 First-Order Approximation of Bias
For each pair of source and target countries, we estimate the oracle bias for the target country, i.e., the left hand side of (5). In specific, we estimate the mean/median of the potential outcome in the target country by either transporting from the source country or using the target data (with the outcome information) itself. The difference between the two estimates is treated as the “oracle bias” from conditional exchangeability since it represent the bias one my occur by directly assuming conditional exchangeability when outcome information in the target is unavailable. The confidence intervals of the oracle bias (i.e., horizontal bars) in Figure 7 are obtained via bootstrap.
Next, we estimate the bias approximated with SLOPE, i.e., the right hand side of (5). In addition to estimating SLOPE as described in Section F.2, we also estimate the sensitivity parameter $\gamma$ as follows. First, suppose the normal assumption holds, i.e., $Y_{i}-\mu_{a}(X_{i},V_{i})$ is conditionally normally distributed with mean zero and variance $\sigma^{2}_{a}(X_{i})$. Then the sensitivity model (3) enlists a location shift in the errors between the source and the target, i.e., $Q_{Y\mid X,V,A=a}\sim N\left(\mu_{a}(X,V)+\gamma\sigma_{a}^{2}(X),\sigma_{a}^{2}(X)\right)$. Therefore, by method of moment, we estiamte $\gamma$ through the following formula,
$\sum_{i=1}^{n_{p}}\left[\widehat{\mu}_{a}(X_{i},V_{i})+\widehat{\gamma}\cdot\widehat{\sigma}_{a}^{2}(X_{i})\right]=\sum_{i=n_{p}+1}^{n}\widehat{\sigma}_{a}^{2}(X_{i}),$
where $\widehat{\mu}_{a}$ and $\widehat{\sigma}_{a}^{2}$ are estimates of $\mu_{a}$ and $\sigma_{a}^{2}$, respectively. Therefore, the approximated bias is the product of $\widehat{\gamma}$ and the estimate of SLOPE. For confidence intervals, we fix $\widehat{\gamma}$ as obtained as above from the original source and target samples, and construct $95\%$ confidence intervals for the SLOPE by bootstrapping the source and the target samples. Therefore, the vertical bars shown in Figure 7 do not include the randomness of estimating $\gamma$. Such construction is to align with the common understanding in sensitivity analysis that the sensitivity parameter is taken as a pre-specified fixed value instead of a random quantity.
The results are shown in Figure 7, where the two panels represent the mean and median, and each panel plots the approximated bias with SLOPE against the oracle bias. These dots roughly lie on the $y=x$ line, thereby validating the bias approximation with SLOPE.
Figure 7: Bias approximation with SLOPE for mean (left) and median (right). Each panel plots the bias approximated with SLOPE in the y-axis and the oracle bias in the x-axis in dots and the corresponding bootstrapped $95\%$ confidence interval in bars. The dashed straight line is $y=x$.
F.4 Auxiliary Data Information for Section 5.3
For the data analysis in Section 5.3, all health variables and the corresponding physical health index were measured at individual level. To keep the sample units as households, we average the individual level measurements over households. Although we changed the outcome variable in terms of the weights, we kept the pre-treatment covariate the same to guarantee fair comparison across weighting schemes. The pre-treatment covariate $X$ is the categorized baseline measurement of the physical health index (i.e., the average of three health variables mentioned in Section 5.3). Table 4 gives the distribution of $X$ across countries on the overlapped sample.
Table 4: Baseline Measurement $X$ of physical health index.
Peru India Ethiopia
Sample size 1307 740 785
Category of $X$
(-1.36, 0.235] 979 (74.9%) 523 (70.7%) 350 (44.6%)
(0.235,0.818] 328 (25.1%) 217 (29.3%) 435 (55.4%)
Appendix G Simulations
In this section, we validate the asymptotic properties of the proposed estimators in synthetic datasets that were generated to mimic the real data.
G.1 Simulation Setting
Suppose $O$ is a continuous variable and $X$ is a random variable that is either binary and continuous. In the case when $X$ is binary, we suppose the support $\mathcal{S}_{X}=\{1,2\}$ and $P_{O\mid X=j}\sim N(\mu_{j},\sigma_{j}^{2})$ with $j=1,2$. In the case when $X$ is continuous, we suppose both $P_{X}$ and $Q_{X}$ are Gaussian and $P_{O\mid X}\sim N(\alpha_{m}+\beta_{m}X,\alpha_{v}+\beta_{m}X^{2})$. Data generation parameters were estimated from real data in Section 5.2. Specifically, the conditional distribution $P_{O\mid X}$ is estimated from the per-capita consumption in the treated group of Pakistan. $Q_{X}$ is estimated from Pakistan (i.e., no covariate shift) and Honduras (i.e., covariate shift). To construct a binary $X$, we dichotomize the log-transformed baseline measurement by whether it is below the median; to construct a continuous $X$ we use the original log-transformed measurement. The exact numbers of the (semi-)synthetic distributions are given in Table 5.
For both data generation procedures, we are interested in the SLOPE for mean and median. When $X$ is continuous, we also consider the regression coefficients in simple linear regression which regresses $O$ on $X$. For SLOPE of the mean and OLS coefficients, we consider three estimators: weighting estimator ((35) and (40)), the regression estimator ((36) and (41)), and the efficient influence function based estimator (38). For the SLOPE of the median, we consider the weighting estimator (44) and the regression estimator (45). During estimation we suppose $P_{O\mid X}$ is Gaussian. For nuisance functions/parameters, $\omega(X)$ is estimated with empirical distributions when $X$ is binary (see Section E.7.3) and entropy balancing when $X$ is continuous (see Section E.7.2), $\mu(X)$ is estimated with linear regression, $\sigma^{2}(X)$ is estimated with subgroup sample variance when $X$ is binary and is estimated by regressing the squared residuals $\{O-\widehat{\mu}(X)\}^{2}$ on $X^{2}$ when $X$ is continuous. The target mean is estimated by a weighted average $\sum_{i=n_{q}+1}^{n}O_{i}\widehat{\omega}(X_{i})/n_{p}$ for weighting estimators and by regression $\sum_{i=1}^{n_{q}}\widehat{\mu}(X_{i})/n_{q}$ for regression estimators. For the target median, when $X$ is binary it is estimated with the qmixnorm function in the R package KScorrect where the group mean and variance of $P_{O\mid X=j}$ are estimated by maximum likelihood estimation; when $X$ is continuous it is estimated by the numerical solution to
$\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\dfrac{1}{\widehat{\sigma}(X_{i})}\varphi\left(\dfrac{m_{1/2}-\widehat{\mu}(X_{i})}{\widehat{\sigma}(X_{i})}\right)=1/2$
via bisection, where $\varphi(\cdot)$ is the p.d.f. of the standard Gaussian. Inference for the weighting and regression estimators is based on bootstrap with 1000 times of resampling and inference for the EIF-based estimator is based on the second moment of the EIF with nuisances plugged in (i.e., (39) and (43)).
Table 5: Data generation in simulated data.
$X$ Cov. shift $Q_{X}$ $P_{X}$ $P_{O\mid X}$
Binary $Q(X=1)$ $Q(X=2)$ $P(X=1)$ $P(X=2)$ $\mu_{1}$ $\mu_{2}$ $\sigma_{1}$ $\sigma_{2}$
No $0.1258$ $0.8742$ $0.1258$ $0.8742$ $4.1816$ $4.4773$ $0.4761$ $0.4524$
Yes $0.6597$ $0.3403$
Continuous $\textnormal{E}_{Q_{X}}(X)$ $\sqrt{\textnormal{Var}_{Q_{X}}(X)}$ $\textnormal{E}_{P_{X}}(X)$ $\sqrt{\textnormal{Var}_{P_{X}}(X)}$ $\alpha_{m}$ $\beta_{m}$ $\alpha_{v}$ $\beta_{v}$
No $4.5803$ $0.5970$ $4.5803$ $0.5970$ $3.1304$ $0.2766$ $0.1924$ 0.0003-0.0003
Yes $3.7054$ $0.5340$
G.2 Simulation Result
Simulations are based on 1000 replicates. In each data setting, we consider $n_{p}=n_{q}\in\{1000,2000\}$ and report the bias (bias), root mean squared error (rMSE), empirical standard deviation (empSD), the average estimated standard error (avgSE) and covarage rate (rate). Tables 6 and 7 list the simulation results for the SLOPE of mean and median. For the OLS coefficients, Table 8 and Table 9 provides results for the slope coefficient and the intercept coefficient ,respectively. As these results show, in all cases, the bias becomes closer to zero and sample sizes increases and the rMSE decays with root $n_{p}$. The average of estimated standard error is close to the empirical standard deviation. The coverage rate is closer to $95\%$. Overall, the simulation results suggest that the estimators are $\sqrt{n}$-consistent and the standard error estimates are consistent.
Table 6: Simulation results for SLOPE of mean and median when $X$ is binary. All numbers have been multiplied with $100$.
Estimand Mean Median
Covariate shift Yes No Yes No
Estimator Regress Weight EIF Regress Weight EIF Regress Weight Regress Weight
$n_{p}=1000$ bias 0.01 -0.10 -0.10 0.03 -0.01 -0.02 0.03 0.01 -0.01 -0.02
rMSE 1.66 1.65 1.65 1.01 1.00 1.04 1.86 1.90 1.01 1.04
empSD 1.66 1.65 1.65 1.01 1.00 1.04 1.86 1.90 1.01 1.04
avgSE 1.67 1.66 1.66 1.01 1.01 1.01 1.78 1.78 1.01 1.01
rate 94.7% 94.6% 94.4% 94.8% 94.8% 94.7% 93.7% 92.0% 94.5% 93.9%
$n_{p}=2000$ bias -0.01 -0.07 -0.07 0.00 -0.02 -0.02 -0.11 -0.13 -0.02 -0.03
rMSE 1.18 1.18 1.18 0.68 0.68 0.68 1.29 1.33 0.72 0.73
empSD 1.18 1.18 1.18 0.68 0.68 0.68 1.29 1.32 0.72 0.73
avgSE 1.19 1.19 1.19 0.72 0.72 0.72 1.26 1.26 0.72 0.72
rate 95.3% 94.9% 94.8% 95.7% 95.8% 95.8% 94.5% 93.6% 94.8% 94.4%
Table 7: Simulation results for SLOPE of mean and median when $X$ is continuous. All numbers have been multiplied with $100$.
Estimand Mean Median
Covariate shift Yes No Yes No
Estimator Regress Weight EIF Regress Weight EIF Regress Weight Regress Weight
$n_{p}=1000$ bias -0.09 -0.15 -0.09 -0.09 -0.09 -0.09 -0.06 -0.07 -0.06 -0.06
rmse 0.86 2.16 2.06 0.86 0.86 0.86 0.87 1.48 0.87 0.89
empSD 0.85 2.16 2.05 0.85 0.85 0.86 0.86 1.48 0.86 0.89
avgSE 0.84 2.24 1.92 0.84 0.85 0.84 0.84 1.46 0.84 0.88
rate 93.9% 91.6% 91.9% 93.9% 94.1% 94.0% 94.3% 93.6% 94.3% 94.5%
$n_{p}=2000$ bias -0.01 -0.14 -0.11 -0.01 -0.01 -0.01 0.01 -0.01 0.01 0.01
rmse 0.61 1.53 1.44 0.61 0.61 0.61 0.59 1.07 0.59 0.62
empSD 0.61 1.53 1.44 0.61 0.61 0.61 0.59 1.07 0.59 0.62
avgSE 0.60 1.44 1.38 0.60 0.60 0.60 0.60 1.04 0.60 0.62
rate 94.2% 92.4% 93.2% 94.2% 94.4% 94.7% 95.3% 93.7% 95.3% 95.5%
Table 8: Simulation results for SLOPE of the slope coefficient in simple linear regression. All numbers have been multiplied with $100$.
Estimand Slope
Covariate shift Yes No
Estimator Regress Weight EIF Regress Weight EIF
$n_{p}=1000$ bias -0.02 -0.03 -0.02 -0.02 -0.02 -0.02
rmse 0.23 0.47 0.47 0.19 0.19 0.19
empSD 0.23 0.47 0.47 0.18 0.19 0.19
avgSE 0.22 0.45 0.45 0.18 0.18 0.19
rate 93.9% 91.0% 92.4% 93.9% 94.1% 95.6%
$n_{p}=2000$ bias 0 -0.03 -0.03 0 0 0
rmse 0.16 0.33 0.33 0.13 0.13 0.13
empSD 0.16 0.33 0.33 0.13 0.13 0.13
avgSE 0.16 0.32 0.32 0.13 0.13 0.14
coverage 94.5% 92.9% 94.4% 94.4% 94.3% 95.9%
Table 9: Simulation results for SLOPE of the intercept coefficient in simple linear regression. All numbers have been multiplied with $100$.
Estimand Intercept coefficient
Covariate shift Yes No
Estimator Regress Weight EIF Regress Weight EIF
$n_{p}=1000$ bias -0.09 -0.12 -0.09 -0.09 -0.09 -0.09
rmse 0.86 2.09 2.06 0.86 0.86 0.86
empSD 0.85 2.08 2.05 0.85 0.86 0.86
avgSE 0.84 1.93 1.94 0.84 0.84 0.90
rate 93.9% 90.6% 92.3% 93.9% 94.0% 95.2%
$n_{p}=2000$ bias -0.01 -0.13 -0.11 -0.01 -0.01 -0.01
rmse 0.61 1.46 1.44 0.61 0.61 0.61
empSD 0.61 1.46 1.44 0.61 0.61 0.61
avgSE 0.60 1.38 1.39 0.60 0.60 0.64
rate 94.2% 92.4% 93.4% 94.2% 94.3% 96.4%
Appendix H Proof for The Derivation of The SLOPE
H.1 Proof of Theorem 2
Since Theorem 2 is a special case of Theorem 5 with $q=1/2$, please find the proof of Theorem 5 in Section H.4.
H.2 Proof of Theorem 3
We recall some notation defined in Section A.1 for the support under $Q_{O,X}^{0}$: denote the support of $Q_{O,X}^{0}$ as $\mathcal{S}_{O,X}$ and the supports of the marginals $Q_{X}$ and $Q_{O}^{0}$ as $\mathcal{S}_{X}$ and $\mathcal{S}_{O}$, respectively.
Proof of Theorem 3. The proof proceeds in two steps. First, we show that fixing $Q_{O,X}=Q_{O,X}^{0}$, the (partial) derivative of $\phi$ with respect to $\gamma$ at $\gamma=0$ is
$\phi^{\prime}_{\gamma}(0)=\textnormal{W}{Q}_{O,X}\in l^{\infty}(\mathcal{S}_{O,X}),\text{ such that }\int_{B}d\textnormal{W}{Q}_{O,X}=\int_{B}\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0},$
for any measurable set $B\in\mathcal{S}_{O,X}$.
Second, we show that for an $H\in l^{\infty}(\mathcal{S}_{O,X})$ such that $\int dH=0$, the Hadamard derivative of $\psi$ with respect to $Q_{O,X}$ at $Q_{O,X}^{0}$ in the direction of $H$ is
$\psi^{\prime}_{Q_{O,X}^{0}}(H)=\int\textnormal{IF}\left(O,X,\psi(Q_{O,X}^{0})\right)dH.$
Then, using chain rule of Hadamard derivative (Theorem 20.4 of Van der Vaart, (2000)), we have the derivative of the composite function $\psi\circ\phi$ with respect to $\gamma$ at $\gamma=0$ being:
$\psi^{\prime}_{Q_{O,X}}\left(\phi^{\prime}_{\gamma}(0)\right)=\int\textnormal{IF}\left(O,X,\psi(Q_{O,X}^{0})\right)\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0}.$
We prove the two steps in order.
For the first step, we note that for any $h\in\mathcal{R}$ and $t\downarrow 0$,
$\int_{B}dQ_{O,X}^{th}-\int_{B}dQ_{O,X}^{0}$ $=\int_{B}\left(\dfrac{dQ_{O,X}^{th}}{dQ_{O,X}^{0}}-1\right)dQ_{O,X}^{0}$
$=\int_{B}\left[\dfrac{\exp(th)}{\textnormal{E}\left\{\exp(thO)\mid X\right\}}\right]dQ_{O,X}^{0}.$
Let $\Delta_{t,h}=\dfrac{dQ_{O,X}^{th}}{dQ_{O,X}^{0}}-1$, then
$\Delta_{t,h}$ $=\dfrac{1+thO+O(t^{2})}{1+th\textnormal{E}_{P_{O\mid X}}(O\mid X)+O(t^{2})}-1$
$=\left\{1+thO+O(t^{2})\right\}\left\{1-th\textnormal{E}_{P_{O\mid X}}(O\mid X)+O(t^{2})\right\}-1$
$=th\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}+O(t^{2}).$
Therefore,
$\dfrac{\int_{B}dQ_{O,X}^{th}-\int_{B}dQ_{O,X}^{0}}{t}=$ $\dfrac{\int_{B}\Delta_{t,h}dQ_{O,X}^{0}}{t}$
$=$ $\dfrac{\int_{B}th\{O-\textnormal{E}_{P_{O\mid X}}\}dQ_{O,X}^{0}+O(t^{2})}{t}$
$\to$ $h\int_{B}\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0}.$
For the second step, we note that under Condition 3, $\textnormal{IF}\left(o,x,\psi(Q_{O,X}^{0})\right)=\psi^{\prime}_{Q_{O,X}^{0}}[\delta_{o,x}-Q_{O,X}^{0}]$. Then
$\int\textnormal{IF}\left(o,x,\psi(Q_{O,X}^{0})\right)dH(o,x)$ $=\int\psi^{\prime}_{Q_{O,X}^{0}}[\delta_{o,x}-Q_{O,X}^{0}]dH(o,x)$
$=\psi^{\prime}_{Q_{O,X}}\left[\int\delta_{o,x}-Q_{O,X}^{0}dH(o,x)\right]$
$=\psi^{\prime}_{Q_{O,X}}[H],$
where the second equality follows from the linearity and continuity of Hadamard derivative, and the last equation follows from the observation that for $B\in l^{\infty}(\mathcal{S}_{O,X})$,
$\int\delta_{o,x}(B)-Q_{O,X}^{0}(B)dH(o,x)=H(B)-Q_{O,X}^{0}(B)\cdot 0=H(B).$
Hence, both two steps have been proved. $\square$
H.3 Proof of Lemma 1
Proof of Lemma 1.
Under sensitivity model (3), the target estimand is
$\psi^{\xi}(Q_{O,X}^{\gamma})=\textnormal{E}_{Q_{X}}\left[\dfrac{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\xi(O,X)\right\}}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\right\}}\right].$
Taking derivative with respect to $\gamma$ at $\gamma=0$, we have
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\xi})$ $=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\left\{O\xi(O,X)\right\}-\textnormal{E}_{P_{O\mid X}}\left\{\xi(O,X)\right\}-\textnormal{E}_{P_{O\mid X}}(O)\right]$
$=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}\left[O,\xi(O,X)\right]\right\}.$
$\square$
H.4 Proof of Theorem 5
Proof of Theorem 5. Let $\kappa(\gamma)=\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}$. Under sensitivity model (3), the $q$-th quantile of $O$ under $Q_{O}^{0}$ can be defined as
$q$ $=F_{Q_{O}^{\gamma}}\left(\psi(F_{Q_{O}^{\gamma}})\right)$
$=\int_{-\infty}^{F_{Q_{O}^{\gamma}}^{-1}(q)}\int\dfrac{\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}}dQ_{X}dP_{O\mid X}$
$=\int\int_{-\infty}^{F_{Q_{O}^{\gamma}}^{-1}(q)}\dfrac{\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X)\}}dP_{O\mid X}dQ_{X}.$
On both sides, we take derivative with respect to $\gamma$. Applying the Leibniz rule, we have
$0$ $=\int\left(\dfrac{\partial}{\partial\gamma}\int_{-\infty}^{F_{Q_{O}^{0}}^{-1}(q)}\left[\dfrac{\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}}\right]dP_{O\mid X}\right)\Bigg|_{\gamma=0}dQ_{X}$
$=\int\left[f_{P_{O\mid X}}(m_{q}\mid X)\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi)+\int_{-\infty}^{m_{q}}\{O-\mu(X)\}dP_{O\mid X}\right]dQ_{X}$
$=\textnormal{SLOPE}(Q_{O,X}^{0},\psi)\cdot\int f_{P_{O\mid X}}(m_{q})dQ_{X}+\textnormal{E}_{Q_{O}^{0}}\{O\mathds{1}(O\leq m_{q}\}-\textnormal{E}_{Q_{X}}\left\{F_{P_{O\mid X}}(m_{q})\mu(X)\right\}$
$=\textnormal{SLOPE}(Q_{O,X}^{0},\psi)f_{Q_{O}^{0}}(m_{q})+\textnormal{E}_{Q_{O}^{0}}\{O\mathds{1}(O\leq m_{q})\}-\textnormal{E}_{Q_{X}}\left\{F_{P_{O\mid X}}(m_{q})\mu(X)\right\}.$
Reorganize terms, we have
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=$ $\dfrac{\textnormal{E}_{Q_{X}}\left\{F_{P_{O\mid X}}(m_{q})\mu(X)\right\}-\textnormal{E}_{Q_{O}^{0}}\{O\mathds{1}(O\leq m_{q})\}}{f_{Q_{O}^{0}}(m_{q})}.$
H.5 Proof of Theorem 6
Proof of Theorem 6. Under the sensitivity model (3), the target estimand is
$\psi^{{\alpha\textnormal{-trim}}}(Q_{O}^{\gamma})$ $=\dfrac{1}{1-2\alpha}\int_{F_{Q_{O}^{\gamma}}^{-1}(\alpha)}^{F_{Q_{O}^{\gamma}}^{-1}(1-\alpha)}OdQ_{O}^{\gamma}$
$=\dfrac{1}{1-2\alpha}\int_{F_{Q_{O}^{\gamma}}^{-1}(\alpha)}^{F_{Q_{O}^{\gamma}}^{-1}(1-\alpha)}\dfrac{O\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}}dQ_{O,X}^{0}.$
Taking derivative with respect to $\gamma$ at $\gamma=0$, we obtain the SLOPE as follows:
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\alpha\textnormal{-trim}}})=$ $\dfrac{\partial\psi^{{\alpha\textnormal{-trim}}}(Q_{O}^{\gamma})}{\partial\gamma}\Bigg|_{\gamma=0}$
$=$ $\dfrac{1}{1-2\alpha}\int\left\{F_{Q_{O}^{0}}^{-1}(1-\alpha)f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(1-\alpha))\left[\dfrac{\partial F_{Q_{O}^{\gamma}}^{-1}(1-\alpha)}{\partial\gamma}\Bigg|_{\gamma=0}\right]\right\}dQ_{X}$
$-\dfrac{1}{1-2\alpha}\int\left\{F_{Q_{O}^{0}}^{-1}(\alpha)f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(\alpha))\left[\dfrac{\partial F_{Q_{O}^{\gamma}}^{-1}(\alpha)}{\partial\gamma}\Bigg|_{\gamma=0}\right]\right\}dQ_{X}$
$+\dfrac{1}{1-2\alpha}\int\left[\int_{F_{Q_{O}^{0}}^{-1}(\alpha)}^{F_{Q_{O}^{0}}^{-1}(1-\alpha)}O\{O-\mu(X)\}dP_{O\mid X}\right]dQ_{X}$
$=$ $\dfrac{1}{1-2\alpha}F_{Q_{O}^{0}}^{-1}(1-\alpha)\textnormal{E}_{Q_{O,X}^{0}}\left[\mathbb{P}(O\leq F_{Q_{O}^{0}}^{-1}(1-\alpha)\mid X)\mu(X)-O\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(1-\alpha))\right]$
$-\dfrac{1}{1-2\alpha}F_{Q_{O}^{0}}^{-1}(\alpha)\textnormal{E}_{Q_{O,X}^{0}}\left[\mathbb{P}(O\leq F_{Q_{O}^{0}}^{-1}(\alpha)\mid X)\mu(X)-O\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(\alpha))\right]$
$+\dfrac{1}{1-2\alpha}\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[O\{O-\mu(X)\}\mathds{1}_{[F_{Q_{O}^{0}}^{-1}(\alpha),F_{Q_{O}^{0}}^{-1}(1-\alpha)]}(O)\mid X\right]\right),$
where the last equality follows from the SLOPE for quantiles (Theorem 5).
$\square$
H.6 Proof of Theorem 7
Proof of Theorem 7. This theorem is a special case of Corollary 3 with $s(Y,\psi^{\textnormal{OLS,Sub}})=X_{\textnormal{Sub}}X_{\textnormal{Sub}}^{\intercal}\psi^{\textnormal{OLS,Sub}}-X_{\textnormal{Sub}}Y$. $\square$
H.7 Proof of Theorem 8
Proof of Theorem 8.
We start with $\tau_{a}$. Suppose $\psi^{{\textnormal{OLS}}}$ satisfies $\textnormal{E}_{Q_{O,X}^{0}}\left(XX^{\intercal}\psi^{{\textnormal{OLS}}}-XY\right)$. From Theorem 7, the SLOPE for $\tau_{a}$ is the second entry of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$, which is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$
$=$ $\left\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\right\}^{-1}\textnormal{E}_{Q_{X}}\left[X\sigma^{2}(X)\right]$
$=$ $\left\{\begin{bmatrix}1&\textnormal{E}_{Q_{A}}(A)&\textnormal{E}_{Q_{Z}}(Z^{\intercal})\\ \textnormal{E}_{Q_{A}}(A)&\textnormal{E}_{Q_{A}}(A^{2})&\textnormal{E}_{Q_{X}}(AZ^{\intercal})\\ \textnormal{E}_{Q_{Z}}(Z)&\textnormal{E}_{Q_{X}}(ZA)&\textnormal{E}_{Q_{Z}}(ZZ^{\intercal})\end{bmatrix}\right\}^{-1}\begin{bmatrix}&\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{A\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{Z\sigma^{2}(X)\}\end{bmatrix}$
$=$ $\begin{bmatrix}1-b^{\intercal}S^{-1}b&{}^{*}&{}^{*}\\ -\frac{\textnormal{E}_{Q_{A}}(A))}{\textnormal{Var}_{Q_{A}}(A)}+\delta^{\intercal}V^{-1}\textnormal{E}_{Q_{A}}(A)\delta-\delta^{\intercal}V^{-1}\textnormal{E}_{Q_{L}}(L)&1/\textnormal{Var}_{Q_{A}}(A)+\delta^{\intercal}V^{-1}\delta&{}^{*}\\ V^{-1}\delta\textnormal{E}_{Q_{A}}(A)-V^{-1}\textnormal{E}_{Q_{L}}(L)&-V^{-1}\delta&-V^{-1}\end{bmatrix}\begin{bmatrix}&\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{A\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{L\sigma^{2}(X)\}\end{bmatrix}$,
where entries as “” are omitted because that matrix is symmetric, $b=[\textnormal{E}_{Q_{A}}(A),\textnormal{E}_{Q_{L}}(L^{\intercal})]$, $\delta=\textnormal{Cov}_{Q_{L,A}}[L,A]/\textnormal{Var}_{Q_{A}}[A]$, $V=\textnormal{Cov}_{Q_{X}}(X)-\delta\delta^{\intercal}\textnormal{Var}_{Q_{A}}(A)$, and $S=\begin{bmatrix}\textnormal{Var}_{Q_{A}}(A)&\textnormal{Cov}_{Q_{X}}[A,L^{\intercal}]\\ \textnormal{Cov}_{Q_{X}}[L,A]&\textnormal{Cov}_{Q_{L}}(L)\end{bmatrix}$. Then $\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})$ in (23) follows by expanding terms in the formula above and taking the second entry.
The SLOPE for $\tau_{\textnormal{u}}$ in the unadjusted model can be obtained similarly and is hence omitted. $\square$
H.8 Proof of Theorem 9
Proof of Theorem 9. Under sensitivity model (3), the MAD satisfies
$1/2=$ $\int_{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})-\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})}^{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})+\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})}dQ_{O,X}^{\gamma}$
$=$ $\textnormal{E}_{Q_{X}}\left[\int_{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})-\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})}^{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})+\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})}\dfrac{\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}}dP_{O\mid X}\right].$
Taking derivative with respect to $\gamma$ at $\gamma=0$, we have
$0=$ $\textnormal{E}_{Q_{X}}\left[f_{P_{O\mid X}}\left(m_{1/2}+{\textnormal{MAD}}\right)\left\{\dfrac{\partial\left\{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})+\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})\right\}}{\partial\gamma}\Bigg|_{\gamma=0}\right\}\right]$
$-\textnormal{E}_{Q_{X}}\left[f_{P_{O\mid X}}\left(m_{1/2}-{\textnormal{MAD}}\right)\left\{\dfrac{\partial\left\{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})-\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})\right\}}{\partial\gamma}\Bigg|_{\gamma=0}\right\}\right]$
$+\textnormal{E}_{Q_{X}}\left[\int_{m_{1/2}-{\textnormal{MAD}}}^{m_{1/2}+{\textnormal{MAD}}}\{O-\mu(X)\}dP_{O\mid X}\right]$
$=$ $\left\{f_{Q_{O}^{0}}(m_{1/2}+{\textnormal{MAD}})+f_{Q_{O}^{0}}(m_{1/2}-{\textnormal{MAD}})\right\}\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{MAD}}})$
$+\left\{f_{Q_{O}^{0}}(m_{1/2}+{\textnormal{MAD}})-f_{Q_{O}^{0}}(m_{1/2}-{\textnormal{MAD}})\right\}\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\mathds{1}_{[m_{1/2}-{\textnormal{MAD}},m_{1/2}+{\textnormal{MAD}}]}(O)\{O-\mu(X)\}\right].$
Then the result follows from re-organizing this equation. $\square$
H.9 Proof of Theorem 10
Proof of Theorem 10. For the L-functional
$\psi(Q_{O})$ $=\int_{0}^{1}h(F_{Q_{O}}^{-1}(p))l(p)dp,$
the corresponding SLOPE is (assuming differentiation and integration can exchange)
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$
$=$ $\dfrac{\partial\psi(Q_{O}^{\gamma})}{\partial\gamma}\Bigg|_{\gamma=0}$
$=$ $\int_{0}^{1}h^{\prime}(F_{Q_{O}^{0}}^{-1}(p))\cdot\frac{\partial F_{Q_{O}^{\gamma}}^{-1}(p)}{\partial\gamma}\Bigg|_{\gamma=0}l(p)dp$
$=$ $\int_{0}^{1}h^{\prime}(F_{Q_{O}^{0}}^{-1}(p))\cdot\left(\dfrac{-\textnormal{E}\{O\mathds{1}(O\leq F_{Q_{O}}^{-1}(p))\}}{f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(p))}+\dfrac{\textnormal{E}_{Q_{X}}\left[\mu(X)\textnormal{E}_{P_{O\mid X}}\left\{\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(p))\mid X\right\}\right]}{f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(p))}\right)l(p)dp,$
where the second line follows from the SLOPE for quantiles (Theorem 5).
$\square$
We note that SLOPE of L-functionals can be alternatively derived from the IF below (Huber,, 1981, (3.11)) using Theorem 3,
$\textnormal{IF}(o,x,\psi(Q_{O,X}^{0}))$ $=\int\frac{ph^{\prime}(F_{Q_{O}^{0}}^{-1}(p))}{f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(p))}l(p)dp-\int_{F_{Q_{O}^{0}}(o)}^{1}\frac{h^{\prime}(F_{Q_{O}^{0}}^{-1}(p))}{f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(p))}l(p)dp.$
Appendix I Proof for the Estimation Theory
I.1 Proof of Theorem 11 and Theorem 12
The proof for Theorems 11 and 12 follows from standard M-estimation theory in Newey and McFadden, (1994) and Van der Vaart, (2000). We next prove Theorem 11 and omit the proof of Theorem 12 since the proof for both theorems follows the same procedure.
Proof of Theorem 11. We begin by proving consistency under Condition 7 using the standard techniques in Newey and McFadden, (1994, Theorem 2.1). Let
$M(\eta)$ $=-E\left[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right]^{\intercal}E\left[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right],\text{ and}$
$\widehat{M}_{n}(\eta)$ $=-\left\{\dfrac{1}{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right\}^{\intercal}\left\{\dfrac{1}{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right\}.$
First, under Condition 7(i), $\eta^{\textnormal{W}}$ uniquely maximizes $M(\eta)$. Next, using Lemma 2.4 of Newey and McFadden, (1994), under Condition 7(ii) and (iii), we have the uniform convergence of $G^{\textnormal{W}}$:
$\sup_{\eta\in\Theta}\left\lVert\dfrac{1}{n}\sum_{i=1}^{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)-\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\}\right\rVert\to_{p}0,$
and that the function $\textnormal{E}[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)]$ is continuous with respect to $\eta$, where $\to_{p}$ denotes convergence in probability. Then $M(\eta)$ is also continuous with respect to $\eta$. Next we show the uniform convergence of $\widehat{M}_{n}$. Note that the compactness of $\Theta$ implies the boundedness of $\textnormal{E}[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)]$. Then for any $\eta\in\Theta$,
$\left|\widehat{M}_{n}(\eta)-M(\eta)\right|$
$\leq$ $\left\lVert\textnormal{E}\left[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right]-\dfrac{1}{n}\sum_{i=1}^{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right\rVert\cdot\ \left\lVert\textnormal{E}\left[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right]+\dfrac{1}{n}\sum_{i=1}^{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right\rVert$
implies the uniform convergence,
$\sup_{\eta\in\Theta}|M_{n}(\eta)-M(\eta)|\to_{p}0.$ (56)
With all regularity conditions checked above, we prove the consistency mimicking the proof of Theorem 2.1 in Newey and McFadden, (1994). For any $\varepsilon>0$. with probability approaching one, we have
$M(\widehat{\eta}^{\textnormal{W}})>\widehat{M}_{n}(\widehat{\eta}^{\textnormal{W}})-\varepsilon/3>\widehat{M}_{n}(\eta^{\textnormal{W}})-\varepsilon/3-\varepsilon/3>M(\eta^{\textnormal{W}})-\varepsilon/3-\varepsilon/3-\varepsilon/3,=M(\eta^{\textnormal{W}})-\varepsilon.$ (57)
where the first inequality and third equality hold by the uniform convergence (56), and the second inequality holds because $\widehat{\eta}^{\textnormal{W}}$ uniquely maximizes $\widehat{M}_{n}(\cdot)$. Let $\mathcal{N}$ be an open subset of $\Theta$ that contains $\eta^{\textnormal{W}}$. By the compactness of $\Theta$ and continuity of $M$, we have $\sup_{\eta\in\Theta\cap\mathcal{N}^{C}}M(\eta)<M(\eta^{\textnormal{W}})$ with probability approaching one. Let $\varepsilon=M(\eta^{\textnormal{W}})-\sup_{\eta\in\Theta\cap\mathcal{N}^{C}}M(\eta)$, then (57) implies that $\widehat{\eta}^{\textnormal{W}}\in\mathcal{N}$ with probability approaching one. Hence, with Condition 8(iv), the consistency of the SLOPE follows from the continuous mapping theorem.
Next, we prove the asymptotic normality. It directly follows from Theorem 5.31 of Van der Vaart, (2000) that under Conditions 7-8, $\widehat{\eta}^{\textnormal{W}}$ is asymptotically normal in that $\sqrt{n}\left(\widehat{\eta}^{\textnormal{W}}-\eta^{\textnormal{W}}\right)\to_{d}N\left(0,V^{\textnormal{W}}(V^{\textnormal{W}})^{\intercal}\right)$. Then the asymptotic normality of the SLOPE estimate follows from delta method (Theorem 3.1 of Van der Vaart, (2000)).
$\square$
I.2 Proof of the Derivation of Efficient Influence Functions
I.2.1 Preliminaries
Let $f_{P_{X}}$, $f_{Q_{X}}$, and $f_{P_{O\mid X}}$ be the density functions of the corresponding random variables on the subscripts. For a generic observation, the log-likelihood of the observed data on the joint population can be written as
$l(T,O,X)=$ $(1-T)\log f_{P_{O\mid X}}(O\mid X)+(1-T)\log f_{P_{X}}(X)+T\log f_{Q_{X}}(X)$
$+(1-T)\log(\textnormal{pr}(T=0))+T\log(\textnormal{pr}(T=1)).$
Consider the Hilbert space $\mathcal{H}$ that contains all one-dimensional zero-mean measurable functions of the observed data with finite variance. Consider $f_{P_{X}}$, $f_{Q_{X}}$, and $f_{P_{O\mid X}}$ as nuisance functions and denote their nuisance tangent spaces as $\mathcal{T}_{P_{X}}$, $\mathcal{T}_{Q_{X}}$, and $\mathcal{T}_{P_{O\mid X}}$, respectively. Then $\mathcal{H}$ can be decomposed as
$\mathcal{H}$ $=\mathcal{T}_{P_{X}}\oplus\mathcal{T}_{Q_{X}}\oplus\mathcal{T}_{P_{O\mid X}},\text{ where}$
$\mathcal{T}_{P_{X}}$ $=\left\{(1-T)a_{1}(O,X):\textnormal{E}_{P_{O\mid X}}\{a_{1}(O,X)\mid X\}=0,\textnormal{Var}_{P_{O\mid X}}\left[a_{1}(O,X)\mid X\right]<\infty\right\},$
$\mathcal{T}_{Q_{X}}$ $=\left\{Ta_{2}(X):\textnormal{E}_{Q_{X}}\{a_{2}(X)\}=0,\textnormal{Var}_{Q_{X}}\left[a_{2}(X)\right]<\infty\right\},$
$\mathcal{T}_{P_{O\mid X}}$ $=\left\{(1-T)a_{3}(X):\textnormal{E}_{P_{X}}\{a_{3}(X)\}=0,\textnormal{Var}_{P_{X}}\left[a_{3}(X)\right]<\infty\right\}.$
We consider parametric submodels $f_{P_{O\mid X}}(O\mid X,\xi_{1})$ and $f_{Q_{X}}(X,\xi_{2})$ where $\xi_{1}=0$ and $\xi_{2}=0$ correspond to the underlying truth. We also let
$\textnormal{SC}_{\xi_{1}}(O,X)=\dfrac{\partial\log f_{P_{O\mid X}}(O\mid X,\xi_{1})}{\partial\xi_{1}},\quad\textnormal{SC}_{\xi_{2}}(X)=\dfrac{\partial\log f_{Q_{X}}(X,\xi_{2})}{\partial\xi_{2}}$
be the score functions.
We revise the notation of the target functional to indicate the dependency on the nuisance parameters $\xi_{1}$ and $\xi_{2}$. Specifically, let $\psi(\xi_{1},\xi_{2})$ be the target functional. Then the efficient influence function for the target functional, $\textnormal{EIF}(T,O,X,\psi)$, satisfies
$\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{1}}\Bigg|_{\xi_{1}=0}$ $=\textnormal{E}\left[(1-T)\cdot\textnormal{EIF}(T,O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right],$ (58)
$\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}\Bigg|_{\xi_{2}=0}$ $=\textnormal{E}\left\{T\cdot\textnormal{EIF}(T,O,X,\psi)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\},$ (59)
Similarly, we revise the notation of the SLOPE and its component: let $\textnormal{SLOPE}(\xi_{1},\xi_{2})$ be the SLOPE which takes the form of
$\textnormal{SLOPE}(\xi_{1},\xi_{2})$ $=-\dfrac{\eta_{1}(\xi_{1},\xi_{2})}{\eta_{2}(\xi_{1},\xi_{2})},\text{ where}$
$\eta_{1}(\xi_{1},\xi_{2})$ $=\textnormal{E}_{Q_{O,X}^{0}}\{s(O,X,\psi)\{O-\mu(X)\},$
$\eta_{2}(\xi_{1},\xi_{2})$ $=\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}.$
Likewise, the efficient influence function for the SLOPE, $\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$, satisfies
$\textnormal{E}\left[(1-T)\textnormal{EIF}(T,O,X,\textnormal{SLOPE})\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]$ $=\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\Bigg|_{\xi_{1}=0}$
$=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}\Bigg|_{\xi_{1}=0},$ (60)
$\textnormal{E}\left\{T\cdot\textnormal{EIF}(T,O,X,\textnormal{SLOPE})\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}$ $=\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\Bigg|_{\xi_{2}=0}$
$=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}\Bigg|_{\xi_{2}=0},$ (61)
I.2.2 Proof of Proposition 1
Proof of Proposition 1. Suppose the efficient influence functions for the target functional and the SLOPE take the following form,
$\textnormal{EIF}(T,O,X,\psi)=(1-T)e_{1}(O,X)+Te_{2}(X),\quad\textnormal{EIF}(T,O,X,\textnormal{SLOPE})=(1-T)a_{1}(O,X)+Ta_{2}(X),$
where $(1-T)e_{1}(O,X),(1-T)a_{1}(O,X)\in\mathcal{T}_{P_{O\mid X}}$ and $Te_{2}(X),Ta_{2}(X)\in\mathcal{T}_{Q_{X}}$. The specific forms of $e_{1}(O,X)$ and $e_{2}(X)$ are given in Proposition 2.
In light of (60) and (61), our goal is to solve the following equations for $a_{1}$ and $a_{2}$,
$\textnormal{E}\left\{(1-T)a_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$ $=\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\Bigg|_{\xi_{1}=0}$
$=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}\Bigg|_{\xi_{1}=0},$ (62)
$\textnormal{E}\left\{Ta_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}$ $=\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\Bigg|_{\xi_{2}=0}$
$=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}\Bigg|_{\xi_{2}=0}.$ (63)
We start with (62). In order to calculate the right hand side (RHS), we first calculate $\partial\eta_{1}(\xi_{1},\xi_{2})/\partial\xi_{1}$ and $\partial\eta_{2}(\xi_{1},\xi_{2})/\partial\xi_{1}$; we have
$\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\Bigg|_{\xi_{1}=0}$
$=$ $\dfrac{\partial}{\partial\xi_{1}}\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\right]$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{1}}$
$-\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\dfrac{\partial}{\partial\xi_{1}}\int Of_{P_{O\mid X}}(O\mid X,\xi_{1})dO\right]$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\textnormal{pr}(T=0)\cdot\textnormal{E}_{Q_{O,X}^{0}}\cdot\left\{\dfrac{1}{\omega(X)}e_{1}(O,X,\psi)\cdot\textnormal{SC}_{\xi}(O,X)\right\}$
$-\textnormal{E}_{Q_{O,X}^{0}}\left[\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}O\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left\{c_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\},$
where
$c_{1}(O,X)=$ $s(O,X,\psi)\{O-\mu(X)\}+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\dfrac{\textnormal{pr}(T=0)}{\omega(X)}e_{1}(O,X,\psi)$
$-\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}O,$
and
$\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{1}}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}\cdot\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\psi_{1}}+\textnormal{E}_{Q_{O,X}^{0}}\left\{\dot{s}(O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}+\textnormal{E}_{Q_{O,X}^{0}}\left\{\dot{s}(O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}\cdot\textnormal{pr}(T=0)\textnormal{E}_{Q_{O,X}^{0}}\left\{\dfrac{1}{\omega(X)}e_{1}(O,X,\psi)\cdot\textnormal{SC}(O,X,\psi)\right\}$
$+\textnormal{E}_{Q_{O,X}^{0}}\left\{\dot{s}(O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\{c_{2}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\},$
where
$c_{2}(O,X)$ $=\dfrac{\textnormal{pr}(T=0)}{\omega(X)}\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}\cdot e_{1}(O,X,\psi)+\dot{s}(O,X,\psi)$
Therefore, the RHS of (62) is
$\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{1}}$ $=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\{\eta_{2}(\xi_{1},\xi_{2})\}^{2}}$
$=\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\left\{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)\right\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}.$
Equalizing both sides of (62), we have
$0=$ $\textnormal{E}\left\{(1-T)a_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}-\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{1}}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[\dfrac{\textnormal{pr}(T=0)}{\omega(X)}a_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]$
$-\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\left\{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)\right\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[\left\{\dfrac{\textnormal{pr}(T=0)}{\omega(X)}a_{1}(O,X)-\dfrac{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right].$
Noticing that $\textnormal{SC}_{\xi_{1}}(O,X)\in\mathcal{T}_{P_{O\mid X}}$ (and hence $\textnormal{E}_{P_{O\mid X}}\{\textnormal{SC}_{\xi}(O,X)\mid X\}=0$), we have
$\dfrac{\textnormal{pr}(T=0)}{\omega(X)}a_{1}(O,X)$
$=$ $\dfrac{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}-\textnormal{E}_{P_{O\mid X}}\left[\dfrac{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\Bigg|X\right].$
Therefore, by plugging in $c_{1}(O,X)$ and $c_{2}(O,X)$ at $\xi_{j}=0$ ($j=1,2$) and reorganizing terms, $a_{1}(O,X)$ can be written as
$a_{1}(O,X)$
$=$ $\dfrac{\omega(X)}{\textnormal{pr}(T=0)}\dfrac{-s(O,X,\psi)\{O-\mu(X)\}+\textnormal{Cov}_{P_{O\mid X}}[s(O,X,\psi),O\mid X]+\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\{O-\mu(X)\}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}$
$-\dfrac{\omega(X)\cdot\textnormal{SLOPE}}{\textnormal{pr}(T=0)}\dfrac{\dot{s}(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}$
$-\textnormal{SLOPE}\cdot\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}e_{1}(O,X,\psi).$
Next, we consider (63) for $a_{2}(O,X)$. For the RHS of (63), we calculate $\partial\eta_{1}(\xi_{1},\xi_{2})/\partial\xi_{2}$ and $\partial\eta_{2}(\xi_{1},\xi_{2})/\partial\xi_{2}$; we have
$\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{2}}=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\cdot\textnormal{SC}_{\xi_{2}}(X)\right]+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\cdot\textnormal{SC}_{\xi_{2}}(X)\right]$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\textnormal{pr}(T=1)\textnormal{E}_{Q_{X}}\{e_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\}$
$=$ $\textnormal{E}_{Q_{X}}\{b_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\},\text{ where}$
$b_{2}(X)=$ $\textnormal{E}_{P_{O\mid X}}\left[s(O,X,\psi)\{O-\mu(X)\}\mid X\right]$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\textnormal{pr}(T=1)\cdot e_{2}(X),$
and
$\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{2}}=$ $\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}\cdot\textnormal{SC}_{\xi_{2}}(X)\right]+\textnormal{E}_{Q_{O,X}^{0}}\left[\ddot{s}(O,X,\psi)\right]\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}$
$=$ $\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}\cdot\textnormal{SC}_{\xi_{2}}(X)\right]$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\ddot{s}(O,X,\psi)\right]\textnormal{pr}(T=1)\textnormal{E}_{Q_{X}}\{e_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\}$
$=$ $\textnormal{E}_{Q_{X}}\{b_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\},\text{ where }$
$b_{2}(X)=$ $\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}+\textnormal{E}_{Q_{O,X}^{0}}\left[\ddot{s}(O,X,\psi)\right]\textnormal{pr}(T=1)e_{2}(X).$
Therefore, the RHS of (63) can be written as
$\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{2}}$ $=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}$
$=\textnormal{E}_{Q_{X}}\left\{\left[\dfrac{-b_{1}(X)-b_{2}(X)\cdot\textnormal{SLOPE}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right]\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}.$
Equalizing both sides of (63), we have
$0=$ $\textnormal{E}_{Q_{X}}\left[\left(\textnormal{pr}(T=1)a_{2}(X)-\left[\dfrac{-b_{1}(X)-b_{2}(X)\cdot\textnormal{SLOPE}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right]\right)\cdot\textnormal{SC}_{\xi_{2}}(X)\right].$
Since $\textnormal{SC}_{\xi_{2}}(X)\in\mathcal{T}_{Q_{X}}$ and hence $\textnormal{E}_{Q_{X}}\{\textnormal{SC}_{\xi_{2}}(X)\}=0$, we have
$a_{2}(X)=$ $\dfrac{\left[\dfrac{-b_{1}(X)-b_{2}(X)\cdot\textnormal{SLOPE}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right]-\textnormal{E}_{Q_{X}}\left\{\left[\dfrac{-b_{1}(X)-b_{2}(X)\cdot\textnormal{SLOPE}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right]\right\}}{\textnormal{pr}(T=1)}$
$=$ $\dfrac{-\textnormal{SLOPE}\cdot\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}-\textnormal{E}_{P_{O\mid X}}\{O-\mu(X)\mid X\}}{\textnormal{pr}(T=1)\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}$
$-e_{2}(X)\cdot\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]+\textnormal{SLOPE}\cdot\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}.$
Finally, by plugging in $a_{1}(O,X)$ and $a_{2}(X)$ into $\textnormal{EIF}(T,O,X,\textnormal{SLOPE})=(1-T)a_{1}(O,X)+Ta_{2}(X)$ and noticing $(1-T)e_{1}(O,X)+Te_{2}(X)=\textnormal{EIF}(T,O,X,\psi)$, we obtain the EIF stated in Proposition 1. $\square$
I.2.3 Proof of Proposition 2
Proof of Proposition 2. Suppose the efficient influence function of the target functional is
$\textnormal{EIF}(T,O,X,\psi)=(1-T)e_{1}(O,X)+Te_{2}(X),$
where $(1-T)e_{1}(O,X)\in\mathcal{T}_{P_{O\mid X}}$ and $Te_{2}(X)\in\mathcal{T}_{Q_{X}}$. Then (58)-(59) can be re-expressed as
$\textnormal{E}\left\{(1-T)e_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$ $=\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{1}},$ (64)
$\textnormal{E}\left\{Te_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}$ $=\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}.$ (65)
First, consider (64). The LHS is
$\textnormal{E}\left\{(1-T)e_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$ $=\textnormal{E}_{Q_{O,X}^{0}}\left[\dfrac{\textnormal{pr}(T=0)}{\omega(X)}e_{1}(O,X)\right],$
and the RHS is
$\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{1}}$ $=\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\textnormal{E}_{Q_{O,X}^{0}}\left\{s(O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}.$
We equalize them and then have
$e_{1}(O,X)=-\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\dfrac{\omega(X)}{\textnormal{pr}(T=0)}\left[s(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\right].$
Next, consider (65). The LHS is
$\textnormal{E}\left\{Te_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}$ $=\textnormal{pr}(T=1)\textnormal{E}_{Q_{X}}\left\{e_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\},$
and the RHS is
$\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}$ $=\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\textnormal{E}_{Q_{O,X}^{0}}\left\{s(O,X,\psi)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}.$
We equalize them and have
$e_{2}(X)=-\dfrac{\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}-\textnormal{E}_{Q_{O,X}^{0}}\{s(O,X,\psi)\}}{\textnormal{pr}(T=1)\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}.$
Therefore, by plugging in $e_{1}(O,X)$ and $e_{2}(X)$ to the EIF of the target functional, we have
$\textnormal{EIF}(T,O,X,\psi)=$ $-\dfrac{(1-T)\omega(X)}{\textnormal{pr}(T=0)}\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\left[s(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\right]$
$-\dfrac{T}{\textnormal{pr}(T=1)}\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}.$
$\square$
Appendix J Extended Remarks
This section details some extended remarks deferred from the main text. Section J.4 discusses defining SLOPE for other types of conditional exchangeability assumptions, including the no unmeasured confounding assumption and the missing at random assumption. Section J.5 comments on the challenge of extending the notion of SLOPE to other bound-based sensitivity models. Section J.6 states the mathematical connection between SLOPE and the marginal interventional effect (Zhou and Opacic,, 2022) with the incremental propensity score intervention (Kennedy,, 2019).
J.1 SLOPE for a Vector Valued $\psi(\cdot)$
When the functional $\psi(\cdot)$ is vector valued with dimension $p$, the corresponding SLOPE, as defined in Definition 1, is also a vector of $p$ elements. Each dimension of SLOPE represents the robustness of the corresponding element of the target estimand when conditional exchangeability is “near” violated. For example, when $X$ is $p$-dimensional and the target estimand is a $p$-dimensional vector of ordinary least squares (OLS) coefficients of regressing $O$ on $X$, then each element of SLOPE describes the robustness of the corresponding coefficient; see Section C.5 for a formal presentation. More generally, for SLOPE as vectors, the connection between SLOPE and IF still holds (i.e., Theorem 3. In addition, the two proposed estimators are still applicable, as mentioned in Remarks 5 and 6.
The main difference between a vector of SLOPE and a scalar SLOPE is on the interpretation, and in particular, the informed guidance in designs. Specifically, as mentioned in Section 3.1, when SLOPE becomes a vector and all elements are relevant, then practitioners need to find an appropriate summary of the vector in order to compare SLOPEs across study designs. This could be a visual summary, such as plotting the SLOPE of each dimension, or a quantitative measure such as a norm of the SLOPE.
J.2 SLOPE Beyond Exponential Tilting
Remark 7 (The Direction of Sensitivity).
In order to understand the re-centering term $O-\mu(X)$ in equation (10) of Theorem 3, we consider a broader class of sensitivity models that tilts the density ratio by a given non-negative function $\rho(O,X,\gamma)$ and incorporates the exponential tilting model (3) as a special case with $\rho(O,X,\gamma)=\exp(\gamma\cdot O)$:
$\dfrac{dQ_{O\mid X}^{\gamma}}{dP_{O\mid X}}(O,X)$ $\propto\rho(O,X,\gamma),$
where $\rho(O,X,0)=1$ and $\int\rho(O,X,\gamma)dP_{O\mid X}\in(0,\infty)$. Then the result in Theorem 3 becomes:
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{O,X}^{0}}\left(\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\}\right]\right),$
where $\dot{\rho}$ is the derivative of $\rho$ with respect to $\gamma$. Here, $\dot{\rho}$ characterizes the belief on how violations to conditional exchangeability may change the conditional distribution $Q_{O\mid X}^{\gamma}$. For example, when $\rho(O,X,\gamma)=\exp(\gamma\cdot O)$ as in (3), the derivative is $\dot{\rho}(O,X,0)=O$. Therefore, the term $\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\}\right]$ (i.e., $O-\mu(X)$ in (10)) can be viewed as the direction of sensitivity that’s unexplained by the observed variable $X$.
J.3 Extended Remarks from Section 6
For the mean, the dependency of its SLOPE on the conditional variance $\sigma^{2}(X)$ advocates source populations with a lower $\sigma^{2}(X)$; this principle holds generally for location parameters in that SLOPE advocates a more homogeneous design. Since $\sigma^{2}(X)$ is smaller when $X$ contains less information, SLOPE seems to advocate a homogeneous $X$ (e.g., $X=x$ almost surely and therefore $\sigma^{2}(X)=0$), which contradicts with the existing understanding that $X$ should be sufficiently rich for the overlap condition (Assumption 1) to hold. To understand why such contradiction appears, we elaborate on the comments noted in Section 6.
First, SLOPE is proposed and applied in cases when the overlap condition holds, meaning that covariates in the source population cannot contain less information than the covariates in the target population. This aligns with most works in sensitivity analysis for the conditional exchangeability assumption in different contexts. We refer readers to Huang, (2025) for a sensitivity analysis solely for the overlap condition. Considering violations to both assumptions is a valuable future direction (see Bonvini and Kennedy, (2022) and Cui and Li, (2025)) and is beyond the scope of this work.
Second, we always prioritize the endeavor to meet conditional exchangeability than to make remedies when the conditional exchangeability is violated. Whenever possible, we echo Tipton and Olsen, (2018) and Degtiar and Rose, (2023) in the importance of a careful data collection for the conditional exchangeability to hold, or, for the violation to be as small as possible. In this sense, it’s more preferable to collect more common characteristics in source and target populations, namely a richer $X$. Secondary to that, SLOPE is a useful tool to assess sensitivity/robustness when it is unrealistic to meet the conditional exchangeability with the observed set of $X$. This happens, for example, when (i) it is infeasible to randomize units into the source or target population, (ii) not all covariates can be measured in both populations, which is common if the investigator opts in after data collection, and/or (iii) there exist unobservable differences between populations (e.g., sites) even under a careful design (Allcott,, 2015; Jin et al.,, 2024).
J.4 SLOPE in Other Contexts
While this paper focuses on the conditional exchangeability assumption in transportability/generalizability, SLOPE can be used to study the sensitivity of other types of conditional exchangeability typed assumptions. With different meanings of $P$ and $Q$, SLOPE can measure the sensitivity/robustness of other types of conditional exchangeability assumptions, including the no unmeasured confounding assumption in causal inference and the missing at random assumption in missing data. We describe these settings in the following subsections.
J.4.1 SLOPE for Unmeasured Confounding
Let $A\in\{0,1\}$ be the treatment and $Y(A)$ be the potential outcome under treatment $A$, where $A=1$ means intervention and $A=0$ means control. We suppose SUTVA (Assumption 4) holds; i.e., when $A=a$, the observed outcome $Y$ is $Y(a)$, for $a=0,1$. Suppose we are interested in the average treatment effect on the treated, a popular estimand in causal inference,
$\psi^{\textnormal{ATT}}$ $=\textnormal{E}\{Y(1)-Y(0)\mid A=1\}$
$=\textnormal{E}\{Y(1)\mid A=1\}-\textnormal{E}\{Y(0)\mid A=1\}$
$=\textnormal{E}\{Y\mid A=1\}-\textnormal{E}\{Y(0)\mid A=1\}.$ (66)
where the third line follows from SUTVA. From (66), the key challenge in identifying $\psi^{\textnormal{ATT}}$ lies in the challenge of identifying the second term, $\textnormal{E}\{Y(0)\mid A=1\}$, since it involves the potential outcome $Y(0)$. One common strategy for identifying $\psi^{\textnormal{ATT}}$ is through the conditional exchangeability assumption (Assumption 2) and the overlap condition (Assumption 1). To demonstrate that, we define $P$ and $Q$ as follows.
Suppose $P$ represents the population of units with $A=0$ and $Q$ represents the population of units with $A=1$. Therefore, $\textnormal{E}_{P}(\cdot)=\textnormal{E}(\cdot\mid A=0)$ and $\textnormal{E}_{Q}(\cdot)=\textnormal{E}(\cdot\mid A=1)$. Suppose $X$ contains pre-treatment covariates (i.e., measured confounders) and $O=Y(0)$ is the potential outcome under condtrol. Then the conditional exchangeability assumption (Assumption 2), i.e., $Q_{O\mid X}(\cdot\mid x)=P_{O\mid X}(\cdot\mid x)$ almost everywhere $Q_{X}$, implies that
$\left(\textnormal{E}_{Q_{O\mid X}}[O\mid X]:=\right)\quad\textnormal{E}\{Y(0)\mid X,A=1\}=\textnormal{E}\{Y(0)\mid X,A=0\}\quad\left(:=\textnormal{E}_{P_{O\mid X}}[O\mid X]\right).$
Then $\psi^{\textnormal{ATT}}$ in (66) can be identified, since
$\textnormal{E}\{Y(0)\mid A=1\}$ $=\textnormal{E}\left[\textnormal{E}\{Y(0)\mid X,A=1\}\mid A=1\right]$
$=\textnormal{E}\left[\textnormal{E}\{Y(0)\mid X,A=0\}\mid A=0\right]$ (by conditional exchangeability)
$=\textnormal{E}\left[\textnormal{E}\{Y\mid X,A=0\}\mid A=0\right]$ (by SUTVA)
no longer involves potential outcomes.
In this context, the conditional exchangeability assumption is also referred to as no unmeasured confounding. In observational studies where unmeasured confounding is highly likely, there is an extensive literature that develops sensitivity analysis methods to study the consequence of unmeasured confounding. With the above defined $P$ and $Q$, SLOPE can be naturally applied to study the sensitivity of violation to no unmeasured confounding. For example, the SLOPE for $\psi^{\textnormal{ATT}}$ as defined above is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{ATT}})$
$=$ $\textnormal{E}_{Q_{X}}\textnormal{Var}_{P_{O\mid X}}(O\mid X)$ (by Theorem 1)
$=$ $\textnormal{E}_{X\mid A=1}\left[\textnormal{Var}_{Y(0)\mid X,A=0}(Y(0)\mid X,A=0)\mid A=1\right]$ $\text{(by definitions of }P\text{ and }Q\text{)}.$
Remark 8 (SLOPE for Other Causal Estimands).
We note that $P$ and $Q$ may need to be defined differently for other causal estimands. For example, when the target estimand becomes the average treatment effect on the control, then the roles of $P$ and $Q$ need to be swapped. Nevertheless, at a high level, the definition of SLOPE remains consistent, i.e., the derivative of the target causal estimand with respect to $\gamma$ at $\gamma=0$, where $\gamma$ is such that
$\dfrac{[(Y(a)\mid X,A=1-a]}{[Y(a)\mid X,A=a]}\propto\exp\{\gamma\cdot Y(a)\}.$ (67)
where we use $\cdot\mid X,A]$ to represent conditional densities (provided exist with respect to some measure) on the underlying population. For more explanations on the sensitivity model (67), see Scharfstein et al., (1999); Franks et al., (2020).
Finally, we make a clarification on this setting. The current section should not be confused with Section D. In Section D, we transport treatment effect from a source population to a target population where we have assumed that the treatment effect can be identified within the source; in other words, the “causal” assumptions already hold in the source population. However, in this section, we no longer consider any transfer learning setting but instead focus on a causal setting with potential unmeasured confounding. Here, may still think of $P$ (i.e., population of control units) as the “source” population and $Q$ (i.e., population of treated units) as the “target” population, bearing in mind the difference in the meaning of conditional exchangeability assumption as well as the different interpretations of the corresponding SLOPEs.
J.4.2 SLOPE for Non-ignorable Missing
Consider a missing data problem where we have observed covariates $X$ on all units, but only observe the outcome variable $O$ on a subset of the units. Let $P$ be the population of units with complete data and $Q$ be the population of units with a missing $O$. Then the conditional exchangeability assumption (Assumption 2) is the missing at random assumption which is commonly adopted in the literature of missing data. In this context, the SLOPE can be defined similarly as the main text and it can be interpreted as the sensitivity/robustness of the target estimand with respect to non-ignorable missingness.
J.5 Remark on Extending SLOPE to Bound-Based Sensitivity Analysis Models
In sensitivity model (3), the sensitivity parameter $\gamma$ quantifies the degree of violation in a parametric way and it elicits a point identification of the target estimand. In contrast, an important line of sensitivity analyses uses the sensitivity parameter(s) to bound the difference between the unobserved distribution (i.e., $Q_{O\mid X}$) and the observed distribution (i.e., $P_{O\mid X}$) nonparametrically and obtain a set identification of the target estimand. A natural question is whether we may extend the concept of SLOPE to sensitivity analyses based on bounds. Unfortunately, this will require a non-trivial extension to the notion of derivatives. We illustrate the challenge through an example below and leave such extensions as important future directions.
We consider a simplified version of the setting in Zeng et al., (2023) and their sensitivity model. Suppose the target estimand is the mean, $\psi^{{\textnormal{mean}}}(Q_{O,X})=\textnormal{E}_{Q_{O,X}}(O)$. Let $\gamma$ be the sensitivity parameter and suppose the target conditional distribution $Q_{O\mid X}^{\textnormal{bias}}$ deviates from the source conditional distribution $P_{O\mid X}$ so that the bias in the conditional mean is no larger than $\gamma$:
$-\gamma+\textnormal{E}_{P_{O\mid X}}(O\mid X)\leq\textnormal{E}_{Q_{O\mid X}^{\textnormal{bias}}}(O\mid X)\leq\gamma+\textnormal{E}_{P_{O\mid X}}(O\mid X).$ (68)
Consequently, the target estimand can be bounded as
$-\gamma+\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right]\leq\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{bias}})\leq\gamma+\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right].$ (69)
For convenience of notation, here we still use $\gamma$ to represent the sensitivity parameter. We use $Q_{O\mid X}^{\textnormal{bias}}$ and $Q_{O,X}^{\textnormal{bias}}$ to denote conditional and joint distributions under sensitivity model (68), bearing in mind that they are not unique since the sensitivity model (3) does not directly identifies/bounds the distribution (but rather, the bias in conditional means). We use $\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{bias}})$ to denote target estimand under the sensitivity model.
Next, we show that it’s not natural to define an analogy to SLOPE. Suppose we focus on the lower bound of (69) where
$\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{L,}\gamma})=-\gamma+\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right].$
Then the derivative of the target estimand $\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{L,}\gamma})$ with respect to $\gamma$ at $\gamma=0$ is 1-1, since
$\lim_{\gamma\to 0}\dfrac{\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{L,}\gamma})-\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{L,}0})}{\gamma}=\lim_{\gamma\to 0}\dfrac{-\gamma}{\gamma}=-1.$
Similarly, consider the upper bound of (69) where
$\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{R,}\gamma})=\gamma+\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right],$
then the derivative is $1$. The two special cases demonstrate the non-uniqueness of the derivative of the target estimand with respect to $\gamma$ in the set of target estimands (i.e., (69)) determined by the sensitivity model (68). Consequently, generalizing the notion of SLOPE to bound-based sensitivity models like (68) is non-trivial.
J.6 Connections to Marginal Interventional Effect
Broadly speaking, the concept and the mathematical format of the SLOPE are connected to the marginal interventional treatment effect (Zhou and Opacic,, 2022) with the incremental propensity score intervention (Kennedy,, 2019). In this section, we elaborate on this connection.
We begin with a brief introduction to the marginal interventional treatment effect; see Zhou and Opacic, (2022) for details. In causal inference (see notation in Section D.1), an interventional effect quantifies the change in outcome when the propensity of a proportion of units changes. This is in contrast to conventional estimands, like the average treatment effect, which compare the outcome when all units receive treatment with the outcome when all receive placebo. To define an interventional effect, suppose the propensity score is $\pi(Z)=P(A=1\mid Z)$ and consider changing the propensity in a way of $\pi^{\gamma}(Z)$ as a function of $\pi(Z)$ with $\pi^{0}(X)=\pi(Z)$. Then under assumptions in Section 3 of Zhou and Opacic, (2022), the interventional effect (IE) is
$\text{IE}^{\gamma}$ $=\textnormal{E}\left[\dfrac{\pi^{\gamma}(Z)-\pi^{0}(Z)}{\textnormal{E}\left[\pi^{\gamma}(X)-\pi^{0}(Z)\right]\tau(X)}\right],$
where $\tau(Z)=\textnormal{E}\{Y(1)-Y(0)\mid X\}$ is the conditional average treatment effect (CATE). The marginal interventional effect (MIE), as proposed by Zhou and Opacic, (2022), describes the marginal change in the interventional effect by taking the limit of $\gamma$ of IE going to zero:
MIE $=\textnormal{E}\left[\dfrac{\dot{\pi}^{0}(Z)}{\textnormal{E}\left\{\dot{\pi}^{0}(Z)\right\}}\tau(Z)\right],$
where $\dot{\pi}^{0}(X)$ is the derivative of $\pi^{\gamma}(Z)$ with respect to $\gamma$ at $\gamma=0$. When the propensity score shift $\pi^{\gamma}(Z)$ is determined by the incremental propensity score interventions (IPSI) proposed by Kennedy, (2019), i.e.,
$\dfrac{\pi^{\gamma}(Z)/\{1-\pi^{\gamma}(Z)\}}{\pi^{0}(Z)/\{1-\pi^{0}(Z)\}}=\exp(\gamma),$ (70)
the MIE is
$\text{MIE}^{\text{IPSI}}$ $=\dfrac{\textnormal{E}\left[\pi^{0}(Z)\{1-\pi^{0}(Z)\}\tau(Z)\right]}{\textnormal{E}\left[\pi^{0}(Z)\{1-\pi^{0}(Z)\}\right]}.$ (71)
The MIE with IPSI is connected to SLOPE for two reasons. First, the IPSI shifts the intervention $\pi^{\gamma}(Z)$ in the same exponential tilting model as in the sensitivity model (3); see below for details. Second, the concept of MIE, which defines a treatment effect in terms of the change in the outcome with an infinitesimal change in propensity, is broadly connected to the concept of SLOPE. To see the connections more explicitly, we next re-express the SLOPE under the regime of incremental propensity score intervention. We note that although the resulting quantity no longer represents a “‘sensitivity to local perturbation from the exchangeability”, we still refer to it as SLOPE for ease of communication, keeping in mind that we temporarily treat SLOPE as a mathematical quantity instead of a sensitivity/robustness measure.
Let $O=A$ be the binary treatment. Suppose $P$ is the population of interest and $P_{Z,A,Y}$ is the joint distribution of pre-treatment covariate $Z$, binary treatment $A$ and outcome $Y$. We adopt the potential outcome framework (Section D.1) and suppose SUTVA (Assumption 4) and strong ignorability (Condition 2) hold.
Let $Q$ be a hypothetical population with the same distribution as $P$ except that the intervention mechanism has been changed by $\gamma$. In specific, let $O=A$, then the sensitivity model (3) implies that the odds of receiving treatment in $Q$ is $\exp(\gamma)$ times the odds of receiving treatment in $P$:
$\exp(\gamma)=\dfrac{Q^{\gamma}(A=1\mid Z)/\{1-Q^{\gamma}(A=1\mid Z\}}{P(A=1\mid Z)/\{1-P(A=1\mid Z)\}}.$
Note that in this case, $X=(A,Z)$ includes the intervention and the pre-treatment covariates. We let $\pi^{\gamma}(Z)=Q^{\gamma}(A=1\mid Z)$ and therefore $\pi^{0}(Z)=Q^{0}(A=1\mid Z)=P(A=1\mid Z)$ and $\pi^{\gamma}(Z)=\dfrac{\exp(\gamma)\pi^{0}(X)}{1-\pi^{0}(Z)+\exp(\gamma)\pi^{0}(Z)}$.
To define SLOPE, let $Q^{\gamma}_{Z,A,Y}=P_{Z}\times Q^{\gamma}_{A\mid Z}\times P_{Y\mid Z,A}$ and let the target functional be the mean outcome, i.e.,
$\psi(Q_{O,X}^{\gamma})$ $=\textnormal{E}_{Q_{Y}^{\gamma}}(Y)$
$=\textnormal{E}_{Q_{X}}\left[\pi^{\gamma}(Z)\textnormal{E}_{P_{Y\mid Z,A=1}}\{Y\mid Z,A=1\}+\{1-\pi^{\gamma}(Z)\}\textnormal{E}_{P_{Y\mid Z,A=0}}\{Y\mid Z,A=0\}\right]$
$=\textnormal{E}_{Q_{X}}\left[\pi^{\gamma}(Z)\tau(Z)\right]+\textnormal{E}_{Q_{X}}\left[\left\{\textnormal{E}_{P_{Y\mid Z,A=0}}(Y(0)\mid Z,A=0\right\}\right],$
where again $\tau(X)$ is the CATE.
Noticing $\partial\pi_{\gamma}(X)/\partial\gamma|_{\gamma=0}=\pi^{0}(Z)\{1-\pi^{0}(Z)\}$, we have the SLOPE as
$\textnormal{E}_{Q_{X}}\left[\dfrac{\partial\pi^{\gamma}(Z)}{\partial\gamma}\Bigg|_{\gamma=0}\tau(Z)\right]$ $=\textnormal{E}_{Q_{X}}\left[\pi^{0}(Z)\{1-\pi^{0}(Z)\}\tau(X)\right].$ (72)
We make two remarks on (72). First, it resembles the general form of the SLOPE of an outcome mean, since
$\textnormal{E}_{Q_{X}}\left[\pi^{0}(X)\{1-\pi^{0}(X)\}\tau(Z)\right]=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}[A,Y(1)-Y(0)\mid Z]\right\},$
where the $\textnormal{Cov}[\cdot\mid X]$ represents conditional covariance under $P$. Note that here SLOPE is the expectation of a conditional covariance instead of a conditional variance because the “shift” is on the treatment (i.e., $O=A$) rather than the outcome. Second, (72) can be viewed as the non-standardized version of the MIE in (71). It is also a non-standardized version of the average treatment effect for the overlap population (ATO) (Li et al.,, 2018) since the two quantities (ATO and MIE) are identical in this specific regime (Zhou and Opacic,, 2022).