### v2

# SLOPE and Designing Robust Studies for Generalization

Abstract
A common task in generalization is to learn about a new target population using data from another source population. This task relies on conditional exchangeability, which assumes that differences between the source and target populations are fully captured by observable variables. However, this assumption is often untenable in practice due to remaining, unobservable differences, and it cannot be verified with data. These limitations warrant the development of robust study designs that are inherently less sensitive to violations of the assumption. We propose SLOPE (Sensitivity of LOcal Perturbations from Exchangeability), a simple and novel measure that quantifies sensitivity to local violations of conditional exchangeability. SLOPE combines ideas from sensitivity analysis in causal inference and derivative-based robustness measure from Hampel’s influence function. To the best of our knowledge, SLOPE is the first metric to quantify the robustness of study designs with respect to violations of conditional exchangeability. Specifically, SLOPE measures the sensitivity of two design-level characteristics: (a) the functional of interest (e.g., the mean or the median) and (b) the study distributions. We demonstrate how SLOPE can guide robust study designs through a re-analysis of a multinational randomized experiment.
Keywords: Generalizability; Conditional exchangeability; Sensitivity analysis; Causal inference; Influence function; Exponential tilting
1 Introduction
1.1 Background and Overview
There has been a growing interest in generalizing or transporting information from an existing source population to a new, target population under the assumption of conditional exchangeability and the setup can be formalized as follows. Suppose each datum is represented as random variables $(O,X)$ and the goal is to learn the target distribution $Q_{O,X}$ given (i) the “full” data $(O,X)$ from the source distribution $P_{O,X}$ and (ii) the “partial” data $X$ from the target distribution $Q_{X}$. Conditional exchangeability states that conditional on $X$, the distribution of $O$ between the target and the source population is identical:
$Q_{O\mid X}(\cdot\mid X=x)=P_{O\mid X}(\cdot\mid X=x)\text{ almost everywhere in $Q_{X}$. }$ (1)
Equation (1) enables learning about the target distribution $Q_{O,X}$ based only on (i) $P_{O,X}$ and (ii) $Q_{X}$ and this phenomenon can be illustrated with a heuristic, yet simple equality:
$Q_{O,X}=Q_{O\mid X}\times Q_{X}=P_{O\mid X}\times Q_{X}.$
The first equality is from the definition of conditional probability and the second equality is from (1). By the same heuristic argument, we can learn a low-dimensional feature of the target distribution $Q_{O,X}$, denoted as $\psi(Q_{O,X})$ and referred to as the target estimand or target functional, by $\psi(Q_{O,X})=\psi(P_{O\mid X}\times Q_{X})$. Some popular target estimands include the mean of $O$ or the average treatment effect in the target population; see Section 2.1 for details and more examples.
Unfortunately, recent works (Allcott,, 2015; Jin et al.,, 2024) argued that conditional exchangeability is likely violated in practice due to unobservable differences between the source and the target population. Worse, conditional exchangeability is inherently untestable because the variable $O$ is unavailable from the target distribution; under the setup above, we only have access to samples from (i) the joint distribution in the source, $P_{O,X}$, and (ii) the marginal distribution in the target, $Q_{X}$ (Dahabreh et al.,, 2023; Zeng et al.,, 2023; Huang,, 2024). Taken together, these challenges underscore the need to have data collection processes and more broadly, study designs that are inherently less sensitive to violation of conditional exchangeability before analyzing data for generalization.
To this end, the main contribution of the paper is to propose a simple and novel tool that helps gauge which study designs are robust to violations of conditional exchangeability and we present a high-level summary of the tool. Let $Q_{O\mid X}^{\gamma}$ be the distribution of $Q_{O\mid X}$ when conditional exchangeability is violated by a degree quantified by a sensitivity parameter $\gamma\in\mathbb{R}$ and let $\gamma=0$ be the case where conditional exchangeability holds (i.e., $Q_{O\mid X}^{0}=P_{O\mid X}$). Then, we propose a metric called SLOPE, which stands for Sensitivity of LOcal Perturbations from Exchangeability:
${\rm SLOPE}(Q_{O,X}^{0},\psi)=\lim_{\gamma\to 0}\frac{\psi(Q_{O,X}^{\gamma})-\psi(Q_{O,X}^{0})}{\gamma},\quad{}\text{ where }Q_{O,X}^{\gamma}=Q_{O\mid X}^{\gamma}\times Q_{X}.$
As its name and definition imply, SLOPE is the slope of the target estimand $\psi(Q_{O,X}^{\gamma})$ at $\gamma=0$ (see Figure 1 for a visual illustration). SLOPE measures how dramatically the target estimand changes when moving from a setting with no violation of conditional exchangeability (i.e., $\gamma=0$) to a setting with a near-violation (i.e., $\gamma\to 0$). Generally, a higher magnitude of SLOPE suggests that the target estimand is more sensitive/less robust to local violations, while a lower magnitude suggests the estimand is less sensitive/more robust; see Section 3.1 for further discussions on interpreting SLOPE.
SLOPE depends on two quantities: (a) the target estimand (i.e., $\psi$) and (b) the target distribution under conditional exchangeability (i.e., $Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$). Importantly, SLOPE does not depend on the estimation procedure of $\psi$. To put it differently, SLOPE is an intrinsic, design-level characteristic about (a) the target estimand (i.e., $\psi$) and (b) the source or target distributions in the setup (i.e., $P_{O\mid X}$ and $Q_{X}$) when there is a local violation of conditional exchangeability.
We briefly answer three common and important questions about SLOPE; see Sections 3 and 4 for detailed discussions. First, SLOPE is a local measure and for small deviations of $\gamma$ from 0, SLOPE provides an accurate reflection about the change in $\psi$. For larger deviations of $\gamma$ from 0, SLOPE may still provide valuable intuition, subject to the usual limitations of linear approximations based on the tangent line. We remark that some well-known robustness measures are local, including Hampel, (1974)’s celebrated influence function (IF), and these local measures yield valuable insights for designing robust estimators and tests (Huber,, 1981). For more discussions behind the motivation for measuring local violations in robust statistics and how SLOPE yields valuable insights about robust study designs, see Sections 3 and 4.1. Second, from the definition of SLOPE, the unit of SLOPE inherits the unit of the target estimand, thereby respecting the investigator’s original choice of units for the target estimand. If the investigator wishes a unit-less SLOPE, a simple solution would be to transform the target estimand to be unit-less (e.g., in z-score units); see Section 3.1 for more discussions on interpreting SLOPE. Third, SLOPE depends on the parametrization of $Q_{O\mid X}^{\gamma}$ or equivalently, the sensitivity model for conditional exchangeability. Our sensitivity model has some benefits, but also carries some limitations (see Section 2.2, Section 4.1, and Remark 2). Regardless of the choice of the sensitivity model, we believe the high-level idea of SLOPE as a derivative-based summary of a sensitivity analysis can provide new and important insights about designing robust studies for generalization.
Figure 1: Left: the data structure of our setup. Right: an illustration of SLOPE. The $y$-axis plots the target estimand/functional $\psi(\cdot)$ and the x-axis plots $\gamma$, which represents the magnitude of violating conditional exchangeability. The point $\gamma=0$ indicates no violation of conditional exchangeability.
1.2 Prior Works
SLOPE fits into the large literature on robust statistics. Specifically, SLOPE and Hampel, (1974)’s IF are related in that both SLOPE and IF use local derivatives to quantify robustness. But, Hampel, (1974)’s IF measures a local change in the target estimand $\psi$ due to contamination of a single data point whereas SLOPE measures a local change in $\psi$ due to violations of conditional exchangeability. Despite these differences, Section 4.1 reveals an interesting analytic and geometric connection between SLOPE and IF.
Our work also fits into the literature on sensitivity analysis of conditional exchangeability for generalizability and transportability (Nguyen et al.,, 2017; Nie et al.,, 2021; Colnet et al.,, 2021; Dahabreh et al.,, 2022, 2023; Duong et al.,, 2023; Ek and Zachariah,, 2023; Huang,, 2024; Jin et al.,, 2024). What differentiates our paper from most existing works on sensitivity analysis for generalizability is that existing works studied sensitivity of estimators or tests for a specific value of the sensitivity parameter $\gamma\neq 0$. In contrast, our work studies sensitivity of study designs; as mentioned in our summary above, SLOPE is agnostic to how the target estimand $\psi$ is estimated and is a population-level characteristic.
There have been some works on presenting simple, numerical summaries about the impact of violating key assumptions in different fields and we highlight some relevant examples. Gupta and Rothenhäusler, (2023) proposed the directional s-value, which quantifies the minimum amount of covariate shift that alters the sign of a target estimand, usually the mean. Rosenbaum, (2004) proposed design sensitivity, which is a scalar, odds-ratio based summary of the power of a test statistic in a sensitivity analysis. Andrews et al., (2017) considered local mis-specification of generalized method of moments (GMM) and their matrix-based measure, denoted as $\Lambda\in\mathbb{R}^{p\times p}$, measured the change in the estimator for GMM parameters. In missing data, Troxel et al., (2004) proposed a measure which quantifies the change in an estimator when there is a local violation of ignorable missingness. Ding and VanderWeele, (2016), Oster, (2019), Zhao, 2019b , and Cinelli and Hazlett, (2020) proposed scalar measures, which summarize the impact of unmeasured confounding in observational studies. Specifically, each proposed metrics that quantify the “minimum unmeasured confounding bias” necessary to alter the study’s conclusion under no unmeasured confounding. Except for Gupta and Rothenhäusler, (2023) and Rosenbaum, (2004) to some extent, all the above works measure robustness of estimators or tests rather than that of study designs. Also, except for Troxel et al., (2004), all these works do not use derivative-based measures of robustness.
Finally, we briefly mention another important line of work on robustness and sensitivity analysis when conditional exchangeability holds, but the distributions of shared characteristics $X$ of the two populations differ dramatically, especially in the observed data; this setting referred to as limited overlap (Stuart et al.,, 2011; Tipton,, 2014; Chen et al., 2023b, ; Huang,, 2025). Except for Huang, (2025), a key distinction between these works and our work is that limited overlap can be, in principle, checked from the observed data since $X$ is observed in both populations. In contrast, conditional exchangeability cannot be checked from the observed data since $O$ is not observed in the target population.
1.3 Organization of Paper and Notation
The paper is organized as follows. Section 2 introduces the setup and the sensitivity model. Section 3 formally introduces SLOPE, its properties, and results on robust study designs. Section 4 discusses more insights and results, including the relationship between SLOPE and IF and estimating SLOPE. Section 5 showcases an application of SLOPE through a re-analysis of a multi-site experiment by Banerjee et al., (2015). Section 6 concludes with a discussion on practical considerations. Proofs and other results are in the Supplement.
We define the notations that we use in the paper. For a population $P$ and random vectors $X_{1}$, $X_{2}$, we let $P_{X_{j}}$ be the marginal distribution of $X_{j}$ for $j=1,2$. Given $P_{X_{j}}$, we let $\textnormal{E}_{P_{X_{j}}}(\cdot)$ and $F_{P_{X_{j}}}$ be the expectation and cumulative distribution function (c.d.f.), respectively, under $P_{X_{j}}$. We let $P_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$ be the conditional distribution of $X_{1}$ given $X_{2}$. Similarly, we let $\textnormal{E}_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$, $F_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$, $f_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$ be the conditional distribution, conditional c.d.f., and conditional probability density function, respectively. We also let $\textnormal{E}_{X_{1}\mid X_{2}}(\cdot\mid x_{2})$ be the conditional expectation given a specific $X_{2}=x_{2}$. Throughout the paper, we assume sufficient regularity conditions for conditional distributions, conditional densities, and conditional expectations to exist; see Shao, (2008, Chapter 1) for the regularity conditions. Finally, for a probability distribution $Q$ defined on the same measurable space of $(X_{1},X_{2})$, we use $P_{X_{1}\mid X_{2}}\times Q_{X_{2}}$ to denote the joint distribution where $P_{X_{1}\mid X_{2}}\times Q_{X_{2}}(A\times B)=\int_{B}P_{X_{1}\mid X_{2}}(A\mid x_{2})dQ_{X_{2}}(x_{2})$ for $A\in\mathcal{B}$ and $B\in\mathcal{B}^{d}$ and $\mathcal{B}$ is a Borel $\sigma$-field with respect to the reals.
2 Setup
2.1 Goal in Generalization and Key Assumptions
Let $P_{O,X}$ and $Q_{O,X}$ be the joint distributions of random variables $(O,X)$ from a source population $P$ and a target population $Q$, respectively, where $O$ is a scalar and $X$ is a vector. The goal in generalization is to learn a functional (i.e., $\psi(\cdot)$) of the target distribution $Q_{O,X}$, which we denote as $\psi(Q_{O,X})$ and refer to as the target estimand or target functional, from (a) “full” data $(O,X)$ from the source distribution $P_{O,X}$ and (b) “partial” data $X$ from the target distribution $Q_{X}$; see Figure 1 for a visual illustration of the data setup.
Before we go any further, we make two brief remarks about the setup. First, while our exposition below considers a scalar functional $\psi$ (e.g., means, medians, average potential outcomes), all of our results extend to a low-dimensional, vector-valued $\psi$; see Section 3.1 and Section J.1 of the Supplement where we discuss the setting when $\psi$ is the regression coefficient of $O$ regressed on $X$ in the target population. Second, almost all results below are agnostic to how the data were sampled within each population, for instance by simple random sampling, i.i.d. sampling, or even adaptive sampling. Specifically, our results remain at the population level until Section 4.2, where we propose estimators of SLOPE.
Under the setup, the two most popular assumptions for identifying the target estimand $\psi$ (e.g., Cole and Stuart, (2010); Tipton, (2014); Kern et al., (2016); Dahabreh et al., (2019); Huang et al., (2023); Zeng et al., (2023); Degtiar and Rose, (2023)) are as follows.
Assumption 1 (Overlap).
$Q_{X}$ is absolutely continuous with respect to $P_{X}$.
Assumption 2 (Conditional Exchangeability).
Equation (1) holds.
Remark 1 (Alternative Formulation of Assumption 2).
When the densities of $P_{O\mid X}$ and $Q_{O\mid X}$ exist with respect to a common measure (e.g., the Lebesgue measure or the counting measure), Assumption 2 can be re-formulated with respect to the corresponding density functions, i.e., $f_{Q_{O\mid X}}(O,X)/f_{P_{O\mid X}}(O,X)=1$ almost everywhere in $P_{O\mid X}\times Q_{X}$.
Assumption 1 states that the support of $Q_{X}$ is within the support of $P_{X}$. Assumption 2 enables replacing $Q_{O\mid X}$ with $P_{O\mid X}$, which can be identified from the “full data” $(O,X)$ in the source population $P_{O,X}$. Under Assumptions 1 and 2, the target estimand can be identified as $\psi(Q_{X}\times P_{O\mid X})$ and some examples of target estimands are listed below.
Example 1 (Mean).
Suppose we are interested in the mean of $O$ in the target distribution, denoted as $\psi^{{\textnormal{mean}}}(Q_{O,X})=\textnormal{E}_{Q_{O}}(O)$. Under Assumptions 1 and 2, the mean is identified via
$\psi^{{\textnormal{mean}}}(Q_{O,X})=\textnormal{E}_{Q_{O}}(O)=$ $\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{Q_{O\mid X}}(O\mid X)\right]=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right].$
Example 2 (Median).
Suppose we are interested in the median of $O$ in the target distribution, denoted as $\psi^{{\textnormal{med}}}(Q_{O,X})=F_{Q_{O}}^{-1}(1/2)$, and $O$ is continuous. Under Assumptions 1 and 2, the median is identified as the solution to the following equation:
$\frac{1}{2}=\int\int_{-\infty}^{\psi^{{\textnormal{med}}}}dQ_{O\mid X}dQ_{X}=\int\int_{-\infty}^{\psi^{{\textnormal{med}}}}dP_{O\mid X}dQ_{X}.$
Example 3 (Z-Estimand).
Suppose $\psi(Q_{O,X})$ is defined as the solution to
$\textnormal{E}_{Q_{O,X}}\{s\left(O,X,\psi(Q_{O,X})\right)\}=0,$ (2)
where $s(O,X,\cdot)$ is a user-specified function, usually a score function of the same dimension as $\psi$. Under Assumptions 1 and 2, the target estimand is identified as the solution to
$0=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{Q_{O\mid X}}\left\{s(O,X,\psi(Q_{O\mid X}\times Q_{X}))\mid X\right\}\right]=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\left\{s(O,X,\psi(Q_{O,X}^{0}))\mid X\right\}\right].$
Example 4 (Mean or Median of Potential Outcomes).
Consider a randomized experiment in the source population to measure the average treatment effect (ATE). Let $Y(a)$ be the potential outcome if, contrary to fact, a study unit was assigned to treatment $a\in\mathcal{A}\subset\mathbb{R}$ where $\mathcal{A}$ is a set of all possible treatment (e.g., $\mathcal{A}=\{0,1\}$) and let $X$ be pre-treatment covariates. The goal is to learn about the ATE in a new, target population based on (a) the randomized experiment in the source population and (b) the distribution of $X$ in the target population. Under Assumptions 1 and 2, the mean and the median of the potential outcome $Y(a)$ in the target population are identified using Examples 1 and 2:
$\psi^{{\textnormal{mean}}}(Q_{Y(a),X})=\mathbb{E}_{Q_{Y(a)}}\{Y(a)\}=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{Y(a)\mid X}}\{Y(a)\mid X\}\right],\text{ and }\quad{}\frac{1}{2}=\int\int_{-\infty}^{\psi^{{\textnormal{med}}}}dP_{Y(a)\mid X}dQ_{X}.$
We remark that both equations involve potential outcomes (i.e., $P_{Y(a)\mid X}$), which is identified under a randomized experiment in the source population; see Section D in the Supplement for details.
2.2 Model for Sensitivity Analysis of Conditional Exchangeability
Suppose we suspect that conditional exchangeability (i.e., Assumption 2) is implausible and we wish to assess how the conclusion of the study may change if conditional exchangeability is violated. A sensitivity analysis addresses this question by supposing that there is a $\gamma$ violation” of conditional exchangeability and quantifying the downstream consequences of this violation. This section presents a model-based sensitivity analysis (Rosenbaum and Rubin, 1983a, ; Robins et al.,, 2000; Franks et al.,, 2020) for quantifying violations of conditional exchangeability based on exponential tilting.
Formally, for each $\gamma\in\mathbb{R}$, let $Q_{O\mid X}^{\gamma}(\cdot\mid X)$ be absolutely continuous with respect to $P_{O\mid X}(\cdot\mid X)$ almost everywhere $Q_{X}$. Suppose the corresponding densities $f_{Q_{O\mid X}^{\gamma}}(O,X)$ and $f_{P_{O\mid X}}(O,X)$ satisfy the following relationship:
$\dfrac{f_{Q_{O\mid X}^{\gamma}}(O,X)}{f_{P_{O\mid X}}(O,X)}\propto\exp(\gamma\cdot O)\text{, almost everywhere in }P_{O\mid X}\times Q_{X}.$ (3)
The notation $\propto$” means “proportional to” and the normalizing constant satisfies $\int\exp(\gamma O)dP_{O\mid X}(O\mid X)<\infty$ almost everywhere $Q_{X}$.
The term $\gamma$ is often referred to as the sensitivity parameter and it measures the difference between $Q_{O\mid X}$ and $P_{O\mid X}$. If $\gamma=0$, the sensitivity model (3) reduces to Assumption 2 where the two distributions are identical, i.e., $Q_{O\mid X}^{0}=P_{O\mid X}$ almost everywhere in $Q_{X}$. As $\gamma$ moves away from zero, the difference between $Q_{O\mid X}$ and $P_{O\mid X}$ becomes larger and conditional exchangeability is violated by a larger amount.
We make some remarks about the sensitivity model (3). First, model (3) was proposed by Scharfstein et al., (1999) and Robins et al., (2000) as a non-parametric (just) identified model for describing selection bias in missing data and has been used by several others (Rotnitzky et al.,, 2001; Birmingham et al.,, 2003; Troxel et al.,, 2004; Linero and Daniels,, 2018; Franks et al.,, 2020; Nabi et al.,, 2024; Dahabreh et al.,, 2022; Miao et al.,, 2024). Second, model (3) can be reformulated as a selection model and under some assumptions, the sensitivity parameter $\gamma$ can be reparameterized to pseudo-$R^{2}$ (Franks et al.,, 2020). Third, model (3) can be extended so that $\gamma$ depends on $X$ and $O$ or the exponential tilting term can be replaced with a non-negative tilting term (see Section 4.3). Fourth, an important caveat of model (3) is the non-collapsibility of the model with respect to $X$ as the model implies a logistic selection model; see Section 7 of Scharfstein et al., (1999). Fifth, model (3) differs from a “bound-based” sensitivity analysis (e.g., Rosenbaum, (1987); Tan, (2006)) and Remark 2 discusses an inherent difficulty in defining SLOPE with such models. In particular, model (3) (i) makes our proposed measure SLOPE tractable in terms of having a unique tangent curve (see Section 4.3), (ii) has an analytic connection to the influence function (see Section 4.1), (iii) posits no testable implications on the data (Franks et al.,, 2020), and most importantly, (iv) leads to simple and empirically validated principles about designing robust study designs for generalizations (see Sections 3.2, 3.3 and 6).
3 SLOPE: Sensitivity to Local Perturbation from Exchangeability
3.1 Definition and Basic Properties
Under Assumption 1 and the sensitivity model (3), with a chosen value of the sensitivity parameter $\gamma$, the target estimand can be identified via $\psi(Q_{O,X}^{\gamma})$ where $Q_{O,X}^{\gamma}=Q_{O\mid X}^{\gamma}\times Q_{X}$ represents the joint distribution of $Q_{O\mid X}^{\gamma}$ induced from (3) and $Q_{X}$. But, $\gamma$ is never known in practice because it represents the magnitude of violating conditional exchangeability (i.e., Assumption 2). Instead, sensitivity analysis seeks to understand how $\psi(Q_{O,X}^{\gamma})$ changes from $\gamma=0$ (i.e., when Assumption 2 holds) to $\gamma\neq 0$ (i.e., when Assumption 2 doesn’t hold). This is typically done by presenting a table or a plot of $\psi(Q_{O,X}^{\gamma})$ for a plausible range of $\gamma$ with the range determined by domain knowledge (Scharfstein et al.,, 1999; Rotnitzky et al.,, 2001; Nabi et al.,, 2024), benchmarking (Huang,, 2024) or calibration (Miao et al.,, 2024).
[…]
Definition 1 (SLOPE).
The sensitivity to local perturbation from exchangeability (SLOPE) of a target functional/estimand $\psi$ with respect to the sensitivity model in (3) is defined as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\lim_{\gamma\to 0}\dfrac{\psi(Q_{O,X}^{\gamma})-\psi(Q_{O,X}^{0})}{\gamma},$ (4)
provided the limit exists.
A large magnitude of SLOPE means that the target estimand $\psi$ will change more drastically if conditional exchangeability is slightly violated (i.e., $\gamma\to 0$). In contrast, a small magnitude of SLOPE means that the target estimand will change less drastically. Note that the magnitude of SLOPE is the absolute value of SLOPE when the target estimand $\psi$ is a scalar. When $\psi$ is a vector, the magnitude of SLOPE corresponds to the researcher’s choice of measuring the magnitude of vectors (e.g., $\ell_{2}$ norm). As remarked in Section 2.1, for expositional purposes, we focus on a scalar $\psi$, and thereby a scalar SLOPE, but our results hold for low-dimensional, vector-valued $\psi$.
When comparing the magnitudes of SLOPE, investigators should note that the unit of SLOPE inherits the unit of the target estimand, which is often defined with scientifically meaningful units. For instance, for the mean $\psi^{{\textnormal{mean}}}$ and the median $\psi^{{\textnormal{med}}}$ in Examples 1 and 2, respectively, the units of SLOPE for both estimands are the unit of $O$ and the two SLOPEs have identical units. If the investigator wishes to change the units of SLOPE, including a unit-less SLOPE, a simple approach is to change the units of the target estimand. More broadly, we recommend interpreting the magnitude of SLOPE with the same caution used for interpreting the magnitude of regression coefficients where the units of the regression coefficients inherit their underlying units from the data.
A key property of SLOPE is that it does not depend on any particular estimation procedure. Instead, SLOPE measures an intrinsic property about the robustness of a study design and is determined by two design quantities: (a) the target estimand $\psi$, and (b) the target distribution under conditional exchangeability ($Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$). For researchers, these choices roughly correspond to answering two questions: (a) “what quantity do I want to study?” and (b) “which dataset should I use to study the quantity?”. Changing the answers to either question can alter the value of SLOPE. Consequently, SLOPE can help researchers pick a robust study design by selecting an estimand (e.g., the mean or the median of $O$ in the target distribution), target distribution (e.g., $Q_{X}$), or the source distribution (e.g., $P_{O,X}$) that leads to a lower magnitude of SLOPE; see Sections 3.2 and 5 for illustrations.
When communicating the meaning of SLOPE, some researchers may find it useful to interpret SLOPE as a measure of the “first-order change” of the estimand when conditional exchangeability is violated. Specifically, a first-order Taylor expansion of $\psi$ yields
$\psi(Q_{O,X}^{\gamma})-\psi(Q_{O,X}^{0})\approx\gamma\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi).$ (5)
The Taylor expansion suggests that for a small $\gamma$ that is near zero, SLOPE provides an accurate measure of the change in $\psi$ when conditional exchangeability is violated. When $\gamma$ is large in magnitude, SLOPE may still provide some intuition about the change in $\psi$, with the usual limitations of first-order linear approximations. We remark, however, that SLOPE cannot identify the bias of an estimator of $\psi$ from violating conditional exchangeability since, as mentioned in Section 2.2, $\gamma$ and $Q_{O,X}^{\gamma}$ are not identifiable.
Finally, we remark that SLOPE does not always exist for every target estimand. For example, if $\psi$ is the sign of the mean of $O$ in the target population, the limit in Definition 1 may not exist when the sign changes near $\gamma=0$. One general condition for SLOPE to exist is to satisfy the conditions for the chain rule under Hadamard differentiability; see Section A of the Supplement for details. For Z-estimands in Example 3, a sufficient condition for the existence of SLOPE is to impose smoothness and boundedness conditions on $s$; note that these conditions are common to establish consistency of Z-estimators.
Condition 1 (Existence of SLOPE for Z-Estimands).
(i) $\textnormal{E}_{Q_{O\mid X}^{\gamma}}[s(O,X,\psi(Q_{O,X}^{0}))]$ is bounded for $\gamma$ in a neighborhood of zero, and $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi(Q_{O,X}^{0}))\left\{O-\mu(X)\right\}\right]$ exists where $\mu(X)=\textnormal{E}_{P_{O\mid X}}[O\mid X]$; (ii) $s(O,X,\cdot)$ is differentiable almost everywhere with the derivative $\dot{s}(O,X,\cdot)$; and (iii) $\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi(Q_{O,X}^{0}))\}$ exists and is non-singular.
3.2 Example: SLOPE for the Mean and Robust Study Designs
This section has two main goals. The first is to show that how researchers can derive SLOPE for a given $\psi(Q_{O,X}^{\gamma})$ using basic calculus. The second is to illustrate how SLOPE can yield useful insights about robust study designs for generalization.
To begin, consider the mean of $O$ in the target population, i.e., $\psi^{{\textnormal{mean}}}$ in Example 1 where $\psi^{\textnormal{mean}}(Q_{O,X})=\textnormal{E}_{Q_{O}}(O)$. From Section 3.1, for a given $\gamma$, the sensitivity model (3) implies the following equality:
$\psi^{{\textnormal{mean}}}(Q_{O,X}^{\gamma})=\textnormal{E}_{Q_{O,X}^{\gamma}}(O)=\textnormal{E}_{Q_{X}}\left\{\textnormal{E}_{Q_{O\mid X}^{\gamma}}\left(O\mid X\right)\right\}=\textnormal{E}_{Q_{X}}\left[\dfrac{\textnormal{E}_{P_{O\mid X}}\left\{O\exp(\gamma O)\mid X\right\}}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\mid X\right\}}\right].$
Then SLOPE for the mean is the derivative of $\psi^{{\textnormal{mean}}}(Q_{O,X}^{\gamma})$ with respect to $\gamma$, evaluated at $\gamma=0$. In principle, researchers can compute this derivative using single-variable calculus and Theorem 1 states the regularity conditions to ensure the existence of this derivative.
Theorem 1 (SLOPE of Mean).
Suppose Condition 1 holds with $s(O,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$. Then the SLOPE of the mean $\psi^{\rm mean}$ from Example 1 is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\},\text{ where }\sigma^{2}(X)=\textnormal{Var}_{P_{O\mid X}}(O\mid X).$ (6)
In words, the SLOPE of the mean is the average variability of $O$ after adjusting for $X$ in the source population and the average is taken over the target’s $Q_{X}$. When this variation is homoskedastic/constant across $X$ (i.e., $\sigma^{2}(X)=\sigma^{2}$), the mean’s SLOPE simplifies to $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\sigma^{2}$. In this case, SLOPE is only determined by the source distribution, specifically $P_{O\mid X}$; the target distribution $Q_{X}$ does not change SLOPE.
Two immediate implications follow from Theorem 1 about robust study designs. First, suppose the shared covariate $X$ explains almost all the variation in $O$ in the source population, then $\sigma^{2}(X)$ will be close to zero. Thus, SLOPE will be close to zero, meaning that the target mean will not change dramatically even if conditional exchangeability is slightly violated.
As a concrete example, consider Example 4 where the goal is to generalize the ATE by letting O=Y(1)Y(0)O=Y(1)-Y(0). If the randomized experiment in the source population suggests that the individual treatment effect is nearly constant (i.e., $Y(1)-Y(0)\approx c$ for some constant $c$), then $\sigma^{2}(X)$ will be close to zero and according to SLOPE, the ATE will not be sensitive even if conditional exchangeability is violated. In short, a near-constant treatment effect is robust in generalization. We briefly remark that Tipton and Olsen, (2018) made a similar observation in the context of generalizing ATEs in empirical works where constant treatment effects generalize better than heterogeneous ones. The main difference between Tipton and Olsen, (2018) and our work is we provide a more theoretical foundation for why constant treatment effects, or more broadly any effects where the conditional variance of O=Y(1)Y(0)O=Y(1)-Y(0) given $X$ is small, generalizes better.
Second, suppose the variation in $O$ is not well-explained by $X$ in some regions of $X$ from the source population. Then, the mean’s SLOPE can be made small by choosing a target distribution $Q_{X}$ that concentrates around a region of $X$ that shows the smallest variation in $O$ in the source population. Figure 2 provides a visual illustration of these examples and Corollary 1 formalizes these observations.
Figure 2: Illustration of designing robust studies for generalization with $\psi^{{\textnormal{mean}}}$. The x-axis represents $X$ and the y-axis represents $O$. For a formal result behind the illustrations, see Corollary 1.
Corollary 1 (Robust Study Design for Learning $\psi^{\rm mean}$).
(i) Consider two source populations $P$ and $\widetilde{P}$ that satisfy $\sigma^{2}(X)\leq\widetilde{\sigma}^{2}(X)$ almost surely for a target distribution $Q_{X}$, where $\sigma^{2}(X)=\textnormal{Var}_{P_{O\mid X}}(O\mid X)$ and $\widetilde{\sigma}^{2}(X)=\textnormal{Var}_{\widetilde{P}_{O\mid X}}(O\mid X)$. Then
$\textnormal{SLOPE}(P_{O\mid X}\times Q_{X},\psi^{{\textnormal{mean}}})\leq\textnormal{SLOPE}(\widetilde{P}_{O\mid X}\times Q_{X},\psi^{{\textnormal{mean}}}),$
(ii) Next, consider two target distributions $Q_{X}$ and $\widetilde{Q}_{X}$ over a common support $\mathcal{S}_{X}$ such that there exists a subset $\mathcal{S}_{X,1}\in\mathcal{S}_{X}$ that satisfies $Q_{X}(\mathcal{S}_{X,1})\leq\widetilde{Q}_{X}(\mathcal{S}_{X,1})$. If there exists a constant $c$ such that $\sigma^{2}(X)<c$ for $X\in\mathcal{S}_{X,1}$ and $\sigma^{2}(x)>c$ for $x\not\in\mathcal{S}_{X,1}$, then
$\textnormal{SLOPE}(P_{O\mid X}\times Q_{X},\psi^{{\textnormal{mean}}})\leq\textnormal{SLOPE}(P_{O\mid X}\times\widetilde{Q}_{X},\psi^{{\textnormal{mean}}}),$
3.3 Example: SLOPE for Median
Similar to the SLOPE of the mean, the SLOPE of the median also depends on the dispersion of the underlying distribution $Q_{O,X}^{0}$. Theorem 2 provides a general formula for the SLOPE of the median along with two special cases.
Theorem 2 (SLOPE of Median).
Suppose Condition 1 holds with $s(O,X,\psi^{{\textnormal{med}}})=\mathds{1}(O\leq\psi^{{\textnormal{med}}})-\mathds{1}(O>\psi^{{\textnormal{med}}})$ and $\psi^{\rm med}$ is unique, where $\mathds{1}(\cdot)$ is the indicator function which is one if the event holds and zero otherwise. Then, the SLOPE of the $\psi^{\rm med}$ from Example 2 is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=$ $\dfrac{\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{1/2}\mid X)\mu(X)\right]-\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{1/2})\right]}{f_{Q_{O}^{0}}(m_{1/2})},$ (7)
where $m_{1/2}=F_{Q_{O}^{0}}^{-1}(1/2)$ is the median of $O$ on the target population under conditional exchangeability and we recall $\mu(X)=\textnormal{E}_{P_{O\mid X}}(O\mid X)$ is the conditional expectation of $O$ given $X$ in the source population.
(i) If $P_{O\mid X}$ is symmetric with respect to O and $\mu(X)=m_{1/2}$ almost surely $Q_{X}$, then (7) simplifies to
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})=\dfrac{m_{1/2}-\textnormal{E}_{Q_{O}^{0}}\left(O\mid O\leq m_{1/2}\right)}{2f_{Q_{O}^{0}}(m_{1/2})}.$ (8)
(ii) If $P_{O\mid X}$ is Gaussian, i.e., $P_{O\mid X}\sim N\left(\mu(X),\sigma^{2}(X)\right)$, then (7) simplifies to
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=\textnormal{E}_{Q_{X}}\left[\sigma^{2}(X)\cdot\dfrac{f_{P_{O\mid X}}(m_{1/2}\mid X)}{\textnormal{E}_{Q_{X}}\left\{f_{P_{O\mid X}}(m_{1/2}\mid X)\right\}}\right].$
Compared with the SLOPE of the mean, the SLOPE of the median depends on the spread of $Q_{O,X}^{0}$ in a way that is more complicated than $\sigma^{2}(X)$. For example, in part (i) which induces symmetry, the SLOPE of the median in (8) depends on two quantities: (a) the difference between the mean and the truncated mean that is lower than the median $\left\{m_{1/2}-\textnormal{E}_{Q_{O}^{0}}\left(O\mid O\leq m_{1/2}\right)\right\}$, and (b) the inverse of the marginal density at the median $1/f_{Q_{O}^{0}}(m_{1/2})$. In essence, (a) measures the spread of $Q_{O}^{0}$ and (b) is called Tukey’s sparsity (Tukey,, 1965), which is designed to measure the inverse of the concentration of $Q_{O}^{0}$. In part (ii) when $P_{O\mid X}$ is Gaussian, the SLOPE of the median becomes a weighted average of the conditional variance $\sigma^{2}(X)$ where the weight is determined by the conditional density $f_{O\mid X}$ evaluated at the median $m_{1/2}$. This form of the median’s SLOPE resembles the mean’s SLOPE (6), which is an (unweighted) average of $\sigma^{2}(X)$. In general, whether the researcher should study the median or the mean as a measure of centrality of $O$ in their study will depend on the conditional variance $\sigma^{2}(X)$ and the shape of the tail of $P_{O\mid X}$.
4 Further Insights and Results
4.1 Relationship Between SLOPE and IF
Another local, derivative-based measure of robustness that precedes and complements our work is Hampel, (1974)’s celebrated influence function (IF) [^1] Hampel, (1974) originally called IF the “influence curve.” In later works, others, including Hampel, referred to the influence curve as the IF due to its generalization to higher dimensions (Hampel et al.,, 2011).. In this section, we show how SLOPE is related to IF. Briefly, the IF of $\psi(\cdot)$ under $Q_{O,X}^{0}$ is a derivative-based local measure of robustness that quantifies the effect of an infinitesimal contamination at the point $(o,x)$ on the estimand $\psi(Q_{O,X}^{0})$,
$\textnormal{IF}(o,x,\psi(Q_{O,X}^{0}))=$ $\lim_{t\downarrow 0}\dfrac{\psi\left((1-t)Q_{O,X}^{0}+t\delta_{o,x}\right)-\psi(Q_{O,X}^{0})}{t}.$ (9)
The term $\delta_{o,x}$ represents the dirac delta function at $(o,x)$. From the definitions, both SLOPE and IFs are local measures of robustness. The main difference is that IFs measure the local change of the target estimand due to contamination of a single data point whereas SLOPE measures the local change of the target estimand due to violation of conditional exchangeability from the entire distribution.
In addition to their definitions, we can also compare the IF and SLOPE from a geometrical perspective. Specifically, both are directional derivatives at the “origin” (i.e., $t=0$ or $\gamma=0$), but they differ in their direction. Theorem 3 shows that SLOPE is equal to IF if the IF it “tilted” towards a particular direction.
Theorem 3 (Connection between SLOPE and IF).
Suppose either (i) $\psi$ is a Z-estimand defined in (2) and Condition 1 hold, or (ii) Conditions 2-3 in the Supplement hold. Then SLOPE can be written as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left\{O-\mu(X)\right\}\mid X\right]\right).$ (10)
Similar to Condition 1, Conditions 2-3 are regularity conditions that ensure the chain rule under Hadamard differentiability holds. In words, (10) states that SLOPE is the expectation (over $Q_{X}$) of the conditional covariance of the IF and $O-\mu(X)$ (over $P_{O\mid X}$). As discussed in Remark 4, the term $O-\mu(X)$ is the “residual variation” of violating conditional exchangeability under the sensitivity model (3) that is unexplained by $X$. If the IF is nearly orthogonal to the residual subspace $O-\mu(X)$ based on violating conditional exchangeability, then SLOPE will be close to zero. For a Z-estimand in (2), its SLOPE will be smaller if the score function $s$ is chosen to be nearly orthogonal to the subspace spanned by $O-\mu(X)$.
In addition to the geometric interpretation of SLOPE, Theorem 3 provides a general formula to derive the SLOPE given an IF. Sections 3.2 and 3.3 derived SLOPEs for the mean and the median using their respective IFs. Also, Section C of the Supplement derives SLOPEs for scale parameters, ordinary least squares (OLS) coefficients, Pearson correlation, L-estimands and other Z-estimands using Theorem 3.
4.2 Estimation
We briefly discuss two estimators of SLOPE, the weighting estimator and the regression estimator, for Z-estimands in Example 3. In short, estimation of SLOPE follows from existing theory on M-estimation (e.g., Chapter 5 of Van der Vaart, (2000)). Details behind the estimators, including implementation, asymptotic properties, and simulations to assess their finite-sample performance, are in Sections E - G of the Appendix. Our data analysis in Section 5 uses the regression estimator below due to better finite-sample performance.
Suppose we have $n_{q}$ independent and identically distributed (i.i.d.) samples $X_{i}\sim Q_{X}$ and another independent $n_{p}$ i.i.d. samples $(X_{i},O_{i})\sim P_{O,X}$ with $i=1,\ldots,n_{q},n_{q}+1,\cdots,n_{p}+n_{q}=n$. Under Condition 1, the SLOPE of a Z-estimand is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=$ $-\left\{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\dot{s}(O,X,\psi)\right]\right)\right\}^{-1}\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}s\left(O,X,\psi\right)\right]\right).$ (11)
Let $\omega(X)=f_{Q_{X}}(X)/f_{P_{X}}(X)$ be the density ratio of $X$ between the two populations. The weighting estimator of (11) re-weighs the source samples to match the target distribution $Q_{X}$:
[…]
The term $\widehat{Q}_{O,X}^{0}$ represents samples from the source and target distributions, and $\widehat{\omega}$, $\widehat{\mu}$, and $\widehat{\psi}$ are estimates of $\omega$, $\mu$, and $\psi(Q_{O,X}^{0})$, respectively, say by parametric or semi-parametric methods. For example, $\widehat{\mu}$ can be from OLS, $\widehat{\omega}$ can be estimated from balancing methods or selection models (see Section E.7 of the Appendix), and $\widehat{\psi}$ can be estimated by a Z-estimator with data from $\widehat{Q}_{O,X}^{0}$.
The regression estimator of (11) works by estimating the corresponding conditional expectations $\textnormal{E}_{P_{O\mid X}}$, denoted as $\widehat{\textnormal{E}}_{P_{O\mid X}}$, using parametric or semi-parametric methods:
[…]
As a concrete example, consider the target mean $\psi^{{\textnormal{mean}}}$ with $s(O,X,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$. The weighting estimator and the regression estimator of the mean’s SLOPE are
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})$ $=\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{O_{i}-\widehat{\mu}(X_{i})\}(O_{i}-\widehat{\psi}^{{\textnormal{mean}}}),$
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})$ $=\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i}).$
Here, $\hat{\psi}^{{\textnormal{mean}}}$ is the estimate of the target mean under conditional exchangeability and $\widehat{\sigma}^{2}(x)$ is an estimate of $\sigma^{2}(x)$. A simple way is to estimate $\widehat{\mu}(X)$ is via OLS and $\widehat{\sigma}^{2}(X)$ is with a log-variance linear model from weighted least squares (e.g., Harvey, (1976); Carroll and Ruppert, (1982); Davidian and Carroll, (1987)).
4.3 Extending SLOPE to Other Types of Sensitivity Analysis
Our development of SLOPE is based on the sensitivity model (3). While our reasons of choosing this model have been discussed in Section 2.2, we discuss two potential extensions of SLOPE to other sensitivity models.
First, consider an extension of (3) where we replace the exponential tilting term with a more general, non-negative function $\rho(O,X,\gamma)$. As we show below, SLOPE can still be well-defined with respect to $\rho$ and it maintains its analytic connection to the IF.
Theorem 4 (SLOPE and IF for $\rho$-Based Sensitivity Model).
Consider a broader class of sensitivity models that tilts the density ratio by a non-negative function $\rho(O,X,\gamma)$:
$\dfrac{f_{Q_{O\mid X}^{\gamma}}(O,X)}{f_{P_{O\mid X}}(O,X)}$ $\propto\rho(O,X,\gamma),$ (14)
where $\rho(O,X,0)=1$ and $\int\rho(O,X,\gamma)OdP_{O\mid X}<\infty$. Suppose $\rho$ is differentiable at $\gamma=0$ with its derivative denoted as $\dot{\rho}(O,X,0)$ and Condition 7 in the Supplement holds. Then
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{X}}\left\{\textnormal{E}_{P_{O\mid X}}\left(\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\mid X\}\right]\right)\right\}.$
Compared to Theorem 3, the relationship between IF and SLOPE defined under the sensitivity model (14) is driven by $\dot{\rho}$ at the point where conditional exchangeability holds (i.e., $\gamma=0$). Specifically, the residual variation in $\dot{\rho}$, as measured by $\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\mid X\}$ can be viewed as the subspace defined in the sensitivity analysis that is not explained by $X$. When $\rho(O,X,\gamma)=\exp(\gamma O)$ as in our original sensitivity model (3), the derivative at $\gamma=0$ is $\dot{\rho}(O,X,0)=O$ and the subspace is defined by $O-\mu(X)$.
Second, an important future direction is to generalize SLOPE to “bound-based” sensitivity analysis (e.g., Rosenbaum, (1987); Tan, (2006); Zeng et al., (2023)) where the upper bound on the difference between $Q_{O\mid X}$ and $P_{O\mid X}$ is some function of $\gamma$. However, we believe such an extension is non-trivial due to the difficulty in (a) generalizing the notion of derivative from a point to a set, and (b) the most natural generalization of this set-based derivative may not be as insightful for guiding robust designs compared to our current approach. We briefly illustrate points (a) and (b) in Remark 2 and defer details to Section J.4 of the Supplement.
Remark 2 (Challenges in Defining SLOPE for Bound-Based Models).
To fix ideas, consider Zeng et al., (2023)’s sensitivity analysis of $\psi^{\textnormal{mean}}$ in the target population. Their sensitivity analysis assumes the target conditional distribution, denoted as $Q_{O\mid X}^{\textnormal{bias}}$, deviate from $P_{O\mid X}$ by at most $\gamma$ where the deviation is measured in terms of conditional means:
$-\gamma+\textnormal{E}_{P_{O\mid X}}(O\mid X)\leq\textnormal{E}_{Q_{O\mid X}^{\textnormal{bias}}}(O\mid X)\leq\gamma+\textnormal{E}_{P_{O\mid X}}(O\mid X).$ (15)
Under (15), Zeng et al., (2023) showed that the target estimand $\psi^{\textnormal{mean}}$ is sharply bounded below and above by $\gamma+\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})$ and $\gamma+\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})$, respectively. Then, an analogous definition of SLOPE where we take the derivative of the upper and lower bounds with respect to $\gamma$ yields 1-1 and $1$, respectively. Since these two numbers disagree, the two-sided limit in Definition 1 no longer exists and the corresponding derivative is not well-defined.
Also, even if we take the maximum magnitudes of the two derivatives to resolve the issue (i.e., the “worst-case” SLOPE), we believe the resulting value (i.e., $1$) cannot be meaningfully interpreted as an intrinsic property of the study design because any source distribution $P_{O\mid X}$ or target distribution $Q_{X}$ will yield the same maximum of $1$. As mentioned in Section 3.2, a measure of robustness that is constant irrespective of the source or the target distribution does not align with some empirical recommendations on robust study designs for generalization. In contrast, our SLOPE based on either exponential tilting (3) or its generalized form (14) depends on $P_{O\mid X}$ and $Q_{X}$, and different choices of these distributions reflect differences in robustness between study designs.
5 Application
5.1 Data Background
We illustrate how to use SLOPE to inform robust study designs by re-analyzing Banerjee et al., (2015)’s multi-national experiment. The goal of the experiment was to evaluate the Graduation program in six countries (Ethiopia, Ghana, Honduras, India, Pakistan, and Peru). The Graduation program provides a holistic set of services, including asset transfers, consumption support and other career and health services, to poor households. Between 2007 and 2014, eligible households in each village were randomized to the intervention or the control group and the experiment lasted for 24 months.
We adopt the potential outcome notation stated in Example 4 where $Y(a)$ denotes the potential outcome under treatment $a$, with $a=1$ denoting participation in the program and $a=0$ denoting otherwise. In subsections below, we use SLOPE to study the violation of conditional exchangeability in transporting the potential outcome of treatment, i.e., $O=Y(1)$, from one country (i.e., the source population $P$) to another country (i.e., the target population $Q$). We remark that identification of the SLOPE requires additional causal assumptions in the source population (i.e., stable unit treatment variable assumption [SUTVA] and strong ignorability), on the source population and these assumptions are satisfied because a randomized experiment was conducted; see Sections D and F.1 in the Supplement for details.
We focus on two types of outcome variables, the per capita consumption and the physical health index, respectively, in Sections 5.2 and 5.3. For each outcome, the baseline covariate corresponds to the same variable measured prior to intervention. Our analyses are based on complete data with overlapped baseline measurement (i.e., Assumption 1 holds). To harmonize $X$ across countries for generalizability, we discretize $X$ into coarse categorical variables; note that due to the randomization of $A$, these transformations of $X$ will not affect the plausibility of strong ignorability of $A$. Finally, for estimation, we have assumed that the conditional expectation of the outcome is linear in the baseline $X$ and the village where the household locates. Moreover, for SLOPE for the median in Section 5.2, the residual of the linear model is assumed to be normal for simplicity and diagnostics in Section F.3 of the Supplement suggest that this is a reasonable assumption. see Sections F.3 and F.4 in the Supplement.
5.2 Which Source Country Is More Robust?
In this section, we study the SLOPE of transporting the per capita consumption across countries. The outcome variable is the log of the average of per capita consumption at two time points and ranges from 1.4 to 6.4. One country (Ethiopia) was excluded from the analysis because their consumption support was substantially different from other countries. For each pair of countries, we treat one as the target population, the other as the source population, and estimate the SLOPE of the mean and median (i.e., $\psi^{\rm{\textnormal{mean}}}$ and $\psi^{{\textnormal{med}}}$).
Results are shown in Table 1. Given a target country, the SLOPE is primarily determined by the data distribution in the source country and does not vary much between the mean and the median. Using India or Peru as the source population yields a lower SLOPE (i.e., a lower sensitivity) compared with using other source countries. Also, there is minimal difference in the SLOPEs of the mean and median, although the median’s SLOPEs are consistently, but slightly, lower than the mean’s SLOPEs. Section F.3.2 in the Supplement further shows that the estimated median and mean themselves (not their corresponding SLOPEs) have comparable estimated variances and thus, there is no (empirical) loss in efficiency when we estimate the median. Therefore, from this analysis based on SLOPE, we generally recommend using the median in transporting the average of per capita consumption.
To explain why India and Peru generally have smaller SLOPEs, Figure 3 plots the the distribution of $Y(1)$ given some discrete values of $X$ across countries. We see that Ghana, which is a source country that almost always has the highest SLOPE, has higher spread of $Y(1)$ conditional on each category of $X$. Conversely, India and Peru, which are the source countries with lower SLOPEs, have more concentrated values of $Y(1)$ relative to others.
Figure 3: Boxplots of $Y(1)$ across countries (in y-axis) and categories of $X$ (in panels).
We conducted some sanity checks of our recommendation and we provide a summary of them. First, because the outcome data is actually measured in the target country, we are in a unique situation where we can assess whether the sensitivity model (3) is reasonable for our analysis. Specifically, we can assess the bias from violating conditional exchangeability empirically and check whether SLOPE can approximate this bias using the first-order approximation in (5). Figure 7 in Section F.3.3 of the Supplement shows that the first-order approximation works well where the bias can be well-approximated by SLOPE. Second, a deeper subject-matter expertise explanation for why India and Peru are robust compared to other countries with respect to violation of conditional exchangeability is beyond the scope of this work. Nevertheless, we present one hypothesis in Section F.3 of the Supplement based on our understanding of Banerjee et al., (2015).
Table 1: The estimated SLOPEs for transporting the counterfactual log-transformed per capita consumption under treatment (i.e., $O=Y(1)$) from a source country (by rows) to a target country (by columns). Bootstrap standard errors are in the parentheses. For each target country and estimand, the lowest SLOPEs among the source countries are bold faced.
Estimand ($\psi$) Source ($P_{O\mid X}$) Target ($Q_{X}$)
Ghana Honduras India Pakistan Peru
Mean Ghana 0.24 (0.01) 0.24 (0.01) 0.24 (0.02) 0.24 (0.02)
Honduras 0.24 (0.01) 0.24 (0.01) 0.25 (0.02) 0.25 (0.02)
India 0.14 (0.01) 0.13 (0.01) 0.20 (0.04) 0.20 (0.04)
Pakistan 0.21 (0.03) 0.21 (0.03) 0.21 (0.04) 0.20 (0.01)
Peru 0.15 (0.01) 0.15 (0.02) 0.15 (0.02) 0.15 (0.01)
Median Ghana 0.24 (0.01) 0.24 (0.01) 0.24 (0.02) 0.24 (0.02)
Honduras 0.24 (0.01) 0.24 (0.01) 0.25 (0.02) 0.25 (0.02)
India 0.13 (0.01) 0.12 (0.01) 0.19 (0.03) 0.19 (0.03)
Pakistan 0.20 (0.03) 0.21 (0.03) 0.21 (0.04) 0.20 (0.01)
Peru 0.15 (0.02) 0.15 (0.02) 0.14 (0.02) 0.15 (0.01)
5.3 Which Health Index Is More Robust for Generalization?
Banerjee et al., (2015) constructed a physical health index to capture the overall physical health of individuals in a household. Specifically, the index is an (equally weighted) average of three standardized variables (z-scores): did not miss work due to illness ($Y_{\text{notMiss}}(1)$), activities of daily living score ($Y_{\text{act}}(1)$), and perception of health status ($Y_{\text{perc}}(1)$). In this section, we ask whether there is another way to define the physical health index so that it’s less sensitive for generalization. Formally, suppose we rewrite the physical health index as a weighted average of three z-scores:
$O=\alpha_{\text{notMiss}}Y_{\text{notMiss}}(1)+\alpha_{\text{act}}Y_{\text{act}}(1)+\alpha_{\text{perc}}Y_{\text{perc}}(1).$
In the original analysis, the weights $\alpha_{\text{notMiss}}$, $\alpha_{\text{act}}$, and $\alpha_{\text{perc}}$ were set to $1/3$ (i.e., equally weighted). Our goal is to find a new vector of weights $\alpha=[\alpha_{\text{notMiss}},\alpha_{\text{act}},\alpha_{\text{perc}}]^{\intercal}$ in a simplex that minimizes the SLOPE of the mean, i.e., $\psi^{{\textnormal{mean}}}=\textnormal{E}_{Q_{O}}(O)$.
We focus on households in three countries (India, Ethiopia, Peru) where all three variables that make up the index were measured. Also, like before, we filtered households so that the overlap was plausible.
Figure 4 presents the SLOPE across different weight combinations where the target country is Ethiopia and the source country is (left) India or (right) Peru. When the source country is India, the SLOPE is minimized at $\alpha_{\text{notMiss}}=0.10$, $\alpha_{\text{act}}=0.55$, $\alpha_{\text{perc}}=0.35$. Upon closer examination, the variable notMiss has the highest variance compared to the other two variables and thus, putting a low weight on it minimizes the overall variance of $O$. Also, when the source country is Peru, the SLOPE is minimized with the weight representing the left bottom vertex where $\alpha_{\text{notMiss}}=1$. Again, upon closer examination, this was because the variance of $Z_{\text{notMiss}}$ in Peru is substantially lower than the other two variables. More broadly, similar to the theoretical discussion from Section 3.2 and Corollary 1, SLOPE prefers distributions of the physical index that are less variable in order to be less sensitive to generalization.
Figure 4: SLOPE for the mean of the physical health index across different weights. In each panel, edges of the triangle represent the weight in percentage scale of the three z-scores that make up the index. The contours represent the SLOPE where a lighter color means a lower SLOPE. The star represents the point in the simplex where SLOPE is minimized.
6 Discussion
This main contribution of this paper is a simple and novel measure, SLOPE, to design robust studies for generalization. Specifically, SLOPE is inspired by principles from robust statistics and is a derivative-based metric that measures the change in the target estimand when there is a local violation of the conditional exchangeability assumption. SLOPE depends on two design-level quantities, (a) the target estimand and (b) the source and target distributions $P_{O\mid X}$ and $Q_{X}$, respectively. Changing either of these quantities can result in a different SLOPE and thus, researchers can assess the robustness of competing study designs for generalization.
Inspired by Tipton and Olsen, (2018)’s recommendations for robust generalization, we summarize some advice for designing robust studies based on SLOPE. We remark that these principles are assuming that Assumption 1 holds:
  1. 1.
    If the target population is fixed and the target estimand is the mean of $O$, a source population will be less sensitive to local violations of conditional exchangeability if the spread of $O$ given $X$ is small. For instance, the mean’s SLOPE becomes smaller when $\sigma^{2}(X)$ becomes smaller (Corollary 1). From our data example in Section 5.2, countries with a smaller SLOPE (e.g., India and Peru as shown in Table 1) also have a less variable distribution of $O$ given $X$ (Figure 3).
  2. 2.
    If the source population is fixed and the target estimand is the mean of $O$, it’s essential to understand which region of $X$ leads to the least amount of variation in $O$ in the source population. Once this region of $X$ is identified, a target population will be robust for generalization if it is homogeneous with respect to its $X$ and its $X$s focus around this region. More concretely, Section 3.2 and Corollary 1 discuss an example where $Q_{X}$ is selected to focus on regions with the least variability in $O$.
  3. 3.
    If both the source and the target populations are fixed, it’s less sensitive to choose a target estimand whose influence function projects more (in proportion) onto the space of shared variable $X$. In a trivial example inspired by Theorem 3, a target functional that concerns $X$ only, i.e., $\psi(Q_{O,X})=\psi(Q_{X})$, has zero SLOPE. Additionally, as shown in Section 5.3, for a weighted average of several physical health variables, increasing weights to variables that are better explained by $X$ will also reduce the magnitude of SLOPE, thereby improving robustness.
We re-emphasize that SLOPE is developed under the validity of the overlap assumption (i.e., Assumption 1), which means that the source population is already sufficiently large compared with the target population, in terms of $P_{X}$ and $Q_{X}$. In addition, we echo Tipton and Olsen, (2018) and Degtiar and Rose, (2023) on the general importance of guaranteeing conditional exchangeability through discussion with domain experts and careful data collection processes (e.g., by collecting a rich $X$). However, when it is infeasible or impractical to plausibly satisfy conditional exchangeability with the observed set of $X$, we believe SLOPE is a useful tool to assess the sensitivity/robustness of the underlying study design, and to guide future designs for generalization.
Finally, we highlight some extensions and future directions. First, while this paper focuses on violations of conditional exchangeability in generalizability, by defining $P$ and $Q$ differently, SLOPE naturally extends to measuring the sensitivity of conditional exchangeability in causal inference and missing data problems; see Section J.3 of the Supplement for these extensions. Second, Section J.5 of the Supplement connects SLOPE to the marginal interventional effect proposed by Zhou and Opacic, (2022) with the incremental propensity score intervention (Kennedy,, 2019), highlighting some potential mathematical connections between SLOPE and incremental treatment effects. Third, while our result has been based on sensitivity model (3), the high level idea of SLOPE as a derivative-based robustness measure for study designs is “generalizable.” As stated in Section 4.3, an important direction is to extend SLOPE to bound-based sensitivity models where a properly defined SLOPE not only exists, but also provides useful insights about robust study designs.
Acknowledgment
The authors gratefully acknowledge Guanhua Chen, Melody Huang, Edward Kennedy, Jae-Kwang Kim, Qingyuan Zhao, and participants of the 2024 International Conference on Statistics and Data Science, 2024 Joint Statistical Meetings, 2025 American Causal Inference Conference, and the statistics seminar at Seoul National University for their valuable feedback. The work of Xinran Miao is supported in part by funding from American Family Funding Initiative and the Morgridge Summer Research Fellowship.
Appendix
The Appendix is organized as follows. Section A discusses the existence of SLOPE, including the definition of SLOPE through Hadamard differentiability and some regularity conditions deferred from the main text. Section B exemplifies the use of SLOPE in choosing a robust location estimand. Section C discusses SLOPE for other estimands, including risks, quantiles, trimmed mean, OLS coefficients, some scale parameters, Pearson correlation coefficient, and general formulas of SLOPEs for L-estimands and Z-estimands. Section D discusses transporting functionals of potential outcomes. Next, Section E and Section F supplement Section 4.2 (estimation) and Section 5 (data application) in the main text, respectively, by providing details, auxiliary results, and some deferred discussions. Section G includes a simulation study that verifies the asymptotic properties of proposed estimators. Afterwards, Sections H and I provide proof for the derivation of the SLOPE (at population level) and the estimation of the SLOPE, respectively. Finally, Section J details some extended remarks, including SLOPE with a vector valued $\psi(\cdot)$, SLOPE for other types of conditional exchangeability assumptions, the challenge of extending SLOPE to bound-based sensitivity models, and the mathematical connection between SLOPE and the marginal interventional effect.
Appendix A Existence of SLOPE
A.1 Notation
We introduce some notation for normed spaces and functional analysis (Van Der Vaart and Wellner,, 1996). For an arbitrary set $T$, let $l^{\infty}(T)$ be the set of all uniformly bounded, real functions on $T$: the set with elements satisfying $z:T\to\mathbb{R}$ such that $\lVert z\rVert_{T}:=\sup_{t\in T}|z(t)|<\infty$. For two normed spaces $\mathbb{D}$ and $\mathbb{E}$, a map $\phi:\mathbb{D}_{\phi}\subset\mathbb{D}\to\mathbb{E}$ is Hadamard differentiable at $\theta\in\mathbb{D}_{\phi}$ tagentially to $\mathbb{D}_{0}$ if there exists a continuous linear map $\phi_{\theta}^{\prime}:\mathbb{D}\to\mathbb{E}$ such that $\{\phi(\theta+t_{n}h_{n})-\phi(\theta)\}/t_{n}\to\phi_{\theta}^{\prime}(h)$ as $n\to\infty$, for all converging sequences $t_{n}\to 0$ and $h_{n}\to h$ such that $\theta+t_{n}h_{n}\in\mathbb{D}_{\phi}$ and $h\in\mathbb{D}_{0}$, where $\mathbb{D}_{\phi}$ and $\mathbb{D}_{0}$ are two subsets of $\mathbb{D}$.
We introduce notations for the supports under $Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$. Let $\mathcal{S}_{O,X}$ be the support under $Q_{O,X}^{0}$ and $\mathcal{S}_{X}$ and $\mathcal{S}_{O}$ be the supports for the marginals $Q_{X}$ and $Q_{O}^{0}$, respectively.
A.2 SLOPE Through Hadamard Differentiability.
We define SLOPE through the derivative of a composite of two functionals. First, define the map from $\gamma$ to $Q_{O,X}^{\gamma}$ as $\phi:\left(\mathbb{R},l^{\infty}(\mathcal{S}_{O,X})\right)\to l^{\infty}(\mathcal{S}_{O,X})$ such that $(\gamma,Q_{O,X}^{0})\mapsto Q_{O,X}^{\gamma}$, with domain $\mathbb{D}_{\phi}=\left([-\varepsilon,\varepsilon],Q_{O,X}^{0}\right)\subset\left(\mathbb{R},l^{\infty}(\mathcal{S}_{O,X})\right)$ for some $\varepsilon>0$. Next, consider the functional $\psi$ as $l^{\infty}(\mathcal{S}_{O,X})\to\mathbb{R}$ with domain $\mathbb{D}_{\psi}\subset l^{\infty}(\mathcal{S}_{O,X})$ which contains probability distributions on $\mathcal{S}_{O,X}$. Then holding $Q_{O,X}^{0}$ fixed, SLOPE as defined in (4) is the Hadamard derivative of the composite function $\psi\circ\phi$ with respect to $\gamma$ at zero. It exists under Condition 2, the standard condition that enables the chain rule of Hadamard differentiability. For Z-estimands, SLOPE exists under Condition 1 of the main text. Part (i) ensures exchanging integration and differentiation and the existence of SLOPE as an integral; parts (ii) and (iii) are analogous to regularity conditions that ensures the existence of the IF of $\psi$ (see Section 4.1).
Condition 2 (Existence of SLOPE).
Suppose $\phi$ as a function of $\gamma$ is Hadamard differentiable at 0 and $\psi$ is Hadamard differentiable at $\phi(0,Q_{O,X}^{0})=Q_{O,X}^{0}$ tangentially to $\phi^{\prime}_{0}(\{0\})$.
Condition 3 is a regularity condition that ensures IF is an evaluation of the Hadamard derivative and therefore its connection to SLOPE follows from the linearity of Hadamard differentiability. We note that Hadamard differentiability (i.e., Condition 3) is stronger than the directional differentiability (i.e., (9) in the main text) since the former requires the limit to exist for every sequence of directions that converges to $\delta_{o,x}-Q_{O,X}^{0}$ whereas the latter only requires the limit to exist in this single direction.
Condition 3 (IF as an Evaluation of a Hadamard Derivative).
Suppose $\psi$ is Hadamard differentiable at $Q_{O,X}^{0}$ tangentially to $\delta_{o,x}-Q_{O,X}^{0}$.
A.3 Regularity Conditions
Condition 4 (Existence of SLOPE for the Mean).
Suppose $\textnormal{E}_{P_{O\mid X}}\{O\exp(\gamma O)\mid X\}/\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}$ is uniformly bounded by an integrable function under $Q_{X}$ for $\gamma$ in a neighborhood of zero, and $\sigma^{2}(X)<\infty$ almost surely on $Q_{X}$.
Condition 5 (Existence of SLOPE for the Median).
(i) Suppose $F_{Q_{O}^{0}}$ is differentiable at $m_{1/2}$ with a positive derivative.
(ii) Suppose $\textnormal{E}_{P_{O\mid X}}\left[O\mathds{1}(O\leq m_{1/2})\exp(\gamma O)\mid X\right]/\textnormal{E}_{P_{O\mid X}}(O\mid X)$ is uniformly bounded by an integrable function under $Q_{X}$ for $\gamma$ in a neighborhood of zero, and $\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{1/2}\mid X)\mu(X)\right]$ and $\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{1/2})\right]$ exist.
Condition 6 (Existence of SLOPE for the $q$-th Quantile).
(i) Suppose $F_{Q_{O}^{0}}$ is differentiable at $m_{q}$ with a positive derivative.
(ii) Suppose $\textnormal{E}_{P_{O\mid X}}\left[\mathds{1}(O\leq m_{q})O\exp(\gamma O)\mid X\right]/\textnormal{E}_{P_{O\mid X}}(O\mid X)$ is uniformly bounded by an integrable function under $Q_{X}$ for $\gamma$ in a neighborhood of zero, and $\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{q}\mid X)\mu(X)\right]$ and $\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{q})\right]$ exist.
Condition 7 (Existence of SLOPE for Sensitivity Model Defined Through $\rho(O,X,\gamma)$).
(i) $\textnormal{E}_{Q_{O\mid X}^{\gamma}}[s(O,X,\psi(Q_{O,X}^{0}))]$ is bounded for $\gamma$ in a neighborhood of zero and
$\textnormal{E}_{Q_{O,X}^{0}}\left(s(O,X,\psi(Q_{O,X}^{0}))\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\mid X\}\right]\right)$
exists; (ii) $s(O,X,\cdot)$ is differentiable almost everywhere with the derivative $\dot{s}(O,X,\cdot)$; (iii) $\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi(Q_{O,X}^{0}))\}$ exists and is non-singular; (iv) $\rho(O,X,\gamma)$ is twice differentiable with respect to $\gamma$ at a neighborhood of zero with derivative $\ddot{\rho}(O,X,\gamma)$.
Appendix B Using SLOPE to Choose A Robust Location Parameter
In this section, we illustrate how SLOPE can guide the choice of a robust location parameter. We focus on comparing two estimands, the mean with functional $\psi^{\textnormal{mean}}$ and the median with functional $\psi^{\textnormal{med}}$. Their SLOPEs have been stated in Theorems 1 and 2 in the main text. When $P_{O\mid X}$ is normal, we recall that in this case the both SLOPEs are weighted average of the conditional variance $\sigma^{2}(X)$. In this case, Theorem 5 lists some sufficient conditions where the SLOPE for the mean is larger than (or equal to) the SLOPE for the median.
Theorem 5 (Comparison in SLOPEs of Mean and Median).
Suppose $P_{O\mid X}\sim N\left(\mu(X),\sigma^{2}(X)\right)$.
(a) If $\sigma^{2}(X)=\sigma^{2}$ almost surely $Q_{X}$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=\sigma^{2}$.
(b) If $\mu(X)=\mu$ almost surely $Q_{X}$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})\geq\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$.
(c) More generally, $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{mean}})\geq\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})$ if and only if
$\textnormal{Corr}_{Q_{X}}[\sigma^{2}(X),f_{P_{O\mid X}}(m_{1/2}\mid X)]\leq 0,\text{ where }\textnormal{Corr}\text{ represents correlation}.$
In part (a) of Theorem 5 where the conditional variance is constant, the SLOPEs for the median and the mean are identical. From part (b) of Theorem 5, roughly speaking, when the conditional means of $P_{O\mid X}$ are sufficiently uniform, the median is more robust than or equally robust with the mean in terms of violation of conditional exchangeability. Intuitively, this is because the conditional density at the median, i.e., $f_{P_{O\mid X=x}}(m_{1/2})$, should be higher for $x$’s with a lower conditional variance ($\sigma^{2}(x)$). Then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$, as a weighted average of $\sigma^{2}(x)$, will assign a higher weight to $x$’s with a lower $\sigma^{2}(x)$. This makes the SLOPE for the median no larger than the SLOPE for the mean, which is an (unweighted) average of $\sigma^{2}(x)$.
While Theorem 5 lists some sufficient conditions for the median to be more robust than the mean, in general, the relationship can be reversed. For example, suppose the target population contains two subgroups indicated by $X\in\{x_{1},x_{2}\}$. In each subgroup, the source distribution is normal with $P_{O\mid X=x_{1}}\sim N(\mu_{1},\sigma_{1}^{2})$ and $P_{O\mid X=x_{2}}\sim N(\mu_{2},\sigma_{2}^{2})$, where we set $\mu_{1}=0$, $\sigma_{1}=0.5$, and $\sigma_{2}=0.6$. Then SLOPE of the median becomes
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})=(1-w_{2})\sigma_{1}^{2}+w_{2}\sigma_{2}^{2},$
where $w_{2}$ is the weight for the more hetereogeneous subgroup ($X=x_{2}$) with the higher variance $\sigma_{2}^{2}$. As shown in Lemma 2, once $\mu_{1}$ and $\sigma_{2}>\sigma_{1}$ are fixed, $w_{2}$ (as a function of $\mu_{2}-\mu_{1}$) is monotonically increasing with $\mu_{2}-\mu_{1}$, the mean difference between the two subgroups. Therefore, as $w_{2}$ increases, the SLOPE for the median assigns an increasing weight for the heterogeneous subgroup, which will eventually exceed the SLOPE for the mean. Intuitively, as $\mu_{2}$ increases, the marginal target distribution under exchangeability $Q_{O}^{0}\sim 0.2N(0,0.5^{2})+0.8N(\mu_{2},0.6^{2})$ becomes more asymmetric and less uni-modal, and thus the usual understanding on median’s robustness no longer holds. Figure 5 provides a visual illustration of different underlying marginal distributions $Q_{O}^{0}$ and how they correspond to different SLOPEs for the median and the mean.
Figure 5: Some toy examples where the SLOPE of the median is more, equal, or less than the SLOPE of the mean.
Appendix C SLOPE for Other Estimands
In this section we provide examples of SLOPE with some other target functionals $\psi(\cdot)$ that were not included in the main text.
C.1 SLOPE for Expectations
Suppose the target estimand is the expectation of a known function $\xi(O,X)$,
$\psi^{\xi}(Q_{O,X})$ $=\textnormal{E}_{Q_{O,X}}\{\xi(O,X)\}.$ (16)
Lemma 1 shows that SLOPE is the expectation (under $Q_{X}$) of the conditional covariance (under $P_{O\mid X})$ of $O$ and $\xi(O,X)$.
Lemma 1.
Suppose Condition 1 holds with $s(O,\psi^{\xi})=\psi^{\xi}-\xi(O,X)$. Then SLOPE for the functional $\psi^{\xi}$ defined in (16) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\xi})$ $=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}\left[O,\xi(O,X)\mid X\right]\right\}.$
For example, if $\xi(O,X)=O$, then $\psi^{\xi}=\psi^{{\textnormal{mean}}}$ and the SLOPE is the expectation of the conditional variance of $O$. Additional examples include the centered moment in Example 5, risks and excess risks in Section C.2.
Example 5 (Centered Moments).
Consider $\psi(Q_{O,X})=\textnormal{E}_{Q_{X}}[\{O-\mu(X)\}^{k}]$ for a positive integer $k$. Then the SLOPE is $\textnormal{E}_{Q_{O,X}^{0}}[\{O-\mu(X)\}^{k+1}]$ provided that the expectation exists.
C.2 SLOPE for Risks
Suppose $\delta(X)$ is a (fixed) decision rule that maps from $X$ to the support of $O$. For a loss function $L(O,\cdot)$, define the risk of $\delta(X)$ on the target distribution as $R(\delta)=\textnormal{E}_{Q_{O,X}}[L(O,\delta(X))]$. In this section we consider SLOPEs for risks and excess risks. Note that they can be viewed as special cases of the SLOPE for expectations in Lemma 1 with different choices of $\xi$.
C.2.1 SLOPE for Mean Squared Error
Consider the squared loss with $L(O,\delta(X))=\{O-\delta(X)\}^{2}$. The target functional is the mean squared error of the decision $\delta(X)$ on the target population, i.e., $\psi=\textnormal{E}_{Q_{O,X}}[\{O-\delta(X)\}^{2}]$. Then the SLOPE is $\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=\textnormal{E}_{Q_{O,X}^{0}}\left[\{\delta(X)-O\}^{2}\{O-\mu(X)\}\right]$.
C.2.2 SLOPE for Excess Risk under Squared Loss
Consider again the squared loss $L(O,\delta(X))=\{O-\delta(X)\}^{2}$. Then $\mu(X)$ minimizes the corresponding risk. For any (fixed) $\delta(X)$, let the target functional be the excess risk, i.e., $\psi(Q_{O,X})=R(\delta)-R(\mu)=\textnormal{E}_{Q_{O,X}}\left[\{\delta(X)-O\}^{2}-\{\mu(X)-O\}^{2}\right]$. SLOPE for this excess risk under squared loss is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=2\textnormal{E}_{Q_{X}}\left[\{\mu(X)-\delta(X)\}\sigma^{2}(X)\right].$
C.2.3 SLOPE for Excess Risk under 0-1 Loss
Suppose $O\in\{1,-1\}$ and let $\eta(X)=P(O=1\mid X)$. Consider the 0-1 loss, $L(O,\delta(X))=\mathds{1}\{\delta(X)\neq O\}$, and the corresponding risk
$R(\delta)=\textnormal{E}_{Q_{O,X}}\{L(O,\delta(X)\}=\textnormal{E}_{Q_{O,X}}\left[\mathds{1}\{\delta(X)\neq 1\}\{2\eta(X)-1\}+1-\eta(X)\right].$
Then one of the Bayes classifiers is $\delta^{*}(X)=\mathds{1}\{\eta(X)\geq 1/2\}-\mathds{1}\{\eta(X)<1/2\}$. For any fixed $\delta(X)$, let the target functional be the excess risk, $\psi(Q_{O,X})=R(\delta)-R(\delta^{*})$. Then the SLOPE for the excess risk under 0-1 loss is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=\textnormal{E}_{Q_{X}}\left[\sigma^{2}(X)\mathds{1}\{\delta(X)\neq\delta^{*}(X)\}\textnormal{sign}\{\eta(X)-1/2\}\right],$
where $\textnormal{sign}(\cdot)$ is the sign function such that $\textnormal{sign}(t)=1$ if $t>0$, $\textnormal{sign}(t)=0$ if $t=0$, and $\textnormal{sign}(t)=-1$ if $t<0$.
C.3 SLOPE for Quantiles
For $q\in(0,1)$, let $\psi(Q_{O})=F_{Q_{O}}^{-1}(q)$ be the $q$ quantile of the marginal distribution $Q_{O}$; for example when $q=1/2$, $\psi=\psi^{{\textnormal{med}}}$. Then SLOPE for the quantile, as a general case of the SLOPE for the median (Theorem 2), is presented in Theorem 6.
Theorem 6 (SLOPE for Quantiles).
Suppose Condition 6 holds. Then SLOPE for the $q$-th quantile is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=$ $\dfrac{\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{q}\mid X)\mu(X)\right]-\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{q})\right]}{f_{Q_{O}^{0}}(m_{q})},$
where $m_{q}$ satisfies $\int_{-\infty}^{m_{q}}dQ_{O}^{0}=q$.
C.4 SLOPE for $\alpha$-Trimmed Mean
Let the target functional be the $\alpha$-trimmed mean, $\psi^{{\alpha\textnormal{-trim}}}$, which is the mean of $Q_{O}$ after trimming off the lower and upper $\alpha$ quantiles. Specifically, the $\psi^{{\alpha\textnormal{-trim}}}$ is defined as
$\psi^{{\alpha\textnormal{-trim}}}(Q_{O})=\dfrac{1}{1-2\alpha}\int_{\alpha}^{1-\alpha}F_{Q_{O}}^{-1}(p)dp,$ (17)
for $\alpha\in(0,1/4)$.
Theorem 7 (SLOPE for $\alpha$-Trimmed Mean).
SLOPE for the $\alpha$-trimmed mean $\psi^{{\alpha\textnormal{-trim}}}$ defined in (17) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\alpha\textnormal{-trim}}})=$ $\frac{F^{-1}_{Q_{O}^{0}}(1-\alpha)}{1-2\alpha}\textnormal{E}_{Q_{O,X}^{0}}\left\{F_{P_{O\mid X}}(F_{Q_{O}^{0}}^{-1}(1-\alpha))\mu(X)-O\mathds{1}_{(-\infty,F_{Q_{O}^{0}}^{-1}(1-\alpha)]}(O)\right\}$
$-\frac{F^{-1}_{Q_{O}^{0}}(\alpha)}{1-2\alpha}\textnormal{E}_{Q_{O,X}^{0}}\left\{F_{P_{O\mid X}}(F_{Q_{O}^{0}}^{-1}(\alpha))\mu(X)-O\mathds{1}_{(-\infty,F_{Q_{O}^{0}}^{-1}(\alpha)]}(O)\right\}$
$+\dfrac{1}{1-2\alpha}\textnormal{E}_{Q_{O,X}^{0}}\left[O\{O-\mu(X)\}\mathds{1}_{\left(F_{Q_{O}^{0}}^{-1}(\alpha),F_{Q_{O}^{0}}^{-1}(1-\alpha)\right)}(O)\right].$
Moreover, if $P_{O\mid X}\sim N\left(\mu(X),\sigma^{2}(X)\right)$, then the SLOPE is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\alpha\textnormal{-trim}}})=\dfrac{1}{1-2\alpha}\textnormal{E}_{Q_{X}}\left[\sigma^{2}(X)\left\{\Phi\left(\frac{F_{Q_{O}^{0}}^{-1}(1-\alpha)-\mu(X)}{\sigma(X)}\right)-\Phi\left(\frac{F_{Q_{O}^{0}}^{-1}(\alpha)-\mu(X)}{\sigma(X)}\right)\right\}\right],$
where $\Phi(\cdot)$ is the cumulative distribution function of standard normal.
C.5 SLOPE for OLS Coefficients
Suppose $O=Y$ is an outcome variable and $X$ is a vector of covariates that includes one. We are interested in the OLS coefficient of regressing $Y$ on $X$ in the target distribution, i.e., $\psi^{{\textnormal{OLS}}}(Q_{Y,X})$ such that
$\textnormal{E}_{Q_{Y,X}}\left[XX^{\intercal}\psi^{{\textnormal{OLS}}}-XY\right]=0.$ (18)
C.5.1 SLOPE for OLS Coefficients with Omitted Variables
We consider transferring OLS coefficients $\psi^{{\textnormal{OLS}}}$ in (18) where $X$ is a vector of covariates and contains one as the intercept, $O=Y$ is a scalar outcome variable. The target estimand $\psi^{{\textnormal{OLS}}}$ can be equivalently defined as the solution to the following least squares problem,
$\psi^{{\textnormal{OLS}}}=\arg\min_{\psi^{{\textnormal{OLS}}}}\textnormal{E}_{Q_{O,X}}\left[\left(Y-X^{\intercal}\psi^{{\textnormal{OLS}}}\right)^{2}\right]$ (19)
For generality, we also consider the OLS coefficient where only a subset $X_{\textnormal{Sub}}\subset X$ that contains the intercept is being modeled:
$\psi^{\textnormal{OLS,Sub}}=\arg\min_{\psi^{\textnormal{OLS,Sub}}}\textnormal{E}_{Q_{O,X}}\left[\left(Y-X_{\textnormal{Sub}}^{\intercal}\psi^{\textnormal{OLS,Sub}}\right)^{2}\right].$ (20)
Note that (20) contains (19) as a special case by setting $X_{\textnormal{Sub}}=X$.
Theorem 8 (SLOPE for OLS Coefficient).
Suppose Condition 1 holds with $s(Y,\psi^{\textnormal{OLS,Sub}})=X_{\textnormal{Sub}}X_{\textnormal{Sub}}^{\intercal}\psi^{\textnormal{OLS,Sub}}-X_{\textnormal{Sub}}Y$. SLOPE for the OLS coefficient with a subset of covariates (20) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{OLS,Sub}})=\left\{\textnormal{E}_{Q_{X}}\left(X_{\textnormal{Sub}}X_{\textnormal{Sub}}^{\intercal}\right)\right\}^{-1}\textnormal{E}_{Q_{X}}\left\{X_{\textnormal{Sub}}\sigma^{2}(X)\right\}.$
In the special case where $X_{\textnormal{Sub}}=1$, coefficient $\psi^{\textnormal{OLS,Sub}}$ becomes the outcome mean (Theorem 1). As expected, SLOPE in Theorem 8 becomes $\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}$, which is identical to the SLOPE for the mean. On the contrary, in another special case where $X_{\textnormal{Sub}}=X$, the SLOPE becomes $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})=\left\{\textnormal{E}_{Q_{X}}\left(XX^{\intercal}\right)\right\}^{-1}\textnormal{E}_{Q_{X}}\left\{X\sigma^{2}(X)\right\}$.
C.5.2 SLOPE in ANCOVA
Suppose $X=[1,A,L^{\intercal}]^{\intercal}$ where $A$ is a binary treatment and $L$ contains pre-treatment covariates. We are interested in the effect of $A$ on $Y$ in two models of analysis of covariance (ANCOVA) in which (21) adjusts for covariates $L$ and (22) does not:
$(\alpha_{\textnormal{a}},\tau_{\textnormal{a}},\beta_{\textnormal{a}})$ $=\arg\min_{(\widetilde{\alpha}_{\textnormal{a}},\widetilde{\tau}_{\textnormal{a}},\widetilde{\beta}_{\textnormal{a}})}\textnormal{E}_{Q_{O,X}^{0}}\left[\left(Y-\widetilde{\alpha}_{\textnormal{a}}-\widetilde{\tau}_{\textnormal{a}}A-\widetilde{\beta}_{\textnormal{a}}^{\intercal}L\right)^{2}\right],$ (21)
$(\alpha_{\textnormal{u}},\tau_{\textnormal{u}})$ $=\arg\min_{(\widetilde{\alpha}_{\textnormal{u}},\widetilde{\tau}_{\textnormal{u}})}\textnormal{E}_{Q_{O,X}^{0}}\left[\left(Y-\widetilde{\alpha}_{\textnormal{u}}-\widetilde{\tau}_{\textnormal{u}}A\right)^{2}\right].$ (22)
Specifically, consider SLOPE for the regression slopes $\tau_{\textnormal{a}}$ and $\tau_{\textnormal{u}}$. We will see that the conditional varaince $\sigma^{2}(X)$ still plays an important role. To reflect its dependency on the treatment $A$ and covariates $L$, we slightly abuse the notation and let $\sigma^{2}(L,A)=\textnormal{Var}_{P_{Y\mid X}}(Y\mid A,L)=\textnormal{Var}_{P_{Y\mid X}}(Y\mid X)$.
Remark 3 (Difference Between Transporting $Y\mid A$ and Transporting $Y(A)$).
Coefficients $\tau_{\textnormal{a}}$ and $\tau_{\textnormal{u}}$ can be interpreted as the (causal) treatment effect of $A$ on $Y$ under modeling assumptions. Nevertheless, they are different from the causal effects in Section D which are based on transporting potential outcomes, even under the same modeling assumptions.
In Section D, the transportation is for the potential outcome $O=Y(A)$, based on the commonly observed pre-treatment covariates in the two populations (i.e., $X$). The SLOPE depends on target population only through the distribution of pre-treatment covariates; notably, the treatment $A$ itself need not be well defined on the target population. The SLOPE depends on the source population only through the distribution of the potential outcome $Y(A)$ given pre-treatment covariates. It does not directly depend on how $A$ has been randomized on the source population, as longs as the identification is guaranteed. This setting is widely adopted in generalizing a causal effect from one population where an experimental has taken place to another population where only baseline covariates are collected.
In this Section, the transportation is for the observed outcome $O=Y$, based on the commonly observed intervention and pre-treatment covariates (i.e., $X$ contains both $A$ and pre-treatment covariates). The intervention $A$ is observed on the target population and the SLOPE depends on the target population not only through covariates, but also through $A$. This setting is less common than the previous one. One example of this setting is in surveys (or in general, missing data) where $P$ contains units whose outcome variables are observed and $Q$ contains units whose outcome variables are not observed.
The following Theorem 9 gives the SLOPE for regression slopes $\tau_{\textnormal{a}}$ and $\tau_{\textnormal{u}}$.
Theorem 9 (SLOPE in ANCOVA).
The SLOPEs for $\tau_{\textnormal{u}}$ and $\tau_{\textnormal{a}}$ are
$\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{u}})$ $=\dfrac{\textnormal{Cov}_{Q_{L,A}}[A,\sigma^{2}(L,A)]}{\textnormal{Var}_{Q_{A}}(A)},$ (23)
$\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})$ $=\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{u}})+\delta^{\intercal}V^{-1}\delta\textnormal{Cov}_{Q_{L,A}}[A,\sigma^{2}(L,A)]-\delta^{\intercal}V^{-1}\textnormal{Cov}_{Q_{L,A}}[L,\sigma^{2}(L,A)],$ (24)
respectively, where $\delta=\textnormal{Cov}_{Q_{X}}[L,A]/\textnormal{Var}_{Q_{A}}[A]$ and $V=\textnormal{Cov}_{Q_{L}}(L)-\delta\delta^{\intercal}\textnormal{Var}_{Q_{A}}(A)$.
From Theorem 9, if $\textnormal{Cov}_{Q_{L,A}}(L,A)=0$ almost surely, e.g., in cases when the intervention $A$ has been randomized, then by definition $\delta=0$ and
$\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})=\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{u}})=\textnormal{Cov}_{Q_{L,A}}[A,\sigma^{2}(L,A)]/\textnormal{Var}_{Q_{A}}(A).$
C.6 SLOPE for Scale Parameters
C.6.1 SLOPE for Variance
Let the target functional be the variance of $O$, i.e., $\psi^{\text{var}}=\textnormal{Var}_{Q_{O}}(O)$. Then the SLOPE is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\text{var}})=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}[O^{2},O\mid X]\right\}-2\textnormal{E}_{Q_{O}^{0}}(O)\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}.$
C.6.2 SLOPE for MAD
Let the target functional be the median absolute deviation from the median (MAD), $\psi^{{\textnormal{MAD}}}(\cdot)$ such that
$Q_{O}\left(\left|O-\psi^{{\textnormal{med}}}(Q_{O})\right|\leq\psi^{{\textnormal{MAD}}}\right)=1/2,$ (25)
where $\psi^{{\textnormal{med}}}$ is the functional that maps to the marginal median which has been defined in Example 2 of the main text. Recall that $m_{1/2}=\psi^{{\textnormal{med}}}(Q_{O}^{0})$ and to ease notation, let ${\textnormal{MAD}}=\psi^{{\textnormal{MAD}}}(Q_{O}^{0})$ be the MAD of $Q_{O}^{0}$.
Theorem 10 (SLOPE for MAD).
Suppose Condition 2 holds for $\psi^{{\textnormal{MAD}}}$. Then SLOPE for the MAD defined through $\psi^{{\textnormal{MAD}}}$ in (25) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{MAD}}})=$ $\dfrac{f_{Q_{O}^{0}}\left(m_{1/2}-{\textnormal{MAD}}\right)-f_{Q_{O}^{0}}\left(m_{1/2}+{\textnormal{MAD}}\right)}{f_{Q_{O}^{0}}\left(m_{1/2}-{\textnormal{MAD}}\right)+f_{Q_{O}^{0}}\left(m_{1/2}+{\textnormal{MAD}}\right)}\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$
$-\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\mathds{1}_{[m_{1/2}-{\textnormal{MAD}},m_{1/2}+{\textnormal{MAD}}]}(O)\{O-\mu(X)\}\right]}{f_{Q_{O}^{0}}\left(m_{1/2}-{\textnormal{MAD}}\right)+f_{Q_{O}^{0}}\left(m_{1/2}+{\textnormal{MAD}}\right)},$
where $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$ is the SLOPE for median (Theorem 2).
C.6.3 SLOPE for $\alpha$-Quantile Range
The $\alpha$-quantile range of $Q_{O}$ is $\psi^{\alpha\text{-range}}(Q_{O})=F_{Q_{O}}^{-1}(1-\alpha)-F_{Q_{O}}^{-1}(\alpha)$ for a given $\alpha\in(0,1/2)$. A common choice is to let $\alpha=1/4$ and the $\alpha$-quantile range is the interquartile range. By the SLOPE for quantiles (Theorem 6), the SLOPE for the $\alpha$-quantile range is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\alpha\text{-range}})=$ $\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[(1-\alpha)O-\mathds{1}\left(O\leq F_{Q_{O}^{0}}^{-1}(1-\alpha)\right)O\right]}{f_{Q_{O}^{0}}\left(F_{Q_{O}^{0}}^{-1}(1-\alpha)\right)}$
$-\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\alpha O-\mathds{1}\left(O\leq F_{Q_{O}^{0}}^{-1}(\alpha)\right)O\right]}{f_{Q_{O}^{0}}\left(F_{Q_{O}^{0}}^{-1}(\alpha)\right)}.$
C.7 SLOPE for Pearson Correlation Coefficient
Suppose $O\in\mathbb{R}$ and let the target functional be the Pearson correlation between $O$ and $X$:
$\psi^{\textnormal{Corr}}(Q_{O,X})=\dfrac{\textnormal{E}_{Q_{O,X}}(XO)-\textnormal{E}_{Q_{X}}(X)\textnormal{E}_{Q_{O}}(O)}{\sqrt{\textnormal{Var}_{Q_{X}}(X)\textnormal{Var}_{Q_{O}}(O)}}.$ (26)
The SLOPE for the Pearson correlation coefficient is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{Corr}})$ $=\dfrac{\textnormal{Cov}_{Q_{X}}[X,\sigma^{2}(X)]}{\sqrt{\textnormal{Var}_{Q_{X}}(X)\textnormal{Var}_{Q_{O}^{0}}(O)}}-\dfrac{\textnormal{Corr}_{Q_{O,X}^{0}}[X,O]}{2\textnormal{Var}_{Q_{O}^{0}}(O)}\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{Var}}),$
where $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{Var}})$ is the SLOPE for the variance of $O$ (Section C.6.1). We also note that the relationship between SLOPE and IF (Theorem 3) can be easily verified, where the IF is
$\textnormal{IF}\left(O,X,\psi^{\textnormal{Corr}}(Q_{O,X}^{0})\right)=$ $\dfrac{X-\textnormal{E}_{Q_{X}}(X)}{\sqrt{\textnormal{Var}_{Q_{X}}(X)}}\cdot\dfrac{O-\textnormal{E}_{Q_{O}^{0}}(O)}{\sqrt{\textnormal{Var}_{Q_{O}^{0}}(O)}}$
$-\dfrac{\textnormal{Corr}_{Q_{O,X}^{0}}(X,O)}{2}\left\{\dfrac{\{X-\textnormal{E}_{Q_{X}}(X)\}^{2}}{\textnormal{Var}_{Q_{X}}(X)}+\dfrac{\{O-\textnormal{E}_{Q_{O}^{0}}(O)\}^{2}}{\textnormal{Var}_{Q_{O}^{0}}(O)}\right\}$
by Devlin et al., (1975).
C.8 SLOPE for L-Estimands
C.8.1 L-Estimands and Their SLOPEs
We discuss one important class of estimands whose sample counterparts correspond to L-estimates (Huber,, 1981, Section 3.3). Specifically, consider a one-dimensional functional $\psi(Q_{O,X})=\psi(Q_{O})$ defined on the marginal distribution of $O$:
$\psi(Q_{O})$ $=\int_{0}^{1}h(F_{Q_{O}}^{-1}(p))l(p)dp.$ (27)
Choosing a particular function $h(\cdot)$ and a density function $l$, both defined over the support $(0,1)$, determines a location parameter $\psi(Q_{O})$. For example, when $l(p)=1$, equation (27) reduces to the mean of $O$ in the target population, i.e., $\psi^{{\textnormal{mean}}}(Q_{O})$. When $l(p)=\mathds{1}_{\frac{1}{2}}(p)$ where $\mathds{1}_{C}(p)$ for a set $C$ is the indicator function such that $\mathds{1}_{C}(p)=1$ if $s\in C$ and $\mathds{1}_{C}(p)=0$ otherwise, equation (27) becomes the marginal median, $\psi^{{\textnormal{med}}}(Q_{O})$. When $l(p)=\mathds{1}_{[\alpha,1-\alpha]}(p)/(1-2\alpha)$, then the target functional becomes the marginal trimmed mean, $\psi^{{\alpha\textnormal{-trim}}}$.
Theorem 11 derives the SLOPE for estimands defined in (27).
Theorem 11 (SLOPE for L-Estimands).
Suppose Condition 2 holds with $\psi$ in (27). Then SLOPE of $\psi$ defined in (27) is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$
$=$ $\int_{0}^{1}\dfrac{-\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(p))\right]+\textnormal{E}_{Q_{O,X}^{0}}\left[\mu(X)\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(p))\right]}{f_{Q_{O}^{0}}\left(F_{Q_{O}^{0}}^{-1}(p)\right)}h^{\prime}\left(F_{Q_{O}^{0}}^{-1}(p)\right)l(p)dp,$
where $h^{\prime}(\cdot)$ is the derivative of $h(\cdot)$.
C.8.2 Lemma 2
Lemma 2 (SLOPE of Median in Two-Component Gaussian Mixtures).
Suppose $X\in\{x_{1},x_{2}\}$ and $q_{1}=Q_{X}(x_{1})$ and $q_{2}=Q_{X}(x_{2})$, $P_{O\mid X=x_{1}}\sim N(\mu_{1},\sigma_{1}^{2})$ and $P_{O\mid X=x_{2}}\sim N(\mu_{2},\sigma_{2}^{2})$ with $\sigma_{2}>\sigma_{1}$ fixed. We denote the SLOPE for the median as a function of $\Delta=\mu_{2}-\mu_{1}$: $\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]$. Then the followings hold.
(i) If $0<q_{1}<q_{2}$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]$ is increasing with $\Delta$.
(ii) If $q_{1}>q_{2}>0$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]$ is decreasing with $\Delta$.
(iii) If $q_{1}=q_{2}=1/2$, then $\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]=\sigma_{1}\sigma_{2}$ does not depend on $\Delta$.
Proof of Lemma 2. We start with parts (i) and (ii). Without loss of generality we suppose $\mu_{1}=0$ and $\sigma_{1}=1$ and therefore $\mu_{2}=\Delta$ and $\sigma_{2}>1$.
The SLOPE of the median can be expressed as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi{\textnormal{med}})[\Delta]=w_{1}(\Delta)(\sigma_{1}^{2}-\sigma_{2}^{2})+\sigma_{2}^{2},$
The weight is
$w_{1}(\Delta)$ $=\dfrac{q_{1}f_{P_{Y\mid X=x_{1}}}(m_{1/2})}{q_{1}f_{Y\mid X=x_{1}}(m_{1/2})+q_{2}f_{P_{Y\mid X=x_{2}}}(m_{1/2})}$
$=\dfrac{1}{1+\dfrac{q_{2}}{q_{1}\sigma_{2}}\dfrac{\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)}{\varphi(m_{1/2}(\Delta))}}$
$=\dfrac{1}{1+\dfrac{q_{2}}{q_{1}\sigma_{2}}\exp\left\{\frac{1}{2}\left[\{m_{1/2}(\Delta)\}^{2}-\{m_{1/2}(\Delta)-\Delta\}^{2}/\sigma_{2}^{2}\right]\right\}}$
$=\dfrac{1}{1+\dfrac{q_{2}}{q_{1}\sigma_{2}}\exp\{h(\Delta)\}}$
where $m_{1/2}=m_{1/2}(\Delta)$ is the marginal median that satisfies
$\int_{-\infty}^{m_{1/2}(\Delta)}\left[q_{1}\varphi(y)+q_{2}\varphi\left(\frac{y-\Delta}{\sigma_{2}}\right)/\sigma_{2}\right]dy=1/2.$ (28)
and $h(\Delta)=\dfrac{1}{2}\left[\{m_{1/2}(\Delta)\}^{2}-\{m_{1/2}(\Delta)-\Delta\}^{2}/\sigma_{2}^{2}\right]$. Since $\dfrac{q_{2}}{q_{1}\sigma_{2}}>0$ and the function $\exp(\cdot)$ is increasing, to prove $w_{1}(\Delta)$ is monotonically decreasing (resp. increasing) with $\Delta$, it’s sufficient to prove $h(\Delta)$ is monotonically increasing (resp. decreasing) with $\Delta$.
We start with case (i) when $0<q_{1}<q_{2}$. From (28) we know that the derivative of $m(\Delta)$ is
$m_{1/2}^{\prime}(\Delta)=\dfrac{\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)}{q_{1}\varphi(m_{1/2}(\Delta))+\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)}$
and it satisfies $m_{1/2}^{\prime}(\Delta)\in(0,1)$. Then the derivative of $h(\Delta)$ is
$h^{\prime}(\Delta)=$ $m_{1/2}(\Delta)\cdot m_{1/2}^{\prime}(\Delta)-\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}^{2}}\left[m_{1/2}^{\prime}(\Delta)-1\right]$
$=$ $\left[q_{1}\varphi(m_{1/2}(\Delta))+\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)\right]\cdot$
$\left[m_{1/2}(\Delta)\dfrac{q_{2}}{\sigma_{2}}\varphi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)+\dfrac{1}{\sigma_{2}^{2}}\left\{m_{1/2}(\Delta)-\Delta\right\}q_{1}\varphi\left(m_{1/2}(\Delta)\right)\right]\cdot$
$=$ $\underbrace{\dfrac{1}{\sigma_{2}}\left[q_{1}\varphi(m_{1/2}(\Delta))+\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)\right]}_{>0}\cdot$
$\left[q_{2}m_{1/2}(\Delta)\varphi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)-\dfrac{q_{1}}{\sigma_{2}}\left\{\Delta-m_{1/2}(\Delta)\right\}\varphi\left(m_{1/2}(\Delta)\right)\right].$
Since the first term $\dfrac{1}{\sigma_{2}}\left[q_{1}\varphi(m_{1/2}(\Delta))+\dfrac{q_{2}}{\sigma_{2}}\varphi\left((m_{1/2}(\Delta)-\Delta)/\sigma_{2}\right)\right]>0$, the sign of $h^{\prime}(\Delta)$ is determined by the sign of the second term. Hence, we have that
$h^{\prime}(\Delta)>0$ $\Longleftrightarrow q_{2}m_{1/2}(\Delta)\varphi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)-\dfrac{q_{1}}{\sigma_{2}}\left\{\Delta-m_{1/2}(\Delta)\right\}\varphi\left(m_{1/2}(\Delta)\right)>0$
$\Longleftrightarrow\dfrac{q_{2}m_{1/2}(\Delta)\varphi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)}{\dfrac{q_{1}}{\sigma_{2}}\left\{\Delta-m_{1/2}(\Delta)\right\}\varphi\left(m_{1/2}(\Delta)\right)}>1$
$\Longleftrightarrow\dfrac{q_{2}}{q_{1}}\cdot\dfrac{\varphi\left(\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}\right)/\left\{\Delta-m_{1/2}(\Delta)\right\}}{\varphi\left(m_{1/2}(\Delta)\right)/m_{1/2}(\Delta)}>1,$
where the second line follows from the fact that both the numerator and the denominator are positive, and the third line follows by re-organizing terms. In order to show $h^{\prime}(\Delta)>0$, it’s sufficient to show $m_{1/2}(\Delta)>\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}$, because
$h^{\prime}(\Delta)>0$ $\Longleftarrow\dfrac{\varphi\left(\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}\right)/\left\{\Delta-m_{1/2}(\Delta)\right\}}{\varphi\left(m_{1/2}(\Delta)\right)/m_{1/2}(\Delta)}>1$
$\Longleftarrow m_{1/2}(\Delta)>\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}},$
where the first line follows from the fact that $q_{2}/q_{1}>1$ and the second line follows since the function $\varphi(x)/x=\exp(-x^{2}/2)/x$ is decreasing when $x>0$.
Therefore, we are left to show $m_{1/2}(\Delta)>\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}$. Recall the definition of the median in (2), we have
$q_{1}\Phi\left(m_{1/2}(\Delta)\right)+q_{2}\Phi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)=1/2$
$\Longrightarrow$ $q_{1}\left[\Phi\left(m_{1/2}(\Delta)\right)-1/2\right]+q_{2}\left[\Phi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)-1/2\right]=0$
$\Longrightarrow$ $q_{1}\left[\Phi\left(m_{1/2}(\Delta)\right)-1/2\right]=-q_{2}\left[\Phi\left(\dfrac{m_{1/2}(\Delta)-\Delta}{\sigma_{2}}\right)-1/2\right]=q_{2}\left[\Phi\left(\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}\right)-1/2\right]$
$\Longrightarrow$ $\dfrac{\Phi\left(m_{1/2}(\Delta)\right)-1/2}{\Phi\left(\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}\right)-1/2}=\dfrac{q_{2}}{q_{1}}>1$
$\Longrightarrow$ $m_{1/2}(\Delta)>\dfrac{\Delta-m_{1/2}(\Delta)}{\sigma_{2}}.$
From the above, we have $h^{\prime}(\Delta)>0$ and in hence $w_{1}^{\prime}(\Delta)<0$. The SLOPE $w_{1}(\Delta)(\sigma_{1}^{2}-\sigma_{2}^{2})+\sigma_{2}^{2}$ is therefore increasing with $\Delta$. We have proven part (i).
The proof of part (ii) follows with a similar argument. When $q_{2}<q_{1}$, we have $m_{1/2}(\Delta)<\left[\delta-m_{1/2}(\Delta)\right]/\sigma_{2}$, and therefore $h^{\prime}(\Delta)<0$. This in turn gives $w_{1}^{\prime}>0$ and the SLOPE is decreasing with $\Delta$.
Part (iii) follows by noticing the median is $m_{1/2}=(\mu_{1}\sigma_{2}+\mu_{2}\sigma_{1})/(\sigma_{1}+\sigma_{2})$.
$\square$
C.9 SLOPE for Z-Estimands
For a Z-estimand defined through (2) in the main text, the corresponding SLOPE is presented in the Corollary 2 below, which is an immediate result of Theorem 3.
Corollary 2 (SLOPE for Z-Estimands).
Under Condition 3, the SLOPE for a Z-estimand is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=-\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi(Q_{O,X}^{0}))\}^{-1}\textnormal{E}_{Q_{O,X}^{0}}\left[s\left(O,X,\psi(Q_{O,X}^{0})\right)\{O-\mu(X)\}\right].$ (29)
Appendix D SLOPE for Functionals of Potential Outcomes
In this section, we discuss how to use SLOPE to quantify the sensitivity of the conditional exchangeability assumption in generalizing parameters in causal inference. We provide a brief review on notation in causal inference in Section D.1 and discuss the SLOPE in Section D.2.
D.1 Notation for Causal Inference
Let $Y$ be the observed outcome variable, $A\in\{0,1\}$ denote the observed treatment, and $X$ denote pre-treatment covariates. Let $Y(a)$ denote the potential outcome if the study unit, contrary to fact, was assigned treatment value $a\in\{0,1\}$. Throughout the paper, we assume that we observe only one of the potential outcomes based on the observed treatment assignment, i.e., $Y=Y(A)$ almost surely (Rubin,, 1980). A central goal is in causal inference to identify some functional of $Y(a)$, for example, the mean of $Y(a)$, for $a\in\{0,1\}$.
D.2 Transporting Functionals of a Potential Outcome
Suppose $P$ is the source population where researchers have completed a randomized experiment to study the effect of treatment $A\in\{0,1\}$ on an outcome $Y$. We adopt notation in Section D.1 and let $Y(a)$ represent the potential outcome under treatment $a$. The goal is to estimate the “average” of the potential outcome $Y(a)$ on the target country, where only the marginal $Q_{X}$ can be identified. In this problem, $O=Y(a)$ is the potential outcome under the treated. To identify its average on the target distribution, in the literature it’s common to impose the overlap condition between $Q_{X}$ and $P_{X}$ and the conditional exchangeability condition concerns the counterfactual outcome between the source and target populations, i.e., Assumptions 1 and 2 with $O=Y(a)$. Specifically, Assumption 2 becomes: $Q_{Y(a),X}(\cdot\mid X)$ is absolute continuous with respect to $P_{Y(a)\mid X}(\cdot\mid X)$ and the Radon-Nikodym derivative satisfies
$dQ_{Y(a),X}(O,X)/dP_{Y(a)\mid X}(O,X)=1\text{ almost everywhere }P_{Y(a)\mid X}\times Q_{X}$ (30)
Accordingly, the sensitivity model (3) becomes
$\dfrac{f_{Q_{Y(a)\mid X}^{\gamma}}(O,X)}{f_{P_{Y(a)\mid X}}(O,X)}\propto\exp(\gamma\cdot O)\text{, almost surely }P_{Y(a)\mid X}\times Q_{X}.$ (31)
Following Theorems 1 and 2, the SLOPEs for the mean potential outcome and median potential outcome are
$\textnormal{SLOPE}(Q_{Y(a),X}^{0},\psi^{{\textnormal{mean}}})=$ $\textnormal{E}_{Q_{X}}\left\{\textnormal{E}_{P_{Y(a)\mid X}}\left[\left\{Y(1)-\textnormal{E}_{P_{Y(a)\mid X}}[Y(a)\mid X]\right\}^{2}\Bigg|X\right]\right\},\text{ and}$ (32)
$\textnormal{SLOPE}(Q_{Y(a),X}^{0},\psi^{{\textnormal{med}}})=$ $\dfrac{\textnormal{E}_{Q_{X}}\{F_{P_{Y(a)\mid X}}(m_{1/2})\mu(X)\}-\textnormal{E}_{Q_{Y(a)}^{0}}\{Y(a)\mathds{1}(Y(a)\leq m_{1/2})\}}{f_{Q_{Y(a)}^{0}}(m_{1/2})},$ (33)
respectively, where we recall that $m_{1/2}=F_{Q_{Y(a)}^{0}}(1/2)$ is the marginal median under conditional exchangeability and $\mu(X)=\textnormal{E}_{P_{Y(a)\mid X}}\{Y(a)\mid X\}$ is the conditional mean function.
We make some remarks on the SLOPEs in (32) and (33) compared with their counterparts for an observed outcome in Theorems 1 and 2. First, the SLOPE for functionals of a potential outcome is no different from in form from the SLOPE for functionals of an observed outcome, except for a change in notation. This is because the identification strategy in transporting/generalizing a causal quantity is based on the couterfactual outcome $Y(a)$ (i.e., conditional exchangeability in (30), and so as its violation (i.e., the sensitivity model (31)). In this perspective, the SLOPE for a potential outcome should exactly mimic the SLOPE for an observed outcome, as we have seen above.
Second, although the forms look identical, we remind the readers that working with potential outcomes requires more discretion because their SLOPEs (e.g., (32) and (33)) usually involve the counterfactual distribution $P_{Y(a)\mid X}$ which is un-identified without additional causal assumptions. Fortunately, if the randomization procedure in the source population is known, then the identification is straightforward with additional justifiable assumptions. We discuss these assumption in the followings.
Suppose the randomization is known to depend on $X$ and another variable $V$ which are both measured in the source population, then by design, the strong ignorability assumption (Assumption 3) holds on the source population, which in turn provides equivalence between $P_{Y(a)\mid X,V}$ and $P_{Y(a)\mid X,V,A=1}$. Further, under SUTVA (Assumption 4), $P_{Y(a)\mid X,V,A=1}$ equals to $P_{Y\mid X,V,A=1}$. Consequently, one can replace $P_{Y(1)\mid X}$ in the SLOPE with $\int P_{Y\mid X,V,A=1}dP_{V\mid X,A=1}$, which no longer involves counterfactual quantities and thus can be identified. Under these identification assumptions, the SLOPEs can be expressed in observed quantities; see (53) and (54) in Section F.1. In the special case when the randomization solely depends on $X$ (i.e., all confounders can been observed in both the source and the target), we have $P_{Y(1)\mid X}=P_{Y(1)\mid X,A=1}$.
Assumption 3 (Strong Ignorability on $P$; Rosenbaum and Rubin, 1983b ).
$P_{X,V}(\cdot)$ is absolute continuous with respect to $P_{X,V\mid A=a}(\cdot)$ and $P_{Y(a)\mid X,V}(\cdot\mid x,v)=P_{Y(a)\mid X,V,A=a}(\cdot\mid x,v)$ almost everywhere $P_{X,V}$.
Assumption 4 (Stable Unit Treatment Variable Assumption (SUTVA)).
$Y=Y(a)$ if $A=a$ on the source population $P$.
Appendix E Supplementary Materials for Estimation
we provide supplementary results and some deferred discussions for the estimation of SLOPE presented in Section 4.2 of the main text. We start with defining some notations in Section E.1. Then we discuss regularity conditions and asymptotic properties of the weighting estimator and the regression estimator in Sections E.2 and E.3, respectively. Next, we detail estimators of SLOPEs for the mean, OLS coefficients, and the median in Sections E.4 to E.6, including the weighting and regression estimators presented in the main text and estimators based on the efficient influence function of the SLOPE (if exists). Then we discuss ways of estimating the nuisance function $\omega(X)$ in Section E.7. Finally, we present a general statement of the efficient influence function for the SLOPE of scalar valued Z-estimands in Section E.8.
E.1 Notation and Setup
As stated in the main, suppose we have i.i.d. samples from the target population with size $n_{q}$ and i.i.d. samples from the source population with size $n_{p}$. We pool the samples together and denote by $T_{i}=1$ if the sample comes from the target distribution $Q$ and $T_{i}=0$ otherwise. Therefore, we have a random sample $\{(X_{i},\textnormal{NA},T_{i}=1),i=1,\cdots,n_{q}\}\cup\{(X_{i},O_{i},T_{i}=0),i=n_{q}+1,\cdots,n\}$ from the pseudo population that combines $Q$ and $P$, where NA means unobserved. We will establish the asymptotic properties for the weighting and regression estimators by re-expressing these estimators as solutions to estimating equations, and then apply standard M-estimation theory (Van der Vaart,, 2000).
After introducing the pseudo population that combines $P$ and $Q$, in this section, we use $\textnormal{pr}(\cdot)$ and $\textnormal{E}(\cdot)$, $f(\cdot)$ without subscripts to denote probabilities, expectations, and densities on this pseudo population, and similarly for conditional quantities. For quantities on a single population (e.g., expectations on $P$ or $Q$ only), there are two equivalent notations. One is to use subscripts, e.g., $\textnormal{E}_{P_{O\mid X}}(O\mid X)$ and $f_{Q_{X}}(X)$, as adopted in the main text. The other is to conditioning on $T$, e.g., $\textnormal{E}(O\mid X,T=0)$ and $f(X\mid T=1)$, We prefer the first approach in order to be consistent with our convention of notation in the main text, while in this section we use the second approach if parts of Section E.7 when the context is clearer.
E.2 Asymptotic Properties for the Weighting Estimator
To start with, we express the SLOPE in equation (11) as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=-\eta_{2}^{-1}\eta_{1},\text{ where}$
$\eta_{2}=\left\{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\dot{s}(O,X,\psi)\mid X\right]\right)\right\},\text{ and }\eta_{1}=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}s\left(O,X,\psi\right)\mid X\right]\right).$
Similarly, we re-express the weighting estimator by the estimates of $\eta_{1}$ and $\eta_{2}$ as follows,
[…]
$\widehat{\eta}_{1}^{\textnormal{W}}$ $=\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{O_{i}-\widehat{\mu}(X_{i})\}s(O_{i},X_{i},\widehat{\psi}),\text{ and}$
[…]
In order to establish the consistency and asymptotic normality of the weighting estimator, we re-express $\widehat{\eta}_{1}^{\textnormal{W}}$ and $\widehat{\eta}_{2}^{\textnormal{W}}$ as solutions to some estimating equations and then apply standard M-estimation theory. Specifically, consider a vector of parameters $\eta^{\textnormal{W}}=[\eta_{1},\eta_{2},\eta_{3}^{\intercal},\eta_{4}^{\intercal},\eta_{5}^{\intercal}]^{\intercal}$, and suppose these parameters are estimated by solving estimating equations $G^{\textnormal{W}}=\left[\left(g_{1}^{\textnormal{W}}\right)^{\intercal},\left(g_{2}^{\textnormal{W}}\right)^{\intercal},\left(g_{3}^{\textnormal{W}}\right)^{\intercal},\left(g_{4}^{\textnormal{W}}\right)^{\intercal},\left(g_{5}^{\textnormal{W}}\right)^{\intercal}\right]^{\intercal}$, i.e., by
$\sum_{i=1}^{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=0$ (34)
Specifically, elements in $\eta^{\textnormal{W}}$ and estimating equations $G^{\textnormal{W}}$ are defined as follows. The first two parameters, $\widehat{\eta}_{1}^{\textnormal{W}}$ and $\widehat{\eta}_{2}^{\textnormal{W}}$, have been defined above, and their corresponding estimating equations are
$g_{1}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})$ $=\dfrac{1-T_{i}}{\textnormal{pr}(T_{i}=0)}\omega(X_{i},\eta_{5})s(O_{i},X_{i},\eta_{3})\{O_{i}-\mu(X_{i},\eta_{4})\}-\eta_{1}=0,\text{and}$
$g_{2}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})$ $=\dfrac{1-T_{i}}{\textnormal{pr}(T_{i}=0)}\omega(X_{i},\eta_{5})\dot{s}(O_{i},X_{i},\eta_{3})-\eta_{2}=0.$
Next, $\eta_{3}=\psi(Q_{O,X}^{0})$ is defined as the target estimand under conditional exchangeability, and is estimated through
$g_{3}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=\dfrac{1-T_{i}}{\textnormal{pr}(T_{i}=0)}\omega(X_{i},\eta_{5})s(O_{i},X_{i},\eta_{3}).$
Finally, suppose parametric models of $\mu(x)$ and $\omega(x)$ are posited by the researcher with parameters $\eta_{4}$ and $\eta_{5}$, respectively, and let $g_{4}^{\textnormal{W}}$ and $g_{5}^{\textnormal{W}}$ be the corresponding estimating equations (e.g., score functions). Denote the two nuisance functions as $\mu(x,\eta_{5})$ and $\omega(x,\eta_{4})$ and hence their estimates are $\widehat{\mu}(x)=\mu(x,\widehat{\eta}_{5})$ and $\widehat{\omega}(x)=\omega(x,\widehat{\eta}_{4})$. Suppose Let $g_{4}^{\textnormal{W}}$ and $g_{5}^{\textnormal{W}}$ are estimating equations (e.g., score functions) that correspond to $\mu(x)$ and $\omega(x)$, respectively, via parametric models posited by the researcher. To indicate the dependencies on nuisance parameters, we denote these two functions as $\mu(x,\eta_{5})$ and $\omega(x,\eta_{4})$, respectively, and hence their estimates are $\widehat{\mu}(x)=\mu(x,\widehat{\eta}_{5})$ and $\widehat{\omega}(x)=\omega(x,\widehat{\eta}_{4})$.
With $G^{W}$ defined as above, the asymptotic properties of the weighting estimator for SLOPE in (12) can be established under standard regularity conditions in M-estimation theory (Newey and McFadden,, 1994; Van der Vaart,, 2000). Below, Condition 8 parallels the condition in Theorem 2.6 of Newey and McFadden, (1994) for consistency, and Condition 9 parallels assumptions in Theorem 3.4 of Newey and McFadden, (1994) and Theorem 5.31 of Van der Vaart, (2000) for asymptotic normality.
Condition 8 (Regularity Conditions for Consistency for Weighting Estimator).
(i) $\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\}=0$ implies $\eta=\eta^{\textnormal{W}}$. (ii) $\eta^{\textnormal{W}}\in\Theta$ where $\Theta$ is compact. (iii) $G^{\textnormal{W}}$ is continuous at each $\eta\in\Theta$ with probability one and $\textnormal{E}\left\{\sup_{\eta\in\Theta}\lVert G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\rVert\right\}<\infty$. (iv) $\eta_{2}=\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O_{i},X_{i},\psi(Q_{O,X}^{0})\right]\neq 0$ with probability one.
Condition 9 (Regularity Conditions for Asymptotic Normality for Weighting Estimator).
(i) $\eta^{\textnormal{W}}$ lies in the interior of $\Theta$. (ii) $\textnormal{E}\lVert G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})\rVert^{2}<\infty$. (iii) The function class $\left\{G^{\textnormal{W}}:\lVert\eta-\eta^{\textnormal{W}}\rVert<\delta\right\}$ is Donsker for some $\delta>0$ and $\textnormal{E}\lVert G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)-G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})\rVert^{2}\to 0$ as $\eta\to\eta^{\textnormal{W}}$. (iv) The map $\eta\mapsto\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\}$ is differentiable at $\eta^{\textnormal{W}}$ with a non-singular derivative matrix $\Omega^{\textnormal{W}}$ with inverse matrix $V^{\textnormal{W}}$.
With Conditions 8 and 9, we establish the asymptotic properties of the weighting estimator.
Theorem 12 (Weighting Estimator).
Let $\widehat{\textnormal{SLOPE}}^{\textnormal{W}}$ be the weighting estimator in (12) where $\widehat{\eta}_{j}$’s are estimated with (34) and we drop the notation in parentheses of the SLOPE for ease of communication. Suppose Condition 8 holds, then $\widehat{\textnormal{SLOPE}}^{\textnormal{W}}$ converges to SLOPE in probability. Additionally suppose Condition 9 holds, then $\sqrt{n}(\widehat{\textnormal{SLOPE}}^{\textnormal{W}}-\textnormal{SLOPE})$ converges in distribution to a normal distribution with mean zero and variance $(\eta_{2})^{2}V_{11}^{\textnormal{W}}/(\eta_{1})^{4}+V_{22}^{\textnormal{W}}/(\eta_{1})^{2}-2\eta_{2}V^{W}_{12}/(\eta_{1})^{3}$, where $V^{\textnormal{W}}$ is inverse of the derivative matrix of $\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})\}$ with respect to $\eta^{\textnormal{W}}$, with $V^{\textnormal{W}}_{ij}$ denoting its entry at the $i$-th row and $j$-th column.
We note that depending on the target estimand, some nuisance parameters listed above may be trivial. For example, for the SLOPE for the mean, $\eta_{2}=-1$ need not be estimated since $\dot{s}=-1$ is constant. In addition, we provide example estimators for other estimands later in this section and discuss ways of estimating $\omega(x)$ in Section E.
Finally, we remark on the weighting estimator for SLOPE when the target functional $\psi(\cdot)$ is vector valued.
Remark 4 (Weighting Estimator with a Vector Valued $\psi$).
With a vector valued target functional $\psi(\cdot)$, the target estimand and the SLOPE become vectors of the same dimension, say $p$. The SLOPE formula derived from the IF (i.e., (11)) still holds and the weighting estimator (12) is still applicable. The difference is that $\eta_{1}$ becomes a vector of length $p$ and $\eta_{2}$ becomes a $p$ by $p$ matrix. To establish the asymptotic properties, we need to adjust the vector of nuisance parameter, $\eta^{\textnormal{W}}$, mainly to include a vectorization of $\eta_{2}$ instead of $\eta_{2}$ itself. More concretely, let $\eta^{\textnormal{W}}=[\eta_{1}^{\intercal},\{\vec{(}\eta_{2})\}^{\intercal},\eta_{3}^{\intercal},\eta_{4}^{\intercal},\eta_{5}^{\intercal}]^{\intercal}$, where $\vec{(}\cdot)$ is vectorization of a matrix, and let $G^{\textnormal{W}}$ be modified such that $g_{2}^{\textnormal{W}}$ corresponds to $\vec{(}\eta_{2})$. Then under the same regularity conditions on the updated $G^{\textnormal{W}}$ and $\eta^{\textnormal{W}}$, the weighting estimator is still consistent and asymptotically normal. The asymptotic variance is $A\Sigma^{\textnormal{W}}A^{\intercal}$, where $\Sigma^{\textnormal{W}}$ is the first $(p^{2}+p)$ diagonal matrix of $V^{\textnormal{W}}$, and $A=\left[-(\textnormal{SLOPE})^{\intercal}\otimes(\eta_{2})^{-1},\eta_{2}^{-1}\right]$.
E.3 Asymptotic Properties for the Regression Estimator
We re-express the regression estimator (13) with $\widehat{\eta}_{1}^{\textnormal{R}}$ and $\widehat{\eta}^{\textnormal{R}}_{2}$, the regression typed estimators for $\eta_{1}$ and $\eta_{2}$ respectively, as follows:
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}(\widehat{Q}_{O,X}^{0},\widehat{\psi})=-\left(\eta_{2}^{\textnormal{R}}\right)^{-1}\eta_{1}^{\textnormal{R}},$ where
$\eta_{1}=\sum_{i=1}^{n_{q}}\widehat{\textnormal{E}}_{P_{O\mid X}}\left[\{O_{i}-\widehat{\mu}(X_{i})\}s(O_{i},X_{i},\psi)\mid X_{i}\right],$ $\text{ and }\eta_{2}=\left[\sum_{i=1}^{n_{q}}\widehat{\textnormal{E}}_{P_{O\mid X}}\{\dot{s}(O_{i},X_{i},\psi)\mid X_{i}\}\right].$
Next, we establish the asymptotic properties of the regression estimator in a similar way as done for the weighting estimator. Specifically, we consider nuisance parameters $\eta^{\textnormal{R}}=[\eta_{1},\eta_{2},\eta_{3},\eta_{4}^{\intercal},\eta_{6}^{\intercal},\eta_{7}^{\intercal},\eta_{8}^{\intercal}]^{\intercal}$ estimated through
$\sum_{i=1}^{n}G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})=0$ (35)
with estimating equations $G^{\textnormal{R}}=\left[g_{1}^{\textnormal{R}},g_{2}^{\textnormal{R}},g_{3}^{\textnormal{R}},\left(g_{4}^{\textnormal{R}}\right)^{\intercal},\left(g_{6}^{\textnormal{R}}\right)^{\intercal},\left(g_{7}^{\textnormal{R}}\right)^{\intercal},\left(g_{8}^{\textnormal{R}}\right)^{\intercal}\right]^{\intercal}$. Specifically, $\eta_{1}$ to $\eta_{4}$ have been defined previously with these new estimating equations based on regressions on the target samples instead of weighting on the source samples,
$g_{1}^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})$ $=\dfrac{T_{i}}{\textnormal{pr}(T_{i}=1)}\textnormal{E}_{P_{O\mid X}}\left[s(O_{i},X_{i},\eta_{3})\{O_{i}-\mu(X_{i})\}\mid X_{i},\eta_{6}\right]-\eta_{1},$
$g_{2}^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})$ $=\dfrac{T_{i}}{\textnormal{pr}(T_{i}=1)}\textnormal{E}_{P_{O\mid X}}\left\{\dot{s}(O_{i},X_{i},\eta_{3})\mid X_{i},\eta_{7}\right\}-\eta_{2},$
$g_{3}^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})$ $=\dfrac{T_{i}}{\textnormal{pr}(T_{i}=1)}\textnormal{E}_{P_{O\mid X}}\{s(O_{i},X_{i},\eta_{3})\mid X_{i},\eta_{8}\},$
and $g_{4}^{\textnormal{R}}=g_{4}^{\textnormal{W}}$ is the estimating equation for the regression function $\mu(x)$ which is now denoted as $\mu(x,\eta_{4})$ as in Section E.2. As denoted in the proceeding formulas, the additional nuisance parameters, $\eta_{6}$, $\eta_{7}$, and $\eta_{8}$ represent the nuisance parameters in parametric models for $\textnormal{E}_{P_{O\mid X}}\{s(O_{i},X_{i},\eta_{3})\{O_{i}-\mu(X_{i})\mid X\}$, $\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O_{i},X_{i},\eta_{3})\mid X\}$ and $\textnormal{E}_{P_{O\mid X}}\{s(O_{i},X_{i},\eta_{3})\mid X\}$, respectively, with estimating equations $g_{6}^{\textnormal{R}}$, $g_{7}^{\textnormal{R}}$ and $g_{8}^{\textnormal{R}}$. Depending on the specific estimand, some nuisance parameters may be trivial and the estimation is simpler than what has been shown. That being said, using regression based estimator typically involves more nuisance functions/parameters than the weighting estimator.
The regularity conditions for the regression estimator are standard and are similar to the conditions for the weighting estimator. Condition 10 parallels the condition in Theorem 2.6 of Newey and McFadden, (1994) for consistency and Condition 11 parallels assumptions in Theorem 3.4 of Newey and McFadden, (1994) and Theorem 5.31 of Van der Vaart, (2000) for asymptotic normality.
Condition 10 (Regularity Conditions for Consistency for Regression Estimator).
(i) $\textnormal{E}\{G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta)\}=0$ implies $\eta=\eta^{\textnormal{R}}$. (ii) $\eta^{\textnormal{R}}\in\Theta$ where $\Theta$ is compact. (iii) $G^{\textnormal{R}}$ is continuous at each $\eta\in\Theta$ with probability one and $\textnormal{E}\left\{\sup_{\eta\in\Theta}\lVert G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta)\rVert\right\}<\infty$. (iv) $\eta_{2}=\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O_{i},X_{i},\psi(Q_{O,X}^{0})\right]\neq 0$ with probability one.
Condition 11 (Regularity Conditions for Asymptotic Normality for Regression Estimator).
(i) $\eta^{\textnormal{R}}$ lies in the interior of $\Theta$. (ii) $\textnormal{E}\lVert G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})\rVert^{2}<\infty$. (iii) The function class $\left\{G^{\textnormal{R}}:\lVert\eta-\eta^{\textnormal{R}}\rVert<\delta\right\}$ is Donsker for some $\delta>0$ and $\textnormal{E}\lVert G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta)-G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})\rVert^{2}\to 0$ as $\eta\to\eta^{\textnormal{R}}$. (iv) The map $\eta\mapsto\textnormal{E}\{G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta)\}$ is differentiable at $\eta^{\textnormal{R}}$ with a non-singular derivative matrix $\Omega^{\textnormal{R}}$ with inverse matrix $V^{\textnormal{R}}$.
Under regularity conditions listed above, the consistency and asymptotic normality of the regression estimator is presented as follows.
Theorem 13 (Regression Estimator).
Let $\widehat{\textnormal{SLOPE}}^{\textnormal{R}}$ be the regression estimator in (13) where $\eta_{j}$’s are estimated with (35) and we drop the notation in parentheses of the SLOPE for ease of communication. Suppose Condition 10 in the Appendix holds, then $\widehat{\textnormal{SLOPE}}^{\textnormal{R}}$ converges to SLOPE in probability. Additionally suppose Condition 11 in the Appendix holds, then $\sqrt{n}(\widehat{\textnormal{SLOPE}}^{\textnormal{R}}-\textnormal{SLOPE})$ converges in distribution to a Gaussian distribution with mean zero and variance $(\eta_{2})^{2}V_{11}^{\textnormal{R}}/(\eta_{1})^{4}+V_{22}^{\textnormal{R}}/(\eta_{1})^{2}-2\eta_{2}V^{\textnormal{R}}_{12}/(\eta_{1})^{3}$, where $V^{\textnormal{R}}$ is the derivative matrix of $\textnormal{E}\{G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})\}$ with respect to $\eta^{\textnormal{R}}$, with $V^{\textnormal{R}}_{ij}$ denoting its entry at the $i$-th row and $j$-th column.
Remark 5 (Regression Estimator with a Vector Valued $\psi$).
When the target functional is vector valued, the regression estimator is still applicable with a similar argument as in Remark 4 for the weighting estimator.
E.4 Estimating SLOPE for the Mean
We elaborate on the estimators of the SLOPE for the mean, $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})$. In Section 4.2, we have presented two estimators, a weighting estimator and a regression estimator,
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})=$ $\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{O_{i}-\widehat{\mu}(X_{i})\}(O_{i}-\widehat{\psi}^{{\textnormal{mean}}}),$ (36)
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})=$ $\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i}).$ (37)
To implement the two estimators, one can resort to Section E.7 for estimating $\omega(x)$ and any regression method for estimating $\mu(x)$. In addition, $\widehat{\psi}^{{\textnormal{mean}}}$ can be obtained by a weighted average of outcomes over source samples, i.e., $\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})O_{i}/{n_{p}}$, and $\widehat{\sigma}^{2}(x)$ can be obtained by regressing the squared residual, $\{O_{i}-\widehat{\mu}(X_{i})\}^{2}$ over $X_{i}$.
Next we motivate the estimator based on the efficient influence function. By expressing the SLOPE explicitly,
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}^{2}\mid X\right]\right),$
an alternative weighting estimator is naturally motivated:
$\widehat{\textnormal{SLOPE}}^{\text{W,Alt}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})=$ $\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{O_{i}-\widehat{\mu}(X_{i})\}^{2}.$ (38)
Moreover, a fourth estimator is motivated by combining properties of the alternative weighting estimator (38) with the regression estimator (37), i.e.,
$\widehat{\textnormal{SLOPE}}^{\text{EIF}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})=$ $\widehat{\textnormal{SLOPE}}^{\text{W,Alt}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})-\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{\sigma}^{2}(X_{i})+\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i})$
$=$ $\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left[\{O_{i}-\widehat{\mu}(X_{i})\}^{2}-\widehat{\sigma}^{2}(X_{i})\right]+\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i}).$ (39)
Since this estimator (39) can be naturally motivated from the efficient influence function (EIF) of the SLOPE (see Section E.8 below), we refer to it as the EIF-based estimator and denote it using the superscript “EIF”. Under standard regularity conditions, the EIF-based estimator is consistent and asymptotically normal,
$\sqrt{n}\left\{\widehat{\textnormal{SLOPE}}^{\text{EIF}}(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{mean}}})-{\textnormal{SLOPE}}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})\right\}\to_{d}N\left(0,\textnormal{E}\{\textnormal{EIF}(T,O,X,\textnormal{SLOPE})\}\right),$
where $\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$ is the efficient influence function of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})$ and it takes the form of
$\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$ $=\dfrac{1-T}{\textnormal{pr}(T=0)}\omega(X)\{O-\mu(X)\}^{2}+\dfrac{T}{\textnormal{pr}(T=1)},$
where $\textnormal{pr}(T=1)$ is the probability limit of $n_{q}/n$ and $\textnormal{pr}(T=0)=1-\textnormal{pr}(T=1)$; see Section E.8 (in particular, Proposition 1) for a general formula of the EIF for the SLOPE. Therefore, the asymptotic variance can be estimated with
$\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\left\{\widehat{\sigma}^{2}(X_{i})-\widehat{\textnormal{SLOPE}}^{\text{EIF}}\right\}+\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left[\{O_{i}-\widehat{\mu}(X_{i})\}^{2}-\widehat{\sigma}^{2}(X_{i})\right],$ (40)
which is consistent under regularity conditions. Consequently, the standard error (SE) of the EIF-based estimator can be estimated with
$\dfrac{1}{\sqrt{n}}\sqrt{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\left\{\widehat{\sigma}^{2}(X_{i})-\widehat{\textnormal{SLOPE}}^{\text{EIF}}\right\}+\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left[\{O_{i}-\widehat{\mu}(X_{i})\}^{2}-\widehat{\sigma}^{2}(X_{i})\right]}.$
E.5 Estimating SLOPE for OLS Coefficients
Suppose $O$ is an outcome variable and $X$ is a vector of covariates that includes $1$ as the first component. We are interested in the OLS coefficient of regressing $O$ on $X$ in the target distribution, i.e., $\psi^{{\textnormal{OLS}}}(Q_{O,X})$ such that $\textnormal{E}_{Q_{O,X}}\left[XX^{\intercal}\psi^{{\textnormal{OLS}}}-XO\right]=0$. Then according to Theorem 8, the SLOPE for the OLS coefficient $\psi^{{\textnormal{OLS}}}$ is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$ $=\left\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\right\}^{-1}\textnormal{E}_{Q_{X}}\{X\sigma^{2}(X)\}.$
To estimate the SLOPE, we consider three estimators based on weighting, regression, and the efficient influence function:
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{OLS}}}\right)=\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{r}_{i}^{2}X_{i},$ (41)
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{OLS}}}\right)=\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}^{2}(X_{i})X_{i}$ (42)
$\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{OLS}}}\right)=\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\left[\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\{\widehat{r}_{i}^{2}-\widehat{\sigma}^{2}(X_{i})\}X_{i}+\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}\widehat{\sigma}^{2}(X_{i})\right],$ (43)
where $\widehat{r}_{i}=O_{i}-\widehat{\mu}(X_{i})$ is the regression residuals on the source data.
In addition, for the EIF-based estimator, the variance can be consistently estimated by
$\dfrac{1}{n}\sum_{i=1}^{n}\widehat{\textnormal{EIF}}\left(T_{i},O_{i},X_{i},\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}}\right)\left\{\widehat{\textnormal{EIF}}(T_{i},O_{i},X_{i},\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}})\right\}^{\intercal},$ (44)
where
$\widehat{\textnormal{EIF}}\left(T_{i},O_{i},X_{i},\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}}\right)=$ $\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\left[\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})X_{i}\{\widehat{r}_{i}^{2}-\widehat{\sigma}^{2}(X_{i})\}\right]$
$+\left(\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}X_{i}X_{i}^{\intercal}\right)^{-1}\left[\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\left\{-X_{i}^{\intercal}\widehat{\textnormal{SLOPE}}^{\textnormal{EIF}}+\widehat{\sigma}^{2}(X_{i})\right\}X_{i}\right].$
E.6 Estimating SLOPE for the Median
In this section, we consider estimating the SLOPE for the median. We start with a simpler case when $P_{O\mid X}$ is Gaussian, i.e., $P_{O\mid X}\sim N\left(\mu(X),\sigma^{2}(X)\right)$. Then by part (ii) of Theorem 2, the SLOPE for the median is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=\dfrac{\textnormal{E}_{Q_{X}}\left[\sigma(X)\varphi(\{m_{1/2}-\mu(X)\}/\sigma(X))\right]}{\textnormal{E}_{Q_{X}}\left[\varphi(\{m_{1/2}-\mu(X)\}/\sigma(X))/\sigma(X)\right]}.$
This motivates the following weighting and regression estimators:
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{\sigma}(X_{i})\varphi\left(\{\widehat{m}_{1/2}-\widehat{\mu}(X_{i})\}/\widehat{\sigma}(X_{i})\right)}{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\varphi\left(\{\widehat{m}_{1/2}-\widehat{\mu}(X_{i})\}/\widehat{\sigma}(X_{i})\right)/\widehat{\sigma}(X_{i})},$ (45)
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{\sigma}(X_{i})\varphi\left(\{\widehat{m}_{1/2}-\widehat{\mu}(X_{i})\}/\widehat{\sigma}(X_{i})\right)}{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\varphi\left(\{\widehat{m}_{1/2}-\widehat{\mu}(X_{i})\}/\widehat{\sigma}(X_{i})\right)/\widehat{\sigma}(X_{i})},$ (46)
where $\varphi(\cdot)$ is the density of the standard normal distribution.
Next, we consider the general case with SLOPE given by (7). Since it involves conditional densities, the efficient influence function does not exist in general. We will consider the weighting estimator and the regression estimator only. Let $\widehat{F}_{P_{O\mid X}}$, $\widehat{f}_{P_{O\mid X}}$ and $\widehat{\textnormal{E}}_{P_{O\mid X}}\left\{O\mathds{1}(O\leq\widehat{m}_{1/2})\right\}$ be estimates of the c.d.f. $F_{P_{O\mid X}}$, the p.d.f. $f_{P_{O\mid X}}$, and the truncated mean $\textnormal{E}_{P_{O\mid X}}\left\{O\mathds{1}(O\leq\widehat{m}_{1/2})\right\}$, respectively, and $\widehat{m}_{1/2}$ be an estimate of $m_{1/2}$ where all estimates are based on parametric models. Then the weighting and regression estimators are
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left\{\widehat{F}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})\widehat{\mu}(X_{i})-O_{i}\mathds{1}(O_{i}\leq\widehat{m}_{1/2})\right\}}{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{f}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})},\text{ and}$
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{F}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})\widehat{\mu}(X_{i})-\widehat{\textnormal{E}}_{P_{O\mid X}}\left\{O_{i}\mathds{1}(O_{i}\leq\widehat{m}_{1/2})\mid X_{i}\right\}}{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{f}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})},$
respectively. In practice, one may impose parametric assumptions on $P_{O\mid X}$ and estimate $F_{P_{O\mid X}}$, $f_{P_{O\mid X}}$ and $\textnormal{E}_{P_{O\mid X}}\left\{O\mathds{1}(O\leq\widehat{m}_{1/2})\right\}$ accordingly. Then the marginal $m_{1/2}$ can be estimated by numerically solving either of the following equations using bisection,
$\sum_{i=1}^{n_{q}}\widehat{F}_{P_{O\mid X}}(m_{1/2}\mid X_{i})$ $=1/2,\text{ or}$
$\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{F}_{P_{O\mid X}}(m_{1/2}\mid X_{i})$ $=1/2.$
E.7 Example Estimating Equations for $\omega(x)$
In this section, we discuss some common approaches to estimating the density ratio $\omega(x)$, including the logistic regression, entropy balancing, and a method based on empirical distributions when the support of $Q_{X}$ (i.e., $\mathcal{S}_{X}$) is finite and discrete. Specifically, following the notation in Section E.2, we build estimating equations $g_{5}^{W}$ for the nuisance parameter $\eta_{5}$ and after obtaining the estimate $\widehat{\eta}_{5}$, we present $\widehat{\omega}(x)$ in terms of $\widehat{\eta}_{5}$. For the sake of exposition, we take the convention that $X$ does not include one (i.e., a constant that can serve as an intercept in regression models).
E.7.1 Logistic Regression
Let $\textnormal{pr}(T=1\mid x)=\textnormal{pr}(T=1\mid X=x)$ be the probability of being included in the target sample. Since the density ratio $\omega(x)$ can be re-formulated in terms of $\textnormal{pr}(T=1\mid x)$ via Bayes rule,
$\omega(x)$ $=\dfrac{f(x\mid T=1)}{f(x\mid T=0)}$ (By definition)
$=\dfrac{\textnormal{pr}(T=1\mid x)}{\textnormal{pr}(T=0\mid x)}\dfrac{\textnormal{pr}(T=0)}{\textnormal{pr}(T=1)}$ $\text{ (By Bayes rule)},$
one common strategy to estimate $\omega(x)$ is to first estimate the conditional probability function $\textnormal{pr}(T=1\mid x)$ using some binary classification/regression model, and then plug in the estimates $\widehat{\textnormal{pr}}(T=1\mid x)$ to obtain $\widehat{\omega}(x)$,
$\widehat{\omega}(x)=\dfrac{\widehat{\textnormal{pr}}(T=1\mid x)}{\widehat{\textnormal{pr}}(T=0\mid x)}\dfrac{\widehat{\textnormal{pr}}(T=0)}{\widehat{\textnormal{pr}}(T=1)}=\dfrac{\widehat{\textnormal{pr}}(T=1\mid x)}{\widehat{\textnormal{pr}}(T=0\mid x)}\dfrac{n_{p}}{n_{q}},$ (47)
where $\widehat{\textnormal{pr}}(T=0)=n_{p}/n$ and $\widehat{\textnormal{pr}}(T=1)=n_{q}/n$.
Next, we list the estimation equations based on maximum likelihood estimation when $\textnormal{pr}(T=1\mid x)$ is modeled by logistic regression, one of the most popular binary regression models. This can be implemented in R using the built-in function glm. To demonstrate the estimating equation for the nuisance parameter $\eta_{5}$, we let $\eta_{5}=[\alpha_{\textnormal{LR}},\beta_{\textnormal{LR}}^{\intercal}]^{\intercal}$ where $\alpha_{\textnormal{LR}}$ and $\beta_{\textnormal{LR}}$ are the intercept and slope coefficients in the logistic regression model. Then the likelihood can be expressed as
$l_{\textnormal{LR}}(\alpha_{\textnormal{LR}},\beta_{\textnormal{LR}})$ $=\sum_{T_{i}=0}(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X_{i})-\sum_{i=1}^{n}\log\left(1+\exp\{\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X_{i}\}\right),$
By setting $\partial l_{\textnormal{LR}}/\partial\alpha_{\textnormal{LR}}=0$ and $\partial l_{\textnormal{LR}}/\partial\beta_{\textnormal{LR}}=0$ and checking the second-order conditions, parameters $\alpha_{\textnormal{LR}}$ and $\beta_{\textnormal{LR}}$ can be estimated by $\widehat{\alpha}_{\textnormal{LR}}$ and $\widehat{\beta}_{\textnormal{LR}}$ which are solutions to the following equations,
$n_{p}-\sum_{i=1}^{n}\dfrac{\exp(\widehat{\alpha}_{\textnormal{LR}}+\widehat{\beta}_{\textnormal{LR}}^{\intercal}X_{i})}{1+\exp(\widehat{\alpha}_{\textnormal{LR}}+\widehat{\beta}_{\textnormal{LR}}^{\intercal}X_{i})}=0,\text{ and}\sum_{T_{i}=0}X_{i}-\sum_{i=1}^{n}\dfrac{X_{i}\exp(\widehat{\alpha}_{\textnormal{LR}}+\widehat{\beta}_{\textnormal{LR}}^{\intercal}X_{i})}{1+\exp(\widehat{\alpha}_{\textnormal{LR}}+\widehat{\beta}_{\textnormal{LR}}^{\intercal}X_{i})}.$
Therefore, the estimating equation $g_{5}^{\textnormal{W}}$ is
$g_{5}^{\textnormal{W}}(T,O,X,\eta^{\textnormal{W}})$ $=\begin{bmatrix}\mathds{1}(T=0)-\dfrac{\exp(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X)}{1+\exp(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X)}\\ {}\\ \mathds{1}(T=0)X-\dfrac{X\exp(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X)}{1+\exp(\alpha_{\textnormal{LR}}+\beta_{\textnormal{LR}}^{\intercal}X)}\end{bmatrix}$,
and the estimate $\widehat{\eta}_{5}=[\widehat{\alpha}_{\textnormal{LR}},\widehat{\beta}_{\textnormal{LR}}^{\intercal}]^{\intercal}$ is obtained by setting $\sum_{i=1}^{n}g_{5}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=0$. After estimating these parameters, by (47), the resulting estimate for the density ratio is
$\widehat{\omega}(x)$ $=\exp\left\{-\left[\widehat{\alpha}_{\textnormal{LR}}-\log(n_{p}/n_{q})\right]-\widehat{\beta}_{\textnormal{LR}}^{\intercal}x\right\}.$
E.7.2 Entropy Balancing
Note that for any measurable function $h(x)$, the density ratio $\omega(x)$ satisfies that
$\textnormal{E}\{\omega(X)h(X)\mid T=0\}=\textnormal{E}\{h(X)\mid T=1\}.$
In other words, $\omega(x)$ reweighs the source population in order to match the target population. This property motivates estimating $\omega(x)$ by balancing functions (usually moments) of the source samples so that they match the target samples.
In this section, we introduce the method of entropy balancing and aim at balancing the first moments of $X$. Specifically, for source samples with $i=1+n_{q},\cdots,n$, suppose $\omega_{i}=\omega(X_{i})$ are the weights. Entropy balancing seeks for $\omega_{i}$’s that maximizes their entropy as well as balances the first moment of $X$:
$\begin{split}&\underset{w_{i}}{\rm argmin}\sum_{T_{i}=0}w_{i}\log(w_{i}),\quad{}\textnormal{s.t. }\dfrac{1}{n_{p}}\sum_{T_{i}=0}w_{i}X_{i}=\dfrac{1}{n_{q}}\sum_{T_{i}=1}X_{i}.\end{split}$ (48)
According to Lee et al., (2023); Chen et al., 2023a , solutions to (48), denoted as $\widehat{\omega}_{i}$, can be expressed by
$\widehat{\omega}_{i}=\widehat{\omega}(X_{i})=\exp\left(-\widehat{\alpha}_{\textnormal{EB}}-\widehat{\beta}_{\textnormal{EB}}^{\intercal}X_{i}\right),$ (49)
where $\widehat{\alpha}_{\textnormal{EB}}$ and $\widehat{\beta}_{\textnormal{EB}}$ satisfy
$\sum_{T_{i}=0}\exp(\widehat{\alpha}_{\textnormal{EB}}-\widehat{\beta}_{\textnormal{EB}}^{\intercal}X_{i})=n_{p},\quad\dfrac{1}{n_{p}}\sum_{T_{i}=0}\exp\left(\widehat{\alpha}_{\textnormal{EB}}-\widehat{\beta}_{\textnormal{EB}}^{\intercal}X_{i}\right)=\dfrac{1}{n_{q}}\sum_{T_{i}=1}X_{i}.$
Following the notation in the main text, let the nuisance parameter be $\eta_{5}=[\alpha_{\textnormal{EB}},\beta_{\textnormal{EB}}^{\intercal}]^{\intercal}$, then the estimating equation $g_{5}^{\textnormal{W}}$ is
$g_{5}^{\textnormal{W}}(T,O,X,\eta^{\textnormal{W}})=\begin{bmatrix}\mathds{1}(T=0)\exp(-\alpha_{\textnormal{EB}}-\beta_{\textnormal{EB}}^{\intercal}X)-\mathds{1}(T=0)\\ \mathds{1}(T=0)X_{i}\exp(-\alpha_{\textnormal{EB}}-\beta_{\textnormal{EB}}^{\intercal}X)-\dfrac{n_{p}}{n_{q}}\mathds{1}(T=1)X\end{bmatrix}$,
and the estimate $\widehat{\eta}_{5}=[\widehat{\alpha}_{\textnormal{EB}},\widehat{\beta}_{\textnormal{EB}}^{\intercal}]^{\intercal}$ is obtained by setting $\sum_{i=1}^{n}g_{5}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=0$. After estimating these parameters, by (49), the resulting estimate for the density ratio is
$\widehat{\omega}(x)$ $=\exp\left(\widehat{\alpha}_{\textnormal{EB}}-\widehat{\beta}_{\textnormal{EB}}^{\intercal}x\right).$
We note that the two methods, entropy balancing and logistic regression, are related in that when models are correctly specified, $\beta_{\textnormal{EB}}=\beta_{\textnormal{LR}}$ and $\alpha_{\textnormal{EB}}=\alpha_{\textnormal{LR}}-\log(\textnormal{pr}(T=0)/\textnormal{pr}(T=1))$, while their estimates are numerically different due to different first-order conditions (Zhao and Percival,, 2017). In practice, we recommend using entropy balancing over logistic regression since by enabling covariate balancing, entropy balancing is more robust when models are slightly mis-specified (Imai and Ratkovic,, 2014; Zhao, 2019a, ).
E.7.3 Estimation for Discrete Covariates
Suppose $X$ is discrete with a finite support, $\mathcal{S}_{X}=\{x_{1},x_{2},\cdots,x_{K}\}$ where $K$ is fixed. Then $\omega(x)$ can be estimated by the ratio of the empirical distributions of $f_{Q_{X}}(x)$ and $f_{P_{X}}(x)$ for $x\in\mathcal{S}_{X}$, i.e.,
$\widehat{\omega}(x)=\dfrac{\sum_{T_{i}=1}\mathds{1}(X_{i}=x)/n_{q}}{\sum_{T_{i}=0}\mathds{1}(X_{i}=x)/n_{p}}.$ (50)
Following the notation in Section E.2, let $\eta_{5}=[w_{1},w_{2},\cdots,w_{K}]^{\intercal}$ where $w_{k}=f_{Q_{X}}(x_{k})/f_{P_{X}}(x_{k})$ for $k=1,2,\cdots,K$. Then by defining the estimating equations as
$g_{5}^{\textnormal{W}}(T,O,X,\eta^{\textnormal{W}})$ $=\begin{bmatrix}\dfrac{T\mathds{1}(X=x_{1})}{n_{q}}-w_{1}\cdot\dfrac{(1-T)\mathds{1}(X=x_{1})}{n_{p}}\\ \ldots\\ \dfrac{T\mathds{1}(X=x_{K})}{n_{q}}-w_{K}\cdot\dfrac{(1-T)\mathds{1}(X=x_{K})}{n_{p}}\end{bmatrix}$,
the nuisance parameter can be estimated by $\widehat{\eta}_{5}=[\widehat{w}_{1},\widehat{w}_{2},\cdots,\widehat{w}_{5}]$ where $\widehat{w}_{k}$’s are estimated by setting $\sum_{i=1}^{n}g_{5}^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})=0$. Then $\widehat{\omega}(x_{k})$ can be estimated with $\widehat{w}_{k}$ for $x_{k}\in\mathcal{S}_{X}$.
E.8 General Statements for the Efficient Influence Function
In this section, we derive the efficient influence function (EIF) of SLOPE when the target functional (and therefore the SLOPE) is scalar valued. Let $\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$ be EIF for $\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$. It is provided by Proposition 1. Additionally, the next Proposition 2 gives the EIF of the target functional, denoted as $\textnormal{EIF}(T,O,X,\psi)$.
Proposition 1 (EIF of SLOPE).
Suppose $\psi$ and $s$ are one-dimensional. The efficient influence function of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ is
$\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\cdot\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$
$=$ $\dfrac{(1-T)\omega(X)}{\textnormal{pr}(T=0)}\left[-s(O,X,\psi)\{O-\mu(X)\}+\textnormal{Cov}_{P_{O\mid X}}[s(O,X,\psi),O\mid X]+\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\{O-\mu(X)\}\right]$
$-\dfrac{T}{\textnormal{pr}(T=1)}\textnormal{E}_{P_{O\mid X}}[s(O,X,\psi)\{O-\mu(X)\}\mid X]$
$-\textnormal{SLOPE}(Q_{O,X}^{0},\psi)\left(\dfrac{(1-T)\omega(X)}{\textnormal{pr}(T=0)}\left[\dot{s}(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}\right]-\dfrac{T}{\textnormal{pr}(T=1)}\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}\right)$
$-\textnormal{EIF}(T,O,X,\psi)\left(\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]+\textnormal{SLOPE}(Q_{O,X}^{0},\psi)\cdot\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}\right),$
where $\textnormal{EIF}(T,O,X,\psi)$ be the EIF of the target functional $\psi(Q_{O,X}^{0})$ (see Proposition 2).
Proposition 2 (EIF of The Target Functional).
The efficient influence function of $\psi(Q_{O,X}^{0})$ is
$\textnormal{EIF}(T,O,X,\psi)=$ $-\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\dfrac{(1-T)\omega(X)}{\textnormal{pr}(T=0)}\left[\{s(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\right]$
$-\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\dfrac{T}{\textnormal{pr}(T=1)}\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}.$
We elaborate these two propositions on a few examples. First, suppose the target functional is $\psi^{{\textnormal{mean}}}$ with $s(O,X,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$. Then $\dot{s}(O,X,\psi)=-1$ and $\ddot{s}(O,X,\psi)=0$. With Proposition 2, the EIF for the target functional is
$\textnormal{EIF}(T,O,X,\psi^{{\textnormal{mean}}})$ $=\dfrac{1-T}{\textnormal{pr}(T=0)}\omega(X)\{O-\mu(X)\}+\dfrac{T}{\textnormal{pr}(T=1)}\left\{\mu(X)-\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})\right\}.$ (51)
This is identical to the EIF derived by Zeng et al., (2023) in their special case when the source and target samples share the same set of covariates. Estimators for the mean based on this EIF were proposed by Dahabreh et al., (2020) and then by Zeng et al., (2023) in a more general setting.
Further, by Proposition 1, the EIF of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})$ is
$\begin{split}\textnormal{EIF}(T,O,X,\textnormal{SLOPE})=&\dfrac{1-T}{\textnormal{pr}(T=0)}\omega(X)\left[\{O-\mu(X)\}^{2}-\sigma^{2}(X)\right]\\ &+\dfrac{T}{\textnormal{pr}(T=1)}\left\{\sigma^{2}(X)-\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})\right\}.\end{split}$ (52)
(52) consists of two parts. The first part is indexed by 1T1-T and it is a weighted, mean-zero term involving the source data. The weight is the density ratio $\omega(X)$ that reweighs the source covariates to match the target population, and the mean zero part can be viewed as the residual of estimating $\sigma^{2}(X)=\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}^{2}\mid X\right]$. The second part is indexed by $T$ and it can be viewed as the conditional variance $\sigma^{2}(X)$ re-centered over the target population. By estimating nuisance functions $\mu(X)$, $\omega(X)$, and $\sigma^{2}(X)$ and setting the summation of the empirical EIFs to zero, we obtain an EIF-based estimator that was presented in (39).
For the second example, suppose the target functional is the OLS coefficient $\psi^{{\textnormal{OLS}}}$ where $s(O,X)=\textnormal{E}_{Q_{O,X}}\left[XX^{\intercal}\psi^{{\textnormal{OLS}}}-XO\right]=0$ with $X$ includes an intercept term as its first component. We notice that although Proposition 1 has been stated for one-dimensional parameters, it is also valid for the $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$ because the corresponding $\ddot{s}(O,X,\psi^{{\textnormal{OLS}}})=0$. Consequently, the EIF for $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$ is
$\textnormal{EIF}(T,O,X,\textnormal{SLOPE})=$ $\left\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\right\}^{-1}\left\{\dfrac{1-T}{\textnormal{pr}(T=0)}\omega(X)\left[\{O-\mu(X)\}^{2}-\sigma^{2}(X)\right]+\dfrac{T}{\textnormal{pr}(T=1)}\sigma^{2}(X)\right\}.$
After presenting these examples of the EIF, finally, we comment on some limitations for estimators based on the EIF. First, the EIF for the SLOPE does not always exist. For example, when the target estimand is the median, $\psi^{{\textnormal{med}}}$, the SLOPE involves the conditional density $f_{P_{O\mid X}}$, and hence the EIF does not exist. Second, unlike many well-known examples in the literature, the SLOPE estimators based on the EIF often do not enjoy the double robustness property. In the simplest case when the target functional is the mean, $\psi^{{\textnormal{mean}}}$, the EIF-based estimator (39) involves three nuisance functions, $\omega(X)$, $\mu(X)$, and $\sigma^{2}(X)$. Roughly speaking, this estimator is conditionally doubly robust in that as long as $\widehat{\mu}(x)$ is consistent, the estimator will be consistent when either $\widehat{\omega}(x)$ or $\widehat{\mu}(x)$ is consistent. A similar result for the conditional rate double robustness holds under standard regularity conditions or cross fitting procedures. Due to these limitations and the complexity of the EIF, in practice we recommend using weighting or regression estimators presented in Section 4.2 of the main text. Meanwhile, we implement all three estimators in the numeric simulations (Section G) and we recognize the development of robust estimators of the SLOPE as an import future direction.
Appendix F Supplementary Materials for Data Application
In this section, we provide supplementary results for the data application in Section 5 of the main text.
F.1 Causal Identification
On the source population, we assume SUTVA (Assumption 4) holds for potential outcome $Y(a)$ under treatment ($a=1$) and control ($a=0$). Let $X$ be the baseline measurement and $V$ be the village that a household belongs to. We assume the strong ignorability assumption holds with $X$ and $V$, i.e., Assumption 3; note that this holds by the design of Banerjee et al., (2015). Under these identification assumptions, the SLOPE for the mean becomes
$\textnormal{SLOPE}(Q_{Y(a),X}^{0},\psi^{{\textnormal{mean}}})$
$=$ $\textnormal{E}_{Q_{X}}\left\{\textnormal{Var}_{P_{Y(a)\mid X}}[Y(a)\mid X]\right\}$
$=$ $\textnormal{E}_{Q_{X}}\textnormal{E}_{P_{V\mid X}}\left\{\textnormal{Var}_{P_{Y\mid X,V,A=a}}[Y\mid X,V,A=a]\mid X\right\}+\textnormal{Var}_{P_{V\mid X}}\left[\mu_{a}(X,V)\mid X\right],$ (53)
where we let $\mu_{a}(X,V)=\textnormal{E}_{P_{Y\mid X,V,A=a}}(Y\mid X,V,A=a)$.
The SLOPE for the median becomes
$\textnormal{SLOPE}(Q_{Y(a),X}^{0},\psi^{{\textnormal{med}}})$
$=$ $\dfrac{\textnormal{E}_{Q_{X}}\left[F_{P_{Y(a)\mid X}}(m_{1/2}\mid X)\mu(X)\right]-\textnormal{E}_{Q_{Y(a),X}^{0}}\left[Y(a)\mathds{1}\{Y(a)\leq m_{1/2}\}\right]}{f_{Q_{Y(a)}^{0}}(m_{1/2})}.$
$=$ $\dfrac{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{V\mid X}}\left[F_{P_{Y\mid X,V,A=a}}(m_{1/2}\mid X,V)\mu_{a}(X,V)-\textnormal{E}_{Y\mid X,V,A=a}\{Y\mathds{1}(Y\leq m_{1/2}\}\Bigg|X\right]\right)}{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{V\mid X}}\left[f_{Y\mid X,V,A=a}(m_{1/2}\mid X,V)\Bigg|X\right]\right)},$ (54)
where $m_{1/2}$ is the marginal median such that
$\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{V\mid X}}\left\{F_{P_{Y\mid X,V,A=a}}(m_{1/2}\mid X,V)\mid X\right\}\right]=1/2.$ (55)
F.2 Estimation of the SLOPE
With a slight abuse of notation, we let $\mu_{a}(X,V)=\textnormal{E}_{P_{Y\mid X,V,A=a}}(Y\mid X,V,A=a)$ and $\mu_{a}(X)=\textnormal{E}_{P_{V\mid X}}\{\mu_{a}(X,V)\mid X\}$. In notation we keep $a\in\{0,1\}$ for generality while in the main text we have focused on $a=1$. For estimation, we assume a linear model whereas for individual $i$ in treatment group $a$, i.e. $A_{i}=a$, we have
$\mu_{a}(x,v)=\alpha_{v}+\beta_{x}+\delta_{a}+(\beta\delta)_{xa},$ (56)
where $v$ is a discrete variable ranges in all villages in the source country and $x\in\mathcal{S}_{X}$ ranges in the categories of the baseline variable, and regression coefficients are constrained to guarantee identification: $\beta_{1}=0$, $\delta_{0}=0$, $(\beta\delta)_{x0}=0$ for $x\in\mathcal{S}_{X}$ and $(\beta\delta)_{0a}=0$ for $a\in\{0,1\}$. In addition, we assume that the outcome variance depends on the baseline measurement $x$ as below.
$\sigma^{2}_{a}(x):=\textnormal{Var}[Y\mid X=x,V=v,A=a].$
In Sections 5.2 and 5.3, the SLOPE for the mean is estimated by a simple plug-in estimator of (53):
$\sum_{x\in\mathcal{S}_{X}}\left[\widehat{Q}_{X}(x)\left\{\widehat{\sigma}^{2}_{a}(x)+\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X=x}(v\mid x)\widehat{\mu}_{a}(x,v)\right\}\right],$
where $\widehat{\mu}_{a}(x,v)$ is estimated by the least squares estimator that regresses $Y$ on $X$ and $V$ as in (56), $\widehat{\sigma}^{2}_{a}(x)$ is estimated by the sample variance of regression residuals, $\widehat{Q}_{X}(x)$ and $\widehat{P}_{V\mid X=x}(v\mid x)$ are estimated by the empirical distribution of the corresponding distributions. Since $X$ is discrete, this plug-in estimator can alternatively be viewed as a weighting estimator with weights estimated by empirical distributions of $X$ in the source and target populations.
To estimate the SLOPE for the median in Section 5.2, we assume $Y_{i}-\mu_{a}(X_{i},V_{i})$ follows a normal distribution. Under this conditional normality assumption, the SLOPE in (54) can be estimated by a plug-in (or equivalently, weighting) estimator as follows:
$\sum_{x\in\mathcal{S}_{X}}\widehat{Q}_{X}(x)\left[\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X}(v\mid x)\Phi\left(\dfrac{\widehat{m}_{1/2}-\widehat{\mu}_{a}(x,v)}{\widehat{\sigma}_{a}^{2}(x)}\right)\right]\left[\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X}(v\mid x)\widehat{\mu}_{a}(x,v)\right]$
$-\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X}(v\mid x)\left[\widehat{\mu}_{a}(x,v)\Phi\left(\dfrac{\widehat{m}_{1/2}-\widehat{\mu}_{a}(x,v)}{\widehat{\sigma}_{a}^{2}(x)}\right)-\widehat{\sigma}_{a}(x)\varphi\left(\dfrac{\widehat{m}_{1/2}-\widehat{\mu}_{a}(x,v)}{\widehat{\sigma}_{a}^{2}(x)}\right)\right],$
where $\varphi(\cdot)$ and $\Phi(\cdot)$ are the probability density functional and cumulative distribution function of the standard normal distribution, and $\widehat{m}_{1/2}$ is the solution of the following equation using bisection search,
$\sum_{x\in\mathcal{S}_{X}}\widehat{Q}_{X}(x)\sum_{v\in\mathcal{S}_{V}}\widehat{P}_{V\mid X}(v\mid x)\Phi\left(\dfrac{\widehat{m}_{1/2}-\widehat{\mu}_{a}(x,v)}{\widehat{\sigma}_{a}^{2}(x)}\right)=1/2.$
F.3 Auxiliary Data Information and Results for Section 5.2
In this section, we provide additional data information and analysis results for Section 5.2 where the outcome variable is the log-transformed per capita consumption.
F.3.1 Auxiliary Data Information
Table 2 gives the distribution of the categorized baseline measurement of the log-transformed per-capita consumption across countries on the overlapped sample.
Table 2: Baseline Measurement $X$ of the log-transformed per-capita consumption.
Peru Pakistan India Honduras Ghana
Sample size 1768 829 771 2152 2379
Category of $X$
(-0.41,3.65] 97 ( 5.5%) 50 ( 6.0%) 450 (58.4%) 1026 (47.7%) 1001 (42.1%)
(3.65,4.24] 449 (25.4%) 170 (20.5%) 260 (33.7%) 817 (38.0%) 832 (35.0%)
(4.24,8] 1222 (69.1%) 609 (73.5%) 61 (7.9%) 309 (14.4%) 546 (23.0%)
Figure 6 provides the normal diagnostics for the conditional normal assumption imposed in Section 5.2 when estimating the SLOPE for the median. From these QQ-plots, the normality assumption is reasonable well for most countries.
Figure 6: $Q$-plots for the residuals of the linear model (56) across source countries. Each plot is generated by qqplot in R where the straight line generated by qqline passes through the first and third quartiles.
F.3.2 Additional Results
First, we present one hypothesis on why SLOPE for the average per capita consumption is lower in India and Peru. For India, the experiment was located at West Bengal, an area abutting Bangladesh and shares a language and a culture. Therefore, the unique cultural and locational features may have caused the uniformity of the underlying population. For Peru, according to Banerjee et al., (2015), there has already been a consumption support program implemented on part of the households. This may has led to a higher homogeneity among the treated households.
Second, as mentioned in the main text, Table 3 presents the mean and median of the transported per-capita consumption.
Table 3: The estimated mean and median (not their SLOPEs) for transporting the counterfactual log-transformed per capita consumption under intervention from a source country (i.e., the rows of the table) to a target country (i.e., the columns of the table). Bootstrap standard errors are in the parentheses.
Estimand ($\psi$) Source ($P_{O\mid X}$) Target ($Q_{X}$)
Ghana Honduras India Pakistan Peru
Mean Ghana 3.47 (0.02) 3.44 (0.02) 3.67 (0.03) 3.67 (0.03)
Honduras 4.23 (0.02) 4.15 (0.02) 4.42 (0.04) 4.42 (0.04)
India 4.07 (0.03) 4.04 (0.02) 4.24 (0.07) 4.24 (0.07)
Pakistan 4.24 (0.05) 4.21 (0.05) 4.17 (0.06) 4.42 (0.02)
Peru 4.77 (0.03) 4.74 (0.03) 4.71 (0.04) 4.93 (0.02)
Median Ghana 3.48 (0.02) 3.44 (0.02) 3.68 (0.03) 3.67 (0.03)
Honduras 4.23 (0.02) 4.16 (0.02) 4.42 (0.04) 4.41 (0.04)
India 4.05 (0.02) 4.03 (0.02) 4.23 (0.07) 4.22 (0.06)
Pakistan 4.24 (0.04) 4.21 (0.05) 4.18 (0.06) 4.42 (0.02)
Peru 4.76 (0.03) 4.73 (0.03) 4.70 (0.04) 4.94 (0.02)
F.3.3 First-Order Approximation of Bias
For each pair of source and target countries, we estimate the oracle bias for the target country, i.e., the left hand side of (5). In specific, we estimate the mean/median of the potential outcome in the target country by either transporting from the source country or using the target data (with the outcome information) itself. The difference between the two estimates is treated as the “oracle bias” from conditional exchangeability since it represent the bias one my occur by directly assuming conditional exchangeability when outcome information in the target is unavailable. The confidence intervals of the oracle bias (i.e., horizontal bars) in Figure 7 are obtained via bootstrap.
Next, we estimate the bias approximated with SLOPE, i.e., the right hand side of (5). In addition to estimating SLOPE as described in Section F.2, we also estimate the sensitivity parameter $\gamma$ as follows. First, suppose the normal assumption holds, i.e., $Y_{i}-\mu_{a}(X_{i},V_{i})$ is conditionally normally distributed with mean zero and variance $\sigma^{2}_{a}(X_{i})$. Then the sensitivity model (3) enlists a location shift in the errors between the source and the target, i.e., $Q_{Y\mid X,V,A=a}\sim N\left(\mu_{a}(X,V)+\gamma\sigma_{a}^{2}(X),\sigma_{a}^{2}(X)\right)$. Therefore, by method of moment, we estiamte $\gamma$ through the following formula,
$\sum_{i=1}^{n_{p}}\left[\widehat{\mu}_{a}(X_{i},V_{i})+\widehat{\gamma}\cdot\widehat{\sigma}_{a}^{2}(X_{i})\right]=\sum_{i=n_{p}+1}^{n}\widehat{\sigma}_{a}^{2}(X_{i}),$
where $\widehat{\mu}_{a}$ and $\widehat{\sigma}_{a}^{2}$ are estimates of $\mu_{a}$ and $\sigma_{a}^{2}$, respectively. Therefore, the approximated bias is the product of $\widehat{\gamma}$ and the estimate of SLOPE. For confidence intervals, we fix $\widehat{\gamma}$ as obtained as above from the original source and target samples, and construct $95\%$ confidence intervals for the SLOPE by bootstrapping the source and the target samples. Therefore, the vertical bars shown in Figure 7 do not include the randomness of estimating $\gamma$. Such construction is to align with the common understanding in sensitivity analysis that the sensitivity parameter is taken as a pre-specified fixed value instead of a random quantity.
The results are shown in Figure 7, where the two panels represent the mean and median, and each panel plots the approximated bias with SLOPE against the oracle bias. These dots roughly lie on the $y=x$ line, thereby validating the bias approximation with SLOPE.
Figure 7: Bias approximation with SLOPE for mean (left) and median (right). Each panel plots the bias approximated with SLOPE in the y-axis and the oracle bias in the x-axis in dots and the corresponding bootstrapped $95\%$ confidence interval in bars. The dashed straight line is $y=x$.
F.4 Auxiliary Data Information for Section 5.3
For the data analysis in Section 5.3, all health variables and the corresponding physical health index were measured at individual level. To keep the sample units as households, we average the individual level measurements over households. Although we changed the outcome variable in terms of the weights, we kept the pre-treatment covariate the same to guarantee fair comparison across weighting schemes. The pre-treatment covariate $X$ is the categorized baseline measurement of the physical health index (i.e., the average of three health variables mentioned in Section 5.3). Table 4 gives the distribution of $X$ across countries on the overlapped sample.
Table 4: Baseline Measurement $X$ of physical health index.
Peru India Ethiopia
Sample size 1307 740 785
Category of $X$
(-1.36, 0.235] 979 (74.9%) 523 (70.7%) 350 (44.6%)
(0.235,0.818] 328 (25.1%) 217 (29.3%) 435 (55.4%)
Appendix G Simulations
In this section, we validate the asymptotic properties of the proposed estimators in synthetic datasets that were generated to mimic the real data.
G.1 Simulation Setting
Suppose $O$ is a continuous variable and $X$ is a random variable that is either binary and continuous. In the case when $X$ is binary, we suppose the support $\mathcal{S}_{X}=\{1,2\}$ and $P_{O\mid X=j}\sim N(\mu_{j},\sigma_{j}^{2})$ with $j=1,2$. In the case when $X$ is continuous, we suppose both $P_{X}$ and $Q_{X}$ are Gaussian and $P_{O\mid X}\sim N(\alpha_{m}+\beta_{m}X,\alpha_{v}+\beta_{m}X^{2})$. Data generation parameters were estimated from real data in Section 5.2. Specifically, the conditional distribution $P_{O\mid X}$ is estimated from the per-capita consumption in the treated group of Pakistan. $Q_{X}$ is estimated from Pakistan (i.e., no covariate shift) and Honduras (i.e., covariate shift). To construct a binary $X$, we dichotomize the log-transformed baseline measurement by whether it is below the median; to construct a continuous $X$ we use the original log-transformed measurement. The exact numbers of the (semi-)synthetic distributions are given in Table 5.
For both data generation procedures, we are interested in the SLOPE for mean and median. When $X$ is continuous, we also consider the regression coefficients in simple linear regression which regresses $O$ on $X$. For SLOPE of the mean and OLS coefficients, we consider three estimators: weighting estimator ((36) and (41)), the regression estimator ((37) and (42)), and the efficient influence function based estimator (39). For the SLOPE of the median, we consider the weighting estimator (45) and the regression estimator (46). During estimation we suppose $P_{O\mid X}$ is Gaussian. For nuisance functions/parameters, $\omega(X)$ is estimated with empirical distributions when $X$ is binary (see Section E.7.3) and entropy balancing when $X$ is continuous (see Section E.7.2), $\mu(X)$ is estimated with linear regression, $\sigma^{2}(X)$ is estimated with subgroup sample variance when $X$ is binary and is estimated by regressing the squared residuals $\{O-\widehat{\mu}(X)\}^{2}$ on $X^{2}$ when $X$ is continuous. The target mean is estimated by a weighted average $\sum_{i=n_{q}+1}^{n}O_{i}\widehat{\omega}(X_{i})/n_{p}$ for weighting estimators and by regression $\sum_{i=1}^{n_{q}}\widehat{\mu}(X_{i})/n_{q}$ for regression estimators. For the target median, when $X$ is binary it is estimated with the qmixnorm function in the R package KScorrect where the group mean and variance of $P_{O\mid X=j}$ are estimated by maximum likelihood estimation; when $X$ is continuous it is estimated by the numerical solution to
$\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\dfrac{1}{\widehat{\sigma}(X_{i})}\varphi\left(\dfrac{m_{1/2}-\widehat{\mu}(X_{i})}{\widehat{\sigma}(X_{i})}\right)=1/2$
via bisection, where $\varphi(\cdot)$ is the p.d.f. of the standard Gaussian. Inference for the weighting and regression estimators is based on bootstrap with 1000 times of resampling and inference for the EIF-based estimator is based on the second moment of the EIF with nuisances plugged in (i.e., (40) and (44)).
Table 5: Data generation in simulated data.
$X$ Cov. shift $Q_{X}$ $P_{X}$ $P_{O\mid X}$
Binary $Q(X=1)$ $Q(X=2)$ $P(X=1)$ $P(X=2)$ $\mu_{1}$ $\mu_{2}$ $\sigma_{1}$ $\sigma_{2}$
No $0.1258$ $0.8742$ $0.1258$ $0.8742$ $4.1816$ $4.4773$ $0.4761$ $0.4524$
Yes $0.6597$ $0.3403$
Continuous $\textnormal{E}_{Q_{X}}(X)$ $\sqrt{\textnormal{Var}_{Q_{X}}(X)}$ $\textnormal{E}_{P_{X}}(X)$ $\sqrt{\textnormal{Var}_{P_{X}}(X)}$ $\alpha_{m}$ $\beta_{m}$ $\alpha_{v}$ $\beta_{v}$
No $4.5803$ $0.5970$ $4.5803$ $0.5970$ $3.1304$ $0.2766$ $0.1924$ 0.0003-0.0003
Yes $3.7054$ $0.5340$
G.2 Simulation Result
Simulations are based on 1000 replicates. In each data setting, we consider $n_{p}=n_{q}\in\{1000,2000\}$ and report the bias (bias), root mean squared error (rMSE), empirical standard deviation (empSD), the average estimated standard error (avgSE) and covarage rate (rate). Tables 6 and 7 list the simulation results for the SLOPE of mean and median. For the OLS coefficients, Table 8 and Table 9 provides results for the slope coefficient and the intercept coefficient ,respectively. As these results show, in all cases, the bias becomes closer to zero and sample sizes increases and the rMSE decays with root $n_{p}$. The average of estimated standard error is close to the empirical standard deviation. The coverage rate is closer to $95\%$. Overall, the simulation results suggest that the estimators are $\sqrt{n}$-consistent and the standard error estimates are consistent.
Table 6: Simulation results for SLOPE of mean and median when $X$ is binary. All numbers have been multiplied with $100$.
Estimand Mean Median
Covariate shift Yes No Yes No
Estimator Regress Weight EIF Regress Weight EIF Regress Weight Regress Weight
$n_{p}=1000$ bias 0.01 -0.10 -0.10 0.03 -0.01 -0.02 0.03 0.01 -0.01 -0.02
rMSE 1.66 1.65 1.65 1.01 1.00 1.04 1.86 1.90 1.01 1.04
empSD 1.66 1.65 1.65 1.01 1.00 1.04 1.86 1.90 1.01 1.04
avgSE 1.67 1.66 1.66 1.01 1.01 1.01 1.78 1.78 1.01 1.01
rate 94.7% 94.6% 94.4% 94.8% 94.8% 94.7% 93.7% 92.0% 94.5% 93.9%
$n_{p}=2000$ bias -0.01 -0.07 -0.07 0.00 -0.02 -0.02 -0.11 -0.13 -0.02 -0.03
rMSE 1.18 1.18 1.18 0.68 0.68 0.68 1.29 1.33 0.72 0.73
empSD 1.18 1.18 1.18 0.68 0.68 0.68 1.29 1.32 0.72 0.73
avgSE 1.19 1.19 1.19 0.72 0.72 0.72 1.26 1.26 0.72 0.72
rate 95.3% 94.9% 94.8% 95.7% 95.8% 95.8% 94.5% 93.6% 94.8% 94.4%
Table 7: Simulation results for SLOPE of mean and median when $X$ is continuous. All numbers have been multiplied with $100$.
Estimand Mean Median
Covariate shift Yes No Yes No
Estimator Regress Weight EIF Regress Weight EIF Regress Weight Regress Weight
$n_{p}=1000$ bias -0.09 -0.15 -0.09 -0.09 -0.09 -0.09 -0.06 -0.07 -0.06 -0.06
rmse 0.86 2.16 2.06 0.86 0.86 0.86 0.87 1.48 0.87 0.89
empSD 0.85 2.16 2.05 0.85 0.85 0.86 0.86 1.48 0.86 0.89
avgSE 0.84 2.24 1.92 0.84 0.85 0.84 0.84 1.46 0.84 0.88
rate 93.9% 91.6% 91.9% 93.9% 94.1% 94.0% 94.3% 93.6% 94.3% 94.5%
$n_{p}=2000$ bias -0.01 -0.14 -0.11 -0.01 -0.01 -0.01 0.01 -0.01 0.01 0.01
rmse 0.61 1.53 1.44 0.61 0.61 0.61 0.59 1.07 0.59 0.62
empSD 0.61 1.53 1.44 0.61 0.61 0.61 0.59 1.07 0.59 0.62
avgSE 0.60 1.44 1.38 0.60 0.60 0.60 0.60 1.04 0.60 0.62
rate 94.2% 92.4% 93.2% 94.2% 94.4% 94.7% 95.3% 93.7% 95.3% 95.5%
Table 8: Simulation results for SLOPE of the slope coefficient in simple linear regression. All numbers have been multiplied with $100$.
Estimand Slope
Covariate shift Yes No
Estimator Regress Weight EIF Regress Weight EIF
$n_{p}=1000$ bias -0.02 -0.03 -0.02 -0.02 -0.02 -0.02
rmse 0.23 0.47 0.47 0.19 0.19 0.19
empSD 0.23 0.47 0.47 0.18 0.19 0.19
avgSE 0.22 0.45 0.45 0.18 0.18 0.19
rate 93.9% 91.0% 92.4% 93.9% 94.1% 95.6%
$n_{p}=2000$ bias 0 -0.03 -0.03 0 0 0
rmse 0.16 0.33 0.33 0.13 0.13 0.13
empSD 0.16 0.33 0.33 0.13 0.13 0.13
avgSE 0.16 0.32 0.32 0.13 0.13 0.14
coverage 94.5% 92.9% 94.4% 94.4% 94.3% 95.9%
Table 9: Simulation results for SLOPE of the intercept coefficient in simple linear regression. All numbers have been multiplied with $100$.
Estimand Intercept coefficient
Covariate shift Yes No
Estimator Regress Weight EIF Regress Weight EIF
$n_{p}=1000$ bias -0.09 -0.12 -0.09 -0.09 -0.09 -0.09
rmse 0.86 2.09 2.06 0.86 0.86 0.86
empSD 0.85 2.08 2.05 0.85 0.86 0.86
avgSE 0.84 1.93 1.94 0.84 0.84 0.90
rate 93.9% 90.6% 92.3% 93.9% 94.0% 95.2%
$n_{p}=2000$ bias -0.01 -0.13 -0.11 -0.01 -0.01 -0.01
rmse 0.61 1.46 1.44 0.61 0.61 0.61
empSD 0.61 1.46 1.44 0.61 0.61 0.61
avgSE 0.60 1.38 1.39 0.60 0.60 0.64
rate 94.2% 92.4% 93.4% 94.2% 94.3% 96.4%
Appendix H Proof for The Derivation of the SLOPE
H.1 Proof of Theorem 2
Since Theorem 2 is a special case of Theorem 6 with $q=1/2$, please find the proof of Theorem 6 in Section H.5.
H.2 Proof of Theorem 3
We recall some notation defined in Section A.1 for the support under $Q_{O,X}^{0}$: denote the support of $Q_{O,X}^{0}$ as $\mathcal{S}_{O,X}$ and the supports of the marginals $Q_{X}$ and $Q_{O}^{0}$ as $\mathcal{S}_{X}$ and $\mathcal{S}_{O}$, respectively.
Proof of Theorem 3. The proof proceeds in two steps. First, we show that fixing $Q_{O,X}=Q_{O,X}^{0}$, the (partial) derivative of $\phi$ with respect to $\gamma$ at $\gamma=0$ is
$\phi^{\prime}_{\gamma}(0)=\widetilde{Q}_{O,X}\in l^{\infty}(\mathcal{S}_{O,X}),\text{ such that }\int_{B}d\widetilde{Q}_{O,X}=\int_{B}\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0},$
for any measurable set $B\in\mathcal{S}_{O,X}$.
Second, we show that for an $H\in l^{\infty}(\mathcal{S}_{O,X})$ such that $\int dH=0$, the Hadamard derivative of $\psi$ with respect to $Q_{O,X}$ at $Q_{O,X}^{0}$ in the direction of $H$ is
$\psi^{\prime}_{Q_{O,X}^{0}}(H)=\int\textnormal{IF}\left(O,X,\psi(Q_{O,X}^{0})\right)dH.$
Then, using chain rule of Hadamard derivative (Theorem 20.4 of Van der Vaart, (2000)), we have the derivative of the composite function $\psi\circ\phi$ with respect to $\gamma$ at $\gamma=0$ being:
$\psi^{\prime}_{Q_{O,X}}\left(\phi^{\prime}_{\gamma}(0)\right)=\int\textnormal{IF}\left(O,X,\psi(Q_{O,X}^{0})\right)\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0}.$
We prove the two steps in order.
For the first step, we note that for any $h\in\mathcal{R}$ and $t\downarrow 0$,
$\int_{B}dQ_{O,X}^{th}-\int_{B}dQ_{O,X}^{0}$ $=\int_{B}\left(\dfrac{dQ_{O,X}^{th}}{dQ_{O,X}^{0}}-1\right)dQ_{O,X}^{0}$
$=\int_{B}\left[\dfrac{\exp(th)}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(thO)\mid X\right\}}\right]dQ_{O,X}^{0}.$
Let $\Delta_{t,h}=\dfrac{dQ_{O,X}^{th}}{dQ_{O,X}^{0}}-1$, then
$\Delta_{t,h}$ $=\dfrac{1+thO+O(t^{2})}{1+th\textnormal{E}_{P_{O\mid X}}(O\mid X)+O(t^{2})}-1$
$=\left\{1+thO+O(t^{2})\right\}\left\{1-th\textnormal{E}_{P_{O\mid X}}(O\mid X)+O(t^{2})\right\}-1+O(t^{2})$
$=th\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}+O(t^{2}).$
Therefore,
$\dfrac{\int_{B}dQ_{O,X}^{th}-\int_{B}dQ_{O,X}^{0}}{t}=$ $\dfrac{\int_{B}\Delta_{t,h}dQ_{O,X}^{0}}{t}$
$=$ $\dfrac{\int_{B}th\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0}+O(t^{2})}{t}$
$\to$ $h\int_{B}\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0}.$
For the second step, we note that under Condition 3, $\textnormal{IF}\left(o,x,\psi(Q_{O,X}^{0})\right)=\psi^{\prime}_{Q_{O,X}^{0}}[\delta_{o,x}-Q_{O,X}^{0}]$. Then
$\int\textnormal{IF}\left(o,x,\psi(Q_{O,X}^{0})\right)dH(o,x)$ $=\int\psi^{\prime}_{Q_{O,X}^{0}}[\delta_{o,x}-Q_{O,X}^{0}]dH(o,x)$
$=\psi^{\prime}_{Q_{O,X}}\left[\int\delta_{o,x}-Q_{O,X}^{0}dH(o,x)\right]$
$=\psi^{\prime}_{Q_{O,X}}[H],$
where the second equality follows from the linearity and continuity of Hadamard derivative, and the last equation follows from the observation that for $B\in l^{\infty}(\mathcal{S}_{O,X})$,
$\int\delta_{o,x}(B)-Q_{O,X}^{0}(B)dH(o,x)=H(B)-Q_{O,X}^{0}(B)\cdot 0=H(B).$
Hence, both two steps have been proved. $\square$
H.3 Proof of Theorem 4
Proof of Theorem 4. The proof follows the same procedure as the proof of Theorem 3 except for replacing the exponential function $\exp(\gamma O)$ with $\rho(O,X,\gamma)$. Specifically, the conclusion of the first step becomes
$\phi^{\prime}_{\gamma}(0)=\widetilde{Q}_{O,X}\in l^{\infty}(\mathcal{S}_{O,X}),\text{ such that }\int_{B}d\widetilde{Q}_{O,X}=\int_{B}\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\mid X)\right]dQ_{O,X}^{0}.$
In the derivation of the first step, we use the Taylor expansion on $\rho(O,X,\gamma)$ around $\gamma=0$,
$dQ_{O,X}^{th}=1+th\dot{\rho}(O,X,0)+O(t^{2}),$
which holds under Condition 7. The rest of the proof follows the same procedure as the proof of Theorem 3.
H.4 Proof of Lemma 1
Proof of Lemma 1.
Under sensitivity model (3), the target estimand is
$\psi^{\xi}(Q_{O,X}^{\gamma})=\textnormal{E}_{Q_{X}}\left[\dfrac{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\xi(O,X)\mid X\right\}}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\mid X\right\}}\right].$
Taking derivative with respect to $\gamma$ at $\gamma=0$, we have
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\xi})$ $=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\left\{O\xi(O,X)\mid X\right\}-\textnormal{E}_{P_{O\mid X}}\left\{\xi(O,X)\mid X\right\}-\textnormal{E}_{P_{O\mid X}}(O\mid X)\right]$
$=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}\left[O,\xi(O,X)\mid X\right]\right\}.$
$\square$
H.5 Proof of Theorem 6
Proof of Theorem 6. Let $\kappa(\gamma)=\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}$. Under sensitivity model (3), the $q$-th quantile of $O$ under $Q_{O}^{0}$ can be defined as
$q$ $=F_{Q_{O}^{\gamma}}\left(\psi(F_{Q_{O}^{\gamma}})\right)$
$=\int_{-\infty}^{F_{Q_{O}^{\gamma}}^{-1}(q)}\int\dfrac{\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}}dQ_{X}dP_{O\mid X}$
$=\int\int_{-\infty}^{F_{Q_{O}^{\gamma}}^{-1}(q)}\dfrac{\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X)\}}dP_{O\mid X}dQ_{X}.$
On both sides, we take derivative with respect to $\gamma$. Applying the Leibniz rule, we have
$0$ $=\int\left(\dfrac{\partial}{\partial\gamma}\int_{-\infty}^{F_{Q_{O}^{0}}^{-1}(q)}\left[\dfrac{\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}}\right]dP_{O\mid X}\right)\Bigg|_{\gamma=0}dQ_{X}$
$=\int\left[f_{P_{O\mid X}}(m_{q}\mid X)\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi)+\int_{-\infty}^{m_{q}}\{O-\mu(X)\}dP_{O\mid X}\right]dQ_{X}$
$=\textnormal{SLOPE}(Q_{O,X}^{0},\psi)\cdot\int f_{P_{O\mid X}}(m_{q}\mid X)dQ_{X}+\textnormal{E}_{Q_{O}^{0}}\{O\mathds{1}(O\leq m_{q}\}-\textnormal{E}_{Q_{X}}\left\{F_{P_{O\mid X}}(m_{q})\mu(X)\right\}$
$=\textnormal{SLOPE}(Q_{O,X}^{0},\psi)f_{Q_{O}^{0}}(m_{q})+\textnormal{E}_{Q_{O}^{0}}\{O\mathds{1}(O\leq m_{q})\}-\textnormal{E}_{Q_{X}}\left\{F_{P_{O\mid X}}(m_{q})\mu(X)\right\}.$
Reorganize terms, we have
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)=$ $\dfrac{\textnormal{E}_{Q_{X}}\left\{F_{P_{O\mid X}}(m_{q})\mu(X)\right\}-\textnormal{E}_{Q_{O}^{0}}\{O\mathds{1}(O\leq m_{q})\}}{f_{Q_{O}^{0}}(m_{q})}.$
H.6 Proof of Theorem 7
Proof of Theorem 7. Under the sensitivity model (3), the target estimand is
$\psi^{{\alpha\textnormal{-trim}}}(Q_{O}^{\gamma})$ $=\dfrac{1}{1-2\alpha}\int_{F_{Q_{O}^{\gamma}}^{-1}(\alpha)}^{F_{Q_{O}^{\gamma}}^{-1}(1-\alpha)}OdQ_{O}^{\gamma}$
$=\dfrac{1}{1-2\alpha}\int_{F_{Q_{O}^{\gamma}}^{-1}(\alpha)}^{F_{Q_{O}^{\gamma}}^{-1}(1-\alpha)}\dfrac{O\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}}dQ_{O,X}^{0}.$
Taking derivative with respect to $\gamma$ at $\gamma=0$, we obtain the SLOPE as follows:
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\alpha\textnormal{-trim}}})=$ $\dfrac{\partial\psi^{{\alpha\textnormal{-trim}}}(Q_{O}^{\gamma})}{\partial\gamma}\Bigg|_{\gamma=0}$
$=$ $\dfrac{1}{1-2\alpha}\int\left\{F_{Q_{O}^{0}}^{-1}(1-\alpha)f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(1-\alpha))\left[\dfrac{\partial F_{Q_{O}^{\gamma}}^{-1}(1-\alpha)}{\partial\gamma}\Bigg|_{\gamma=0}\right]\right\}dQ_{X}$
$-\dfrac{1}{1-2\alpha}\int\left\{F_{Q_{O}^{0}}^{-1}(\alpha)f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(\alpha))\left[\dfrac{\partial F_{Q_{O}^{\gamma}}^{-1}(\alpha)}{\partial\gamma}\Bigg|_{\gamma=0}\right]\right\}dQ_{X}$
$+\dfrac{1}{1-2\alpha}\int\left[\int_{F_{Q_{O}^{0}}^{-1}(\alpha)}^{F_{Q_{O}^{0}}^{-1}(1-\alpha)}O\{O-\mu(X)\}dP_{O\mid X}\right]dQ_{X}$
$=$ $\dfrac{1}{1-2\alpha}F_{Q_{O}^{0}}^{-1}(1-\alpha)\textnormal{E}_{Q_{O,X}^{0}}\left[\mathbb{P}(O\leq F_{Q_{O}^{0}}^{-1}(1-\alpha)\mid X)\mu(X)-O\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(1-\alpha))\right]$
$-\dfrac{1}{1-2\alpha}F_{Q_{O}^{0}}^{-1}(\alpha)\textnormal{E}_{Q_{O,X}^{0}}\left[\mathbb{P}(O\leq F_{Q_{O}^{0}}^{-1}(\alpha)\mid X)\mu(X)-O\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(\alpha))\right]$
$+\dfrac{1}{1-2\alpha}\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[O\{O-\mu(X)\}\mathds{1}_{[F_{Q_{O}^{0}}^{-1}(\alpha),F_{Q_{O}^{0}}^{-1}(1-\alpha)]}(O)\mid X\right]\right),$
where the last equality follows from the SLOPE for quantiles (Theorem 6).
$\square$
H.7 Proof of Theorem 8
Proof of Theorem 8. This theorem is a special case of Corollary 2 with $s(Y,\psi^{\textnormal{OLS,Sub}})=X_{\textnormal{Sub}}X_{\textnormal{Sub}}^{\intercal}\psi^{\textnormal{OLS,Sub}}-X_{\textnormal{Sub}}Y$. $\square$
H.8 Proof of Theorem 9
Proof of Theorem 9.
We start with $\tau_{a}$. Suppose $\psi^{{\textnormal{OLS}}}$ satisfies $\textnormal{E}_{Q_{O,X}^{0}}\left(XX^{\intercal}\psi^{{\textnormal{OLS}}}-XY\right)$. From Theorem 8, the SLOPE for $\tau_{a}$ is the second entry of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$, which is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$
$=$ $\left\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\right\}^{-1}\textnormal{E}_{Q_{X}}\left[X\sigma^{2}(X)\right]$
$=$ $\left\{\begin{bmatrix}1&\textnormal{E}_{Q_{A}}(A)&\textnormal{E}_{Q_{L}}(L^{\intercal})\\ \textnormal{E}_{Q_{A}}(A)&\textnormal{E}_{Q_{A}}(A^{2})&\textnormal{E}_{Q_{X}}(AL^{\intercal})\\ \textnormal{E}_{Q_{L}}(L)&\textnormal{E}_{Q_{X}}(LA)&\textnormal{E}_{Q_{L}}(LL^{\intercal})\end{bmatrix}\right\}^{-1}\begin{bmatrix}&\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{A\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{L\sigma^{2}(X)\}\end{bmatrix}$
$=$ $\begin{bmatrix}1-b^{\intercal}S^{-1}b&{}^{*}&{}^{*}\\ -\frac{\textnormal{E}_{Q_{A}}(A))}{\textnormal{Var}_{Q_{A}}(A)}+\delta^{\intercal}V^{-1}\textnormal{E}_{Q_{A}}(A)\delta-\delta^{\intercal}V^{-1}\textnormal{E}_{Q_{L}}(L)&1/\textnormal{Var}_{Q_{A}}(A)+\delta^{\intercal}V^{-1}\delta&{}^{*}\\ V^{-1}\delta\textnormal{E}_{Q_{A}}(A)-V^{-1}\textnormal{E}_{Q_{L}}(L)&-V^{-1}\delta&-V^{-1}\end{bmatrix}\begin{bmatrix}&\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{A\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{L\sigma^{2}(X)\}\end{bmatrix}$,
where $b=[\textnormal{E}_{Q_{A}}(A),\textnormal{E}_{Q_{L}}(L^{\intercal})]$, $\delta=\textnormal{Cov}_{Q_{L,A}}[L,A]/\textnormal{Var}_{Q_{A}}[A]$, $V=\textnormal{Cov}_{Q_{X}}(X)-\delta\delta^{\intercal}\textnormal{Var}_{Q_{A}}(A)$, $S=\begin{bmatrix}\textnormal{Var}_{Q_{A}}(A)&\textnormal{Cov}_{Q_{X}}[A,L^{\intercal}]\\ \textnormal{Cov}_{Q_{X}}[L,A]&\textnormal{Cov}_{Q_{L}}(L)\end{bmatrix}$, and entries as “” are omitted because that matrix is symmetric. Then $\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})$ in (24) follows by expanding terms in the formula above and taking the second entry.
The SLOPE for $\tau_{\textnormal{u}}$ in the unadjusted model can be obtained similarly and is hence omitted. $\square$
H.9 Proof of Theorem 10
Proof of Theorem 10. Under sensitivity model (3), the MAD satisfies
$1/2=$ $\int_{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})-\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})}^{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})+\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})}dQ_{O,X}^{\gamma}$
$=$ $\textnormal{E}_{Q_{X}}\left[\int_{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})-\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})}^{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})+\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})}\dfrac{\exp(\gamma O)}{\textnormal{E}_{P_{O\mid X}}\{\exp(\gamma O)\mid X\}}dP_{O\mid X}\right].$
Taking derivative with respect to $\gamma$ at $\gamma=0$, we have
$0=$ $\textnormal{E}_{Q_{X}}\left[f_{P_{O\mid X}}\left(m_{1/2}+{\textnormal{MAD}}\mid X\right)\left\{\dfrac{\partial\left\{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})+\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})\right\}}{\partial\gamma}\Bigg|_{\gamma=0}\right\}\right]$
$-\textnormal{E}_{Q_{X}}\left[f_{P_{O\mid X}}\left(m_{1/2}-{\textnormal{MAD}}\mid X\right)\left\{\dfrac{\partial\left\{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})-\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})\right\}}{\partial\gamma}\Bigg|_{\gamma=0}\right\}\right]$
$+\textnormal{E}_{Q_{X}}\left[\int_{m_{1/2}-{\textnormal{MAD}}}^{m_{1/2}+{\textnormal{MAD}}}\{O-\mu(X)\}dP_{O\mid X}\right]$
$=$ $\left\{f_{Q_{O}^{0}}(m_{1/2}+{\textnormal{MAD}})+f_{Q_{O}^{0}}(m_{1/2}-{\textnormal{MAD}})\right\}\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{MAD}}})$
$+\left\{f_{Q_{O}^{0}}(m_{1/2}+{\textnormal{MAD}})-f_{Q_{O}^{0}}(m_{1/2}-{\textnormal{MAD}})\right\}\cdot\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\mathds{1}_{[m_{1/2}-{\textnormal{MAD}},m_{1/2}+{\textnormal{MAD}}]}(O)\{O-\mu(X)\}\right].$
Then the result follows from re-organizing this equation. $\square$
H.10 Proof of Theorem 11
Proof of Theorem 11. For the L-estimand
$\psi(Q_{O})$ $=\int_{0}^{1}h(F_{Q_{O}}^{-1}(p))l(p)dp,$
the corresponding SLOPE is (assuming differentiation and integration can exchange)
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$
$=$ $\dfrac{\partial\psi(Q_{O}^{\gamma})}{\partial\gamma}\Bigg|_{\gamma=0}$
$=$ $\int_{0}^{1}h^{\prime}(F_{Q_{O}^{0}}^{-1}(p))\cdot\frac{\partial F_{Q_{O}^{\gamma}}^{-1}(p)}{\partial\gamma}\Bigg|_{\gamma=0}l(p)dp$
$=$ $\int_{0}^{1}h^{\prime}(F_{Q_{O}^{0}}^{-1}(p))\cdot\left(\dfrac{-\textnormal{E}\{O\mathds{1}(O\leq F_{Q_{O}}^{-1}(p))\}}{f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(p))}+\dfrac{\textnormal{E}_{Q_{X}}\left[\mu(X)\textnormal{E}_{P_{O\mid X}}\left\{\mathds{1}(O\leq F_{Q_{O}^{0}}^{-1}(p))\mid X\right\}\right]}{f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(p))}\right)l(p)dp,$
where the second line follows from the SLOPE for quantiles (Theorem 6).
$\square$
We note that SLOPE of L-estimands can be alternatively derived from the IF below (Huber,, 1981, (3.11)) using Theorem 3,
$\textnormal{IF}(o,x,\psi(Q_{O,X}^{0}))$ $=\int\frac{ph^{\prime}(F_{Q_{O}^{0}}^{-1}(p))}{f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(p))}l(p)dp-\int_{F_{Q_{O}^{0}}(o)}^{1}\frac{h^{\prime}(F_{Q_{O}^{0}}^{-1}(p))}{f_{Q_{O}^{0}}(F_{Q_{O}^{0}}^{-1}(p))}l(p)dp.$
Appendix I Proof for the Estimation Theory
I.1 Proof of Theorem 12 and Theorem 13
The proof for Theorems 12 and 13 follows from standard M-estimation theory in Newey and McFadden, (1994) and Van der Vaart, (2000). We next prove Theorem 12 and omit the proof of Theorem 13 since the proof for both theorems follows the same procedure.
Proof of Theorem 12. We begin by proving consistency under Condition 8 using the standard techniques in Newey and McFadden, (1994, Theorem 2.1). Let
$M(\eta)$ $=-E\left[G^{W}(T_{i},O_{i},X_{i},\eta)\right]^{\intercal}E\left[G^{W}(T_{i},O_{i},X_{i},\eta)\right],\text{ and}$
$\widehat{M}_{n}(\eta)$ $=-\left\{\dfrac{1}{n}G^{W}(T_{i},O_{i},X_{i},\eta)\right\}^{\intercal}\left\{\dfrac{1}{n}G^{W}(T_{i},O_{i},X_{i},\eta)\right\}.$
First, under Condition 8(i), $\eta^{W}$ uniquely maximizes $M(\eta)$. Next, using Lemma 2.4 of Newey and McFadden, (1994), under Condition 8(ii) and (iii), we have the uniform convergence of $G^{W}$:
$\sup_{\eta\in\Theta}\left\lVert\dfrac{1}{n}\sum_{i=1}^{n}G^{W}(T_{i},O_{i},X_{i},\eta)-\textnormal{E}\{G^{W}(T_{i},O_{i},X_{i},\eta)\}\right\rVert\to_{p}0,$
and that the function $\textnormal{E}[G^{W}(T_{i},O_{i},X_{i},\eta)]$ is continuous with respect to $\eta$, where $\to_{p}$ denotes convergence in probability. Then $M(\eta)$ is also continuous with respect to $\eta$. Next we show the uniform convergence of $\widehat{M}_{n}$. Note that the compactness of $\Theta$ implies the boundedness of $\textnormal{E}[G^{W}(T_{i},O_{i},X_{i},\eta)]$. Then for any $\eta\in\Theta$,
$\left|\widehat{M}_{n}(\eta)-M(\eta)\right|$
$\leq$ $\left\lVert\textnormal{E}\left[G^{W}(T_{i},O_{i},X_{i},\eta)\right]-\dfrac{1}{n}\sum_{i=1}^{n}G^{W}(T_{i},O_{i},X_{i},\eta)\right\rVert\cdot\ \left\lVert\textnormal{E}\left[G^{W}(T_{i},O_{i},X_{i},\eta)\right]+\dfrac{1}{n}\sum_{i=1}^{n}G^{W}(T_{i},O_{i},X_{i},\eta)\right\rVert$
implies the uniform convergence,
$\sup_{\eta\in\Theta}|M_{n}(\eta)-M(\eta)|\to_{p}0.$ (57)
With all regularity conditions checked above, we prove the consistency mimicking the proof of Theorem 2.1 in Newey and McFadden, (1994). For any $\varepsilon>0$. with probability approaching one, we have
$M(\widehat{\eta}^{W})>\widehat{M}_{n}(\widehat{\eta}^{W})-\varepsilon/3>\widehat{M}_{n}(\eta^{W})-\varepsilon/3-\varepsilon/3>M(\eta^{W})-\varepsilon/3-\varepsilon/3-\varepsilon/3,=M(\eta^{W})-\varepsilon.$ (58)
where the first inequality and third equality hold by the uniform convergence (57), and the second inequality holds because $\widehat{\eta}^{W}$ uniquely maximizes $\widehat{M}_{n}(\cdot)$. Let $\mathcal{N}$ be an open subset of $\Theta$ that contains $\eta^{W}$. By the compactness of $\Theta$ and continuity of $M$, we have $\sup_{\eta\in\Theta\cap\mathcal{N}^{C}}M(\eta)<M(\eta^{W})$ with probability approaching one. Let $\varepsilon=M(\eta^{W})-\sup_{\eta\in\Theta\cap\mathcal{N}^{C}}M(\eta)$, then (58) implies that $\widehat{\eta}^{W}\in\mathcal{N}$ with probability approaching one. Hence, with Condition 9(iv), the consistency of the SLOPE follows from the continuous mapping theorem.
Next, we prove the asymptotic normality. It directly follows from Theorem 5.31 of Van der Vaart, (2000) that under Conditions 8-9, $\widehat{\eta}^{W}$ is asymptotically normal in that $\sqrt{n}\left(\widehat{\eta}^{W}-\eta^{W}\right)\to_{d}N\left(0,V^{W}(V^{W})^{\intercal}\right)$. Then the asymptotic normality of the SLOPE estimate follows from delta method (Theorem 3.1 of Van der Vaart, (2000)).
$\square$
I.2 Proof of the Derivation of Efficient Influence Functions
I.2.1 Preliminaries
Let $f_{P_{X}}$, $f_{Q_{X}}$, and $f_{P_{O\mid X}}$ be the density functions of the corresponding random variables on the subscripts. For a generic observation, the log-likelihood of the observed data on the joint population can be written as
$l(T,O,X)=$ $(1-T)\log f_{P_{O\mid X}}(O\mid X)+(1-T)\log f_{P_{X}}(X)+T\log f_{Q_{X}}(X)$
$+(1-T)\log(\textnormal{pr}(T=0))+T\log(\textnormal{pr}(T=1)).$
Consider the Hilbert space $\mathcal{H}$ that contains all one-dimensional zero-mean measurable functions of the observed data with finite variance. Consider $f_{P_{X}}$, $f_{Q_{X}}$, and $f_{P_{O\mid X}}$ as nuisance functions and denote their nuisance tangent spaces as $\mathcal{T}_{P_{X}}$, $\mathcal{T}_{Q_{X}}$, and $\mathcal{T}_{P_{O\mid X}}$, respectively. Then $\mathcal{H}$ can be decomposed as
$\mathcal{H}$ $=\mathcal{T}_{P_{X}}\oplus\mathcal{T}_{Q_{X}}\oplus\mathcal{T}_{P_{O\mid X}},\text{ where}$
$\mathcal{T}_{P_{X}}$ $=\left\{(1-T)a_{1}(O,X):\textnormal{E}_{P_{O\mid X}}\{a_{1}(O,X)\mid X\}=0,\textnormal{Var}_{P_{O\mid X}}\left[a_{1}(O,X)\mid X\right]<\infty\right\},$
$\mathcal{T}_{Q_{X}}$ $=\left\{Ta_{2}(X):\textnormal{E}_{Q_{X}}\{a_{2}(X)\}=0,\textnormal{Var}_{Q_{X}}\left[a_{2}(X)\right]<\infty\right\},$
$\mathcal{T}_{P_{O\mid X}}$ $=\left\{(1-T)a_{3}(X):\textnormal{E}_{P_{X}}\{a_{3}(X)\}=0,\textnormal{Var}_{P_{X}}\left[a_{3}(X)\right]<\infty\right\}.$
We consider parametric submodels $f_{P_{O\mid X}}(O\mid X,\xi_{1})$ and $f_{Q_{X}}(X,\xi_{2})$ where $\xi_{1}=0$ and $\xi_{2}=0$ correspond to the underlying truth. We also let
$\textnormal{SC}_{\xi_{1}}(O,X)=\dfrac{\partial\log f_{P_{O\mid X}}(O\mid X,\xi_{1})}{\partial\xi_{1}},\quad\textnormal{SC}_{\xi_{2}}(X)=\dfrac{\partial\log f_{Q_{X}}(X,\xi_{2})}{\partial\xi_{2}}$
be the score functions.
We revise the notation of the target functional to indicate the dependency on the nuisance parameters $\xi_{1}$ and $\xi_{2}$. Specifically, let $\psi(\xi_{1},\xi_{2})$ be the target functional. Then the efficient influence function for the target functional, $\textnormal{EIF}(T,O,X,\psi)$, satisfies
$\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{1}}\Bigg|_{\xi_{1}=0}$ $=\textnormal{E}\left[(1-T)\cdot\textnormal{EIF}(T,O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right],$ (59)
$\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}\Bigg|_{\xi_{2}=0}$ $=\textnormal{E}\left\{T\cdot\textnormal{EIF}(T,O,X,\psi)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\},$ (60)
Similarly, we revise the notation of the SLOPE and its component: let $\textnormal{SLOPE}(\xi_{1},\xi_{2})$ be the SLOPE which takes the form of
$\textnormal{SLOPE}(\xi_{1},\xi_{2})$ $=-\dfrac{\eta_{1}(\xi_{1},\xi_{2})}{\eta_{2}(\xi_{1},\xi_{2})},\text{ where}$
$\eta_{1}(\xi_{1},\xi_{2})$ $=\textnormal{E}_{Q_{O,X}^{0}}\{s(O,X,\psi)\{O-\mu(X)\},$
$\eta_{2}(\xi_{1},\xi_{2})$ $=\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}.$
Likewise, the efficient influence function for the SLOPE, $\textnormal{EIF}(T,O,X,\textnormal{SLOPE})$, satisfies
$\textnormal{E}\left[(1-T)\textnormal{EIF}(T,O,X,\textnormal{SLOPE})\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]$ $=\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\Bigg|_{\xi_{1}=0}$
$=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}\Bigg|_{\xi_{1}=0},$ (61)
$\textnormal{E}\left\{T\cdot\textnormal{EIF}(T,O,X,\textnormal{SLOPE})\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}$ $=\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\Bigg|_{\xi_{2}=0}$
$=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}\Bigg|_{\xi_{2}=0},$ (62)
I.2.2 Proof of Proposition 1
Proof of Proposition 1. Suppose the efficient influence functions for the target functional and the SLOPE take the following form,
$\textnormal{EIF}(T,O,X,\psi)=(1-T)e_{1}(O,X)+Te_{2}(X),\quad\textnormal{EIF}(T,O,X,\textnormal{SLOPE})=(1-T)a_{1}(O,X)+Ta_{2}(X),$
where $(1-T)e_{1}(O,X),(1-T)a_{1}(O,X)\in\mathcal{T}_{P_{O\mid X}}$ and $Te_{2}(X),Ta_{2}(X)\in\mathcal{T}_{Q_{X}}$. The specific forms of $e_{1}(O,X)$ and $e_{2}(X)$ are given in Proposition 2.
In light of (61) and (62), our goal is to solve the following equations for $a_{1}$ and $a_{2}$,
$\textnormal{E}\left\{(1-T)a_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$ $=\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\Bigg|_{\xi_{1}=0}$
$=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}\Bigg|_{\xi_{1}=0},$ (63)
$\textnormal{E}\left\{Ta_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}$ $=\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\Bigg|_{\xi_{2}=0}$
$=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}\Bigg|_{\xi_{2}=0}.$ (64)
We start with (63). In order to calculate the right hand side (RHS), we first calculate $\partial\eta_{1}(\xi_{1},\xi_{2})/\partial\xi_{1}$ and $\partial\eta_{2}(\xi_{1},\xi_{2})/\partial\xi_{1}$; we have
$\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\Bigg|_{\xi_{1}=0}$
$=$ $\dfrac{\partial}{\partial\xi_{1}}\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\right]$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{1}}$
$-\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\dfrac{\partial}{\partial\xi_{1}}\int Of_{P_{O\mid X}}(O\mid X,\xi_{1})dO\right]$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\textnormal{pr}(T=0)\cdot\textnormal{E}_{Q_{O,X}^{0}}\cdot\left\{\dfrac{1}{\omega(X)}e_{1}(O,X,\psi)\cdot\textnormal{SC}_{\xi}(O,X)\right\}$
$-\textnormal{E}_{Q_{O,X}^{0}}\left[\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}O\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left\{c_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\},$
where
$c_{1}(O,X)=$ $s(O,X,\psi)\{O-\mu(X)\}+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\dfrac{\textnormal{pr}(T=0)}{\omega(X)}e_{1}(O,X,\psi)$
$-\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}O,$
and
$\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{1}}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}\cdot\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\psi_{1}}+\textnormal{E}_{Q_{O,X}^{0}}\left\{\dot{s}(O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}+\textnormal{E}_{Q_{O,X}^{0}}\left\{\dot{s}(O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}\cdot\textnormal{pr}(T=0)\textnormal{E}_{Q_{O,X}^{0}}\left\{\dfrac{1}{\omega(X)}e_{1}(O,X,\psi)\cdot\textnormal{SC}(O,X,\psi)\right\}$
$+\textnormal{E}_{Q_{O,X}^{0}}\left\{\dot{s}(O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\{c_{2}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\},$
where
$c_{2}(O,X)$ $=\dfrac{\textnormal{pr}(T=0)}{\omega(X)}\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}\cdot e_{1}(O,X,\psi)+\dot{s}(O,X,\psi)$
Therefore, the RHS of (63) is
$\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{1}}$ $=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{1}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\{\eta_{2}(\xi_{1},\xi_{2})\}^{2}}$
$=\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\left\{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)\right\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}.$
Equalizing both sides of (63), we have
$0=$ $\textnormal{E}\left\{(1-T)a_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}-\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{1}}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[\dfrac{\textnormal{pr}(T=0)}{\omega(X)}a_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]$
$-\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\left\{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)\right\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right]}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[\left\{\dfrac{\textnormal{pr}(T=0)}{\omega(X)}a_{1}(O,X)-\dfrac{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right\}\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right].$
Noticing that $\textnormal{SC}_{\xi_{1}}(O,X)\in\mathcal{T}_{P_{O\mid X}}$ (and hence $\textnormal{E}_{P_{O\mid X}}\{\textnormal{SC}_{\xi}(O,X)\mid X\}=0$), we have
$\dfrac{\textnormal{pr}(T=0)}{\omega(X)}a_{1}(O,X)$
$=$ $\dfrac{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}-\textnormal{E}_{P_{O\mid X}}\left[\dfrac{-c_{1}(O,X)-\textnormal{SLOPE}(\xi_{1},\xi_{2})\cdot c_{2}(O,X)}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\Bigg|X\right].$
Therefore, by plugging in $c_{1}(O,X)$ and $c_{2}(O,X)$ at $\xi_{j}=0$ ($j=1,2$) and reorganizing terms, $a_{1}(O,X)$ can be written as
$a_{1}(O,X)$
$=$ $\dfrac{\omega(X)}{\textnormal{pr}(T=0)}\dfrac{-s(O,X,\psi)\{O-\mu(X)\}+\textnormal{Cov}_{P_{O\mid X}}[s(O,X,\psi),O\mid X]+\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\{O-\mu(X)\}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}$
$-\dfrac{\omega(X)\cdot\textnormal{SLOPE}}{\textnormal{pr}(T=0)}\dfrac{\dot{s}(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}$
$-\textnormal{SLOPE}\cdot\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}e_{1}(O,X,\psi).$
Next, we consider (64) for $a_{2}(O,X)$. For the RHS of (64), we calculate $\partial\eta_{1}(\xi_{1},\xi_{2})/\partial\xi_{2}$ and $\partial\eta_{2}(\xi_{1},\xi_{2})/\partial\xi_{2}$; we have
$\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{2}}=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\cdot\textnormal{SC}_{\xi_{2}}(X)\right]+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}$
$=$ $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi)\{O-\mu(X)\}\cdot\textnormal{SC}_{\xi_{2}}(X)\right]$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\textnormal{pr}(T=1)\textnormal{E}_{Q_{X}}\{e_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\}$
$=$ $\textnormal{E}_{Q_{X}}\{b_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\},\text{ where}$
$b_{2}(X)=$ $\textnormal{E}_{P_{O\mid X}}\left[s(O,X,\psi)\{O-\mu(X)\}\mid X\right]$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]\cdot\textnormal{pr}(T=1)\cdot e_{2}(X),$
and
$\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{2}}=$ $\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}\cdot\textnormal{SC}_{\xi_{2}}(X)\right]+\textnormal{E}_{Q_{O,X}^{0}}\left[\ddot{s}(O,X,\psi)\right]\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}$
$=$ $\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}\cdot\textnormal{SC}_{\xi_{2}}(X)\right]$
$+\textnormal{E}_{Q_{O,X}^{0}}\left[\ddot{s}(O,X,\psi)\right]\textnormal{pr}(T=1)\textnormal{E}_{Q_{X}}\{e_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\}$
$=$ $\textnormal{E}_{Q_{X}}\{b_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\},\text{ where }$
$b_{2}(X)=$ $\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}+\textnormal{E}_{Q_{O,X}^{0}}\left[\ddot{s}(O,X,\psi)\right]\textnormal{pr}(T=1)e_{2}(X).$
Therefore, the RHS of (64) can be written as
$\dfrac{\partial\textnormal{SLOPE}(\xi_{1},\xi_{2})}{\partial\xi_{2}}$ $=-\dfrac{\dfrac{\partial\eta_{1}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{2}(\xi_{1},\xi_{2})-\dfrac{\partial\eta_{2}(\xi_{1},\xi_{2})}{\partial\xi_{2}}\cdot\eta_{1}(\xi_{1},\xi_{2})}{\left\{\eta_{2}(\xi_{1},\xi_{2})\right\}^{2}}$
$=\textnormal{E}_{Q_{X}}\left\{\left[\dfrac{-b_{1}(X)-b_{2}(X)\cdot\textnormal{SLOPE}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right]\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}.$
Equalizing both sides of (64), we have
$0=$ $\textnormal{E}_{Q_{X}}\left[\left(\textnormal{pr}(T=1)a_{2}(X)-\left[\dfrac{-b_{1}(X)-b_{2}(X)\cdot\textnormal{SLOPE}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right]\right)\cdot\textnormal{SC}_{\xi_{2}}(X)\right].$
Since $\textnormal{SC}_{\xi_{2}}(X)\in\mathcal{T}_{Q_{X}}$ and hence $\textnormal{E}_{Q_{X}}\{\textnormal{SC}_{\xi_{2}}(X)\}=0$, we have
$a_{2}(X)=$ $\dfrac{\left[\dfrac{-b_{1}(X)-b_{2}(X)\cdot\textnormal{SLOPE}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right]-\textnormal{E}_{Q_{X}}\left\{\left[\dfrac{-b_{1}(X)-b_{2}(X)\cdot\textnormal{SLOPE}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}\right]\right\}}{\textnormal{pr}(T=1)}$
$=$ $\dfrac{-\textnormal{SLOPE}\cdot\textnormal{E}_{P_{O\mid X}}\{\dot{s}(O,X,\psi)\mid X\}-\textnormal{E}_{P_{O\mid X}}\{O-\mu(X)\mid X\}}{\textnormal{pr}(T=1)\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}$
$-e_{2}(X)\cdot\dfrac{\textnormal{E}_{Q_{O,X}^{0}}\left[\dot{s}(O,X,\psi)\{O-\mu(X)\}\right]+\textnormal{SLOPE}\cdot\textnormal{E}_{Q_{O,X}^{0}}\{\ddot{s}(O,X,\psi)\}}{\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}.$
Finally, by plugging in $a_{1}(O,X)$ and $a_{2}(X)$ into $\textnormal{EIF}(T,O,X,\textnormal{SLOPE})=(1-T)a_{1}(O,X)+Ta_{2}(X)$ and noticing $(1-T)e_{1}(O,X)+Te_{2}(X)=\textnormal{EIF}(T,O,X,\psi)$, we obtain the EIF stated in Proposition 1. $\square$
I.2.3 Proof of Proposition 2
Proof of Proposition 2. Suppose the efficient influence function of the target functional is
$\textnormal{EIF}(T,O,X,\psi)=(1-T)e_{1}(O,X)+Te_{2}(X),$
where $(1-T)e_{1}(O,X)\in\mathcal{T}_{P_{O\mid X}}$ and $Te_{2}(X)\in\mathcal{T}_{Q_{X}}$. Then (59)-(60) can be re-expressed as
$\textnormal{E}\left\{(1-T)e_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$ $=\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{1}},$ (65)
$\textnormal{E}\left\{Te_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}$ $=\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}.$ (66)
First, consider (65). The LHS is
$\textnormal{E}\left\{(1-T)e_{1}(O,X)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}$ $=\textnormal{E}_{Q_{O,X}^{0}}\left[\dfrac{\textnormal{pr}(T=0)}{\omega(X)}e_{1}(O,X)\right],$
and the RHS is
$\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{1}}$ $=\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\textnormal{E}_{Q_{O,X}^{0}}\left\{s(O,X,\psi)\cdot\textnormal{SC}_{\xi_{1}}(O,X)\right\}.$
We equalize them and then have
$e_{1}(O,X)=-\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\dfrac{\omega(X)}{\textnormal{pr}(T=0)}\left[s(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\right].$
Next, consider (66). The LHS is
$\textnormal{E}\left\{Te_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}$ $=\textnormal{pr}(T=1)\textnormal{E}_{Q_{X}}\left\{e_{2}(X)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\},$
and the RHS is
$\dfrac{\partial\psi(\xi_{1},\xi_{2})}{\partial\xi_{2}}$ $=\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\textnormal{E}_{Q_{O,X}^{0}}\left\{s(O,X,\psi)\cdot\textnormal{SC}_{\xi_{2}}(X)\right\}.$
We equalize them and have
$e_{2}(X)=-\dfrac{\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}-\textnormal{E}_{Q_{O,X}^{0}}\{s(O,X,\psi)\}}{\textnormal{pr}(T=1)\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}}.$
Therefore, by plugging in $e_{1}(O,X)$ and $e_{2}(X)$ to the EIF of the target functional, we have
$\textnormal{EIF}(T,O,X,\psi)=$ $-\dfrac{(1-T)\omega(X)}{\textnormal{pr}(T=0)}\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\left[s(O,X,\psi)-\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}\right]$
$-\dfrac{T}{\textnormal{pr}(T=1)}\left[\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi)\}\right]^{-1}\textnormal{E}_{P_{O\mid X}}\{s(O,X,\psi)\mid X\}.$
$\square$
Appendix J Extended Remarks
This section details some extended remarks deferred from the main text. Section J.1 discusses vector valued SLOPE. Section J.2 includes some extended remarks from Section 6. Section J.3 discusses defining SLOPE for other types of conditional exchangeability assumptions, including the no unmeasured confounding assumption and the missing at random assumption. Section J.4 details Remark 2 on the challenge of extending the notion of SLOPE to other bound-based sensitivity models. Section J.5 states the mathematical connection between SLOPE and the marginal interventional effect (Zhou and Opacic,, 2022) with the incremental propensity score intervention (Kennedy,, 2019).
J.1 SLOPE for a Vector Valued $\psi(\cdot)$
When the functional $\psi(\cdot)$ is vector valued with dimension $p$, the corresponding SLOPE, as defined in Definition 1, is also a vector of $p$ elements. Each dimension of SLOPE represents the robustness of the corresponding element of the target estimand when conditional exchangeability is “near” violated. For example, when $X$ is $p$-dimensional and the target estimand is a $p$-dimensional vector of ordinary least squares (OLS) coefficients of regressing $O$ on $X$, then each element of SLOPE describes the robustness of the corresponding coefficient; see Section C.5 for a formal presentation of the SLOPE and Section E.5 for the estimation. More generally, for SLOPE as vectors, the connection between SLOPE and IF still holds (i.e., Theorem 3). In addition, the two proposed estimators are still applicable, as mentioned in Remarks 4 and 5.
The main difference between a vector of SLOPE and a scalar SLOPE is on the interpretation, and in particular, the informed guidance in designs. Specifically, as mentioned in Section 3.1, when SLOPE becomes a vector and all elements are of scientific interest, then practitioners need to find an appropriate summary of the vector in order to compare SLOPEs across study designs. This could be a visual summary, such as plotting the SLOPE of each dimension, or a quantitative measure such as a norm of the SLOPE.
Example 6 (SLOPE for OLS Coefficient).
Suppose $X$ is a $p$-dimensional vector and $O$ is an outcome variable of interest. The target estimand is defined as the regression coefficients of regressing $O$ on $X$. In other words, consider $\psi^{{\textnormal{OLS}}}$ such that
$\textnormal{E}_{{\textnormal{OLS}}}\left(XX^{\intercal}\psi^{{\textnormal{OLS}}}-XO\right)=0,$
which we suppose the uniqueness of $\psi^{{\textnormal{OLS}}}(Q_{O,X})$. Following Section C.5, we know the SLOPE for $\psi^{{\textnormal{OLS}}}$ is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$ $=\left\{\textnormal{E}(XX^{\intercal})\right\}^{-1}\textnormal{E}_{Q_{X}}\left\{\sigma^{2}(X)X\right\},$
which is a $p$-dimensional vector.
To compare SLOPE across designs, we define the magnitude of SLOPE in two scenarios. First, suppose $X$ has been standardized within the source and the target population. Then we can define the norm of SLOPE as the $L_{2}$ norm of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$. With this notion of the magnitude, interpretations of SLOPE based on its magnitude that were discussed in Section 4.1 are applicable. Second, suppose $X$ has not been standardized, then we consider the Mahalanobis distance with covariance matrix $\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\}^{-1}$. More specifically, we define the magnitude of SLOPE as
$\left\{\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})\right\}^{\intercal}\textnormal{E}_{Q_{X}}(XX^{\intercal})\left\{\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})\right\}$
$=$ $\left[\textnormal{E}_{Q_{X}}\left\{\sigma^{2}(X)X\right\}\right]^{\intercal}\left\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\right\}^{-1}\textnormal{E}_{Q_{X}}\left\{\sigma^{2}(X)X\right\}.$
In practice, the SLOPE needs to be estimated. Example estimators of the SLOPE for OLS coefficients are provided Section E.5.
J.2 Extended Remarks from Section 6
For the mean, the dependency of its SLOPE on the conditional variance $\sigma^{2}(X)$ advocates source populations with a lower $\sigma^{2}(X)$; this principle holds generally for location parameters in that SLOPE advocates a more homogeneous design. Since $\sigma^{2}(X)$ is smaller when $X$ contains less information, SLOPE seems to advocate a homogeneous $X$ (e.g., $X=x$ almost surely and therefore $\sigma^{2}(X)=0$), which contradicts with the existing understanding that $X$ should be sufficiently rich for the overlap condition (Assumption 1) to hold. To understand why such contradiction appears, we elaborate on the comments noted in Section 6.
First, SLOPE is proposed and applied in cases when the overlap condition holds, meaning that covariates in the source population cannot contain less information than the covariates in the target population. This aligns with most works in sensitivity analysis for the conditional exchangeability assumption in different contexts. We refer readers to Huang, (2025) for a sensitivity analysis solely for the overlap condition. Considering violations to both assumptions is a valuable future direction (see Bonvini and Kennedy, (2022) and Cui and Li, (2025)) and is beyond the scope of this work.
Second, as mentioned in the main text, we echo Tipton and Olsen, (2018) and Degtiar and Rose, (2023) in the importance of a careful data collection for the conditional exchangeability to hold, or, for the violation to be as small as possible. In this sense, it’s more preferable to collect more common characteristics in source and target populations, namely a richer $X$. Secondary to that, SLOPE is a useful tool to assess sensitivity/robustness when it is unrealistic to meet the conditional exchangeability with the observed set of $X$. This happens, for example, when (i) it is infeasible to randomize units into the source or target population, (ii) not all covariates can be measured in both populations, which is common if the investigator opts in after data collection, and/or (iii) there exist unobservable differences between populations (e.g., sites) even under a careful design (Allcott,, 2015; Jin et al.,, 2024).
J.3 SLOPE in Other Contexts
While this paper focuses on the conditional exchangeability assumption in transportability/generalizability, SLOPE can be used to study the sensitivity of other types of conditional exchangeability typed assumptions. With different meanings of $P$ and $Q$, SLOPE can measure the sensitivity/robustness of other types of conditional exchangeability assumptions, including the no unmeasured confounding assumption in causal inference and the missing at random assumption in missing data. We describe these settings in the following subsections.
J.3.1 SLOPE for Unmeasured Confounding
Let $A\in\{0,1\}$ be the treatment and $Y(A)$ be the potential outcome under treatment $A$, where $A=1$ means intervention and $A=0$ means control. We suppose SUTVA (Assumption 4) holds; i.e., when $A=a$, the observed outcome $Y$ is $Y(a)$, for $a=0,1$. To fix ideas, suppose we are interested in the average treatment effect on the treated, a popular estimand in causal inference,
$\psi^{\textnormal{ATT}}$ $=\textnormal{E}\{Y(1)-Y(0)\mid A=1\}$
$=\textnormal{E}\{Y(1)\mid A=1\}-\textnormal{E}\{Y(0)\mid A=1\}$
$=\textnormal{E}\{Y\mid A=1\}-\textnormal{E}\{Y(0)\mid A=1\}.$ (67)
where the third line follows from SUTVA. From (67), the key challenge in identifying $\psi^{\textnormal{ATT}}$ lies in the challenge of identifying the second term, $\textnormal{E}\{Y(0)\mid A=1\}$, since it involves the potential outcome $Y(0)$. One common strategy for identifying $\psi^{\textnormal{ATT}}$ is through the conditional exchangeability assumption (Assumption 2) and the overlap condition (Assumption 1). To demonstrate that, we define $P$ and $Q$ as follows.
Suppose $P$ represents the population of units with $A=0$ and $Q$ represents the population of units with $A=1$. Therefore, $\textnormal{E}_{P}(\cdot)=\textnormal{E}(\cdot\mid A=0)$ and $\textnormal{E}_{Q}(\cdot)=\textnormal{E}(\cdot\mid A=1)$. Suppose $X$ contains pre-treatment covariates (i.e., measured confounders) and $O=Y(0)$ is the potential outcome under condtrol. Then the conditional exchangeability assumption (Assumption 2), i.e., $Q_{O\mid X}(\cdot\mid x)=P_{O\mid X}(\cdot\mid x)$ almost everywhere $Q_{X}$, implies that
$\left(\textnormal{E}_{Q_{O\mid X}}[O\mid X]:=\right)\quad\textnormal{E}\{Y(0)\mid X,A=1\}=\textnormal{E}\{Y(0)\mid X,A=0\}\quad\left(:=\textnormal{E}_{P_{O\mid X}}[O\mid X]\right).$
Then $\psi^{\textnormal{ATT}}$ in (67) can be identified, since
$\textnormal{E}\{Y(0)\mid A=1\}$ $=\textnormal{E}\left[\textnormal{E}\{Y(0)\mid X,A=1\}\mid A=1\right]$
$=\textnormal{E}\left[\textnormal{E}\{Y(0)\mid X,A=0\}\mid A=0\right]$ (by conditional exchangeability)
$=\textnormal{E}\left[\textnormal{E}\{Y\mid X,A=0\}\mid A=0\right]$ (by SUTVA)
no longer involves potential outcomes.
In this context, the conditional exchangeability assumption is also referred to as no unmeasured confounding. In observational studies where unmeasured confounding is highly likely, there is an extensive literature that develops sensitivity analysis methods to study the consequence of unmeasured confounding. With the above defined $P$ and $Q$, SLOPE can be naturally applied to study the sensitivity of violation to no unmeasured confounding. For example, the SLOPE for $\psi^{\textnormal{ATT}}$ as defined above is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{ATT}})$
$=$ $\textnormal{E}_{Q_{X}}\textnormal{Var}_{P_{O\mid X}}(O\mid X)$ (by Theorem 1)
$=$ $\textnormal{E}_{X\mid A=1}\left[\textnormal{Var}_{Y(0)\mid X,A=0}(Y(0)\mid X,A=0)\mid A=1\right]$ $\text{(by definitions of }P\text{ and }Q\text{)}.$
Remark 6 (SLOPE for Other Causal Estimands).
We note that $P$ and $Q$ may need to be defined differently for other causal estimands. For example, when the target estimand becomes the average treatment effect on the control, then the roles of $P$ and $Q$ need to be swapped. Nevertheless, at a high level, the definition of SLOPE remains consistent, i.e., the derivative of the target causal estimand with respect to $\gamma$ at $\gamma=0$, where $\gamma$ is such that
$\dfrac{[(Y(a)\mid X,A=1-a]}{[Y(a)\mid X,A=a]}\propto\exp\{\gamma\cdot Y(a)\}.$ (68)
where we use $\cdot\mid X,A]$ to represent conditional densities (provided exist with respect to some measure) on the underlying population. For more explanations on the sensitivity model (68), see Scharfstein et al., (1999) and Franks et al., (2020).
Finally, we make a clarification on this setting. The current section should not be confused with Section D. In Section D, we transport treatment effect from a source population to a target population where we have assumed that the treatment effect can be identified within the source; in other words, the “causal” assumptions already hold in the source population. However, in this section, we no longer consider any transfer learning setting but instead focus on a causal setting with potential unmeasured confounding. Here, one may still think of $P$ (i.e., population of control units) as the “source” population and $Q$ (i.e., population of treated units) as the “target” population, while keeping in mind the difference in the meaning of conditional exchangeability assumption as well as the different interpretations of the corresponding SLOPEs.
J.3.2 SLOPE for Non-Ignorable Missingness
Consider a missing data problem where we have observed covariates $X$ on all units, but only observe the outcome variable $O$ on a subset of the units. Let $P$ be the population of units with complete data and $Q$ be the population of units with a missing $O$. Then the conditional exchangeability assumption (Assumption 2) is the missing at random assumption which is commonly adopted in the literature of missing data. In this context, the SLOPE can be defined similarly as the main text and it can be interpreted as the sensitivity/robustness of the target estimand with respect to non-ignorable missingness.
J.4 Remark on Extending SLOPE to Bound-Based Sensitivity Analysis Models
This section provides details of Remark 2 in the main text.
In sensitivity model (3), the sensitivity parameter $\gamma$ quantifies the degree of violation in a parametric way and it elicits a point identification of the target estimand. In contrast, an important line of sensitivity analyses uses the sensitivity parameter(s) to bound the difference between the unobserved distribution (i.e., $Q_{O\mid X}$) and the observed distribution (i.e., $P_{O\mid X}$) nonparametrically and obtain a set identification of the target estimand. A natural question is whether we may extend the concept of SLOPE to sensitivity analyses based on bounds. Unfortunately, this will require a non-trivial extension to the notion of derivatives. We illustrate the challenge through an example below and leave such extensions as important future directions.
We consider a simplified version of the setting in Zeng et al., (2023) and their sensitivity model. Suppose the target estimand is the mean, $\psi^{{\textnormal{mean}}}(Q_{O,X})=\textnormal{E}_{Q_{O,X}}(O)$. Let $\gamma$ be the sensitivity parameter and suppose the target conditional distribution $Q_{O\mid X}^{\textnormal{bias}}$ deviates from the source conditional distribution $P_{O\mid X}$ so that the bias in the conditional mean is no larger than $\gamma$, as sated in (15). Consequently, the target estimand can be bounded as
$-\gamma+\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right]\leq\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{bias}})\leq\gamma+\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right].$ (69)
For convenience of notation, here we still use $\gamma$ to represent the sensitivity parameter. We use $Q_{O\mid X}^{\textnormal{bias}}$ and $Q_{O,X}^{\textnormal{bias}}$ to denote conditional and joint distributions under sensitivity model (15), bearing in mind that they are not unique since the sensitivity model (3) does not directly identifies/bounds the distribution (but rather, the bias in conditional means). We use $\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{bias}})$ to denote target estimand under the sensitivity model.
Next, we show that it’s not natural to define an analogy to SLOPE. Suppose we focus on the lower bound of (69) where
$\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{L,}\gamma})=-\gamma+\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right].$
Then the derivative of the target estimand $\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{L,}\gamma})$ with respect to $\gamma$ at $\gamma=0$ is 1-1, since
$\lim_{\gamma\to 0}\dfrac{\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{L,}\gamma})-\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{L,}0})}{\gamma}=\lim_{\gamma\to 0}\dfrac{-\gamma}{\gamma}=-1.$
Similarly, consider the upper bound of (69) where
$\psi^{{\textnormal{mean}}}(Q_{O,X}^{\textnormal{R,}\gamma})=\gamma+\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right],$
then the derivative is $1$. The two special cases demonstrate the non-uniqueness of the derivative of the target estimand with respect to $\gamma$ in the set of target estimands (i.e., (69)) determined by the sensitivity model (15). Consequently, generalizing the notion of SLOPE to bound-based sensitivity models like (15) is non-trivial.
J.5 Connections to Marginal Interventional Effect
Broadly speaking, the concept and the mathematical format of the SLOPE are connected to the marginal interventional treatment effect (Zhou and Opacic,, 2022) with the incremental propensity score intervention (Kennedy,, 2019). In this section, we elaborate on this connection.
We begin with a brief introduction to the marginal interventional treatment effect; see Zhou and Opacic, (2022) for details. In causal inference (see notation in Section D.1), an interventional effect quantifies the change in outcome when the propensity of a proportion of units changes. This is in contrast to conventional estimands, like the average treatment effect and the quantile treatment effect, which compare the outcome when all units receive treatment with the outcome when all receive placebo. To define an interventional effect, suppose the propensity score is $\pi(Z)=P(A=1\mid Z)$ and consider changing the propensity in a way of $\pi^{\gamma}(Z)$ as a function of $\pi(Z)$ with $\pi^{0}(X)=\pi(Z)$. Then under assumptions in Section 3 of Zhou and Opacic, (2022), the interventional effect (IE) is
$\text{IE}^{\gamma}$ $=\textnormal{E}\left[\dfrac{\pi^{\gamma}(Z)-\pi^{0}(Z)}{\textnormal{E}\left[\pi^{\gamma}(X)-\pi^{0}(Z)\right]\tau(X)}\right],$
where $\tau(Z)=\textnormal{E}\{Y(1)-Y(0)\mid X\}$ is the conditional average treatment effect (CATE). The marginal interventional effect (MIE), as proposed by Zhou and Opacic, (2022), describes the marginal change in the interventional effect by taking the limit of $\gamma$ of IE going to zero:
MIE $=\textnormal{E}\left[\dfrac{\dot{\pi}^{0}(Z)}{\textnormal{E}\left\{\dot{\pi}^{0}(Z)\right\}}\tau(Z)\right],$
where $\dot{\pi}^{0}(X)$ is the derivative of $\pi^{\gamma}(Z)$ with respect to $\gamma$ at $\gamma=0$. When the propensity score shift $\pi^{\gamma}(Z)$ is determined by the incremental propensity score interventions (IPSI) proposed by Kennedy, (2019), i.e.,
$\dfrac{\pi^{\gamma}(Z)/\{1-\pi^{\gamma}(Z)\}}{\pi^{0}(Z)/\{1-\pi^{0}(Z)\}}=\exp(\gamma),$ (70)
the MIE is
$\text{MIE}^{\text{IPSI}}$ $=\dfrac{\textnormal{E}\left[\pi^{0}(Z)\{1-\pi^{0}(Z)\}\tau(Z)\right]}{\textnormal{E}\left[\pi^{0}(Z)\{1-\pi^{0}(Z)\}\right]}.$ (71)
The MIE with IPSI is connected to SLOPE for two reasons. First, the IPSI shifts the intervention $\pi^{\gamma}(Z)$ in the same exponential tilting model as in the sensitivity model (3); see below for details. Second, the concept of MIE, which defines a treatment effect in terms of the change in the outcome with an infinitesimal change in propensity, is broadly connected to the concept of SLOPE. To see the connections more explicitly, we next re-express the SLOPE under the regime of incremental propensity score intervention. We note that although the resulting quantity no longer represents the “sensitivity to local perturbation from the exchangeability”, we still refer to it as SLOPE for ease of communication, keeping in mind that we temporarily treat SLOPE as a mathematical quantity instead of a sensitivity/robustness measure.
Let $O=A$ be the binary treatment. Suppose $P$ is the population of interest and $P_{Z,A,Y}$ is the joint distribution of pre-treatment covariate $Z$, binary treatment $A$ and outcome $Y$. We adopt the potential outcome framework (Section D.1) and suppose SUTVA (Assumption 4) and strong ignorability (Assumption 3) hold.
Let $Q$ be a hypothetical population with the same distribution as $P$ except that the intervention mechanism has been changed by $\gamma$. In specific, let $O=A$, then the sensitivity model (3) implies that the odds of receiving treatment in $Q$ is $\exp(\gamma)$ times the odds of receiving treatment in $P$:
$\exp(\gamma)=\dfrac{Q^{\gamma}(A=1\mid Z)/\{1-Q^{\gamma}(A=1\mid Z\}}{P(A=1\mid Z)/\{1-P(A=1\mid Z)\}}.$
Note that in this case, $X=(A,Z)$ includes the intervention and the pre-treatment covariates. We let $\pi^{\gamma}(Z)=Q^{\gamma}(A=1\mid Z)$ and therefore $\pi^{0}(Z)=Q^{0}(A=1\mid Z)=P(A=1\mid Z)$ and $\pi^{\gamma}(Z)=\dfrac{\exp(\gamma)\pi^{0}(X)}{1-\pi^{0}(Z)+\exp(\gamma)\pi^{0}(Z)}$.
To define SLOPE, let $Q^{\gamma}_{Z,A,Y}=P_{Z}\times Q^{\gamma}_{A\mid Z}\times P_{Y\mid Z,A}$ and let the target functional be the mean outcome, i.e.,
$\psi(Q_{O,X}^{\gamma})$ $=\textnormal{E}_{Q_{Y}^{\gamma}}(Y)$
$=\textnormal{E}_{Q_{X}}\left[\pi^{\gamma}(Z)\textnormal{E}_{P_{Y\mid Z,A=1}}\{Y\mid Z,A=1\}+\{1-\pi^{\gamma}(Z)\}\textnormal{E}_{P_{Y\mid Z,A=0}}\{Y\mid Z,A=0\}\right]$
$=\textnormal{E}_{Q_{X}}\left[\pi^{\gamma}(Z)\tau(Z)\right]+\textnormal{E}_{Q_{X}}\left[\left\{\textnormal{E}_{P_{Y\mid Z,A=0}}(Y(0)\mid Z,A=0\right\}\right],$
where again $\tau(X)$ is the CATE.
Noticing $\partial\pi_{\gamma}(X)/\partial\gamma|_{\gamma=0}=\pi^{0}(Z)\{1-\pi^{0}(Z)\}$, we have the SLOPE as
$\textnormal{E}_{Q_{X}}\left[\dfrac{\partial\pi^{\gamma}(Z)}{\partial\gamma}\Bigg|_{\gamma=0}\tau(Z)\right]$ $=\textnormal{E}_{Q_{X}}\left[\pi^{0}(Z)\{1-\pi^{0}(Z)\}\tau(X)\right].$ (72)
We make two remarks on (72). First, it resembles the general form of the SLOPE of an outcome mean, since
$\textnormal{E}_{Q_{X}}\left[\pi^{0}(X)\{1-\pi^{0}(X)\}\tau(Z)\right]=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}[A,Y(1)-Y(0)\mid Z]\right\},$
where the $\textnormal{Cov}[\cdot\mid X]$ represents conditional covariance under $P$. Note that here SLOPE is the expectation of a conditional covariance instead of a conditional variance because the “shift” is on the treatment (i.e., $O=A$) rather than the outcome. Second, (72) can be viewed as the non-standardized version of the MIE in (71). It is also a non-standardized version of the average treatment effect for the overlap population (ATO) (Li et al.,, 2018) since the two quantities (ATO and MIE) are identical in this specific regime (Zhou and Opacic,, 2022).